A Complete Bibliography of Electronic Communications in Probability

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

20 October 2023 Version 1.15

Title word cross-reference

(1 + 1) [CB10]. (d, α, β) [Zho10]. (∇ + Δ) [CB10]. 0 [Sch12, Wag16]. 1 [Duc19, HL15b, Jac14, Li14, Sch12, SK15, Sim00, Uch18]. 1/2 [KV15]. 1/4 [JPR19]. 2 [BDT11, GH18b, Har12, Li14, RSS18, Sab21, Swa01, VZ11, vdBN17]. 2D [DXZ11]. 2M − X [Bau02, HMO01, MY99]. 3 [AB14, PZ18, SK15]. [0, t] [MLV15]. α [BGT21, DXZ11, Pat07]. α ∈ [0, 1/2) [Sch12]. BES_{0}(d) [Win20]. β [AS22, Ven13]. d [Häg02, Mal15, Van07, Zho20]. d = 2 [KO06]. d > 1 [Sal15]. d_{l_{2}} [Wan14]. d ≥ 2 [BR07]. d ≥ 3 [ST20]. d ∈ (0, 1) [Win20]. E = Z ⊗ E [MM21]. f [DGG^{+}13]. \frac{∂u}{∂t} = κ_{m} \frac{∂^{m}u}{∂x^{m}} [OD12]. G [NY09, BCH^{+}00, FGM11]. H [Woj12, WP14]. k [AV12, BKR06, Gao08, GRS03, Wan21]. k(n) [dBJP13]. k^{α} [Sch12]. L^{1} [CV07, LMV22, MR01]. L^{1}([0, 1]) [FP11]. L^{2} [HN09]. l^{∞} [MHC13]. L^{p} [CGR10]. L_{1} [EM14]. Λ [Fou13, Fou14, Lag07, Zho14]. Lu = u^{α} [Kuz00]. m(n) [dBJP13]. C^{n} [Tko11]. R^{2} [Kri07]. R^{d} [MN09]. Z [Sch12]. Z^{2} [Gla15]. Z^{d} [DV23, DP14, SBS15, BC12]. Z_{n}^{d} [ST20]. R^{2} [HL15a]. R^{3} [Far98].

Z^{d} [BS96]. C^{∞} [DCF06]. N [AGN21, Han21]. N × N × 2 [BF11]. p [Eva06, GL14, Man05]. p_{c} < p_{u} [NP12a]. ϕ [Kie97]. Φ_{2}^{2n} [Hua21]. q [Bar14]. r [SC09]. S [RS07]. S_{2}(δ) [Ost14]. SU(3) [Ras10]. T [IR10, GG04, Gri02, MLV15]. T^{2} [DS06]. U [Arc98, RW02]. \underline{□}_{λ} [Loe13]. Ξ [Fre12]. Z [DJ06, RS11a, Sin14]. Z^{d} [Pet08, Zer06]. Z_{+}^{2} [LK08]. ζ(2) [Wäs09]. ζ(2n) [BFY07]. Z × {0, 1} [HLSX21].

* [KNN15].

-adic [Eva06]. -balayages [SC09]. -coalescent [Möh18]. -coalescents [Fre12, Lag07]. -connected [AB14]. -cut [Wan21]. -D [BDT11]. -dependent [HL15b]. -dimensional [CB10, Häg02, Jac14, Mal15, Van07, Zho20, Swa01]. -distance [EM14]. -divisible [AV12]. -Expectation [BCH^{+}00]. -Hahn [Bar14]. -interaction [CB10]. -Martin [IR10]. -measures [FGM11]. -metric [Wan14, Loe13]. -mixing [Kie97]. -Monotone [Gao08]. -Norm [MR01]. -projective [LMV22]. -selfadjoint [WP14]. -selfadjont [Woj12]. -smooth [VZ11]. -spaces [CV07]. -spine [RSS18]. -Stable [Pat07, BGT21, DXZ11]. -Statistic [GG04, Gri02]. -Statistics [DS06, Arc98, RW02]. -step [GRS03]. -superprocess [Zho10]. -th [Han21]. -transform [RS07]. -valued [BC12, MHC13]. -Variation [Man05]. -variations [GL14]. -vectors [DGG^{+}13]. -Watanabe [Cri21]. -wise [BKR06].

10 [MZ05a].

2D [Kis14, vdBJV07, vdBC12, vdBC13, GPL08]. 2d-random [GPL08].

A. [KDV17]. a.k.a. [Jon13]. a.s [KSY06]. Abelian [Pri17]. Above [Jan97]. absence [IS17]. Absolute [Arc98, DM18, DGK19, SSS15, SV11b, Wag16]. Absolutely [Mac02, Mik02, AF19]. absorption [HH07]. Account [Ros02]. accuracy [MR13]. achieved [MLV15]. Acknowledgement [PN16]. Acknowledgment [MZ05a]. Activated [GGA10]. adaptive [AA07]. add [Tri19]. add-one [Tri19]. additive [EM14, NX13, Pel20]. adic [Eva06]. adjacency [BS07a]. admits [BT17]. ageing [BP10]. Aggregation [Hus08, DCLYY13]. Airy [BL13, CS14, TW03]. Aldous [BP10, War99]. algebra [LR16]. algebraic [AF06]. Algebraically [BFT13]. Algorithm [Mac02]. algorithms [AA07, RR15, RU13]. allocation [FGG21, Kri07]. allocations [HM15]. Almeida [Che21a]. Almost [App02, AN19, BS18a, Fre12, LZ23, Lin09, Res01, Zho10, HR14, Li20]. Almost-sure [BS18a]. along [Hil12, KT11, NS13]. always [FGM11]. amenable [AST14, Hus08, NP12a, Tim19]. amnesia [Lau22]. among [HK16, KSW12, McV08]. Analysis [KLS05, Kub11, FLP20]. analytic [BT20, DP18a, Unt10, BT21]. Analyticity [PS20]. Anchored [Dem20]. Andersen [ES21]. Anderson [BCM22, Duc19, Dun20, GH18b, HL15a, MNX20]. Animals [Ham05, MS11]. Annealed [CG05, VM13]. annihilating [BK11, ST17, TYZ12]. Anomalous [Buc13]. Answers [BS96]. anti [Zha22]. anti-derivative [Zha22]. anticipated [YE13]. antisymmetry [Blo23]. Application [DS10, HCA17, Ruf15b, BP10, Car22, CP19b, CK12, Gau16, Jac23, JK13, Sai07]. Applications [GG11, Jan97, Li99, BKS16, MU10, Pri09, TM15, dlPP09]. Approach [DZ96, Lon04, BC15, Ber17, DW12, DL09a, GH20, HvdHS08, Lac10, Lâm22, Lau20, Led17, LS13b]. approaching [DC13]. Approximating [BBF18, YLW15]. Approximation [BZ06, DP18b, DL08, KS97, BRT10, BC15, BJ18, BN08, CK12, Dal13, Dal17, FT07, FM12b, JTT18, JPR19, KV13, LY13, McV22, MR13, Pri15, SŽ17, Sio14, Tou22, VM13, WYY13]. Approximations [Pec07, CNPP16, Fan16, FM12a, KDV17, MZ18, Sab13]. Arbitrage [SV11a]. Arbitrage-free [SV11a]. Arbitrary [JK04, Bec22, KK15]. arboreal [Eas22]. arc [VY12a]. arc-sine [VY12a]. arcsine [Pan21]. area [Tod22]. arise [MAPS14]. arising [Che21c, KNN15]. Arithmetic [BYZ07, BYZ12]. Arrays [BL10, CK12, MAPS14, Van07]. assembly [GM22]. asset [Rok07]. assignment [Wäs09]. Associated [BL10, Gao03, Mar10, Tan06, FF12, HRKU11, Haj15, HCS08, Ili19, Lab13, MR15a]. association [Dal13]. assumption [Ose11]. asymmetric [GRS03, MPP15]. Asymptotic [BB07, BDN10, CY13, DHR18, DHI11, DC15, EM16, FH22, GWZ21, Kar08, Rio11, SB07, Spr07, Tuc11, BS07a, GMT15, Lan22, MP16, BB22]. asymptotically [GN06, Zha12]. Asymptotics [HM14, Jun11b, Li20, RW02, Rev03, Aid10, ADS20, Aou21, BS18a, Che21c, DW12, EP17, Fre12, Fuk09, Jac14, Jeg09, Maś18, Mec09, PR15, RRZ11, Thi20, Wat12, Yin15, de 06]. Atlas [Tsa18]. attachment [CJ13, HJ18, MP14a, Mal20, Tam07]. Attracted [PV05, Uch18]. attracting [Gau16]. attraction [LC22]. average [BL13]. averaged [LY16]. averages [LPP15]. averaging [BHM21, CD17a, Ruf15b, Wai13]. Avoid [ABV03]. Avoidance [AHM^{+}13]. avoiding [BP21, Gla15, Hut18]. away [Eri16]. Axis [BM05]. Azéma [Çet12]. Azuma [Rio13a].

backward [AED13, Ban15, CD13, Duc19, Owo15, YE13, YRE16]. Baik [Zha22]. Baik-Rains [Zha22]. Balanced [BB10, FGG21, Ste13]. balancing [JK13]. balayages [SC09]. Ball [Gao08, Li99, RŻ98, Nua18, Sap10]. Ballistic [AAK01, Zer02, Flu08]. Balls [CP20, Jun11b, BBMT09, CP19a]. Banach [CK18a, Cha10, Kie97, MN08, VZ11]. band [BGP14, Kar09, Tod22]. band-limited [Tod22]. bank [Nie22]. Barbour [KDV17]. Barnes [NY09, Ost13]. Barrier [GKH03, BS18b, MV14]. barycentric [Hou09]. based [BZ16, CL14, Rey15]. be [BDKS19, FF12]. Becker [HY21]. becomes [Liu22]. before [MLV15]. Behavior [Lim99, CY13, FJR20, HK17, LM17, Tod22, Xio04]. behaviors [KP22, MYZ21]. behaviour [Bas20, Buc13, FH22, KLL18]. Behavioural [RR14]. belong [BLL16]. Benford [DL08]. Berg [Mar10]. Bergomi [Gas19].

Bernoulli [AST14, Bob08, Cer14, Lat08, MAPS14, NP20, Pal08]. Bernstein [DS16]. Berry [GS09, NPY19, Sal22]. Bertoin [Win20]. Bessel [AM18, Alt20, ESY08, HM14, KR19, KO01, Pan21, PW18, PŻ16]. Best [Mar98]. bet [HW19]. Beta [MA17, PR15, DS15, LM17, Möh18, Ost13]. Beta-gamma [PR15]. between [BG21, BCW20, HT05, Le 08, Lou04, MW09, MW12, Pin17, SV08, vdBC13]. Beyond [BS96, Kov09]. Bi [Sau19, BT17]. bi-infinite [BT17]. Bi-log-concavity [Sau19]. bias [Gol22, LG09]. biased [BCNP19, BT20, BT21, GG11]. biasing [LX22]. big [MHC13]. Biggin [BM18]. Biham [AHM05]. binary [CM22, RW21, vdBKN12]. binomial [Hil12]. bins [CP19a, CP20]. biorthogonal [But17]. births [CM22]. Bisection [Har21, Hua18]. bivariate [MAPS14]. BK [Jon13]. BKKKL [Ros08]. block [KLS21, LM17, LS18b, MP14b, Möh18, Ora07]. block-matrices [Ora07]. blocks [Ora07]. blowup [LX15]. Bluetooth [BDL15]. Bobkov [KPS96]. Body [LW05]. Bohman [Sen16]. Bolthausen [KP15, MP14b]. bond [Ald16, Dem20]. bonds [ARS22]. Boolean [For20, For21, Pen18, Sal22, VM13]. bootstrap [GP14, HS22, dBJP13]. Borel [OS16]. Bouchaud [Mui15]. Boué [HW22]. Bougerol [Ass18, HM23]. Bound [Han98, JK04, RŻ98, CG15, Fan15, Han99, Hut18, IM16, KO06, KF09, PS08, Sap10, vdBD20]. boundaries [AM18, BBF18, ESY08, Lyo19]. Boundary [CKS99, Has05, MYZ21, Wag16, vdBN17, Baz21, BA14, BFP^{+}09, BJ18, Bur19, CM22, Dok15, FJR20, GHJ16, Hon19, IR10, KM17, Kua16]. Bounded [Dem96, ESvRS09, RR14, HM15, Mar18, Ose08, Rio15, Ruf17]. Boundedly [LR16]. boundedness [Lat08]. Bounding [Roc05]. Bounds [BCG12b, BP20, CG05, DL08, GS09, GP19, Mar98, PR11, SV04, Wer96, BGHK08, BN08, CM13, Cho18, Doh13, FM12a, FM12b, GP14, KM09, Kis14, LR15, NW15, NPY19, Nua18, QM17, RR15, RS06, Tan17, Yas14, Yas15]. Bovier [GKS18]. box [Cer14, Mor08]. branched [Yan22]. Branching [CGP19, FZ10, HN11, KS97, Mor05, Wan02, Aid10, BM18, BK11, BH16, CM22, CM18b, CHL21, Cra13, CP11, DGK19, DV23, DG17, DP18b, DGH20, EP17, FF12, FJR20, GH18a, GM13, Gho22, GKS18, HH07, HR14, HK15, Hut11, JL08, KM17, Kri21, LT20, MYZ21, Mai13, MM21, Mal15, Mül08, Öz20, PW11, SK15, WH19]. Brascamp [Har14]. breaking [DGdHM21, vdBN17]. Breuer [NPY19]. Bridge [Gao03, Li17, Li16]. Bridges [Ali01, CLMR15, Con16, Alt20, BCP03]. Brownian [MW12, AG15, Ald98, Ali01, Aur11, Bar05, BBB97, BBKM00, BB06, BA14, BCSW18, BPR99, BBM22, BDE13, BH16, BGT07, BFP^{+}09, Bor10, BZ18, Bor22, BN20, BM20, BO03, BN08, BC98, Bur19, CM12a, CK18a, CC98, CGS21, CM22, CK08, CHL21, CSS99, DeB07, DM09, Far98, FV21, Gao03, Gao08, GKS18, GL14, GT11, GHJ16, HCA17, Han21, HH07, HK15, Has05, HT05, Hon19, Hon21, Hoo99, HN09, HN10, HLN13, HSY15, IS17, Ist05, Jan19a, Jan13, JV09, KT03, KT13, Ken09, KLS05, Law96, LM06, Li17, Mai13, Mal15, Mar11, MY99, MLV15, MMB18, MS19b, MW09, NP13, NR18, NX15, NX13, NS13, Oça20, Oka14, Olo18, ORV22, Owo15, Öz20, Par17, PW11, Per22, PW18, QR11, Ros20, SV04, SW02, Spi13, Tan06]. Brownian [Tud09, Unt10, VA06, VY12b, VY12a, Wag16, WYY13, Wan15, Wan21, ZN03]. BSDEs [Bah02, BCH^{+}00, BDM01, CEK11, FPZ16, JKL17, XFT20]. buffers [AS16]. bullet [DKJ^{+}19]. Burdzy [CP23]. Burkholder [HHJ22]. BV [Tre13a]. BV-regularity [Tre13a].

C [AS22, KS10]. càdlàg [CC18]. càdlàguity [MX19]. Calculus [GKH03, TM15]. cancellative [Swa13]. Cannibal [Kub11]. Canonical [Ali01]. Cantor [Mar17, Zhu14]. Capacity [CKS99, LLN09]. caring [BK11]. Carlo [BA01, Hol21]. cascades [DW15, Nik20]. case [AN19, BI15, CM22, Def12, Gau16, IM16, Maś18, RS07, SW10]. cases [MYZ21]. catalytic [BH16, SS06]. catastrophes [FS19]. Cauchy [Ber00, BFY07, GN06, LS18a, MZ14a, Rie18]. Cayley [Mar17, NP12a, Pri17]. cells [BDM07, HM15]. cellular [BP10]. censoring [DM14]. Center [JC04]. centered [Che21a]. Central [AS22, HN10, NX13, Rok15, RX20, AN19, Fér20, GV14, GMP18, GPPdS14, HN09, LZ23, LMV22, Rio11, Tri19, BLL08, BR07, Cha10, DV11, DBGP03, Kar07]. Certain [Ali01, Häg02, MR01, BB06, JK13, MP16, RR15]. Chafee [DHI11]. Chain [BA01, Gui99, HR07, Yad09, CD17b, Hol21, Lac15, LW09, SS21, YRE16]. Chains [BLL08, Ros02, Tel00, ADOS11, CP17b, CP14, DG15, Die15, HLMV20, KF09, Mül08, NW15, Pel20, Pel23, Rao19, Rao21, RR97, SB07]. chambers [KS10]. Chance [DeB07]. Chandra [KT03]. change [BY13, FI21, GG14]. changed [HRKU11, KP22]. Changes [MY99, AF19]. Chaos [GH18b, AN19, Ber17, Bou16, CP19a, GHSS18, Lac18, NP12b, NP12c]. chaoses [CNPP16]. Chaoticity [Rey15]. Characterisation [Die15, JR11]. Characterising [BMMU20]. Characteristic [AGN21, Kös08, BCG12b, Har04]. characteristics [vSW19]. Characterization [AP16, FJ00, LG09, Arg07, BMV07, DN07, Ejs13, KNN15, Neh14]. charge [DHR18]. Chase [BHJR22, DJT20, HJCW20]. Chase-escape [BHJR22, HJCW20]. Cheeger [Mon07]. Chen [ST20]. Chervonenkis [Pan02]. chi [Jou12, MR22a]. chi-square [MR22a]. choice [MP14a]. choices [Mal20]. Choquet [Men14, Men13]. chordal [AK08, Doh13]. Chung [Hil06]. CIR [Aly13, ARS22]. circulant [BHS10]. circular [MZ14a]. claims [PRT13]. Clark [MN08]. Class [DCF06, Mor05, NY10, Wan02, Aym21, BBCG08, EN20, Eri16, Kaz18, Kli12a, Kub11, KZ13, Led17, LST15, MYZ21, MG16a, MG16b, O’R12, San13, Swa13, Woj12, WP14]. classes [CLMR15, GP11, LMK03, MN09]. classical [Li17]. Classification [Wan02]. Close [BM05, Mar05]. Closed [Val19, Jab17]. Closed-form [Val19]. Closeness [HM16]. Closure [Mar10]. CLT [Fan15, NPY19, Pel20, Pel23]. CLTs [HR07]. clumpy [JM15]. cluster [Aou21, BS17, Dem20, GJ09b, Hua19, Kis14, Li14, Mic19, Pet08, PR12b, Sap11, vB15, vdBC12]. clustering [FLP20, vdBHH10]. Clusters [KS03, PV05, vdBKN12, vdBC13, vdBC16]. co [HJ18]. co-existing [HJ18]. coagulation [Ber10]. coalescence [Nic06, Zha20b]. Coalescent [Möh11, Sch99, DPS15, FH22, GWZ21, KP15, MP14b, Möh18]. coalescents [Fre12, Lag07, LM17]. Coalescing [FHJ18, HT05, BFGG^{+}16, TYZ12]. Coarsening [DKNS16]. coefficient [Bah02, ES21, MY20]. coefficients [AF06, Böt11, GŁM23, Gru23, Jab17, Owo15, Thi20, Zha20a]. Coexistence [DJT20, RS21, TT20]. collector [Ili19]. Collide [KP04]. Colliding [KO01, OY01, BFP^{+}09]. Collisions [HP15, IS17]. colored [Fan15, Gri11, HHN16]. coloring [HHL18]. colorings [HL15b]. coloured [BP09]. colouring [BMMU20]. colours [Jan19b]. Column [Men18]. combinatorial [CP23]. combinatorics [WP14]. Combine [BA01]. combined [MP14a, Mal20]. Come [Sch99]. Comm [MZ05a]. Comment [AB14, Tót13]. common [HM16, LW15]. Commutative [Kar07, FdM07]. compact [BMV07, GZ19, MP13, Nie22, RX20]. compactification [Ras10]. Comparison [Ang19, BCH^{+}00, Mar18, YRE16, Pin16]. competing [IS17, RS21]. Competition [DV23, LK08, TT20]. complementary [MR15b]. Complete [DS16, Wäs08]. completely [ADK22, McV08]. completion [LV20]. Complex [KT13, DM18, HK15, IM16, KM17, LP19, WH19]. component [Ald16, Bor22, DM21, Pen22, Rát18, Sen16, Squ21]. Compositions [MU10]. Compound [CK12, Dal13, KM06, BJ20, Dal17, Möh11]. Computation [GKH03]. concave [MG16a, MG16b]. Concavity [Hil12, Kul16, Sau19]. Concentration [BT12, Del10, DZ96, FV14, FM12a, FM12b, GZ19, GLP18, GZ00, GL09, KM06, Mar05, NX15, Pan01, Pol19, PR12b, TM15, Wan14, BT19, BK13, CM12b, DG15, GG11, Lan22, NW15, Rio13b, RV13]. Concerning [War99, Dub23]. Condition [Bir04, CLS05, Sch99, BHS18, BJ18, CPS12, GV14, Hoe09, IS17, KV13, LMV22, Liu15, LG20, RW21, San13, Wan17, Yas15]. Conditional [GLY14, HYZ19]. Conditioned [Ald98, PR12a, War99, Baz21, Car18, DJ12, FF12, JL08, Jan21, KS10, Mil08, Tou22]. conditioning [ALW14]. Conditions [AAGL19, BL10, BDT11, HR07, Lou04, BCM22, FV21, JR19, MR15a, Yas16]. conductance [Buc13]. conductances [ADS19, ADS20, Ave12, BHM21, HK16, KSW12]. cone [VY12b]. cones [Dur14]. configuration [BHJR22, Cha20]. confined [EM21]. confining [Har12]. Conformal [MP20, BM20, Gro19]. Conjecture [KPS96, CP23, Duq09]. Conjectures [PW96]. Connected [Kri07, AB14, HM15, LL20, Mar11]. Connectedness [ABIT22]. connection [Har14, MW09, MW12, PS16, vdBD20]. connections [DJ06, Pin17]. Connective [Gla15, Mar17]. Connectivity [BDL15, FvdHH16]. consensus [LL20]. conservative [Gou18]. Consistent [Cra13, FZ10]. Constant [Kes96, AF06, Gla15, JPR19, JL18, Pim06, Pin19, PR12b, Ruf17]. constants [LR15, Mar17, MS11, Rio11, Rio17]. Constrained [CG05, BC14, CM13, DGdHM21, FPZ16, Har21, Yin15]. Constrains [Lon04]. constraints [DGO20, Fra13, Rok14]. Construction [CEG11, Böt11, HL15a, Hon21, Hua21, Nut12, Win20]. Constructions [BPR99]. Contact [Can15, GMT15, HD20, HD23, SV16]. Containing [DCLYY13]. contains [Mar17]. content [KP22]. context [Ruf15a]. Continuity [DZ19, GMT15, DM18, DGK19, Hon19, LX19, LD21, MR15a, MR22a, MP13, SSS15, SV11b]. continuou [Vid14]. Continuous [KS05b, Mac02, Mik02, RR14, Vov08, vZ02, AP16, AF19, Ban15, Bec22, Dem20, DP18b, FF12, FG13, FJR20, FJ21, GŁM23, JR11, KK15, KO06, MYZ21, Owo15, PR19, SŽ17, Zha20a]. continuous-state [DP18b, FF12, FJR20, MYZ21]. Continuous-Time [RR14, Vov08, FJ21, SŽ17]. Continuum [Stu13, ATT18, AG15, Ang19, Dun20, HL15a, MWW11, Wan21]. contour [HL13]. contours [HS22]. contracting [GLP18]. contraction [PS17]. contractive [BI15]. Control [Mik02, Wee06, GS12, JPR19, Rok14, Sir14]. Controlled [Ale13, BLY15, Gor15, GGPZ14]. Controller [Wee06]. Converge [MZ05a, MZ05b, CD17a]. Convergence [AJ14, AR18, BL10, BC98, GG04, Hua18, Ili19, KM08, KM17, MU12, NP12b, Pec04, Ros02, SS21, Tho16, Wan17, WH19, ALW14, BM18, CGR10, Cha15, CL14, DT18, Ern21, GL08, GM16, GM17, HJ18, JPR19, KNN15, Kri14, Lan22, LR16, MV14, NX15, NS13, Oça20, Stu16, NP12c]. convergent [ARS22, HK11]. Converse [BCH^{+}00, YRE16]. Convex [APRB11, ABV03, Ber00, MW16, BJ20, DJR16, DT18, Dur14, FV14, LMK03, Rei13, Teh17, Tko11, Vys21]. Convexity [Kes96, Lal03, Ken09]. convolution [D’O10, DZ19, MM21, NY09, Sal22]. Convolutions [AS08, Kar08, Hil12, VZ11]. cookie [RS11a]. cooling [Xie20]. Coordinate [GS09]. Coordinates [Spr07]. coprime [Mar22]. core [Häg02]. Cores [Hol21]. corner [Emr16, GMP18]. Corners [Gne08]. Correction [Bas11]. Correlated [Tuc11, BS22, CGP19, LD21, Par17, dHP14]. Correlation [Han98, Kös08, LP99, Li99, MR11, Aou21, Car22, DY21, Han99, Ren22, Wei03]. Corrigendum [HD23]. Cosiness [ST99]. cost [AGS14, Goz06, Hue16, KLL18, PS19, Tri19]. Costs [SV11a, DGO20]. Coulomb [GZ19, Har12]. Countable [RW09, CH21, LS18a, Owo15]. counter [GV14]. counterexample [GJ12, KS07, RS16]. Counting [DV11, CLMR15, Con16, Fan15, LM17, MP14b, Möh18, Squ21]. coupled [Bur19]. Coupling [AHM^{+}13, BA14, CK18a, BCY22, BJ20, FGM10, KS07, NP20, Pos09, Szn19]. Couplings [Sch20, GG11, Gol22, Ken09]. coupon [Ili19]. Covariance [SP00, Sep03, HCS08, McV08, Yas14]. Covariation [DMPARA13]. Cover [JS00]. covering [PZ18]. covers [HS12, Pin17, SHH14]. Cox [HJM22]. Coxeter [Fér20]. Cramér [BT11, CP17a]. Cranston [KS07]. Credit [Lon04]. criteria [Goz06, Sok13, Zho14]. criterion [Ern21, Nou11]. Critical [APS19, FXA18, Gra16, Ham05, Kah03, Mic19, Aid10, BCW20, CP17a, Due06, GM13, JJ16, Ker17, Kis14, MYZ21, Maś18, Pim06, RSS18, Sub12, Yao14, vdBC12, vdBC13, vdBC16, vdBD20, vdHKM09]. criticality [DC13, vdBN17]. crossing [AV12, BS17, BB01]. Crossings [CKS99, Has05, Gan14]. Crumbs [ABP00]. Crystal [DC13]. CSBP [Lab13]. Cubes [ST99]. cubic [NS13]. cumulants [BHS11, FM21, HS11]. cup [Ili19]. Curie [CP17a]. current [VR10]. currents [LW16]. curves [LR15]. cut [Wan21]. Cutoff [BHP18, Her17, CP22, Lac15, Oza20]. cuts [IM07]. cycle [HHL18, KM20]. Cycles [Mar99, Bjö15, MNZ12, NP20, Pin17]. cylinder [Sil20]. cylindrical [Rie18].

D [BDT11, Duc19, GH18b, Har12, KS10, Sab21, vdBN17]. D-Vertex [Sab21]. Damped [App02]. data [BCP21, Che21b, Hua17, Zha20a]. David [War99]. Davis [HHJ22]. dealer [JM15]. Dean [KLvR19]. Decay [BLY15, FV21, dBM15]. decaying [DZ13]. decomposition [Agu18, CJK18, KK15, KP15, MP14b, Möh18, Möh19, MR15b, RSS18]. Decompositions [Ali01]. decreasing [GRS03]. defaultable [ARS22]. Deffuant [LL20]. deficiency [JSS22]. defined [Fra13]. deflated [HJT12]. degeneracy [HLN13]. Degenerate [Wee06, ADS19, FMP17, Luo14, Men11]. Degree [Tam07, Her17, JJ16, JL18]. degrees [Bac11, DJ06, Dei09]. Delange [Aym21]. Delay [AK04, CR05, Buo21, RM16]. delayed [Ore21]. delays [CY13]. Delocalization [JL18]. delocalized [ADK22]. Densities [BBB97, BC14, DOS16, DM09, Jeg09, LT20, Sim11]. Density [GJ09b, Sch16, Alt17, BM22, FH19, Fun07, GL08, HK13, HKST18, HK11, JJ16, LS13b, PŻ16, Stu16, Tho16]. Deny [Men14, Men13]. depend [HKZ12a]. dependence [Dok15, Squ21, Zha20a]. dependencies [Ada15, Wei03]. Dependent [Lin09, Wan02, AS22, FH22, GPL08, GL09, HCS08, HL15b, HHL18, HK16, O’R12, Oli10b, Pol19, Thi20, Wan22]. Deposition [AAK01]. depth [LPN22]. Derivative [Rin98, Tan06, Tre13a, YY18, Zha22]. derive [NW15]. derived [Möh11, RW09]. Derrida [SK15]. descent [Fér20]. destruction [Gra16]. destructive [AST14]. detection [Buo21, FY15]. Determinantal [Pet10, Lyo18]. determinants [Har04]. determined [PS20]. Deterministic [Eri16, Hol21, Ste08]. Deviation [GJ09a, Oli10b, CP17a, DL18, DJ12, Dzi13, FGL12, FJ21, Gan14, Kis14, Rei13, WY08, dHP14]. Deviations [Big04, Dem96, DS06, FX02, KS03, BP09, BG20, But17, BZ17, Car22, Com08, DF16, DJR16, DPS15, DGH20, ES09, EMR15, GRR14, GH18a, GJ09a, GWZ21, GG20, Gho22, Har12, HJT20, Jan19a, KLM15, KLS21, KSW12, LY22, LW09, MPY14, Nik20, Puh19, Tsi13, Yin15, Zhu14]. Diaconis [Hil06]. diagonal [FG13, HM16]. Diagonals [Sch16]. diameter [Wan15]. Dichotomy [Fun07]. dies [GM13]. diffeomorphisms [Att10]. Difference [CP05, CV07]. different [BDE13, BP20, NS13]. differentiability [AP14, Pan08]. differentiable [Luo14]. Differential [AK04, Bar05, BLY15, CR05, FW00, TW03, AF06, AAS21, Ban15, Beg14, CD13, Cri21, CY13, D’O10, DOS16, DC15, FT07, Fra13, FI21, Hoe09, Jac23, Kaz18, Ose08, Ose10, Ose11, Ose20, ORV22, Owo15, Ruf15a, RM16, Tap13, Tap15, Unt10, YE13, YRE16, Yan22, Zha20a]. differential-algebraic [AF06]. differentially [Osȩ16]. Diffusion [Hus08, Jan96, Rin98, Wan09, AS16, BBF18, BR16, Buo21, Car22, CSC13, Cla14, DCLYY13, ES21, Eth14, GG14, GM16, GM17, HLWZ15, KDV17, KTT17, LST15, PW11, RBS15, Sai07]. Diffusion-Limited [Hus08]. Diffusions [Sam10, Win20, BC14, CK14, DN07, Fan16, Gau16, GG20, Hut11, Jab17, KSY06, Kli12a, MU12, MR22b, PS19, Pin19, Rey15, Ruf15b, RW09, Tug16]. Diffusive [NY10, dBM15]. dilations [Gri11, LS18a, Tko11]. Dimension [BR07, Sim00, Bec22, BCSW18, Bor13, BJ22, Buo21, Far98, GP19, Hol15, HKZ12a, Hue16, Jou12, Law96, Le 08, PW18, Sal15, SXY18, Yan06, vB15]. Dimensional [Spr07, Swa01, Abe15, ADS20, AS11, Att10, BFRH15, BDZ11, BS07b, BR16, Can15, CB10, CEG11, DFK18, DG17, Due06, EK08, FI21, GJ18, GPHS13, Häg02, HM09, Har04, HM23, Hon19, JTT18, Jac14, Kab19, Kli12b, KTT17, Mal15, MU12, OdS16, Pet15, Pin19, PR19, PZ18, Roi05, RW09, San13, Sau19, ST17, TYZ12, TYZ15, Van07, Wag16, Wan09, Wat12, WZ19, Yuk08, Zha20a, Zho20]. Dimensions [Law98, Baz21, BDKS19, EMR15, MS11, NZ20, Sap11, SŽ17]. Directed [Bir04, FJ21, GK22, HS09, HS12, Jan19a, Liu22, SHH14, Wat12]. direction [Cou11]. directional [OdS16]. Dirichlet [Arg07, Aym19, AFM20, Fen23, JK08, LD21, MA17, Oui18, RW09, Uem07]. Disaggregation [DCF06]. disc [FKV23]. disc-polygon [FKV23]. disc-polygons [FKV23]. disconnectedness [Zho14]. Disconnection [PW96, Wer96]. discontinuities [BKS16]. Discontinuity [Jan97]. discontinuous [AP14, Att10, GS12, Lej11, LST15]. discrepancy [DY21]. Discrete [Fan16, Kaz18, MS19c, SBS15, Van08, Ang19, BDZ11, CF20, Cra13, Dol14, DKJ^{+}19, KZ13, Lup16, Mak08, PR18, Rok07, RS21, SB07, Win08]. discrete-time [SB07]. discretely [BZ18]. discretisation [Kaz18]. discretization [FI21]. discretized [BBMT09]. Disjoint [Gan14, Wei03]. disk [Gaá14]. Disorder [Bir04, BS22, DW15, FJ21, Lac10]. Disordered [BD02, CdH13, KO06]. Displacement [FZ10, Mal15]. Dissipative [LT11]. distance [CD20, EM14, Liu15, Rei13, Rio11]. distances [GP19]. distant [Uch15]. Distributed [EZ99, vdBHH10]. Distribution [CK08, DFN00, Jan97, JK04, JK08, MZ05a, MZ05b, Ost14, Spr07, Wan22, Bac11, Bar14, BD13, Bas15, BH16, Bor10, BW08, But17, DR12, HY21, Jan15, JL18, KV22, Kli12a, KLL18, Led16, MZ14a, Mic13, MPP15, MP16, Péc19, Pol19, Puh19, Rey15, SV11b, Sub12, Tam07, Val19, Zha22]. distribution-valued [Led16]. Distributional [FLP20]. distributionally [BEK21]. Distributions [Jun11b, KM06, Res01, Arg07, ALW14, BC15, BJ20, Bob08, Dem11, DMPARA13, Gra15, GJ18, GPHS13, HRS20, HM14, HK14, KV11, Lab13, LC22, LG09, MG16a, MG16b, MN09, MU10, MA17, MAPS14, Mar14, Ost13, PR19, Tan17, Tsa18]. Divergence [CP05, Nak19, BD15, Oto09]. divergent [BI15]. diversity [FPRW20]. Divisibility [MR08, AJ14, VY12b]. Divisible [MR01, AV12, DMPARA13, MN09, MU10, Wat12]. Dixie [Ili19]. DLA [RS11a, Sil20]. do [EHW15]. Dobrushin [MY20, Wan17]. Does [MZ05a, MZ05b, BLL16, BS18b, Sap11, Wei03]. Domain [DeB07, Dok15, LC22]. domains [BB06, BBM22, Fra13, Mar11, MS19c]. dominate [Sap11]. Dominating [McV22]. Domination [Lin99, GS21]. Doney [PN15, PN16]. Donsker [BRT10, BDM01, CD20, Par17]. Donsker-Type [BDM01, Par17]. Doob [Rio18]. Döring [HY21]. Double [Wai13, BK11]. double-branching [BK11]. doubly [AED13, HRS20, Tre18, Val19]. doubly-intractable [HRS20]. doubly-stochastic [Val19]. Dovbysh [Pan10]. Down [Sch99]. drawdown [HSY15]. drawn [dHP14]. Drazin [SB07]. drift [AP14, Att10, BFJ^{+}19, BO03, DN07, DP14, GN06, GJ09a, GM16, GM17, Jan13, JV09, KSS11, Le 08, LST15, Luo14, Oça20, Pan21, PW11]. drifted [BA14]. Drifts [LT11]. Driven [AR18, HT05, Ban15, BO03, Cri21, DV23, DXZ11, GŁM23, HHN16, KLL18, MNX20, MY13, ORV22, Owo15, Tap15, Unt10]. dual [Cri21]. dualities [JK13]. Duality [HT05, Jan96, CLMR15, HA07, Kua21, Lin19]. Dupuis [HW22]. duration [BBM22]. dyadic [Bia13, Sio14]. Dyck [KM09]. Dykema [But17]. dynamic [Blo23, Buc13, GS12, OdS16]. Dynamical [AABP22, DN07, Arg07, Ave12, BLMZ12]. dynamically [BHM21]. Dynamics [DE23, CP20, Gor15, KLvR19]. Dyson [BFP^{+}09, KT13, Per22].

easy [Wäs09]. Edge [Gui99, MR11, ACT19, ABIT22, Cra13, Geo10, Law14, Mon07, RRZ11, SV16]. Edge-Reinforced [MR11, ACT19]. edges [Fan15]. Edgeworth [MY13, Pri15]. efficiency [AA07]. Eigenfunctions [AB02]. Eigenvalue [DV11, Wan09, BCY22, SV11b, Woj12, Yas14]. Eigenvalues [KO01, Sos04, BGP14, BG20, MS19a, Wan18]. eigenvector [CD17b]. Eigenvectors [BGZ18, BGP14, BG20, Dub23]. elastic [BBF18, HM22]. Elect. [MZ05a]. elementary [Ber17, Duq09]. elephant [KV22, Lau22]. elliptic [Ave12]. embedded [PW18]. embedding [BM20, Gro19, Har14, YY13]. embeddings [CK14, MP20]. emergence [Vov08]. empirical [BB07, BB22, Mar18, Pol19]. ended [GGNS17, Tim19]. Energy [Tan17, Che21b, HM09, HK15, MR13, SK15, Wat12]. Engel [Zhu14]. enhanced [Gol22]. Ensemble [BDN10, CS14, DGdHM21]. Ensembles [Def11, HCS08, But17, DMPARA13, DS15, LT20, Wan18]. entire [KN19]. Entries [Tuc11, BHS10, FG13, GL09, HCS08, O’R12, PS08, Pol19]. entropic [Oza20]. Entropy [Gao08, BCG12b, CG15, DY21, Fat18, Fen23, Hil12, KF09, Le 08, Rio18]. envelope [Agu18]. Environment [Bir04, HN11, RA05, Zer02, BKS16, Baz21, Blo23, CGP19, DGK19, DFK18, DL09a, ER09, GN14, GM13, Hut11, JL21, Jul19, Kri21, Pet15, WH19, Wat12, Xie20, dBM15, JL21]. Environments [Roi05, BHS18, EP17, HS09, KTT17, LS13a, OdS16, Ste13, Zer07]. equal [AS16, Def12]. Equation [HT05, TM06, AG15, DHI11, DR12, DXZ11, DT18, FP11, HCA17, HHN16, Hua17, KSS11, LX19, MM21, MW09, MW12, Nie22, QR11, Ren22, Sal15, Sim00, YY18, Zha22]. Equations [App02, AK04, Bar05, BDT11, BLY15, CR05, CP05, FW00, LT11, TW03, AF06, AAS21, Ban15, Beg14, BKP22, CD13, Cri21, CY13, D’O10, DOS16, DM18, DC15, FT07, Fra13, FI21, GIK22, HRKU11, HM22, Hoe09, KL14, Kaz18, Lac18, LP12, LZ23, LY13, Luo14, Men11, Nua18, NZ20, ORV22, Owo15, Ruf15a, RM16, Tap13, Tap15, Unt10, Wan22, XFT20, YE13, YRE16, Yan22, Zha20a]. Equidistant [JV09]. equilibrium [CPS12, DT18, JF19, MR22b]. equiprobable [BDM07]. Equivalence [Jou12, Loe13, Lou04, BG21, DGdHM21, vZ08]. equivalent [BCG12a]. Erased [Law98, Mar99, Law14]. Erdos [ADK22, DL18, GM22, JL18, Kif15, Kif16, Rát18]. Ergodic [BLL08, DG15, HLMV20, HLWZ15, Jos07, PY17]. Ergodicity [BKP22, CS14, Ken04, Lou04, BLMZ12, BHS18, BCP21, KS14, NZ20, RR97]. Erratum [BYZ12, BB22, Don22, Fou14, FM12b, GM17, Kif16, KS19, MG16a, Men14, MW12, NP12c, SHH14, TYZ15]. Error [BN08, DL08, Fan15, McV08]. errors [JL21]. Escape [Kor05, Yad09, vdBHH10, BHJR22, Cha20, DJT20, HJCW20, Lâm22]. Esseen [NPY19, Sal22, GS09]. Estimate [Gao08, Yan22]. Estimates [BLY15, CKS99, Rin98, Tel00, ADS19, AJ14, BCY22, Con16, DFK18, EM21, Gan14, GL08, HK16, Hue16, KZK19, LPP15, LP08, MX19, OW22, PR12b, Unt10]. estimation [Che21b, CK12, Mak08]. Estimator [Gao03, Stu16]. Estimators [PR11, GJ09a]. Euclidean [BBM22, Bor13]. Euler [BFY07, GL08, Kaz18, Sab13]. events [GH21, HR14, Wei03]. Eventual [EP98]. Evolution [LT11, BGT21, BLZ18]. Ewens [Zha12]. Exact [BA01, DZ13, KM08, LX19, NX15, Yin15]. exactly [CM18b]. example [GV14]. examples [DJR16, KMiS06]. exceptional [Neu11]. Exchangeable [EZ99, PY17]. Excited [BW03, BR07, DK12, Hol15, Pil17, Zer06]. exclusion [Ave12, BCP21, Che17, CCGS19, GRS03, VR10, Zho20, KS02]. Excursion [Ald98, CH04, Hoo99, Jan97, War99, SW10, Tod22]. Excursions [Jan96, MPY14]. exist [EHW15]. Existence [ATT18, Bah02, Ban15, Fit06, LT11, AS16, BJT17, CEK11, Jac23, LST15, OS13, PR19]. existing [HJ18]. Exit [Lyo19, Aur11, CGPPS13, DHI11, Dok15, Hon21, Mar11, Pat07, VY12b]. expander [Sar21]. Expansion [CSS99, EM16, GH18b, HvdHS08, MY13]. expansions [GWZ21, Mar09, Pri15]. Expectation [BCH^{+}00, HW19]. expectations [GHJL17, HLWZ15]. expected [EM16, Eva06, FV14, JL18, Mar11, NX15, Sap10, vdBC16]. Explicit [BL10, D’O10, Dub23, DL08, HKST18, Mic13]. exploding [KT11]. explosion [BS16, Mau20]. explosive [Lab13]. exponent [KV15, Uch18, Ven13, Zha20b]. Exponential [Bau02, BGHK08, BY01, DS10, GM12, IM10, Jun11b, KS14, MY20, Ngu22, PR11, Rio15, RS06, TM06, dlPP09, AI12, BLL16, BHS18, Buo21, CP19b, CSC13, DM09, DG17, DGH20, Emr16, FGL12, GRR14, HW22, KM09, KMiS06, Pet08, Sch09, SC09, Yin15, Zha20b]. exponentially [DZ13]. Exponents [Ham05, PW96, Wer96, NP12a]. extended [TYZ12, TYZ15, VY12a]. Extension [MR15b, Hoe09, HvdHS08, Pin16, Pos09, Sau19, Uem07]. Extensions [BGT07, BN20, Fit06, Pan02, Rio13a, HM23]. external [Che21a]. extinct [FF12]. extinction [BK11, JL08, SV16]. extra [FLP20]. extra-clustering [FLP20]. extrema [CM18a]. Extremal [CJK18, Dzi13, CCH15, GKS18, MP20]. Extreme [KLS05, CHL21, LC22]. Extreme-Value [KLS05]. extremes [Oui18].

factor [RR15]. Factorial [BLY15]. factorization [Pan21]. Factorizations [BY01]. Factors [Bal05, Cho18, Dal17, MG16a, MG16b, Tim04]. fail [CD17a]. family [Bac11, BLL16, BN20, Neu11, Ngu22]. Fast [BA01, JS18, BR16, CM18c, Wai13]. Fastest [Roc05]. favorable [EKR22]. FBSDE [Jac23]. feedback [HM22]. Feller [Böt11, NP20, PW11, Zha20a]. Ferguson [JK08]. Few [BS96]. Feynman [Bal09, Tak10]. field [BZ16, BDE13, BDZ11, CD13, Che21a, JTT18, LW15, LW16, NP12a, Pel23, CCH15]. Fields [CGXM96, Ist06, BMV07, BCW20, CGXM97, Eis21, ER19, Kli12b, KZ13, LP19, LMV22, MZ18, MHC13, Tod22, WZ19]. Fill [Mac02]. filter [MY20]. Filtered [Çet12]. Filtering [Mak08, CL06, FMP17]. filters [Van08]. filtrations [KK15, Lau13]. finding [MU10]. Fine [Fit00]. Finetti [GK21]. finitary [CZ21]. Finite [Har04, LX15, AGN21, Ald16, BC15, Cer14, DJ06, Der16, Die15, GK21, HM09, JV09, KS14, LL20, LR16, MLV15, Mon07, Rok07, Sio14, Tho16, Zha20b, vdBKN12]. finite- [AGN21]. Finitely [Gne10, HHL18, KP04]. finiteness [AAGL19, DR12, KSY06, Zho10]. fire [Due06, Gra16]. First [And06, BT17, DG17, IM10, Kes96, Lal03, Sch09, Wan09, Ang19, BCY22, BCP03, BB01, CSC13, CEG11, DHI11, Dok15, Gan14, GM21, Gne21, GP19, HM14, Vid14, Yao14, Nak19]. First-Passage [Kes96, Lal03, Sch09, GM21, Yao14]. Fisher [Fou14, Fou13, HT05]. Fit [Sam10]. fitness [BLZ18, GMRC16]. fix [Pin17]. Fixation [GGA10, Möh18]. Fixed [FJ00, AG15, ALW14, BKS16, LS13a]. FK [LW16]. FKG [Bar05]. flat [PSY13]. Flatness [Tap15]. Fleming [Zho14, FX02, LSY99]. Flow [Tha98, AP14, Led17, Xio04]. Flows [CSS99, Att10, Die15, EMR15, Haj15, HCA17, MR15b, VA06]. fluctuation [KO06, Nak19]. Fluctuations [ES16, KZ13, LS18b, BGT10, Dok15, GJ09b, GK22, Sil20]. Fokker [DT18, HRKU11, Luo14, Ren22]. foliated [Ruf15b]. foliations [MR15b]. forced [Wai13]. forest [Due06, Gra16, ABIT22]. forest-fire [Due06, Gra16]. Form [CP05, BW08, Fat18, Lac18, Val19]. Forms [Tha98, HKZ12b, Uem07]. Formula [Bal09, Sch01, BY13, EW09, HW22, HKST18, KT03, MN08, Mic13, Oto09, Pan08, Rát18, YY18, Yan06]. Formulae [Han05, Alt20, BFY07, DP18a]. formulas [CLMR15, Def12, Dub23, Val19]. formulation [BZ16, But17, Sir14]. forward [CD13, Duc19]. forward-backward [CD13, Duc19]. Four [Law98]. Fourth [DP18a, AJ14, Yas15]. fractal [BS18b, BJT17, GHSS18]. fraction [AHM06]. Fractional [BDE13, BGT07, CC98, GG14, Ist05, Ist06, NR18, Aur11, BRT10, Beg14, BCSW18, BZ18, BO03, BN08, D’O10, DOS16, Han21, HLN13, Jun11a, MW09, MW12, NX13, NS13, Nua18, ORV22, Par17, PS16, Tud09, Unt10, WYY13, YY18, ZN03]. fragmentation [Ber10, BS16]. fragmentation-coagulation [Ber10]. Fragmenting [GMS08]. Free [AS08, CCH15, HMO01, Kar08, AV12, ALW14, BDZ11, Bou16, DY21, Dem11, DZ19, DMPARA13, Eis21, Ejs12, Ejs13, GŁM23, HK14, KZ13, LP19, LW16, MR13, RS07, SV11a, Wat12, ABIT22]. Freedman [Tro11]. friendly [BHS18]. Frieze [Sen16]. fringe [DJ14]. Frobenius [CD17b]. frog [BFJ^{+}19, DR20, DP14, JJ16, JR19, RS21]. Front [GJ18, Bjo09, HHN16]. frontier [Law96]. frozen [DKNS16, vdBKN12, vdBN17]. Fuk [Rio17]. Function [DV11, Kös08, Ost14, ABIT22, Gri11, Kin08, LR15, LW09, Liu15, NY09, Oui18, Vys21]. Functional [BI15, ESvRS09, Pan05, Tan06, Xie20, DC15, GPL08, GPPdS14, IK18, JK08, Kri14, MO22, NX13, Pil17]. functionals [CSC13, GG14, Har04, HJT20, KSY06, MU12, MP16, Pel20, SB07, Tri19]. Functions [Gao08, SV04, Tha98, Agu18, Alt20, Aou21, Aym21, BCG12b, CC18, CF20, DS16, Dur14, Dzi13, ES16, For20, For21, FJ21, Gal20, Har04, HL13, Jan09, KN19, KNN15, Li20, LR16, MS19c, Nob20, OW22, Pat07, PSY13, Ras10, SBS15]. fundamental [OD12, SSS15]. Further [AM18, Bau02].

G [HLWZ15]. G-diffusion [HLWZ15]. Gains [RR14]. Galton [BT21, PS20, BT20, Duq09, GP14, HL13, HMSH15, Jan21, KF09, Mic19, PS17, RSS18, Tas10]. Game [Wee06, BZ16, Dol14, Jon04]. games [Jac23, LW15, dBJP13]. Gamma [AS08, BT11, D’O10, KLL18, Mar14, PR15, NY09]. gap [Mor08]. gaps [FW17, Jac14, vdBC13]. gas [Eas22, Har12]. Gases [Han98, GZ19, Han99]. Gauss [Oto09]. Gaussian [BMV07, BCW20, Ber17, BDN10, BG20, BDZ11, BW04, CNPP16, Car22, Che21a, CCH15, DS15, Eis21, Fat18, GLY14, GHSS18, Gri11, GK22, HRKU11, HW22, HJT12, HS09, HK16, HT20, Jos07, KM08, KN19, Kli12b, KZ13, LP19, Li99, Lin09, Liu15, LW16, MR08, Pec07, Pin16, RV13, Sim17, Teh17, Thi20, Tko11, Tuc11, Val17, WCS19, Yor15, vZ08]. Gaussianity [MA17]. GCD [Mar22]. Ge [GJ12]. Gebelein [NPY19]. genealogy [CM18b]. General [CLS05, JL08, Tim04, Ald16, AF14, ADS19, CdH13, CJK18, EK08, Gor15, Jan09, Jan21, KV22, Kur14, RW21, Sai07, SW10]. generalisation [HHJ22]. Generalization [Bar97, Bau02, Bar98, PS16]. Generalized [AS08, Ali01, Def11, Nob20, PP20, BN17, CL14, D’O10, DP18b, KLL18, NY09, Pat07, Sch20, YE13]. generated [FLP20, RS11a]. generating [Rát18]. generator [EK08]. generators [JKL17, Völ20, XFT20]. genus [ACCR13]. geodesic [BCNP19, GP19, Liu22]. geodesic-biased [BCNP19]. Geodesics [BD02, BT17, Cou11, Zha20b]. Geometric [Beg14, Ken04, LLN09, LX22, MY99, RR97, Emr16, HJT20, Kua16, Par17, Pen22]. geometrically [DG15, HLMV20]. Geometry [CR05, PT11, RM16, KTA17, Tri19, VA06]. germ [AHM06]. get [CM18c].

Giant [Pen22, Ald16]. Gibbs [CJK18, Der16, MWW11, Stu13, Wan14, Wan17]. Gibbsian [FGM11]. Gierer [LX15]. Gilbert [HJCW20]. Ginzburg [WZ19]. girth [NP12a]. Given [BPR99, Jan97, Kuz00, Kli12a, Law14, Luo14, PR19]. Gladyshev [Han21]. glass [BS22]. glasses [Kli12b]. glassy [HK15]. Global [Jac23, VA06]. Gluing [BPR15]. GMS [GMRC16]. Godbillon [Led17]. Goldie [Kev16]. Gorin [Har16]. gradient [DN07, Die15, EMR15, Luo14]. Graham [Hil06]. grain [AHM06]. Gram [Alt17]. Graph [AR18, ADK22, BM22, BT17, CL14, Che17, DHS14, DL18, FvdHH16, GH21, HJ18, OS16, Pen22, Pri17, Tam07, Wäs08, vdHKM09]. Graphical [HA07, JK13, Mec09]. graphon [Mar22]. Graphs [JS00, KP04, AST14, Ald16, BCG12a, BP09, BFRH15, CJ13, Che21c, CMT23, DJ06, Dei09, DM21, DGdHM21, Doh13, Fan15, FHJ18, Gan21, GM21, GM22, Her17, HJCW20, Hua18, Hus08, HS12, HP15, JL18, KZ13, LL20, LP19, Mar17, NP12a, OW13, Oza20, RBS15, Sar21, SHH14, Sil20, Tim19, Yin15]. greedy [GT18]. Greeks [GKH03]. Green [ADS20, LR15, Ras10, SV04]. Grid [Tim04]. Grincevicius [Kev16]. Gromov [Loe13]. Gross [BM20]. Grounded [PSY13]. Group [Rev03, AAS21]. groups [BMV07, BFT13, CGR10, Fér20, KV15, MP13]. Growing [AB14, Hua19, Mal20]. Growth [Gou18, Kar08, BS16, CMT23, DZ19, Emr16, GMP18, Mil08, MS11, PW11, TT20]. growth-fragmentation [BS16]. Grushin [BN20]. Grushin-type [BN20]. Gundy [HHJ22].

Haagerup [But17]. Hahn [Bar14]. Half [BHS11, LLN09, ACT19, CM22, Dun20, Gra16, IR10, KLM15, Lup16]. half-Laplacian [Dun20]. half-line [ACT19, CM22, KLM15]. Half-Plane [LLN09, Gra16, Lup16]. half-space [IR10]. halfspace [LPN22]. Hamiltonian [Mar99]. Hammersley [Hut18]. Hamming [FvdHH16]. Hankel [BS07b]. Hanson [Ada15, RV13]. Hard [Häg02, RRZ11, Gaá14]. Hard-core [Häg02]. Harish [KT03]. Harish-Chandra [KT03]. harmonic [Dur14, Le 08, MS19c, SBS15]. harmonics [FXA18]. Harnack [ESvRS09, Wag16, YY18]. Hartung [GKS18]. Hastings [JS18]. Hausdorff [BCSW18, SXY18, Yan06]. Hawkes [DZ13, Duq09, RL20]. heaps [BS18a]. Heat [ADS19, HT05, Rev03, Tha98, Buc13, FP11, HHN16, HK16, Hua17, KP22, Led17, MW09, MW12, Nie22, Nua18, Pri17, YY18]. Heavy [FdLS04, Lim99, Sos04, BHS10, But18, Cha10, DHI11]. heavy-tailed [DHI11]. Hedging [Dol14, Sai07]. Height [Wan15]. Heights [CH04, Gra16, Uch18]. Heisenberg [Bjö15]. helix [Kab19]. Hermite [BN17, BN08, Law08]. Hermitian [AGN21, Bor11, DMPARA13, DE23, Oli10a, Ora07]. Heuristic [BA01]. Hexagone [AB14]. Hierarchical [Fen23]. High [GKS18, Kli12b, Spr07, WCS19, ACCR13, Bor13, CEG11, MS11, NP12a]. High-dimensional [Kli12b, CEG11]. higher [Baz21, Kua21]. Hilbert [BGT21, GV14, Jeg09, LG20]. historical [MO22]. hit [BS18b, Cla14, RU13]. hit-and-run [RU13]. Hits [And06, BR16]. Hitting [Cho18, JK04, Pes08, Bas10, Bas11, Bor10, Bur19, CK08, GHJ16, HM14]. HJB [XFT20]. Hoeffding [IM16, Maj06, Rao19, Rio13a]. Hölder [Hon19, LD21, Zha20a]. holding [DFK18]. holomorphic [Kin08]. Homogeneous [KS97, Kor05, BMV07, BGT10, EM14, GMW18, Li14, MP13, MPY14]. Homogenization [CP05]. homozygosity [DF16]. horizon [JV09]. Horn [DeB07]. Horn-Shaped [DeB07]. Hotelling [DS06]. Houdré [KPS96]. Hsu [Tud09]. Hua [Ass18]. Hull [ABV03, FV14]. hulls [LMK03, MW16]. Hurst [BZ18, NR18]. hybrid [RR97]. Hydrodynamic [Zho20]. hyper [CSC13]. hyper-exponential [CSC13]. Hyperbolic [BP21, Ist05, BT17, DM21, Le 08, PŻ16]. Hypercontractivity [ST99, HW22]. hypercube [BHP18, KMU16]. hypergeometric [Alt20, Pat07]. hypoelliptic [GM16, GM17, Li16]. hypothesis [JR11].

i.i.d [Zer07]. I.I.D. [Zer02]. Identically [EZ99]. Identifiability [EZ99]. Identification [LW09]. identities [GM12, Pri09]. identity [Ass18, Bar14, BB01, Har16, HM23]. IDLA [Hua19]. II [FXA18, NX13, NS13]. iid [DJ06, Dei09, EP17]. ill [KLvR19]. ill-posedness [KLvR19]. illiquid [NR21]. Illustration [YY13]. images [SXY18]. immigration [FJR20]. impact [Fou13, Fou14]. impatient [HPS14]. implicit [Kaz18]. implies [GKS21, Swa13]. imply [Wei03]. imprimitive [AEK14]. Improved [Hon19, JPR19, QM17]. improvement [Tan17]. impulses [DC15]. incipient [Ald16, BS17, Mic19, Sap11, vB15]. include [Völ20]. Inclusion [KS02]. Inclusion-Exclusion [KS02]. incomplete [Olo18]. Increasing [RZ13, BY13, FY15, Geo10, Wei03]. increment [Glo14]. increment-stationary [Glo14]. Increments [HMO01, KM08, BKR06, GHJL17, HN10, Kli12b, MV14]. Independence [BDN10, BHS11, HS11, Jan20]. Independent [KP04, Pan21, Sch16, Tuc11, BKR06, BB07, BB22, CM12b, GHJL17, Jul19, McV22, PS08, Yor15]. Index [JC04, Mar05]. indexed [BH12, CP14, Ist06]. Indicator [Jun11a]. indices [LL07]. individual [GMRC16]. inductive [HvdHS08]. Inequalities [CHL97, DZ96, MR11, Teh17, BT19, BT12, DG15, Del10, Dzi13, FGL12, GJ09a, GLP18, Goz06, HHJ22, HKZ12a, KLS21, LS13b, MWW11, MZ14a, Mar18, Mon07, Oli10b, Ose10, Ose11, PS19, Rei13, Rio13a, Rio15, Rio17, TM15, Wan14, WY08, dlPP09]. Inequality [Bar05, ESvRS09, Kah03, Li99, Mar10, Pan01, Pan02, Ada15, BBCG08, CdH13, Car22, CGR10, CM12b, Fat18, GS21, Haj14, Har14, HLMV20, HKZ12b, Jon13, Jou12, Maj06, NPY19, Oli10a, Ose08, Ose09, Ose14, Osȩ16, Ose20, Pos09, Pri17, Rao19, Rio13b, Rio18, RV13, Tro11, VZ11, YY18]. Infante [DHI11]. Infinite [BZ06, MR08, AJ14, BS17, BC12, BT17, CY13, DKW14, Dem20, Due06, HJT12, Kab19, Li14, Mic19, RW09, Sap11, SŽ17, VY12b, vB15]. infinite-dimensional [Kab19]. Infinitely [Law98, MR01, DMPARA13, Jan19b, MN09, MU10, Wat12]. infinitesimal [Cla14]. Infinity [Sch99, AP16]. inflation [Her17]. Influence [CPS12]. Information [Val17, GK21, LS13b, MWW11, Olo18]. information-theoretic [GK21]. informational [BD15]. informative [Van08]. inhibition [RL20]. Inhomogeneous [GP01, RT08, Emr16, EM21, GT18, Gan21, NP20]. initial [BCM22, BCP21, BGT10, CPS12, DR20, FV21, Hua17, JR19, Zha20a]. insights [BT19]. Instability [BJ22, HK16]. integer [AAHT22]. integers [DK12, ER09, HL15b]. integrability [Liu15]. Integral [Goz06, HN09, KSY06, Luo14, MU12]. Integrals [MR01, Pec07, KMiS06, MY12, Ngu22, Nut12, Ose09, Ose14, Pri09, Pri15, YLW15]. Integrated [Gao08, MY13]. Integration [Alt20, KS09]. integro [DC15]. integro-differential [DC15]. intensities [Ern21]. intensity [DZ13]. Interacting [Def12, Wan02, Che21b, CGS21, CP17b, Def11, Kua16, Rey15]. interaction [CB10, DHS14, Der16, DP18b]. interactions [CL14]. interchange [KMU16, Mor08]. Interface [AS11, Nie22]. interfaces [SS08]. interlacement [PT11]. interlacements [CZ21, RS11b, Rát15, Szn12, Tas10, Win08]. Intermittency [MNX20, HHN16]. Internal [Hus08, RS11a, Sil20, DCLYY13]. Interpretation [LLN09, Har16, XFT20]. Intersect [Law98]. intersecting [vdBC16]. Intersection [AK08, EP98, BCG12a, ORV22]. Interval [Tan06, BA14, FPRW20, KS14, MLV15, Win20]. intervals [Eri16]. intractable [HRS20]. Intrinsic [KTA17, ADS19, HKZ12a, Sap10]. Introducing [Lau22]. Invariance [CH04, BHM21, DL09b, GMW18, HZ07, Jab17, LS18a, Ore21]. Invariant [DXZ11, GN06, HLWZ15, LT11, Ald21, GH21, Kab19, KM20, LW15, Pin16, Sch20, Tap15]. invasion [Sap11, vdBJV07]. inverse [SXY18, SB07]. inverses [SW10]. inversion [CDNX17]. investment [NR21]. Investor [RR14]. investors [Sai07]. inviscid [Bia13]. involving [Eis21]. irreducible [LW09]. irregular [SSS15]. irrelevance [Lac10]. irrelevant [DR20]. irreversible [RBS15]. ISE [JC04]. Ising [GLP18, LS18b, LW16]. Ising-FK [LW16]. isolate [IM07]. isomorphism [Szn12]. isoperimetric [Dem20, Pet08, PR12b]. Isotropic [CSS99, BMV07, CM12b, Kli12b, Kul16, VA06]. Itô [App02, BBMT09, EW09, GG20, LY16, Pri15]. Iterated [Sep03, AFM20, GG20].

Jacobi [DS15]. Jammed [AHM05]. Joint [HS11, MP16, NS13, GL14]. Jump [Sab21, BY13, CSC13, FPZ16, HKST18, Jab17, Kua16, Lau13, LP12, Mak08, MX19, Sai07, Uem07]. jump-diffusion [CSC13]. jump-diffusions [Jab17]. jump-type [LP12, Uem07]. Jumping [SV04]. Jumps [Dem96, Pes08, Dji18, Hoe09, Mil08, Nob20, Ruf17, RM16, Sok13, Val19].

Kac [Bal09, BZ17, But18, Tak10]. Kac-Type [Bal09]. Kähler [BCY22].

Kanter [Dem11]. Kardar [DDT07]. Kawasaki [KLvR19]. Keller [JTT18]. Kemeny [Pin19]. Kendall [BLR17, BB01, KS07]. Kernel [Rev03, ADS19, ADS20, Buc13, HK16, Pri17, Stu16]. Kernels [Mac02, Mon07, Nic06]. Kesten [Men14, And06, BS17, Kev16, Mar10, Men13]. Killed [Ros20, Sam10, SV04, Aid10, DW16, Dur14, KS14, Ras10, SV08, Tou22, Wag16]. Kinetic [Per22]. kinetically [CM13, Har21]. Kingman [DKW14, DPS15, GWZ21, KP15]. Kirkpatrick [Che21a]. KKL [OW13]. Kob [ES21]. Kolmogorov [Car22, HRKU11, Men11]. KP [Zha22]. KPZ [FV21, QR11]. Kramers [Tug16]. Krein [LG20].

Labouchere [HW19]. lace [HvdHS08]. ladder [HLSX21, Uch18]. làglàd [KT11]. Lagrangian [BJ18]. Laguerre [Def11, KO01, Law08]. Laha [Ejs12]. Lambda [Sch99]. Lambda-Coalescent [Sch99]. laminations [CP11]. Lamplighter [Rev03]. Landau [WZ19]. landscape [Liu22]. landscapes [DT18]. Langevin [CP17b, MR22b, RBS15]. Laplace [Dji18]. Laplacian [Dun20]. Large [ABP00, Arc98, Big04, BP09, BG20, Bjö15, But17, CLS05, DF16, DS06, DJ12, DGH20, FGL12, FX02, GRR14, GH18a, GG20, Gho22, GZ00, Jac14, Kis14, KLM15, KLS21, KP22, KSW12, KS03, LY22, MPY14, Puh19, RA05, SP00, WY08, Yin15, Zer02, dHP14, BBCG08, BS07b, BZ17, Che21c, Com08, DJR16, DL18, DPS15, EMR15, EW17, FJ21, GL09, Har12, HR14, HJT20, Jan19a, LW09, Ora07, Rei13, Sai07, Tsi13, Van07, Yao14, Zhu14, vdBKN12, vdBC13]. Large-time [KP22]. Largest [BGP14, Sos04, BG20, But18, DM21, HW19, Kis14, Sen16, vdBC12]. Last [BM05, Oka14, RT08, CPS12, Gho22, MLV15, Pan21, Zha20b]. Last-Passage [BM05, CPS12]. Lattice [Ham05, Han98, DC13, Geo10, Han99, LTV18, MS11, Yao14]. Lattices [Häg02]. Law [Arc98, AFM20, CLS05, DL08, KDN05, Pes08, RA05, SP00, Sep03, Yao14, Zer02, AEK14, ALW14, BLR17, GG20, HM23, HR14, IM07, Kif15, Kif16, NP12b, NP12c, O’R12, OdS16, Pan21, Sch12, Spi13, Tug16, VY12a, Van07, ZN03, Zer07]. Laws [Mar05, Alt20, Kie97, Ora07, Uch18]. layered [Lej11]. lazy [AH18]. least [EKR22, Yas15]. leaves [Ruf15b]. left [Ban15]. lemma [Che17, FdM07, Men13, Men14]. lemmas [BPR15]. Lenglart [GS21]. length [DKW14, LG09]. length-bias [LG09]. Lengths [CH04, Ald21, Cha15, Cra13]. letter [dHP14]. Level [Jan97, AAGL19, CHL21, Szn19]. Levine [AHM05]. Lévy [Cri21, ME21, APRB11, AI12, BGT21, BM18, Bor22, CM18a, DHI11, Dji18, DMPARA13, DL09b, EK08, EW09, EM14, EP98, GJ18, KS14, KMiS06, Kul16, Lau17, LP08, Mar09, MR15a, MY13, MPP15, Mil08, MW16, PP20, Pes08, PS16, Rie18, San13, SW10, SC09, SXY18, Tap15, Yan06]. Lévy-type [BGT21]. Lie [AAS21]. Lieb [Har14]. Lifetime [DeB07]. LIFO [Lim99]. Lifshitz [Fuk09]. liftings [OS16]. lifts [BD21]. Liggett [CGS21]. Likelihood [Gao03, Che21b]. LIL [KLS05, LL07]. Limit [Bac11, BLL08, BFV10, BR07, Cha10, CCGS19, DV11, DFN00, DBGP03, DC13, Emr16, Kar07, KDN05, Lal03, Lin09, NY10, Roi05, Yuk08, AS22, ACCR13, APS19, AN19, BN17, BI15, CP14, DGK19, Fér20, Flu08, Fun07, GV14, Gou18, GMP18, GPL08, GPPdS14, Hol09, HN09, HN10, IM07, IK18, JTT18, LZ23, LMV22, Mar22, MR22b, NX13, Oça20, Pil17, Rio11, Rok15, RX20, Sal22, Sar21, Ste08, Ste13, Tri19, VR10, Wäs09, Xie20, Zho20, vdHKM09]. Limited [Hus08, DCLYY13, Tod22]. Limiting [BD13, BH16, KLL18, Tod22, FH19, GPHS13, HYZ19, Jan15, JL18, KNN15, SV11b]. Limits [BEK21, Bas15, vZ02, AS16, Bou16, DK12, Gri11, HL13, MU10, Mie08, OS16, Tre13b]. Lindeberg [Kar07]. Lindley [CK18b]. line [ACT19, BVBP20, Che21a, CM22, CS14, KLM15, Möh18, vdBC16]. Linear [AF06, App02, CSS99, KS03, NY10, PR11, Bob08, GJ09a, GIOQ20, GM16, GM17, Unt10]. linearity [CL14]. linearization [GS12]. Lines [AB02]. link [Sch09]. Liouville [Ang19, APS19, BCG12a, GP19, KV15]. Lipschitz [Bah02, BB06, Dem20, DDG^{+}10, Dzi13, GŁM23, Jab17, MS19c, PSY13]. Littlewood [Rao21]. Local [AEK14, Ald98, BS16, DS10, LX19, MR22a, PR18, QR11, Abe15, AP16, ACCR13, DW16, ER19, Fan16, Hon21, HN09, HN10, HK17, HKST18, JR11, KSW12, Mar22, PW11, PW18, Ruf17, Sar21, Sok13, Ste13, Tre13b, vdHKM09]. localisation [Mui15]. Locality [CMT23]. Localization [Baz21, CB10, GGPZ14, HN11, Sab21]. locally [Bah02]. Location [Wan02, Jan13]. Log [Roi05, MG16a, MG16b, Sau19]. log-concave [MG16a, MG16b]. Log-scale [Roi05]. Logarithm [Sep03, AFM20, GG20]. Logarithmic [CHL97, MZ14a, KLS21, Zha22]. logistic [PW11]. lognormal [Nik20]. Long [DCF06, DeB07, FdLS04, Gaá14, Xio04, Aou21, Bas20, Can15, NR21, Puh19]. Long-range [Gaá14, Aou21]. Long-term [Xio04, NR21]. long-time [Bas20]. longest [AS16, Geo10, HM16]. look [DR12]. Lookback [GKH03]. Loop [Law98, Lup16, Mar99, Law14, LW16]. Loop-Erased [Law98, Mar99, Law14]. loop-soups [LW16]. loops [Cha15]. loss [Val17]. Lower [Doh13, GP14, Han98, HJT20, Yas14, Agu18, CG15, Han99, KO06, PS08, Yas15, vdBD20]. Lukacs [Ejs12]. Lyapunov [AGN21, Liu15]. Lyons [Mar98].

M. [Tót13]. M1 [Led16]. Macroscopically [RT08]. magic [Dal17]. Major [NPY19]. Malliavin [GKH03, Lau17, Tan06, TM15, Tre13a]. Mandelbrot [LL15, Nik20]. Manhattan [LTV18, Li21]. manifolds [BCY22, GZ19, KTA17, RM16, Tap15]. Many [BS96, Gne10, Jan19b]. mappings [MU10]. maps [AB14, AAS21, ACCR13, CM18c, Gou18, SW16]. Marcenko [BLR17]. Marchal [DS16]. Marchenko [O’R12, Yas16]. Marcus [Sim00]. Marginals [BS22, GH20, GPL08]. Marked [DGP11]. Markets [SV11a, NR21]. Markov [ADOS11, BA01, BLL08, BY01, BW08, CNPP16, CC98, CD17b, CLMR15, Con16, Cra13, CP14, DG15, Die15, Fit06, Gui99, HR07, HLMV20, HYZ19, Hol21, HKST18, Kua21, KF09, Lin19, LW09, MPY14, Mon07, Mül08, NW15, Pel20, Pel23, Puh19, Rao19, Rao21, RR97, Ros02, SS21, SB07, Tel00, Tou22, Völ20, WY08, YRE16]. Markovian [AGS14, Cha15, DL09a, FPZ16, HMO01, Le 19, Man05]. Martin [IR10, Ras10]. Martingale [AF19, CHL97, DW12, GKS18, Lac10, MZ18, Rok07, TM06, BM18, BJ22, CV07, DGK19, FGM10, Gas19, Kin08, KV13, KZK19, Lau20, Men11, Ose11, OW22, Ruf15a, WH19, Yor15]. Martingales [Dem96, Pec04, vZ02, AP16, BT19, Çet12, DM18, FM21, GŁM23, GGPZ14, JR11, KM17, LL15, MO22, Ose09, Osȩ16, Ose20, PR12a, PR18, Ruf17, Sok13, Tro11]. Maruyama [Kaz18]. Mass [JC04, Mik02, BK13, RS16]. massless [KO06]. matching [GH20, Wäs08]. Matchings [HP03, Gan21, Sal15]. Matrices [DV11, DBGP03, GZ00, Sch16, Sos04, Tuc11, AGN21, Alt17, BD21, BB10, BD13, BGP14, BGZ18, BG20, Bor11, Bor13, BS07b, BHS10, Del10, Dub23, DS15, EM16, ES16, FG13, GL09, HCS08, HKZ12a, HT20, JSS22, Kar09, Lan22, LPP15, MM13, MS19a, O’R12, Oli10a, Ora07, PS08, Pol19, Sim17, Ste08, Tko13, Ven13, Wan18, Woj12, WP14, vSW19]. Matrix [Kös08, Law08, SP00, Sep03, AEK14, Ass18, BS07a, DE23, ES09, GKS21, KS18, KS19, Mec07, Men18, SV11b, SC09, Tro11, Yas14, Yas15]. matrix-exponential [SC09]. Matrix-Valued [Law08]. Matsumoto [KV11]. max [AP16, RS06]. max-continuous [AP16]. max-recursive [RS06]. Maxima [BDM07, Fit00, Lin09, CP19b, KZ13]. Maximal [BYZ07, DY21, Mal15, Ose14, Haj14, HSY15, Kaz18, Ose09, Osȩ16, OW22, VZ11, BYZ12]. maximizing [LS13a]. Maximum [Abe15, BFP^{+}09, Che21b, Gao03, KM08, AI12, Bob08, BDZ11, BZ18, Car18, Cha20, DGH20, GH18a, GT11, Hua18, Jan13, JV09, KM09, LC22, MLV15, Oka14, Tre13a, WZ19]. maximums [Maś18]. Maxwell [Tót13]. McDiarmid [Rio13b]. McDiarmid’s [HLMV20]. McKean [Bas20, Gru23, HM22, Lac18]. McKean’s [GKS18]. MCMC [AA07, RR15]. Mean [CD13, Gao03, JTT18, MP13, BZ16, BBMT09, DJ06, DS15, Dzi13, LW15, NP12a, RS07, Thi20, Uch15]. Mean-field [JTT18, NP12a]. Mean-Square [MP13]. meander [HSY15]. Means [MR08, CP19b]. Measurability [FGM10, Bas10, Bas11]. measurable [OS13, PRT13]. Measure [Gui99, GZ00, KM06, Mar05, RŻ98, ADS19, APS19, Ban15, BHN22, Del10, DGP11, ER09, EM16, Fat18, Fun07, GG14, GL09, Hon21, Hua17, LM06, Le 08, Loe13, NW15, SW10, Teh17, Li17]. measurement [Men18]. Measures [LP99, LT11, Mar10, APS19, Ass18, BBCG08, BB07, BB22, CJK18, DXZ11, DS15, EM14, FGM11, FGM10, GN06, GG11, Gho16, Hue16, LS18a, LR16, MWW11, MM13, Neu11, Sch20, Tko11, Zha12]. mechanics [Gaá14]. Media [RT08]. medium [Lej11]. meeting [Che21c]. Meinhardt [LX15]. Meixner [Ejs13]. membrane [BDKS19]. Memory [DCF06, ESvRS09, Lau22, NR21]. mesoscopic [AS22]. Metastability [FS19, HY21]. metastates [CJK18]. Method [Kar07, ARS22, Dal13, Dal17, FI21, Haf18, KDV17, NW15, Rok14, ST20]. methods [BCY22, GS12, YY13]. Metric [Ist06, ADS19, DGP11, Loe13, RX20, Wan14]. Metrics [FPRW20, BG21]. Metropolis [JS18]. Mézard [Sal15]. Microscopic [GRS03]. Mid [Kul16].

Mid-concavity [Kul16]. Middleton [AHM05]. Mild [FW00, Tap13]. Mineka [Pos09]. minima [WCS19]. Minimal [FZ10, CK14, Rio11]. Minimax [KDN05]. minimization [DGO20]. Minimum [BPR99, Abe15, BLL16]. Minorant [Ber00]. minorants [APRB11]. minus [GT11]. missing [BK13]. mixed [Mon07]. Mixer [Yad09]. Mixing [BZ06, CP20, ČS16, CM13, DM14, Lou04, Ram14, RA05, Roc05, Wil03, Cho18, DP13, ER09, GV14, GJ12, JS18, JM15, Kie97, Kov10, QM17, Sil20]. Mixtures [Big04, MAPS14]. Mod [KNN15]. Mod-* [KNN15]. Model [AHM05, AS11, Gne10, Kor05, MR13, AABP22, AHM06, ARS22, Ave12, BCM22, BS22, BFJ^{+}19, BAMR11, BHJR22, BLZ18, BP10, Bia13, Bjö15, BDKS19, Buc13, CP17a, CP19a, Che21a, Che21c, DDT07, Def11, DR20, DP14, Dol14, Dun20, ES09, ES21, Fou13, FLP20, Gaá14, GŁM23, Gas19, Gra16, GMRC16, GH18b, HL15a, HK15, HY21, JTT18, JF19, JJ16, JR19, KT13, KLS21, KO06, Lag07, LK08, LL20, Li14, Lin19, MNX20, Möh19, Mui15, Pen18, Ram14, RV17, SK15, Sir14, SS06, SS08, Tsa18, TT20, Wan21, Fou14]. model-free [GŁM23]. Models [CG05, Häg02, SV11a, Aou21, Bec22, BJT17, CB10, Car18, CPS12, CM13, DHR18, Def12, Due06, Emr16, FH22, FV21, GLP18, Har21, HJCW20, HA07, JK13, Kua21, Kur14, Lac10, LS18b, Möh11, Pim06, RS21, BJ18]. Moderate [Dem96, DS06, ES09, BM22, CP17a, GJ09a, LY22, Tsi13]. modifications [OS13]. modified [DDT07, Gho22]. moduli [MR22a, Wan18]. modulus [HN09, LX19]. Moment [CLS05, LP08, Nua18, PR11, Pri09, Unt10, AAS21, AJ14, DP18a, HN10, JK13, KM09, MX19, Rát18, Yas15]. moment-generating [Rát18]. Moments [ADOS11, DS10, IM10, Jan13, KV22, MY12, AAGL19, GHSS18, LL15]. monochromatic [Fan15]. Monotone [Bal05, DHS14, Gao08, Mac02, AAHT22, BVBP20, BO03, CDNX17, DM14, Gal20, GS21, McV08]. Monotonicity [Häg02, DY21, DGG^{+}13, Hol15, RS16]. Monte [BA01, Hol21]. most [CM12a, Gho22, KV15]. Motion [Bar05, BBKM00, BGT07, CC98, DeB07, Has05, Ist05, KLS05, LW05, MY99, Aur11, BB06, BA14, BCSW18, BBM22, BH16, BFP^{+}09, Bor10, BZ18, Bor22, BN20, BM20, BO03, BN08, CM12a, CM22, CK08, CHL21, DM09, Far98, Flu08, GKS18, GT11, HCA17, Han21, HH07, HK15, Hon19, Hon21, HLN13, Jan13, JV09, KT13, Law96, LM06, Mai13, Mal15, Mar11, MLV15, MMB18, MS19b, MW09, MW12, NP13, NR18, NX13, NS13, Oça20, ORV22, PW11, Per22, Ros20, Spi13, Tud09, Unt10, VY12b, VY12a, Wag16, ZN03]. Motions [BC98, Gao08, SV04, SW02, Bur19, CK18a, CGS21, GHJ16, Jun11a, KT03, Oka14, Owo15, Par17]. moving [Che17]. mult [ES09]. mult-matrix [ES09]. multi [Due06, FJR20, Kua21, KTT17, RS21, Sau19, Yan07, Yuk08]. multi-dimensional [Due06, KTT17, Sau19, Yuk08]. multi-parameter [Yan07]. multi-species [Kua21]. multi-type [FJR20, RS21]. multiclass [AGS14]. Multidimensional [JKL17, AP14, CEK11, CK18b, LL07, Lyo19, Nee14, Pin16]. multifractal [HK17]. multinomial [BDM07]. multiparameter [GP11]. Multiple [Cou11, Pec07, EMR15, Fra13, YLW15]. Multiplication [RS07]. Multiplicative [Kar08, Aym21, Ber17, DZ19, DW15, GHSS18, Sim11]. multiplicativity [BS17]. multiply [GGNS17]. multiply-ended [GGNS17]. multisets [Pin17]. multitype [CP11]. Multivariate [CNPP16, GG04, MG16b, Mar14, BGHK08, Jan09, Ker17, LL20, LC22, Maj06, MG16a]. mutation [Möh19]. mutations [FH22, Lag07].

Nagaev [Rio17]. narrow [QR11]. natural [HS11]. Navier [BDT11, BKP22, DXZ11, Wan22]. near [Hon19, Jan97, Ker17, Kua16]. near-critical [Ker17]. nearby [Tam07]. nearest [Kov09, LY22]. nearly [Vid14]. Necessary [MR15a, Sch99, Yas16]. needed [IM07]. Negative [GHSS18, Pat07, Pes08, CM22, Dal13, Mic13, NR21, PW18, Völ20]. Negatively [Mar10]. neighbor [Kov09, LY22]. Nemytskii [Gru23]. Nemytskii-type [Gru23]. ness [Ken09]. nested [DS16, MU10]. networks [Ald21, BDL15, CH21, Le 19, VM13]. Neumann [AB02]. neutral [CY13, DC15, Lag07]. never [FF12]. next [Lau13]. next-jump [Lau13]. No [BCW20, CP22, Pes08, Mil08, Nob20]. Nodal [AB02]. nodes [DJ14]. Noise [HT05, ST99, AABP22, BCM22, BC14, DHI11, Dun20, EKR22, FMP17, Gal20, LW15, LD21, MNX20, Mau20, Nee14, PP20, Tre13b, YY18, YRE16]. noises [Gri11, HHN16]. noisy [Ram14]. Non [AK04, CP05, DL09a, HJ18, HLN13, Kar07, Kes96, KO01, KV15, Man05, MR08, Mau20, NP12a, OY01, Pen18, Zer02, AST14, AGN21, AHM06, AV12, AS16, Ave12, AN19, Ban15, BGZ18, Bor11, BFP^{+}09, BN08, CGS21, CF20, DJR16, Def12, Dub23, DT18, FdM07, FPZ16, GMW18, GIOQ20, HPS14, Hus08, Jab17, Kaz18, KT11, KF09, LY22, LS18a, MPY14, MR22b, Nak19, Neu11, ORV22, Puh19, RR15, WCS19, Yor15]. non- [CF20]. Non-amenable [NP12a, AST14, Hus08]. Non-Ballistic [Zer02]. non-central [AN19]. Non-Colliding [KO01, OY01, BFP^{+}09]. Non-Commutative [Kar07, FdM07]. Non-convergence [HJ18]. non-convex [DT18]. Non-Convexity [Kes96]. non-crossing [AV12]. Non-degeneracy [HLN13]. Non-Divergence [CP05]. non-elliptic [Ave12]. non-equilibrium [MR22b]. non-existence [AS16]. non-exploding [KT11]. Non-explosion [Mau20]. non-Hermitian [AGN21, Bor11]. non-homogeneous [GMW18, MPY14]. non-interacting [CGS21]. non-linear [GIOQ20]. Non-Liouville [KV15]. non-Lipschitz [Jab17]. Non-Markovian [Man05, FPZ16]. non-nearest [LY22]. non-normal [BN08]. Non-oscillation [AK04]. non-overlapping [AHM06]. Non-perturbative [DL09a]. non-random [Nak19]. non-reconstruction [KF09]. non-reversible [RR15]. non-self-intersection [ORV22]. non-semimartingales [Yor15]. non-smooth [WCS19]. non-stopping [HPS14]. Non-triviality [Pen18]. non-uniform [Kaz18, Neu11]. non-unitary [Dub23, LS18a]. Non-zero [MR08]. Noncentral [BN17, BS07b]. noncoexistence [Swa13]. Noncolliding [KT03]. Noncommutative [Ejs13]. Nonconventional [KS18, Haf18, Kif15, Kif16, KS19]. nondeterminism [LX19].

Noninvadability [Swa13]. Nonlinear [AK04, FMP17, Mor05, GHJL17, HLWZ15, MYZ21, Ren22, Van08]. nonnegative [Ose10, Ose20]. nonpositive [Woj12]. nonstandardness [Lau13]. Norm [MR01, Tan06, BD21, BS07b, BHS10, Mec07, PS08]. Normal [GG04, BGZ18, BN08, MR13]. normalization [Men18]. normalized [BT19, BGHK08, CD17b, DW15, MZ05a, MZ05b, Spi13, dlPP09]. normalizing [Gou18]. norms [HLN13]. Note [Bal09, Bjo09, CF20, KDV17, KS05a, NY10, Ost14, Pan01, ST17, SW02, Ada15, AF14, AS16, BCY22, Blo23, Bob08, Bor22, CEK11, CSC13, CCH15, DV11, DS16, DGO20, Flu08, GS12, Gor15, HK13, Har12, HL13, HS09, HZ07, HLSX21, Jos07, Kev16, KSY06, Kua21, KS09, Lau17, LMK03, LW16, Lyo18, MN09, Mai13, Men13, Men14, O’R12, PS19, Pan07, Péc19, Pet08, Sab13, Sar21, SŽ17, SC09, Sir14, Tko13, Vid14]. Notes [Car05]. Novikov [Sok13, KS05b]. Novikov-type [Sok13]. Nualart [Nou11]. number [BS18a, CM22, Eva06, Fre12, IM07, Jan20, Mal20, MNZ12, MV14, Oka14, Thi20, Uch15, vdBC16]. Numbers [Arc98, CLS05, RA05, SP00, Zer02, BBMT09, HR14, LY22, Ruf17, Van07, Yao14].

observation [CL09, FG13]. observations [Mak08, Van08]. obtaining [BM20]. occupancy [McV22]. Occupation [Dji18, Gui99, Hoo99, KS05a, LT20, BGT10, LM06, Zho10]. occurring [Mai13]. Ocone [MN08, Pec04, vZ02]. ODEs [Mau20, Mor05]. Offord [Rao21]. Often [KP04, Law98]. old [AC10]. once [HLSX21, Ros08]. One [AS11, RA05, Roi05, Tim19, TYZ12, Wan09, Att10, Aur11, BFRH15, BR16, Can15, CGPPS13, CM18a, DFK18, DE23, EK08, FI21, GJ18, GPHS13, Hol15, Hon19, Hue16, JTT18, Jou12, KSS11, MU12, OdS16, Pet15, Pin19, PR19, San13, ST17, Tri19, Wag16, Woj12, Zer07, Zha20a, TYZ15]. One-dimensional [AS11, Roi05, Wan09, Att10, BFRH15, BR16, Can15, DFK18, EK08, GJ18, GPHS13, Hon19, JTT18, MU12, OdS16, Pet15, Pin19, PR19, San13, ST17, Wag16, Zha20a]. One-ended [Tim19]. one-sided [Aur11, CGPPS13, KSS11]. open [Zho20]. operator [Jac14]. operators [Duc19]. opinions [BP20]. Optimal [BD15, GM16, HHZ18, KS05b, Mik02, MS19b, NR21, Sok13, Zha20b, BJ22, CF20, EW09, FGM10, FPZ16, GIOQ20, Jon04, ME21, Olo18, Rok14, GM17]. optimality [KS07]. Optimisation [RR14]. Optimising [McV08]. Optimization [GL14, Tan17]. Optimizing [Cla14]. Option [Kle02]. Optional [KK15, KT11, Agu18, KS09]. Options [GKH03, Dol14]. Order [Kös08, Ren22, Wan18, BT12, Gaá14, Han21, JPR19, LPP15, LM17, Sch12]. ordered [FW10]. ordering [BJ20]. ordinary [FT07]. organized [Due06, vdBN17]. oriented [CM13, GM21, LK08, LTV18]. origin [Lâm22, Tam07]. origins [Eri16]. Ornstein [BH12, CGXM96, CGXM97, CM18b, CKS99, GJ09a, GP11, Jeg09, MY13, Pat07]. orthant [EM21, SBS15]. orthogonal [BD13]. Orthogonality [Kov09, Kov10]. oscillating [MS19b]. Oscillation [AK04]. Oscillator [BW04].

Overdamped [MR22b]. overlap [GK22]. overlapping [AHM06].

P. [DS16]. pair [AS22]. Palm [Gho16]. PAM [Bec22]. paper [DS16, KDV17]. parabola [GT11]. Parabolic [BCM22, Dun20, EK08, GH18b, HL15a, HK17, JTT18, Jan13, MNX20]. parabolic-parabolic [JTT18]. Parameter [Aly13, BZ18, Mak08, NR18, RW09, Yan07]. parameters [GJ09a, Neu11]. parametrises [Kli12a]. Parametrix [Men11]. paraproducts [KZK19]. parent [FH22]. Parisi [DDT07, Pan05, Pan08, Sal15]. parsimony [RW21]. Partial [CP05, EZ99, CY13, Kaz18, Kri14, LL07, Tap13, Tap15, dBJP13]. Partially [FW10, Kua16, PY17]. Particle [BGT07, BGT10, BJ18, Ven13, BH16, Che17, CCGS19, Che21b, Def12, JTT18, Kua16, TYZ12, TYZ15, VR10, dBM15]. Particles [HT05, Def11, IS17, ST17]. Partitions [Pet10, AV12, BMMU20, FPRW20, Li20, PY17, Win20, Zha12]. partly [YY13]. parts [Alt20]. Passage [BM05, IM10, Kes96, Lal03, Nak19, RT08, Ang19, BT17, BBMT09, BCP03, CPS12, CSC13, CEG11, DG17, Gan14, GM21, Gne21, GP19, Sch09, Vid14, Yao14, Zha20b]. Past [ABV03]. Pastur [BLR17, O’R12, Yas16]. Path [BPR99, BCP03, CHL97, Tan06, AK08, CEG11, FW17, KM09, MP16, de 06]. Paths [BM05, Bas20, CDNX17, Geo10, PZ18, RZ13, Tou22]. Pathwise [BB06, Nut12, CC18, Har16, Hoe09, LY16]. PCA [Lou04]. PDEs [HK17]. Peano [PP20]. Peccati [Nou11]. peeling [CM18c]. Penalisations [Tak10]. penalty [Buo21]. Pennington [Péc19]. Percolation [BS96, BM05, Far98, FdLS04, GM21, Ham05, Kah03, Kes96, KS03, Lal03, PV05, RT08, Sch01, AST14, ATT18, Ald16, Ang19, BT17, BFRH15, Can15, CPS12, Cer14, CMT23, CEG11, Dem20, DG17, DDG^{+}10, DC13, Gan14, GJ09b, GP14, GP19, HM09, HS22, HJM22, Jan19a, Kis14, Lup16, Mar22, Mic19, Pet08, Pim06, PR12b, RV17, Sap11, Sch09, Stu13, Szn19, Yao14, Zha20b, vdBJV07, vdBKN12, vdBC12, vdBC13, vdBC16, vdBN17, vdBD20, Nak19]. percolations [Eis21]. Perfect [AH18, DFN00, Ken04, Ste08]. Periodic [BDT11, Dun20, Gau16, HD20, HD23]. periodically [Wai13]. permanental [MR15a]. permutation [MNZ12, Pin17]. permutations [Bjö15, GMS08, KM20, NP20, Pin17]. perpetual [KSY06]. perpetuities [BI15]. Perron [CD17b, Rok14]. persistence [GLY14]. persistent [Eri16]. Perturbation [LW05, DE23]. Perturbations [App02, KL14]. perturbative [DL09a]. perturbed [DK18]. Petersburg [dBJP13]. Petrov [Eth14]. Pfaffian [TYZ12, TYZ15]. Phase [ACT19, AHM05, HJCW20, BJT17, CZ21, HK15, HJM22, KLS21, Pen18, Rát15]. phenomena [Gho16, Ros08]. phenomenon [Ber10]. Phi [WY08]. phylogenetic [FLP20]. Pickands [AC10]. Pickrell [Ass18]. Piecewise [Ruf17, Hol21, JPR19]. piecewise-deterministic [Hol21]. Pieri [Def12]. Pieri-type [Def12]. pinball [Li21]. pinned [Uch15]. Pinning [CG05, CB10, Car18, Lac10]. Pitman [Bas15, Bau02, CP23, HMO01, MY99]. pivotal [Jon13]. Pivotality [Gal20]. Planar [AAK01, JS00, BB06, BCW20, BCG12a, BM20, CM12a, CM18c, Geo10, GPPdS14, GGNS17, Law96, Law14, Mar11, Mie08, SW16, VY12b, Zer07]. Planck [HRKU11, DT18, Luo14, Ren22]. Plane [LLN09, Gra16, GHJ16, Lup16]. Poincaré [CGR10, BBCG08, GKS21, Jou12, MZ14a]. Point [BL10, HP03, Pet10, Tim04, AG15, Der16, Ern21, Gho16, Glo14, Ili19, LR15, Lyo18, Mai13, Rei13, Stu13, TYZ12, TYZ15, Uch15, vdBD20]. point-to-point [vdBD20]. Points [And06, FJ00, Pes08, ALW14, BKS16, BCW20, BR16, GKS18, Kri07, Sub12]. poisoning [SS06]. Poisson [Arg07, Bal05, BDE13, BFRH15, Bou16, BJT17, CK12, Dal13, Dal17, Dei09, DP18a, GT18, Gne08, Gne21, HJCW20, HM15, KM06, Kri07, MAPS14, Möh11, NP20, Neh14, Oui18, PS20, PS16, Rei13, RW09, Sos04, Val19]. policy [JPR19]. Pólya [CL14, CH21, Jan19b, Lau20]. polygon [FKV23]. polygons [FKV23]. polymer [FJ21]. Polymers [Bir04, Jan97, CdH13, GK22, HS09, Wat12]. Polynomial [Kös08, Oça20, Sab21, CMT23, Sub12, Wan18]. Polynomials [Law08, AGN21, BZ17, But18, Eva06, GLP18, Pin17, Thi20]. polytopes [DGG^{+}13, FV14]. pond [vdBJV07]. population [HA07, JK13, Lag07, Möh11]. Portfolio [RR14]. posedness [KLvR19]. position [Gho22]. Positive [Wei03, BDKS19, Dem11, DJ12, ER09, EM21, Gan14, HM22, Lan22, Mil08, Nob20, PW18, Ren22]. Positivity [BBB97, RU13, BC14, Pan07]. possibly [Ban15]. Potential [FX02, Jan97, BHN22, Che21b, Flu08, Har12, Yan07]. potentials [Bec22]. Potts [Aou21, KLS21]. power [ARS22, ALW14, BN08, FGG21, KT11, MP14a]. powers [HT20, MM13]. Precise [Che21c, GWZ21]. predictable [Sio14]. predicting [AI12]. prediction [McV08]. predictor [CL06]. Preemptive [Lim99]. preferential [HJ18, MP14a, Mal20, Tam07]. prescribed [DJ06, Dei09]. preservation [Ren22]. Price [Kle02, ARS22]. prices [NR21]. pricing [Rok07]. Principle [And06, Sam10, BHM21, DL18, Der16, GMW18, HZ07, Ore21, Pan07, Ruf15b, Wag16, Wai13]. Principles [CH04, GS12, WY08, dHP14]. Priority [MZ05a, AGS14, PN16]. Probab. [MZ05a]. Probabilistic [OD12, XFT20, Hua21, IM07, RSS18]. Probabilities [And06, Kah03, Li99, MY99, AK08, BEK21, BS17, BF11, FH22, HH07, Mar09, Pim06, SS21, vdBD20]. Probability [BDKS19, Gao08, Jun11b, LL20, MLV15, Tel00, AF19, BBCG08, BP20, CG15, CP19b, Doh13, GHJ16, HM14, Jan09, KV15, Kov09, Kov10, Kul16, Law14, PS17, Spi13]. Problem [BDT11, LSY99, Mik02, Aur11, Buo21, CGPPS13, DKJ^{+}19, Fuk09, Gne21, GIOQ20, Li21, LD21, Olo18, Pat07, PP20, Rao21, Rie18, Rok07, Sai07, Wäs09, YY13]. Problems [KS05b, FPZ16, GN06, GS12, Ili19, JPR19, MP20, Men11, Rok14, Sir14, Wäs08]. Process [Ber00, CKS99, FX02, Gne08, KO01, Man05, Mor05, Pat07, Pes08, Sab21, TW03, de 06, Aly13, BL13, BH12, BN17, BS18b, BCP21, Ber10, Bor10, Bor22, Can15, Che17, CCH15, CM18b, CM18c, CK18b, DKW14, DZ13, Dei09, DMPARA13, DP18b, Duc19, EK08, Fen23, GT18, GJ09a, GMT15, GKS18, Gne21, GJ12, GJ09b, GRS03, HK13, Hil06, HD20, HD23, HK11, JK08, Jon13, KMU16, KLL18, Lej11, LM17, Mak08, MY13, Mic13, MP16, MP14b, Möh18, Mor08, Neh14, PS16, Puh19, PŻ16, RW09, Sen16, Squ21, SV16, Tre13a, Val17, VR10, Völ20, YLW15, Zho20, vdBHH10, vdBKN12]. Processes [BL10, BY01, DCF06, EP98, Fit06, HP03, Jan96, KO01, KS05a, Law08, LSY99, LP08, Mik02, Pet10, Tak10, Wan09, APRB11, AM18, AF14, AI12, BKS16, Bas15, BGT21, Beg14, BBMT09, BY13, BS16, BM18, Bob08, Böt11, Buo21, BW08, CGPPS13, CM18a, CSC13, CGP19, CCGS19, CLMR15, Con16, CP11, DGK19, Der16, Dji18, DXZ11, EW09, Ejs12, Ejs13, ESY08, EN20, Eri16, Ern21, EM14, FF12, FJR20, GHJL17, GG14, Gho16, GP11, GJ18, HM09, HRKU11, HM14, HR14, HYZ19, Hol21, HLWZ15, HKST18, Ili19, JL08, Jos07, KR19, Ker17, KLM15, KP22, KS14, KMiS06, Kri14, KT11, Kul16, KTT17, Lat08, Lau20, Led16, Lyo18, MYZ21, Mai13, Mar09, Mar18, MR15a, MR22a, McV22, MPP15, MX19, Mil08, MPY14, Möh11, MW16, ME21]. processes [Nob20, Ore21, Pan21, Par17, PW18, PS16, PR19, Pro18, RL20, Rei13, Rie18, San13, SW10, SC09, Sio14, SV08, SXY18, Stu13, Tap15, Tim04, Tou22, TYZ12, TYZ15, Tsi13, Val19, WY08, WCS19, Yan06, dlPP09, vZ08]. Product [DZ96, GHJL17, Bor11, BW08, DP18a, HZ07, KM20, Lac15, Ros08, Sim17, Tko13, Wan18]. product-form [BW08]. Products [AV12, DBGP03, RW02, Tuc11, AGN21, BFY07, EHW15, HJT12, KS18, LS18a, NY09, Pro18, Ste08, Zhu14, KS19]. profile [Dem20, IK18, Mar22, Pet08]. progeny [Aid10, Gho22]. programming [GS12]. Progressions [BYZ07, BYZ12]. progressively [OS13]. Prohorov [Loe13]. Prohorov- [Loe13]. projection [Mec09]. Projections [GS09, KR19, MMB18, Kab19]. projective [LMV22]. Proof [And06, BLL08, CHL97, Kar07, KPS96, AC10, BBCG08, Bar14, BJ20, CGS21, CP23, Duq09, GK21, IM07, MW09, MW12, Nou11, Oza20, Rát15, RSS18, Sim11, Tug16, VY12a, Wäs09, Yan06]. proofs [CM18a]. Propagation [CP19a, GJ18, Lac18]. Properties [BCH^{+}00, Mar10, BCG12a, BS07a, DN07, Ejs12, KMiS06, KTT17, PS20, Sau19, VY12b, WP14, YE13]. Property [BM05, CC98, DL09b, Eth14, Gas19, HYZ19, KV11, LG09, Nie22, QR11, Ruf15a, Zha20a]. prophets [HPS14]. proportions [HJ18]. Protected [DJ14]. Prudent [BFV10]. Pseudo [HLN13, Sal15]. pseudo-dimension [Sal15]. Pseudo-norms [HLN13]. Pure [BY13, NP13]. pursuit [Mec09].

quadrangulations [Mie08]. Quadrant [Gne08]. Quadratic [Tan06, CC18, DKW14, Har04, HKZ12b, Jac23, JKL17]. quantiles [BCG12b]. Quantitative [BLMZ12, BCP21, Mec09, Ros02, AJ14, Con16, HLMV20]. Quantities [IM10]. Quantization [Jun11b]. quantizers [Yuk08]. Quantum [FW00, NP13]. Quasi [FX02, HY21, Lab13, LS18a, ORV22, Ban15, BS17, BC15, BGT10, GJ18]. quasi-homogeneous [BGT10]. Quasi-invariance [LS18a]. quasi-left [Ban15]. quasi-multiplicativity [BS17]. Quasi-Potential [FX02]. Quasi-stationary [HY21, Lab13, BC15, GJ18]. Quasi-sure [ORV22]. quaternion [BCY22]. Quenched [BHM21, GMP18, Fuk09, GPHS13, GPPdS14]. Question [Pan05]. Questions [BS96]. queue [AGS14, AS16]. Queues [Lim99]. quickest [Buo21]. QuickSort [FH19, DFN00, FJ00, Jan15].

Rademacher [Pin16]. Radial [Jun11b, GMW18]. radically [BP20]. Radii [Jun11b]. radius [FXA18]. Radix [EW17]. Rains [Zha22]. Ramanujan [Her17, Oza20]. Random [ABV03, Aym21, BW03, BKR06, BYZ07, BR07, Bir04, Bjo09, BD02, CJ13, CP11, DBGP03, FZ10, GP01, GG04, GS09, GGA10, HMO01, HN11, IM10, Jan19b, JK04, Jul19, KS10, KP04, Lau22, LTV18, LW05, Mar99, MR01, MR13, MR11, NP13, OY01, Pet10, PV05, RA05, Roi05, Sch16, Sos04, Tas10, Tre18, Tuc11, Wäs08, Win08, Zer02, vSW19, AAHT22, Abe15, APRB11, Ada15, AB14, Aid10, ACT19, AG15, Ale13, Alt17, ADS19, ADS20, Aou21, AV12, Ave12, Aym19, AFM20, Bac11, BM22, BD21, Ban15, BEK21, BB10, BS18a, Baz21, Bec22, BCW20, BCNP19, BHP18, BGP14, BGZ18, BYZ12, BFT13, BFGG^{+}16, BS18b, BB07, BB22, BF11, BLZ18, BHM21, BP09, BG20, Bjö15, BMMU20, BK11, Blo23, Bor11]. random [Bor13, BS07a, BS07b, BT11, BT20, BT21, BJT17, Buc13, But18, CZ21, Car18, CGPPS13, ČS16, Cha10, CGP19, Che21c, CJK18, DK18, DGK19, DJ06, Dei09, DV23, Del10, DL18, DFK18, Dem11, DW12, DW16, DGG^{+}13, DJ14, DM21, DG17, DGdHM21, DL09a, DK12, DJ12, Duc19, Dur14, DS15, DGH20, ER09, EP17, EM21, Eva06, ER19, FvdHH16, FXA18, Flu08, FKV23, FS19, FHJ18, FV14, FG13, FY15, Gan21, GRR14, GH18a, GLY14, GM13, GM22, GH21, GMW18, Gho22, Glo14, GMS08, Gri11, GPL08, GPHS13, GPPdS14, GGPZ14, Hil06, HCS08, Hol09, Hol15, HM16, HKZ12b, HKZ12a, HZ07, HK16, Hua18, HLSX21, Hue16, HP15, Hut11, IR10, IM07, IK18, Jac14, JSS22, JL21, Jan20, JK08, Jon04, KM20, Kar09, Kie97]. random [KS18, KS19, KV22, Kli12a, KM17, KSW12, KMU16, Kri21, KTA17, KO06, KTT17, LT20, Lâm22, LS13a, Lan22, Law14, Le 19, Le 08, LY22, LMV22, LW16, MZ18, MNZ12, MV14, Mec07, MM13, MS19a, Men18, Mie08, MHC13, Nak19, O’R12, Oli10a, Ora07, OdS16, PS08, Pen22, Pet15, Pil17, PY17, Pol19, PT11, PZ18, QM17, RS11a, RS07, Ras10, RS11b, Rát15, RS16, Rio15, Sch12, SK15, SV11b, Sin14, ST20, Squ21, Ste13, SW16, Szn12, Thi20, Tko13, Tod22, TT20, Uch15, Uch18, Van07, Ven13, Vid14, Vys21, WYY13, Wan17, Wan18, WH19, Wan21, Wäs09, Woj12, WP14, Xie20, Yas15, Yin15, Yuk08, Zer06, Zer07, vdHKM09, JL21]. random-cluster [Aou21]. Random-to-front [Bjo09]. random-to-random [QM17]. randomized [HHZ18, Oui18]. Randomly [MZ05a, MZ05b, BHS18, BP09, LTV18]. randomly-oriented [LTV18]. Range [FdLS04, Aou21, Can15, ČS16, Der16, Gaá14, GJ09b, Sav14, Spi13]. ranges [DS16]. Rank [BF11, JSS22, BZ16, BS07a, DE23, Rey15]. rank-based [BZ16, Rey15]. rank-one [DE23]. Ranked [CH04]. Rapid [JM15, GJ12]. rapidly [PS20]. Rate [KM08, Yad09, Cha20, Cla14, CP17b, FJ21, Lâm22, LW09, Mil08, Oça20, Vys21]. Rates [Ros02, Kua16, NX15, Völ20, Wan17]. ratio [Hua18, Ose20]. rational [Dji18]. re [DL09b]. re-rooting [DL09b]. Real [Aym19, Kös08, BF11, Sim17]. Realized [FM21, KT11]. rearrangement [BVBP20]. Reciprocal [CLMR15, VY12b]. Reconstructing [GN14, JL21]. Reconstruction [PR11, KF09, RW21]. record [BCSW18, LY22]. Records [Gne08, GM12]. Recovering [LY16, Bor22]. Recurrence [BC12, BD02, Car05, DP14, GGNS17, KTT17, Mül08, Sin14, Zer06, ADOS11, BFGG^{+}16, CK18b, DHS14, HS12, Ker17, SHH14]. Recurrent [Fit06, GP01, KP04, BFT13, DK12, Hua19, JL21, KZ13]. Recursions [BDZ11]. recursive [IM07, IK18, RS06]. redistribution [BA14]. reduction [Buo21, RBS15]. Redux [KS02]. Reflected [DM09, Owo15, Bur19, CEK11, DW16, Fan16, IR10, MLV15, MP16]. Reflecting [BC98, SW02, BB06, Kua16]. Reflection [BDT11]. reflections [BBF18, PS19]. Regeneration [BLL08]. regenerative [Ore21]. regime [AS22, KF09, Sen16]. region [ADK22]. Regular [FGM11, CCGS19, Che21c, Eas22, RZ13]. Regularity [Arc98, BGT21, Kaz18, Tre13a]. Regularization [Bec22]. Reinforced [MR11, Sab21, ACT19, CH21, Hol09, HLSX21, Sch12, Sin14]. reinforcements [Sin14]. Related [BGT07, BCH^{+}00, IM10, MY99, Bec22, Beg14, Bjö15, DOS16, DS15, GG20, GM12, Gra15, Haj15, MP20, Par17, Ras10, RU13]. Relation [Jan96, Le 08]. relations [HA07, NY09]. relationship [SV08]. relative [CG15, Fen23, RS16]. relaxation [CP17b]. relaxing [Ose11]. remainder [Don20, Don22, Yan22]. Remark [HN11, Pec04, HT20, vZ08]. Remarks [BM20, FW17, Tha98, CV07, Sau19]. removal [SV16]. Renewal [DS10, ESY08, Bas15, DK18, Don20, Don22, EN20, MZ14b, Tsi13]. renewal-reward [Tsi13]. Rényi [ADK22, BD15, DL18, GM22, JL18, Kif15, Kif16, Rát18]. repeated [Hou09, CP20]. replacements [Jan19b]. replica [GK22]. Representation [AED13, CHL97, FPZ16, Mor05, OY01, BPR13, CD17b, HK13, HN09, HK11, HA07, JK13, Jan09, KT13, MO22, OD12, Pan10, RW09, Tho16, Völ20, XFT20]. Representations [GP11, AF19, BPR15, DMPARA13]. repulsion [BCW20, Pet08, Ven13]. repulsive [GIK22]. Rescaled [vZ02]. resembling [BDE13]. resources [vdBHH10]. respect [CG15, KS09]. Respondent [AR18]. Respondent-Driven [AR18]. Result [And06, Häg02, War99, AC10, Cha15, MW09, MW12, Pan10, PN15, PN16, SXY18]. results [AGN21, BB07, BB22, CZ21, CP14, DHR18, DJ12, Geo10, Gor15, LL07, SS06]. Resume [Lim99]. Resummed [Mor05]. Retraction [BT21]. return [KV15]. reversal [BL13]. Reversibility [LSY99]. Reversible [Gui99, Tel00, Blo23, HP15, RR15]. Revisited [AS11, BP10, BB01, Har21, Hut18, Neh14, Ros08, Zer07]. reward [Tsi13]. rewards [HPS14]. richness [LMK03]. Riemann [Ost14, Oui18]. Riemannian [FW17]. Right [SW10, Gho22, Vid14]. right-continuou [Vid14]. right-most [Gho22]. rightmost [BH16]. Rigid [LW05, Liu22]. Rigidity [KN19, Gho16]. rigorous [CZ21]. Ring [Jan97]. risk [AGS14]. risk-sensitive [AGS14]. Robbins [Tud09]. robust [BEK21, PRT13]. role [CL06]. Root [KDN05, AM18, But18, ESY08, IM07, RW21]. rooting [DL09b]. Rosenblatt [YLW15]. Rotatable [ER19]. rotationally [Tko11]. Rotor [HMSH15, Cha20, HS12, SHH14]. rotor-router [HS12, SHH14]. Rotor-routing [HMSH15]. rotors [CP17b]. Rough [Ore21, BCM22, Gas19, MNX20, NR18, ORV22, Ros20, Yan22]. Route [Ald21]. router [HS12, SHH14]. routing [HMSH15]. Rowlinson [Häg02]. Rudelson [Oli10a]. Rudvalis [Wil03]. ruin [CP19b]. rules [vdBN17]. Rumor [FY15]. run [RU13]. runs [MAPS14]. Russell [Mar98]. RWRE [GN14].

SABR [GHJ16]. same [Cou11, KM20]. Sample [SP00, Sep03, Tan06, BLL16, CP19b, HCS08, Yas14, de 06]. Sampled [DBGP03, BZ18]. sampler [Wan17]. samplers [RBS15]. Sampling [AR18, BA01, Gne10, Han05, FH22, JV09, Jon13, Ste08, Wan14]. Sandpiles [SW16]. Sanov [MZ14b]. Sasamoto [JF19]. Sasamoto-Spohn [JF19]. satisfies [Zha22]. Sausage [ABP00, Öz20]. Scalar [GKS21]. Scale [ALW14, Ber10, Nob20, Roi05]. Scale-free [ALW14]. scaled [KR19]. Scaling [BFV10, DK12, JF19, Lan22, NY10, Fun07, HL13, Hol09, LG09, Mie08, VY12a]. scan [Wan17]. sceneries [GPPdS14]. scenery [CGPPS13, GPHS13, Tre18]. scheme [FT07, FI21, GL08, Kaz18, LY13]. Schreier [OW13]. Schrödinger [Jac14]. Schur [Li20]. SDE [AP14, Att10, BO03, FF12, Haj15, Swa01]. SDEs [BRT10, ESvRS09, GŁM23, Gru23, LST15, Lyo19, Wan22]. search [KLL18]. Second [Kös08, LM17, NP12b, NP12c]. Second-Order [Kös08]. secretaries [FW10]. seed [Nie22]. seen [GN14, Kli12b]. Segel [JTT18]. segment [vdBC16]. Selberg [Aym21]. Selection [Han05, CM18b, FH22, Fou13, Fou14, Gne21, JK13, Rok07]. Self [BY01, CD17a, ES21, Fit06, Gau16, MZ05a, MZ05b, AST14, BP21, BT19, BGHK08, BS16, Due06, Gla15, Hut18, Jos07, Neu11, ORV22, Tug16, dlPP09, vdBN17]. Self-averaging [CD17a]. self-avoiding [Gla15, Hut18]. self-destructive [AST14]. Self-diffusion [ES21]. Self-normalized [MZ05a, MZ05b, BT19, BGHK08, dlPP09]. self-organized [Due06, vdBN17]. Self-Similar [BY01, Fit06, BS16, Jos07, Neu11]. self-stabilizing [Tug16]. selfadjoint [WP14]. selfadjont [Woj12]. selfsimilar [KTT17]. Semi [Li17, CGR10, KTT17]. Semi-classical [Li17]. semi-groups [CGR10]. semi-selfsimilar [KTT17]. semicircle [AEK14, AK08]. semicircular [Bou16]. semigroup [GN06]. Semigroups [Rin98, DZ19, Völ20]. Semimartingale [Kle02]. semimartingales [GL14, KK15, KS09, Yor15]. s’enfuir [AB14]. senile [Hol09]. sensitive [AGS14]. Sensitivity [DP13, JR19, Aly13, AABP22]. separability [BPR13]. separated [JKL17]. separation [LR16]. sequence [BS18a, For20, Sar21]. Sequences [BZ06, EP98, EZ99, Lin09, CD17a, CV07, ER19, For21, Gou18, Kie97, McV08, RS06, dHP14]. sequential [CP19b, Gne21]. series [ARS22, Aym19, AFM20, ESY08, HK13, HK11, Oli10b, Zhu14]. serve [AS16]. serve-the-longest-queue [AS16]. Set [FJ00, ATT18, BH12, BCSW18, Cla14, DR20, Kin08, Mar17, Pin17, Rát15, ST20, Szn19]. set-indexed [BH12]. Sets [GP01, JK04, AAGL19, CHL21, FPRW20, GHSS18, Glo14, Jab17, KN19, Lat08, Teh17, Tko11]. setting [BRT10, GMP18]. sewing [FdM07]. shadow [BHN22, LX15]. Shape [Lal03, GMT15]. Shaped [DeB07]. shapes [Emr16]. Sharp [Aou21, GL08, HJM22, Mon07, Ose08, Ose09, Ose10, SV04, Wat12, Yas15, Fat18, GS21]. Sharpness [GS21, OW13]. sheet [WYY13]. Sherrington [Che21a]. Shields [BP10]. Shiryaev [KS05b]. Shkolnikov [Har16]. shock [GRS03]. Short [CM18a, Bar14, CEG11, CP17b, Kua21, Rát15, Sim11, Yan06]. Shotgun [GM22]. shrinking [Öz20]. Shuffle [Wil03, QM17]. shuffles [Bjo09, JM15]. shuffling [AH18]. shy [Ken09]. shy-ness [Ken09]. sided [Aur11, CGPPS13, CM18a, Fér20, KSS11, ME21, Pat07]. sigma [Pel23]. Signature [CDNX17, DGdHM21]. signatures [NX15]. signed [NS13]. Similar [BY01, BS16, Fit06, Jos07, Neu11]. Simple [CHL97, LPN22, Ros02, BBCG08, Fre12, FG13, HL15a, KO06, Möh19, MW09, MW12, Tug16]. simply [Mar11]. Simulation [DFN00, Ken04, Lej11, DZ13]. Simulations [PW96]. Simultaneous [Bur19]. sine [VY12a]. single [Che21b, FPZ16, Gri11, JL21, SV16]. single-trajectory [Che21b]. singular [HT20, LST15, Luo14, Yas15]. Singularities [Alt17, BN20]. Site [BFGG^{+}16, Ale13, DC13, McV22, Mui15]. sites [CM12a, Uch15]. six [Lin19]. size [AS16, GG11, HW19, LX22, Sap10, vdBJV07, vdBC12]. sizes [Rát18, vdBC13]. Skew [Sch16, EHW15]. Skew-Diagonals [Sch16]. skew-products [EHW15]. Skewed [BBKM00]. Skip [HMO01]. Skip-Free [HMO01]. Skorohod [BPR13, BPR15, BM20, MNX20, Ngu22, Pri09, Pri15, Tho16]. Skorokhod [Gro19, Har14, Led16, LV20, MP20, YY13]. SLE [AK08, Dub03, LR15]. sliding [AF14]. Slow [BA01, Wai13]. slow-fast [Wai13]. Slowdown [BCNP19, DFK18]. slowly [Mui15]. Small [Gao08, Jeg09, Li99, Mar09, Nik20, RŻ98, BBMT09, BZ18, BC14, Car22, DHI11, GL08, KM20, KP22, LC22]. small-time [KP22]. smallest [HT20, Yas14]. Smooth [Sam10, Agu18, Lau22, VZ11, WCS19]. smoothing [CdH13, DM18, JTT18]. Smoothness [ZN03, CF20, GG14, Lau17]. snake [BC12]. Snell [Agu18]. Sobolev [CHL97, HLN13, KLS21, MZ14a, WY08]. Soccer [Bar97, Bar98]. Soft [Geo10, Pen22]. Solution [Kuz00, AG15, LY16, QR11, SSS15]. Solutions [AK04, FW00, Swa01, Bah02, CEK11, D’O10, DM18, Gru23, Kur14, LST15, LP12, LD21, LY16, OD12, Tap13, Unt10]. solvable [CM18b, DHR18, EN20]. solving [YY13]. Some [BY01, BGT07, CZ21, Car05, CV07, CP14, DZ96, DPS15, Eis21, HM09, HYZ19, KMiS06, LL07, MY99, Pan02, Pin17, SS06, Tha98, VY12b, YE13, BGT21, BLMZ12, Che21c, DOS16, DJ14, Ejs12, HLN13, HA07, Jan20, KTA17, LY22, MU10, Mar09, Men11, MX19, Möh11, Nua18, Ose11, PS16, Sau19]. sort [EW17]. soups [LW16]. source [FY15]. Space [CKS99, BCM22, BBM22, DP18a, FW17, GHJL17, HW22, HN10, IR10, Jan09, Lau17, Le 08, LV20, Loe13, Mar17, PS16, Pri09, Win20]. space-fractional [PS16]. Spaces [CHL97, DZ96, Ist06, AF19, BMV07, BGT21, BC15, CK18a, CV07, DGP11, GV14, HCS08, Jeg09, Kie97, Loe13, MN08, MP13, MR15b, Oto09, RX20, Ros08, VZ11]. spacings [NP20]. Spanning [ABIT22, Mar99, Tim19]. Sparse [AR18, BGP14, BDL15, Hua18]. Spatial [Lou04, NZ20, Ald21, Dei09]. Spatially [KS97, LD21]. SPDEs [AED13, LD21, SSS15]. special [BS07a]. Species [Gne10, BAMR11, BLZ18, Kua21]. Spectral [BS07b, BHS10, DOS16, GZ00, MM13, Mor08, RR15, Sch16, BD21, BD13, Del10, DGdHM21, DS15, EM16, FW17, GL09, JL18, KP22, KP15, Mec07, MP14b, Möh18, Möh19, Ora07, PS08, Pol19, WP14]. Spectrally [Pat07, CM18a, Mic13, MPP15]. Spectrum [BDN10, Kar09, Bor11, Duc19, Sim17]. specular [BJ18]. Speed [KS03, ADS19, Blo23, BT20, BT21, Gan14, GM16, GM17, Hol15, LS13a]. speeds [DKJ^{+}19]. Spent [Jan97]. sphere [Gau16, JK08, Tan17]. Spheres [Spr07]. Spherical [Ist05, BS22, FXA18, Kab19, MMB18, Tod22]. spherically [CP22, Pin16]. sphericity [Mie08]. spiked [BG20]. spin [BS22, CM13, Kli12b, KLS21, Kua21, KO06, LS18b]. spine [HR14, RSS18]. Spitzer [Tud09]. splitting [GH21]. Spohn [JF19]. spread [RV17]. spread-out [RV17]. spreading [FY15]. Square [MP13, Tan06, AM18, ESY08, Jou12, MR22a]. square-root [AM18, ESY08]. Squared [KO01, PW18]. Squares [MR08]. Srinivasan [Jon13]. St. [dBJP13]. Stability [App02, FP11, HRS20, HK16, RL20, ST99, CL06, DC15, Gal20, MY20, Nee14]. stabilizing [Tri19, Tug16]. Stable [Mar05, Pat07, Pes08, Rie18, RŻ98, Tak10, BGT21, Beg14, BFRH15, CGPPS13, Dem11, DXZ11, Gra15, HK13, HK14, HK11, Jun11a, KP22, Mai13, Mic13, Pro18, Sim11, SXY18, Uch18, Van08]. staircase [Li20]. Standard [GG04, Jan09, Jan97, Nic06, Nob20]. Standardized [KM08]. Standardness [Lau13]. State [Lon04, Wan02, Die15, DP18b, FF12, FJR20, MYZ21, Rok14]. statements [DP18a]. states [Alt17, BGT10, NP13]. stationarity [MR22b]. Stationary [BL10, DJ06, Dei09, KS05a, Tsa18, BC15, BCP21, BW08, DM09, Glo14, GJ18, HY21, KLL18, Lab13, LMV22, Pel23, Rey15]. Statistic [GG04, Gri02]. statistical [DR12, Gaá14]. Statistics [DS06, Sos04, AS22, Arc98, BT12, Fér20, LPP15, Oui18, RW02, Wan18]. stay [DJ12, KS10, Mil08]. stays [BBM22]. Stefankovic [GJ12]. Stein [Cho18, DY21, Dal13, Dal17, Gol22, Haf18, KDV17, LS13b, MG16a, MG16b, Pri15, ST20]. step [GRS03, KV22]. sticky [HCA17]. still [JL08]. Stochastic [AAS21, App02, AK04, Att10, Bar05, BDT11, BJ18, CR05, FW00, Fra13, HN09, Jab17, KL14, Law08, Lin99, LT11, MR01, Mik02, Rok14, Tou22, AF06, AP14, Ban15, BC15, BLZ18, Bia13, BKP22, CD13, Cri21, CY13, DOS16, DC15, DGO20, DXZ11, Ern21, FT07, FP11, FM12a, FM12b, FI21, GN06, GS12, Gor15, HCA17, HY21, Hoe09, HHN16, HK17, Hua17, HA07, Jac23, JPR19, Kaz18, Ker17, Kua21, KS09, Kur14, LX19, Lej11, LP12, LX15, LZ23, Lin19, LY13, Mak08, MX19, MR15b, MW09, MW12, Nic06, Nie22, Nua18, NZ20, Nut12, Ose09, Ose14, Owo15, Rie18, Ruf15a, RM16, Sai07, SŽ17, Sir14, Ste08, Tap13, Tap15, Tri19, Unt10, Val19, VZ11, Wai13, Xio04, YY18, YE13, YRE16, Zha20a]. Stochastically [Mac02, Sap11]. Stock [Kle02]. Stokes [BDT11, BKP22, DXZ11, LY13, Wan22]. Stone [Com08]. Stopper [Wee06]. Stopping [KS05b, CF20, FPZ16, GIOQ20, HM22, HHZ18, HPS14, KT11, MS19b, ME21]. Strassen [Lin99]. strategies [HHZ18]. strategy [Jon04]. stratified [BHP18, Jul19]. Stratonovich [LY16, Mau20]. stretched [DGH20, GRR14]. stretches [Sch09]. Stretching [CGXM96, CGXM97]. Strict [Fit00, Lal03, Pet10, GIOQ20, Hol15]. strictly [DJR16, Vys21]. string [Lâm22]. Strip [AAK01]. strips [Zer06]. Strong [BZ06, BG21, CLS05, Gan21, Gru23, Kie97, LP12, Pet15, Zha20a, BZ16, DW15, FGM10, GV14, HR14, Kur14, Lac18, Ose10, Sir14, Van07]. strongest [GMRC16]. Strongly [Lin09, MG16a, MG16b, Thi20]. structure [CHL21, GRS03, KM20, Kar09]. Student [GG04, Gri02]. studies [AM18]. Sub [BGT07, MA17, RV13]. Sub-Fractional [BGT07]. sub-Gaussian [RV13]. sub-Gaussianity [MA17]. subclasses [MU10]. Subcritical [HS22, ATT18, Aou21, APS19, DM21, GMRC16, Sen16]. Subdiagonal [Res01]. subdivision [Hou09]. subexponential [DW12]. Subgaussian [DG15, HKZ12b]. subgraphs [FvdHH16]. Sublinear [Mal20]. submartingales [Ose08, Ose10]. submatrices [CL09]. Subordinate [SV04, Osȩ16, SV08, Wag16]. subordinates [Ose08, Ose10, Ose20]. subordination [Ose11]. Subordinator [And06]. Subordinators [BY01, DS16, MY12, Sav14]. subsequences [HM16, NS13]. Subsequential [Kri21, WZ19]. Subsets [BYZ07, BYZ12]. subtrees [DJ14, Jan21]. Sudakov [Pan10]. sufficiency [HHZ18]. Sufficient [RW21, Sch99, MR15a, Yas16]. Sum [MZ05a, MZ05b, BD13, Bor11, DM09, Kri14, Nic06, Yor15]. sum-of-exponential [DM09]. summability [Kie97]. Sums [EZ99, Oli10a, RW02, AAHT22, BB07, BB22, BGHK08, CM12b, GRR14, GH20, Gri11, Haf18, HJT12, HKZ12a, HZ07, Kif15, Kif16, LL07, LX22, NY09, Oli10b, Rio15, dBJP13]. super [Hon19, Hon21, LM06, Ros20]. super-Brownian [Hon19, Hon21, LM06, Ros20]. Supercritical [Hut11, Cer14, CCH15, Dem20, GMT15, PR12b]. Supercriticality [VM13]. superdiffusive [KV22]. superhedging [PRT13]. supermartingales [FGL12, Haj14]. Superprocess [KS97, Zho10]. Superprocesses [Wan02, Xio04]. Support [Kar08, Sim00, DR12, Zho14]. suprema [Mar18]. supremum [DK18, FXA18, HK13, HK11, Mic13, MPP15, Pro18, ZN03]. Sure [App02, Lin09, AN19, BS18a, Fre12, HR14, LZ23, ORV22, Zho10].

Surface [CGXM96, CGXM97]. Survival [BK11, EP17, HH07, BAMR11, Fuk09, Kul16, SV16, Tou22]. swallowed [CM18c]. swaps [AH18]. switched [BLMZ12, BHS18]. switching [Olo18]. Sylvester [Zhu14]. Symmetric [Ave12, GP01, GG04, GS09, Han05, HL15b, Kös08, Tak10, BCP21, CG15, CP22, HYZ19, HCS08, KF09, PS08, Pro18, San13, Tko11, Uem07, VR10, Zho20]. Symmetrization [Pal08]. symmetrized [Fat18]. symmetry [Bar14]. symplectic [Def12]. System [LX15, TW03, Eva06, HW19, Kua16, ST17]. Systems [BD02, BGT07, NY10, BLMZ12, BGT10, BFP^{+}09, Che21b, Ern21, Jac23, Swa13, TYZ12, TYZ15, Ven13, Wai13]. Sznitman [KP15, MP14b].

tagged [CCGS19]. Tail [Aid10, HKZ12a, Maś18, FV21, Fuk09, GT11, HKZ12b, Jan19a, Lyo18, Ose10, Pel23, PR15, RRZ11, RS06]. tailed [BHS10, But18, Cha10, DHI11]. Tails [FdLS04, Jun11b, Sos04, DGH20, FH19, Jan15]. Talagrand [MW12, Fat18, MW09, Pan01, Pan07]. tame [For20, For21]. tamed [Sab13]. Tameness [Lon04]. Tanaka [Haj15, KSS11]. tangent [CV07]. TAP [GIK22]. target [Ale13, Cla14, DGO20, Sai07]. targets [JS18]. tau [BLR17]. Tauberian [Gor15]. Tauberian-type [Gor15]. Tchebichef [Rio18]. technique [BM20]. techniques [Men11]. tempered [Gra15]. Temporal [Lou04]. tensor [Tko13]. tensors [BF11]. term [NR21, Puh19, Xio04]. terms [BCG12b]. Tesselation [Hou09]. test [JR11]. Testing [DR12]. th [Han21]. Their [ABV03, BGT10, Eri16, Ern21, HRKU11, Ose08, Ose10, Ose20]. Theorem [Bau02, BCH^{+}00, BDM01, HMO01, Lin09, Lin99, Mar98, MY99, Roi05, AS22, BN17, BPR13, BT11, CP17a, CGS21, CD20, Cri21, DK18, Don20, Fér20, GK21, GV14, GMP18, GPL08, GPPdS14, Han21, HN09, HN10, IK18, Kev16, LMV22, MZ14b, NX13, Pil17, PY17, RSS18, Rio11, Rok15, Sal22, Ste13, Szn12, Tap13, Tót13, Tsi13, VR10, XFT20, Yan07, Yas16, vdHKM09, BLL08, BR07, Cha10, DV11, Kar07, Don22]. Theorems [DS10, AED13, AN19, BI15, CCGS19, Com08, DP18a, Gou18, HYZ19, HCS08, LZ23, Par17, RX20, Tri19, Tud09, YRE16, Yuk08, DBGP03]. theoretic [GK21]. Theory [Ost13, Yan07]. thermostats [CP17b]. thin [Lat08]. Thinning [Bal05, ALW14, FGG21]. third [HN10]. Thouless [Che21a]. Three [Gra15, HJ18]. Threshold [Ros08, CK12, FvdHH16, HHZ18, RV17]. Tightness [BKS16, Gri02, SS08, BDZ11, Car18, HYZ19, Kri21, Ngu22, WZ19]. Time [Ald98, BGT21, Han98, Hoo99, Jan97, JK04, JS00, Kes96, KS05b, RR14, Wil03, BL13, Bas20, Ber10, BGT10, BB01, CP20, Cer14, ČS16, CM13, Cho18, CF20, Cla14, Dji18, Dol14, Eri16, EM14, FI21, FJ21, GH21, GL08, HRKU11, Han99, Hon21, HN09, HN10, HK16, Hua19, Jeg09, Kaz18, Kli12a, KP22, Lau13, LX15, Mar09, Mar11, MR15b, Oka14, Oli10b, PW11, Pim06, PR18, QM17, Rok07, RS21, SŽ17, SB07, Tho16, VY12b, Van08, Vov08, WH19, Zho10]. time-changed [HRKU11, KP22]. time-dependent [HK16]. time-homogeneous [EM14]. time-invariant [GH21]. Times [DS10, IM10, KS05a, Pes08, Abe15, ADOS11, Bas10, Bas11, BBMT09, Che21c, DHI11, DFK18, DW16, DP13, Dok15, Fan16, HM14, HM22, HPS14, HKST18, KSW12, Kov10, KT11, PW18, Tre13b, Vid14]. Toeplitz [BB10, BS07b, Kar09, Mec07, SV11b]. Toom [HS22]. Topological [Car05]. topology [Led16, Ruf15b, Stu16]. tori [ČS16]. torus [Win08]. Total [CM22, Aid10, Hon21, Zho10]. Touchard [Pin17]. Trace [Kuz00]. traces [ES09]. trading [Vov08]. Traffic [AHM05, Lim99]. trajectory [Che21b, JL21]. Transaction [SV11a]. transform [LG20, RS07]. Transformation [FJ00]. transformations [ALW14, BCP03, Jos07]. transforms [Dji18, Gra15]. Transience [HS12, SHH14, DHS14, Ker17, KTT17, OdS16, Pet15, RS11b, San13, Zer06]. Transient [GP01, MR15a]. Transition [BBB97, PŻ16, Tel00, BEK21, BJT17, BC14, CZ21, Che21a, FH22, HJM22, KLS21, Mar09, Pen18, Rát15, SS21]. transitions [ACT19, HJCW20, Kov09]. transitive [For20, Gal20]. Translation [LW15, Bec22]. translations [Gri11]. Transport [Hue16, BD15, BJ22, Fat18, Jou12]. transport-chi-square [Jou12]. transport-entropy [Fat18]. Transportation [DZ96, MWW11, Mik02, FGM10, Goz06, PS19]. Transportation-information [MWW11]. trap [Mui15]. trapping [Fuk09]. traps [Mui15]. travel [Cer14]. Tree [Kor05, Tim04, AG15, Ald21, Bia13, BT20, BT21, DKW14, IM07, Jan21, JJ16, RSS18, Wan15, Wan21, vdBKN12]. Trees [HP03, JC04, KDN05, Mar99, PR11, ST99, Tas10, War99, Bac11, BFJ^{+}19, Bjo09, CCGS19, CP22, Cra13, CP14, DJ14, Duq09, DL09b, Eas22, EW17, FY15, FLP20, GP14, HL13, HD20, HMSH15, IK18, Jan20, KF09, Maś18, Mic19, MS11, PSY13, PS20, RZ13, RW21, Tim19, HD23]. trials [NP20]. triangle [Hou09]. Triangles [Dub03]. triangular [CK12, DC13, Yao14]. triangulations [GGNS17]. trifractional [Han21]. trigonometric [Thi20]. trimmed [LL07]. trivial [Pel23]. triviality [KLvR19, Lyo18, Pen18]. truncated [Cha10]. truncations [MS19a]. Tsirelson [DR12]. Two [DJR16, Kah03, KP04, ME21, Mui15, Pat07, Abe15, ADS20, Baz21, Ber10, BDZ11, BJT17, DHR18, DG17, Fér20, GHJL17, HM09, HM23, LR15, RR15, RW09, Sap11, Sch09, Tko13, WZ19, Yor15]. two-charge [DHR18]. two-dimensional [Abe15, ADS20, DG17, HM09, HM23]. two-factor [RR15]. two-parameter [RW09]. two-point [LR15]. Two-sided [ME21, Pat07, Fér20]. Two-site [Mui15]. two-time-scale [Ber10]. Type [Bal09, BDT11, BDM01, BG21, BGT21, BN20, BHS10, BKP22, CP17a, Com08, CM18b, Def12, FJR20, Gor15, Gru23, KS10, LP12, LL07, MNX20, Mon07, Ose14, OW22, Oto09, Par17, Pri15, RS21, Sal22, Sok13, Tug16, Uem07, Wan22, Woj12, dBJP13]. Types [Gne10, Fre12, HJ18]. typically [PSY13].

Uhlenbeck [BH12, CGXM96, CGXM97, CM18b, CKS99, GJ09a, GP11, Jeg09, MY13, Pat07]. ultimate [AI12]. UMD [MN08]. Unbounded [Man05, ATT18, Böt11]. uncertainty [Dol14, Rok15, Sir14]. uncovered [ST20]. underlying [WP14]. unicellular [ACCR13]. unicycles [SW16]. Uniform [ABIT22, BLY15, KM09, LPP15, Res01, RŻ98, SXY18, BC14, DL18, GM12, KM20, Kaz18, MR22a, Neu11, Squ21, Stu16, Tsi13]. Uniformly [BLL08, Fan15, FLP20]. unimodal [JS18]. unimodality [Sim11]. unimodular [FHJ18, Tim19]. unimodularity [BC12]. union [Doh13]. Unique [Swa01, AG15]. Uniqueness [Bia13, Due06, ER09, Kuz00, Li14, LT11, Luo14, Bah02, Ban15, BB06, CEK11, Hoe09, Sch20, Wan17]. unit [JK08]. Unitary [BDN10, Def11, BD13, Dub23, LS18a, MS19a, NP13, Tko13]. Universal [BZ17, Fan15, GH20]. Universality [BM05, DW16, Zha12, AGN21]. unstable [GN06]. Up-to-constants [LR15]. Upper [EM21, FV21, Jan19a, RŻ98, Sap10]. upsilon [Gra15]. Urbanik [GP11]. urn [CL14, Lau20]. urns [CH21, Jan19b, Kub11]. used [Hol21]. user [BHS18]. user-friendly [BHS18]. uses [Law14]. Using [GKH03, FI21, Tre13b]. Utility [RR14].

vacancy [Pen18]. vacant [ATT18, Rát15, Szn19]. Value [KDN05, KLS05, CF20, GIOQ20, HT20, LC22, Wag16, Yas15]. Valued [Law08, AAS21, BC12, Bou16, Cha10, Led16, MHC13]. vanishing [AP16, RS07]. Vapnik [Pan02]. Variable [BY01, Cla14, Dem11, Squ21]. variable-rate [Cla14]. variables [AAHT22, AV12, BB07, BY13, BFY07, BT11, GRR14, HZ07, Kie97, NY09, RS07, Rio15, Van07, BB22]. Variably [BBKM00]. Variance [BZ06, HR07, RBS15, Wee06, AEK14, Rok15, SB07]. Variation [Man05, CC18, DKW14, KT11, NS13, Sio14]. Variational [Der16, KZK19, But17]. variations [BN08, GL14, Tud09]. various [YY13]. varying [BKS16, Mui15]. Vasicek [ARS22]. Vasicek-CIR [ARS22]. VC [LMK03]. Vector [Bou16, Vys21]. Vector-valued [Bou16]. Vectors [GG04, GS09, MR01, Ada15, Cha10, CM12b, DGG^{+}13, HKZ12b]. Velocity [CGXM96, CGXM97]. Version [MY99, GS21, Maj06, MZ14b, Nee14]. versus [DHS14, Gal20]. Vertex [Sab21, DKNS16, Kua21, Lin19, Mon07, Sch12, Sin14]. vertex-reinforced [Sch12, Sin14]. Vey [Led17]. via [Alt20, AED13, ALW14, BC12, D’O10, Dal13, DP18a, Doh13, Gao08, Ham05, HS22, Her17, MY20, NP13, Pet08, SB07, TM15, WY08]. viewed [dBM15]. viewpoint [NR18]. violating [Mar10]. Viot [FX02, LSY99, Zho14]. visibility [Mar22]. visited [CM12a, Uch15]. Vlasov [Bas20, DT18, Gru23, HM22, Lac18]. volatility [NR18, Vov08]. Volterra [App02, Jos07, KL14, Val17]. volume [AHM06, AAGL19, Due06, Kis14, Öz20]. Voter [AS11, AABP22, Che21c, Ram14, RV17, SS08]. vs [KLvR19].

Wald [HPS14]. Walk [BFV10, BW03, BR07, DBGP03, JL21, JK04, KS97, MR11, PV05, RA05, Abe15, Aid10, Ale13, BCNP19, BHP18, BFGG^{+}16, BP21, BK11, Blo23, BT20, BT21, CJ13, ČS16, DK18, DV23, DHS14, DG17, DL09a, DGH20, ER09, FS19, FHJ18, GT18, GH18a, GM13, Gho22, Gla15, GPL08, GPHS13, Hol09, HLSX21, Hut18, Jul19, KV22, KM17, KSW12, Kri21, LS13a, Law14, Le 08, LTV18, MV14, Pet15, PZ18, ST20, Uch15, WH19, Win08, Lau22]. Walks [ABV03, BD02, FZ10, GP01, GGA10, HMO01, HN11, IM10, KP04, Law98, Mar99, OY01, Roi05, Zer02, APRB11, ACT19, ADS19, ADS20, BEK21, Baz21, BKR06, BFT13, BHM21, Car18, DFK18, DW12, DW16, DK12, DJ12, Dur14, EP17, EM21, GLY14, GMW18, GPPdS14, GGPZ14, Hol15, HK16, HS12, HP15, IR10, KS10, LT20, LY22, OdS16, Pil17, RS11a, Ras10, RS16, SHH14, Sch12, SK15, Sin14, Ste13, Tre18, Uch18, Vid14, WYY13, Xie20, Zer06, Zer07]. wall [Def11]. Walsh [VY12a]. Wasserstein [BG21, CD20, EMR15]. Wasserstein-1 [CD20]. Watanabe [Cri21, Hoe09, Tap13]. Watson [BT21, BT20, Duq09, GP14, HL13, HMSH15, Jan21, KF09, Mic19, PS20, PS17, RSS18, Tas10]. wave [LX19, LZ23, NZ20]. way [Bor10, CK08]. Weak [BVBP20, BRT10, Bir04, BC98, Kur14, MV14, OW22, Pec04, SP00, WYY13, vZ02, FJ21, IM07, Kri14, MO22, Ose14, Sar21, Sin14, Stu16, Xie20]. Weak-type [OW22, Ose14]. Weakly [CH21, LT11, Dub23, Har12, Luo14, Oli10b]. wedge [DM09, QR11]. Weierstrass [Com08]. weight [Sch12]. Weighted [LP19, LL15, MZ05a, MZ05b, Osȩ16, Ald16, BB07, BB22, Che17, Gan21, GRR14, Gla15, Rio15]. weights [ABIT22, Def12, Emr16, RW21, Sch09]. Weiss [CP17a]. Welch [Tan17]. Welsh [Hut18]. Wetting [CG05]. Weyl [KS10]. Where [BS18b, KP04, vdBKN12]. Which [KV11, And06, CD17a, Eri16, Yor15]. White [Gri11]. Whose [Swa01, Luo14, Völ20]. Widom [Häg02]. width [Sch09]. width-two [Sch09]. Wiener [NP12c, ABP00, AN19, Bar97, Bar98, CKS99, Fun07, Har04, JC04, MY12, NP12b, Oto09, Pri09, Tan06, Tre13a, YLW15]. Wiener/Wigner [NP12c, NP12b]. Wigner [NP12c, AGN21, DV11, EM16, ES16, HCS08, Kös08, NP12b, Sos04, vSW19]. winding [Oka14]. window [AF14]. windows [Tho16]. wired [Eas22]. wise [BKR06]. Wishart [GL09]. without [BPR13, JTT18, Lac15, Yas15]. Woodroofe [KV13, Tót13]. Worah [Péc19]. words [ES09, HM16, dHP14]. works [Haj15]. Wright [Fou13, Fou14, Ada15, HT05, Pat07, RV13]. Wronskian [Kli12a]. Wulff [DC13].

Xi [Ost14].

Yaglom [Oça20, RSS18]. Yamada [Cri21, Hoe09, Tap13]. Yor [Bas15, KV11]. Yule [de 06].

Zero [Gol22, OdS16, RA05, Tre13b, Bor10, CK08, GJ09b, Kin08, KN19, LX22, MR08, MV14, MLV15, NR18, Oka14, Pan21, Zer07]. Zero-one [OdS16, Zer07]. zero-range [GJ09b]. zero-set [Kin08]. zeros [Aym19, Eva06, Thi20]. zeta [Oui18]. Zhang [DDT07]. Zhao [KV13].

References

Andrieu:2007:EAM

[AA07] Christophe Andrieu and Yves Atchade. On the efficiency of adaptive MCMC algorithms. Electronic Communications in Probability, 12:33:336–33:349, 2007. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1320.

Amir:2022:DNS

[AABP22] Gideon Amir, Omer Angel, Rangel Baldasso, and Ron Peretz. Dynamical noise sensitivity for the voter model. Electronic Communications in Probability, 27(??):1–7, 2022. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-27/issue-
none/Dynamical-noise-sensitivity-for-the-voter-model/
10.1214/22-ECP483.full.

Armentano:2019:CFM

[AAGL19] D. Armentano, J-M. Azaïs, D. Ginsbourger, and J. R. León. Conditions for the finiteness of the moments of the volume of level sets. Electronic Communications in Probability, 24(??): 17:1–17:8, ???? 2019. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1553220033.

Aamand:2022:SMR

[AAHT22] Anders Aamand, Noga Alon, Jakob Bæk Tejs Houen, and Mikkel Thorup. On sums of monotone random integer variables. Electronic Communications in Probability, 27(??):1– 8, 2022. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-27/issue-none/On-sums-of-monotone-
random-integer-variables/10.1214/22-ECP500.full.

Atar:2001:BDP

[AAK01] Rami Atar, Siva Athreya, and Min Kang. Ballistic deposition on a planar strip. Electronic Communications in Probability, 6: 3:31–3:38, 2001. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1032.

Alekseev:2021:SDE

[AAS21] Anton Alekseev, Elizaveta Arzhakova, and Daria Smirnova. Stochastic differential equations for Lie group valued moment maps. Electronic Communications in Probability, 26(??):1–9, 2021. CODEN ???? ISSN 1083-589X.

Atar:2002:NLN

[AB02] Rami Atar and Krzysztof Burdzy. On nodal lines of Neumann eigenfunctions. Electronic Communications in Probability, 7: 14:129–14:139, 2002. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1055.

Addario-Berry:2014:GRC

[AB14] Louigi Addario-Berry. Growing random 3-connected maps or, comment s’enfuir de l’hexagone. Electronic Communications in Probability, 19:54:1–54:12, 2014. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/3314.

Abe:2015:MML

[Abe15] Yoshihiro Abe. Maximum and minimum of local times for twodimensional random walk. Electronic Communications in Probability, 20(??):22:1–22:14, ???? 2015. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/3877.

Alexy:2022:CFU

[ABIT22] Marcell Alexy, Márton Borbényi, András Imolay, and Ádám Timár. Connectedness of the Free Uniform Spanning Forest as a function of edge weights. Electronic Communications in Probability, 27(??):1–12, 2022. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-27/issue-
none/Connectedness-of-the-Free-Uniform-Spanning-Forest-
as-a-function/10.1214/22-ECP453.full.

Angel:2000:LWS

[ABP00] Omer Angel, Itai Benjamini, and Yuval Peres. A large Wiener sausage from crumbs. Electronic Communications in Probability, 5:7:67–7:71, 2000. CODEN ???? ISSN 1083-589X. URL http:

//ecp.ejpecp.org/article/view/1019.

Angel:2003:RWA

[ABV03] Omer Angel, Itai Benjamini, and Bálint Virág. Random walks that avoid their past convex hull. Electronic Communications in Probability, 8:2:6–2:16, 2003. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1065.

Albin:2010:NPO

[AC10] J. M. P. Albin and Hyemi Choi. A new proof of an old result by Pickands. Electronic Communications in Probability, 15:32:339– 32:345, 2010. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1566.

Angel:2013:LLU

[ACCR13] Omer Angel, Guillaume Chapuy, Nicolas Curien, and Gourab Ray. The local limit of unicellular maps in high genus. Electronic Communications in Probability, 18:86:1–86:8, 2013. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3037.

Akahori:2019:PTE

[ACT19] Jiro Akahori, Andrea Collevecchio, and Masato Takei. Phase transitions for edge-reinforced random walks on the half-line. Electronic Communications in Probability, 24(??):39:1–39:12, ???? 2019. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1561169055.

Adamczak:2015:NHW

[Ada15] Radoslaw Adamczak. A note on the Hanson–Wright inequality for random vectors with dependencies. Electronic Communications in Probability, 20(??):72:1–72:13, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3829.

Alt:2022:CDR

[ADK22] Johannes Alt, Raphaël Ducatez, and Antti Knowles. The completely delocalized region of the Erdős–Rényi graph. Electronic Communications in Probability, 27(??):1–9, 2022. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
journals/electronic-communications-in-probability/volume-
27/issue-none/The-completely-delocalized-region-of-the-
Erd%c5%91s-R%c3%a9nyi-graph/10.1214/22-ECP450.full.

Aurzada:2011:MRT

[ADOS11] Frank Aurzada, Hanna Döring, Marcel Ortgiese, and Michael Scheutzow. Moments of recurrence times for Markov chains. Electronic Communications in Probability, 16:28:296–28:303, 2011. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1632.

Andres:2019:HKE

[ADS19] Sebastian Andres, Jean-Dominique Deuschel, and Martin Slowik. Heat kernel estimates and intrinsic metric for random walks with general speed measure under degenerate conductances. Electronic Communications in Probability, 24(??):5:1–5:17, ???? 2019. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1549357292.

Andres:2020:GKA

[ADS20] Sebastian Andres, Jean-Dominique Deuschel, and Martin Slowik. Green kernel asymptotics for two-dimensional random walks under random conductances. Electronic Communications in Probability, 25(??):1–14, 2020. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-25/issue-
none/Green-kernel-asymptotics-for-two-dimensional-random-
walks-under-random/10.1214/20-ECP337.full.

Aman:2013:RTS

[AED13] Auguste Aman, Abouo Elouaflin, and Mamadou Diop. Representation theorems for SPDEs via backward doubly. Electronic Communications in Probability, 18:64:1–64:15, 2013. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2223.

Ajanki:2014:LSL

[AEK14] Oskari Ajanki, Lászlo Erdős, and Torben Krüger. Local semicircle law with imprimitive variance matrix. Electronic Communications in Probability, 19:32:1–32:9, 2014. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/3121.

Alabert:2006:LSD

[AF06]  Aureli Alabert and Marco Ferrante.  Linear stochastic

differential-algebraic equations with constant coefficients. Electronic Communications in Probability, 11:32:316–32:335, 2006. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1236.

Alon:2014:NGS

[AF14] Noga Alon and Ohad Noy Feldheim. A note on general sliding window processes. Electronic Communications in Probability, 19 (??):66:1–66:7, ???? 2014. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/3341.

Aksamit:2019:MSR

[AF19] Anna Aksamit and Claudio Fontana. Martingale spaces and representations under absolutely continuous changes of probability. Electronic Communications in Probability, 24(??):62:1– 62:13, ???? 2019. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1570500302.

Aymone:2020:LIL

[AFM20] Marco Aymone, Susana Frómeta, and Ricardo Misturini. Law of the iterated logarithm for a random Dirichlet series. Electronic Communications in Probability, 25(??):1–14, 2020. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-25/issue-none/Law-of-the-iterated-
logarithm-for-a-random-Dirichlet-series/10.1214/20-ECP340.
full.

Albenque:2015:BCR

[AG15] Marie Albenque and Christina Goldschmidt. The Brownian continuum random tree as the unique solution to a fixed point equation. Electronic Communications in Probability, 20(??):61:1– 61:14, ???? 2015. CODEN ???? ISSN 1083-589X. URL http:/

/ecp.ejpecp.org/article/view/4250.

Akemann:2021:CPP

[AGN21] Gernot Akemann, Friedrich Götze, and Thorsten Neuschel. Characteristic polynomials of products of non-Hermitian Wigner matrices: finite-N results and Lyapunov universality. Electronic Communications in Probability, 26(??):1–13, 2021. CODEN ???? ISSN 1083-589X.

Atar:2014:RSC

[AGS14] Rami Atar, Anindya Goswami, and Adam Shwartz. On the risksensitive cost for a Markovian multiclass queue with priority. Electronic Communications in Probability, 19:11:1–11:13, 2014. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2905.

Aguilar:2018:LSE

[Agu18] Erick Trevino Aguilar. The lower Snell envelope of smooth functions: an optional decomposition. Electronic Communications in Probability, 23(??):12:1–12:10, ???? 2018. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1519722242.

Angel:2018:PSL

[AH18] Omer Angel and Alexander E. Holroyd. Perfect shuffling by lazy swaps. Electronic Communications in Probability, 23(??): 47:1–47:11, ???? 2018. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1532657019.

Angel:2005:JPB

[AHM05] Omer Angel, Alexander Holroyd, and James Martin. The jammed phase of the Biham–Middleton–Levine traffic model. Electronic Communications in Probability, 10:17:167–17:178, 2005. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1148.

Andersson:2006:VFN

[AHM06] Jenny Andersson, Olle Häggström, and Marianne Månsson. The volume fraction of a non-overlapping germ–grain model. Electronic Communications in Probability, 11:8:78–8:88, 2006. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1197.

Angel:2013:AC

[AHM^{+}13] Omer Angel, Alexander Holroyd, James Martin, Peter Winkler, and David Wilson. Avoidance coupling. Electronic Communications in Probability, 18:58:1–58:13, 2013. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/2275.

Ano:2012:PUM

[AI12] Katsunori Ano and Roman Ivanov. On predicting the ultimate maximum for exponential Lévy processes. Electronic Communications in Probability, 17:46:1–46:9, 2012. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/1805.

Aidekon:2010:TAT

[Aid10] Elie Aidekon. Tail asymptotics for the total progeny of the critical killed branching random walk. Electronic Communications in Probability, 15:47:522–47:533, 2010. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1583.

Arizmendi:2014:CFM

[AJ14] Octavio Arizmendi and Arturo Jaramillo. Convergence of the fourth moment and infinite divisibility: quantitative estimates. Electronic Communications in Probability, 19:25:1–25:12, 2014. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3354.

Appleby:2004:ONO

[AK04] John Appleby and Conall Kelly. Oscillation and non-oscillation in solutions of nonlinear stochastic delay differential equations. Electronic Communications in Probability, 9:12:106– 12:118, 2004. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1115.

Alberts:2008:IPC

[AK08] Tom Alberts and Michael Kozdron. Intersection probabilities for a chordal SLE path and a semicircle. Electronic Communications in Probability, 13:43:448–43:460, 2008. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1399.

Aldous:1998:BEC

[Ald98] David J. Aldous. Brownian excursion conditioned on its local time. Electronic Communications in Probability, 3:10:79–10:90, 1998. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/996.

Aldous:2016:IGC

[Ald16] David Aldous. The incipient giant component in bond percolation on general finite weighted graphs. Electronic Communications in Probability, 21(??):68:1–68:9, ???? 2016. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1474462208.

Aldous:2021:RLI

[Ald21] David Aldous. Route lengths in invariant spatial tree networks. Electronic Communications in Probability, 26(??):1–12, 2021. CODEN ???? ISSN 1083-589X.

Alexander:2013:CRW

[Ale13] Kenneth Alexander. Controlled random walk with a target site. Electronic Communications in Probability, 18:43:1–43:6, 2013. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/2763.

Alili:2001:CDC

[Ali01] Larbi Alili. Canonical decompositions of certain generalized Brownian bridges. Electronic Communications in Probability, 7:3:27–3:35, 2001. CODEN ???? ISSN 1083-589X. URL http:

//ecp.ejpecp.org/article/view/1044.

Alt:2017:SDS

[Alt17] Johannes Alt. Singularities of the density of states of random Gram matrices. Electronic Communications in Probability, 22 (??):63:1–63:13, ???? 2017. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1511233247.

Altman:2020:IPF

[Alt20] Henri Elad Altman. Integration by parts formulae for the laws of Bessel bridges via hypergeometric functions. Electronic Communications in Probability, 25(??):46:1–46:11, ???? 2020. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
euclid.ecp/1593569166.

Arratia:2014:SFP

[ALW14] Richard Arratia, Thomas Liggett, and Malcolm Williamson. Scale-free and power law distributions via fixed points and convergence of (thinning and conditioning) transformations. Electronic Communications in Probability, 19:38:1–38:10, 2014. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2923.

Aly:2013:PSC

[Aly13] Sidi Mohamed Ould Aly. Parameter sensitivity of CIR process. Electronic Communications in Probability, 18:34:1–34:6, 2013. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/2035.

Alili:2018:FSS

[AM18] Larbi Alili and Hiroyuki Matsumoto. Further studies on squareroot boundaries for Bessel processes. Electronic Communications in Probability, 23(??):39:1–39:9, ???? 2018. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1529460064.

Azmoodeh:2019:ASL

[AN19] Ehsan Azmoodeh and Ivan Nourdin. Almost sure limit theorems on Wiener chaos: the non-central case. Electronic Communications in Probability, 24(??):9:1–9:12, ???? 2019. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1550199821.

Andrew:2006:PFP

[And06] Peter Andrew. A proof from ‘first principles’ of Kesten’s result for the probabilities with which a subordinator hits points. Electronic Communications in Probability, 11:6:58–6:63, 2006. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1193.

Ang:2019:CDC

[Ang19] Morris Ang. Comparison of discrete and continuum Liouville first passage percolation. Electronic Communications in Probability, 24(??):64:1–64:12, ???? 2019. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1572509098.

Aoun:2021:SAC

[Aou21] Yacine Aoun. Sharp asymptotics of correlation functions in the subcritical long-range random-cluster and Potts models. Electronic Communications in Probability, 26(??):1–9, 2021. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-26/issue-none/Sharp-asymptotics-
of-correlation-functions-in-the-subcritical-long-range/
10.1214/21-ECP390.full.

Aryasova:2014:DSF

[AP14] Olga Aryasova and Andrey Pilipenko. On differentiability of stochastic flow for a multidimensional SDE with discontinuous drift. Electronic Communications in Probability, 19:44:1–44:17, 2014. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/2886.

Acciaio:2016:CMC

[AP16] Beatrice Acciaio and Irina Penner. Characterization of maxcontinuous local martingales vanishing at infinity. Electronic Communications in Probability, 21(??):71:1–71:10, ???? 2016. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1476368439.

Appleby:2002:ASS

[App02] John Appleby. Almost sure stability of linear Itô–Volterra equations with damped stochastic perturbations. Electronic Communications in Probability, 7:22:223–22:234, 2002. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1063.

Abramson:2011:CMR

[APRB11] Josh Abramson, Jim Pitman, Nathan Ross, and Geronimo Uribe Bravo. Convex minorants of random walks and Lévy processes. Electronic Communications in Probability, 16:38:423– 38:434, 2011. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1648.

Aru:2019:CLM

[APS19] Juhan Aru, Ellen Powell, and Avelio Sepúlveda. Critical Liouville measure as a limit of subcritical measures. Electronic Communications in Probability, 24(??):18:1–18:16, ???? 2019. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
euclid.ecp/1553306557.

Athreya:2018:RDS

[AR18] Siva Athreya and Adrian Röllin. Respondent-driven sampling and sparse graph convergence. Electronic Communications in Probability, 23(??):3:1–3:12, ???? 2018. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1517626933.

Arcones:1998:LLN

[Arc98] Miguel A. Arcones. The law of large numbers for U-statistics under absolute regularity. Electronic Communications in Probability, 3:2:13–2:19, 1998. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/988.

Arguin:2007:DCP

[Arg07] Louis-Pierre Arguin. A dynamical characterization of Poisson– Dirichlet distributions. Electronic Communications in Probability, 12:28:283–28:290, 2007. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1300.

Antonelli:2022:CPS

[ARS22] Fabio Antonelli, Alessandro Ramponi, and Sergio Scarlatti. On a convergent power series method to price defaultable bonds in a Vasicek-CIR model. Electronic Communications in Probability, 27(??):1–12, 2022. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-27/issue-
none/On-a-convergent-power-series-method-to-price-defaultable-
bonds/10.1214/22-ECP458.full.

Abreu:2008:FGG

[AS08] Victor Perez Abreu and Noriyoshi Sakuma. Free generalized gamma convolutions. Electronic Communications in Probability, 13:50:526–50:539, 2008. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1413.

Athreya:2011:ODV

[AS11] Siva Athreya and Rongfeng Sun. One-dimensional voter model interface revisited. Electronic Communications in Probability, 16:70:792–70:800, 2011. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1688.

Atar:2016:NNE

[AS16] Rami Atar and Subhamay Saha. A note on non-existence of diffusion limits for serve-the-longest-queue when the buffers are equal in size. Electronic Communications in Probability, 21(??): 2:1–2:10, ???? 2016. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1454514622.

Aguirre:2022:CLT

[AS22] Ander Aguirre and Alexander Soshnikov. Central limit theorem for CβE pair dependent statistics in mesoscopic regime. Electronic Communications in Probability, 27(??):1– 10, 2022. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-27/issue-none/Central-limit-theorem-
for-C%ce%b2E-pair-dependent-statistics-in-mesoscopic/
10.1214/22-ECP481.full.

Assiotis:2018:MBI

[Ass18] Theodoros Assiotis. A matrix Bougerol identity and the Hua– Pickrell measures. Electronic Communications in Probability, 23 (??):7:1–7:11, ???? 2018. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1519182082.

Ahlberg:2014:BSD

[AST14] Daniel Ahlberg, Vladas Sidoravicius, and Johan Tikesson. Bernoulli and self-destructive percolation on non-amenable graphs. Electronic Communications in Probability, 19:39:1–39:6, 2014. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/2611.

Attanasio:2010:SFD

[Att10] Stefano Attanasio. Stochastic flows of diffeomorphisms for onedimensional SDE with discontinuous drift. Electronic Communications in Probability, 15:20:213–20:226, 2010. CODEN

????  ISSN 1083-589X. URL http://ecp.ejpecp.org/
article/view/1545.

Ahlberg:2018:EUV

[ATT18] Daniel Ahlberg, Vincent Tassion, and Augusto Teixeira. Existence of an unbounded vacant set for subcritical continuum percolation. Electronic Communications in Probability, 23(??): 63:1–63:8, ???? 2018. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1536977436.

Aurzada:2011:OSE

[Aur11] Frank Aurzada. On the one-sided exit problem for fractional Brownian motion. Electronic Communications in Probability, 16: 36:392–36:404, 2011. CODEN ???? ISSN 1083-589X. URL http:

//ecp.ejpecp.org/article/view/1640.

Arizmendi:2012:PFR

[AV12] Octavio Arizmendi and Carlos Vargas. Products of free random variables and k-divisible non-crossing partitions. Electronic Communications in Probability, 17:11:1–11:13, 2012. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1773.

Avena:2012:SEM

[Ave12] Luca Avena. Symmetric exclusion as a model of non-elliptic dynamical random conductances. Electronic Communications in Probability, 17:44:1–44:8, 2012. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/2081.

Aymone:2019:RZR

[Aym19] Marco Aymone. Real zeros of random Dirichlet series. Electronic Communications in Probability, 24(??):54:1–54:8, ???? 2019. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1568253716.

Aymone:2021:RMF

[Aym21] Marco Aymone. Random multiplicative functions: the Selberg– Delange class. Electronic Communications in Probability, 26 (??):1–8, 2021. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-26/issue-none/Random-multiplicative-
functions-the-Selberg-Delange-class/10.1214/21-ECP396.
full.

Bandyopadhyay:2001:HCF

[BA01] Antar Bandyopadhyay and David Aldous. How to combine fast heuristic Markov chain Monte Carlo with slow exact sampling. Electronic Communications in Probability, 6:8:79–8:89, 2001. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1037.

Ben-Ari:2014:CDB

[BA14] Iddo Ben-Ari. Coupling for drifted Brownian motion on an interval with redistribution from the boundary. Electronic Communications in Probability, 19:16:1–16:11, 2014. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/2756.

Backhausz:2011:LDD

[Bac11] Agnes Backhausz. Limit distribution of degrees in random family trees. Electronic Communications in Probability, 16:4:29–4:37, 2011. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1598.

Bahlali:2002:EUS

[Bah02] Khaled Bahlali. Existence and uniqueness of solutions for BSDEs with locally Lipschitz coefficient. Electronic Communications in Probability, 7:17:169–17:179, 2002. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1058.

Ball:2005:PTM

[Bal05] Karen Ball. Poisson thinning by monotone factors. Electronic Communications in Probability, 10:7:60–7:69, 2005. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1134.

Balan:2009:NFK

[Bal09] Raluca Balan. A note on a Feynman–Kac-type formula. Electronic Communications in Probability, 14:25:252–25:260, 2009. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1468.

Ben-Ari:2011:SSM

[BAMR11] Iddo Ben-Ari, Anastasios Matzavinos, and Alexander Roitershtein. On a species survival model. Electronic Communications in Probability, 16:22:226–22:233, 2011. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1625.

Bandini:2015:EUB

[Ban15] Elena Bandini. Existence and uniqueness for backward stochastic differential equations driven by a random measure, possibly non quasi-left continuous. Electronic Communications in Probability, 20(??):71:1–71:13, ???? 2015. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/4348.

Baryshnikov:1997:WSG

[Bar97] Yuliy Baryshnikov. Wiener soccer and its generalization. Electronic Communications in Probability, 3:1:1–1:11, 1997. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/987.

Baryshnikov:1998:WSG

[Bar98] Yuliy Baryshnikov. Wiener soccer and its generalization. Electronic Communications in Probability, 3:1–11, 1998. CODEN ???? ISSN 1083-589X.

Barbato:2005:FIB

[Bar05] David Barbato. FKG inequality for Brownian motion and stochastic differential equations. Electronic Communications in Probability, 10:2:7–2:16, 2005. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1127.

Barraquand:2014:SPS

[Bar14] Guillaume Barraquand. A short proof of a symmetry identity for the q-Hahn distribution. Electronic Communications in Probability, 19:50:1–50:3, 2014. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/3674.

Bass:2010:MHT

[Bas10] Richard Bass. The measurability of hitting times. Electronic Communications in Probability, 15:10:99–10:105, 2010. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/ article/view/1535. See correction [Bas11].

Bass:2011:CMH

[Bas11] Richard Bass. Correction to “The measurability of hitting times”. Electronic Communications in Probability, 16:18:189– 18:191, 2011. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1627. See [Bas10].

Basrak:2015:LRP

[Bas15] Bojan Basrak. Limits of renewal processes and Pitman–Yor distribution. Electronic Communications in Probability, 20(??): 51:1–51:13, ???? 2015. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/4080.

Bashiri:2020:LTB

[Bas20] K. Bashiri. On the long-time behaviour of McKean–Vlasov paths. Electronic Communications in Probability, 25(??):1– 14, 2020. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-25/issue-none/On-the-long-time-
behaviour-of-McKean-Vlasov-paths/10.1214/20-ECP330.full.

Baudoin:2002:FEG

[Bau02] Fabrice Baudoin. Further exponential generalization of Pitman’s 2M − X theorem. Electronic Communications in Probability, 7: 4:37–4:46, 2002. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1045.

Bazaes:2021:LBC

[Baz21] Rodrigo Bazaes. Localization at the boundary for conditioned random walks in random environment in dimensions two and higher. Electronic Communications in Probability, 26(??):1–13, 2021. CODEN ???? ISSN 1083-589X.

Borovkov:2001:KIF

[BB01] Konstantin Borovkov and Zaeem Burq. Kendall’s identity for the first crossing time revisited. Electronic Communications in Probability, 6:9:91–9:94, 2001. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1038.

Bass:2006:PUR

[BB06] Richard Bass and Krzysztof Burdzy. Pathwise uniqueness for reflecting Brownian motion in certain planar Lipschitz domains. Electronic Communications in Probability, 11:18:178– 18:181, 2006. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1213.

Bercu:2007:ARE

[BB07] Bernard Bercu and Włodek Bryc. Asymptotic results for empirical measures of weighted sums of independent random variables. Electronic Communications in Probability, 12:19:184– 19:199, 2007. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1273. See erratum [BB22].

Basak:2010:BRT

[BB10] Aniran Basak and Arup Bose. Balanced random and Toeplitz matrices. Electronic Communications in Probability, 15:14:134– 14:148, 2010. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1537.

Bercu:2022:EAR

[BB22] Bernard Bercu and Wlodzimierz Bryc. Erratum: Asymptotic results for empirical measures of weighted sums of independent random variables. Electronic Communications in Probability, 27 (??):??, 2022. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-27/issue-none/Erratum--Asymptotic-
results-for-empirical-measures-of-weighted-sums/10.1214/
22-ECP464.full. See [BB07].

Barlow:1997:PBT

[BBB97] Martin T. Barlow, Richard F. Bass, and Krzysztof Burdzy. Positivity of Brownian transition densities. Electronic Communications in Probability, 2:4:43–4:51, 1997. CODEN ???? ISSN

1083-589X. URL http://ecp.ejpecp.org/article/view/983.

Bakry:2008:SPP

[BBCG08] Dominique Bakry, Franck Barthe, Patrick Cattiaux, and Arnaud Guillin. A simple proof of the Poincaré inequality for a large class of probability measures. Electronic Communications in Probability, 13:7:60–7:66, 2008. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1352.

Becherer:2018:ADR

[BBF18] Dirk Becherer, Todor Bilarev, and Peter Frentrup. Approximating diffusion reflections at elastic boundaries. Electronic Communications in Probability, 23(??):40:1–40:12, ???? 2018. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
euclid.ecp/1529546623.

Barlow:2000:VSB

[BBKM00] Martin Barlow, Krzysztof Burdzy, Haya Kaspi, and Avi Mandelbaum. Variably skewed Brownian motion. Electronic Communications in Probability, 5:6:57–6:66, 2000. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/1018.

Betsakos:2022:DSB

[BBM22] Dimitrios Betsakos, Maher Boudabra, and Greg Markowsky. On the duration of stays of Brownian motion in domains in Euclidean space. Electronic Communications in Probability, 27(??): 1–12, 2022. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-27/issue-none/On-the-duration-
of-stays-of-Brownian-motion-in-domains/10.1214/22-ECP498.
full.

Bernardin:2009:MNP

[BBMT09] Frédéric Bernardin, Mireille Bossy, Miguel Martinez, and Denis Talay. On mean numbers of passage times in small balls of discretized Itô processes. Electronic Communications in Probability, 14:30:302–30:316, 2009. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1479.

Burdzy:1998:WCR

[BC98] Krzysztof Burdzy and Zhen-Qing Chen. Weak convergence of reflecting Brownian motions. Electronic Communications in Probability, 3:4:29–4:33, 1998. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/990.

Benjamini:2012:RVI

[BC12] Itai Benjamini and Nicolas Curien. Recurrence of the Z^{d}-valued infinite snake via unimodularity. Electronic Communications in Probability, 17:1:1–1:10, 2012. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1700.

Budhiraja:2014:UPT

[BC14] Amarjit Budhiraja and Zhen-Qing Chen. On uniform positivity of transition densities of small noise constrained diffusions. Electronic Communications in Probability, 19:1:1–1:9, 2014. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2967.

Benaim:2015:SAA

[BC15] Michel Benaïm and Bertrand Cloez. A stochastic approximation approach to quasi-stationary distributions on finite spaces. Electronic Communications in Probability, 20(??):37:1–37:13, ???? 2015. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/3956.

Benjamini:2012:LIP

[BCG12a] Itai Benjamini, Nicolas Curien, and Agelos Georgakopoulos. The Liouville and the intersection properties are equivalent for planar graphs. Electronic Communications in Probability, 17:42:1– 42:5, 2012. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1913.

Bobkov:2012:BCF

[BCG12b] Sergey Bobkov, Gennadiy Chistyakov, and Friedrich Götze. Bounds for characteristic functions in terms of quantiles and entropy. Electronic Communications in Probability, 17:21:1–21:9, 2012. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/2053.

Briand:2000:CCT

[BCH^{+}00] Philippe Briand, François Coquet, Ying Hu, Jean Mémin, and Shige Peng. A converse comparison theorem for BSDEs and related properties of g-expectation. Electronic Communications in Probability, 5:13:101–13:117, 2000. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1025.

Balan:2022:PAM

[BCM22] Raluca Balan, Le Chen, and Yiping Ma. Parabolic Anderson model with rough noise in space and rough initial conditions. Electronic Communications in Probability, 27(??):1– 12, 2022. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-27/issue-none/Parabolic-Anderson-
model-with-rough-noise-in-space-and-rough/10.1214/22-
ECP506.full.

Beliayeu:2019:SGB

[BCNP19] Mikhail Beliayeu, Petr Chmel, Bhargav Narayanan, and Jan Petr. Slowdown for the geodesic-biased random walk. Electronic Communications in Probability, 24(??):73:1–73:8, ???? 2019. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1573722086.

Bertoin:2003:PTF

[BCP03] Jean Bertoin, Loic Chaumont, and Jim Pitman. Path transformations of first passage bridges. Electronic Communications in Probability, 8:17:155–17:166, 2003. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1096.

Bertini:2021:QES

[BCP21] Lorenzo Bertini, Nicoletta Cancrini, and Gustavo Posta. Quantitative ergodicity for the symmetric exclusion process with stationary initial data. Electronic Communications in Probability, 26(??):1–9, 2021. CODEN ???? ISSN 1083-589X.

Benigni:2018:HDR

[BCSW18] Lucas Benigni, Clément Cosco, Assaf Shapira, and Kay Jörg Wiese. Hausdorff dimension of the record set of a fractional Brownian motion. Electronic Communications in Probability, 23 (??):22:1–22:8, ???? 2018. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1522375381.

Beliaev:2020:NRB

[BCW20] Dmitry Beliaev, Valentina Cammarota, and Igor Wigman. No repulsion between critical points for planar Gaussian random fields. Electronic Communications in Probability, 25(??):1– 13, 2020. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-

in-probability/volume-25/issue-none/No-repulsion-between-
critical-points-for-planar-Gaussian-random-fields/10.
1214/20-ECP362.full.

Baudoin:2022:NFE

[BCY22] Fabrice Baudoin, Gunhee Cho, and Guang Yang. A note on first eigenvalue estimates by coupling methods in Kähler and quaternion Kähler manifolds. Electronic Communications in Probability, 27(??):1–8, 2022. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-27/issue-
none/A-note-on-first-eigenvalue-estimates-by-coupling-
methods-in/10.1214/22-ECP452.full.

Boivin:2002:GRR

[BD02] Daniel Boivin and Jean-Marc Derrien. Geodesics and recurrence of random walks in disordered systems. Electronic Communications in Probability, 7:11:101–11:115, 2002. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1052.

Basak:2013:LSD

[BD13] Anirban Basak and Amir Dembo. Limiting spectral distribution of sum of unitary and orthogonal matrices. Electronic Communications in Probability, 18:69:1–69:19, 2013. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2466.

Bobkov:2015:OTR

[BD15] Sergey G. Bobkov and Ying Ding. Optimal transport and Rényi informational divergence. Electronic Communications in Probability, 20(??):4:1–4:12, ???? 2015. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/3431.

Bandeira:2021:SNR

[BD21] Afonso S. Bandeira and Yunzi Ding. The spectral norm of random lifts of matrices. Electronic Communications in Probability, 26(??):1–10, 2021. CODEN ???? ISSN 1083-589X.

Bierme:2013:FPF

[BDE13] Hermine Biermé, Yann Demichel, and Anne Estrade. Fractional Poisson field and fractional Brownian field: why are they resembling but different? Electronic Communications in Probability, 18:11:1–11:13, 2013. CODEN ???? ISSN 1083-589X. URL http:

//ecp.ejpecp.org/article/view/1939.

Buchholz:2019:PPM

[BDKS19] Simon Buchholz, Jean-Dominique Deuschel, Noemi Kurt, and Florian Schweiger. Probability to be positive for the membrane model in dimensions 2 and 3. Electronic Communications in Probability, 24(??):44:1–44:14, ???? 2019. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1562292105.

Broutin:2015:CSB

[BDL15] Nicolas Broutin, Luc Devroye, and Gabor Lugosi. Connectivity of sparse Bluetooth networks. Electronic Communications in Probability, 20(??):48:1–48:10, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3644.

Briand:2001:DTT

[BDM01] Philippe Briand, Bernard Delyon, and Jean Mémin. Donskertype theorem for BSDEs. Electronic Communications in Probability, 6:1:1–1:14, 2001. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1030.

Bose:2007:MCE

[BDM07] Arup Bose, Amites Dasgupta, and Krishanu Maulik. Maxima of the cells of an equiprobable multinomial. Electronic Communications in Probability, 12:11:93–11:105, 2007. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1260.

Bianchi:2010:AIS

[BDN10] Pascal Bianchi, Mérouane Debbah, and Jamal Najim. Asymptotic independence in the spectrum of the Gaussian unitary ensemble. Electronic Communications in Probability, 15:35:376– 35:395, 2010. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1568.

Barbu:2011:RTP

[BDT11] Viorel Barbu, Giuseppe Da Prato, and Luciano Tubaro. A reflection type problem for the stochastic 2-D Navier–Stokes equations with periodic conditions. Electronic Communications in Probability, 16:29:304–29:313, 2011. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1633.

Bolthausen:2011:RTM

[BDZ11] Erwin Bolthausen, Jean-Dominique Deuschel, and Ofer Zeitouni. Recursions and tightness for the maximum of the discrete, two dimensional Gaussian free field. Electronic Communications in Probability, 16:11:114–11:119, 2011. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1610.

Bechtold:2022:RRT

[Bec22] Florian Bechtold. Regularization by random translation of potentials for the continuous PAM and related models in arbitrary dimension. Electronic Communications in Probability, 27(??): 1–13, 2022. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-27/issue-none/Regularization-by-
random-translation-of-potentials-for-the-continuous-
PAM/10.1214/22-ECP490.full.

Beghin:2014:GSP

[Beg14] Luisa Beghin. Geometric stable processes and related fractional differential equations. Electronic Communications in Probability, 19:13:1–13:14, 2014. CODEN ???? ISSN 1083-589X. URL http:

//ecp.ejpecp.org/article/view/2771.

Bartl:2021:LRW

[BEK21] Daniel Bartl, Stephan Eckstein, and Michael Kupper. Limits of random walks with distributionally robust transition probabilities. Electronic Communications in Probability, 26(??):1– 13, 2021. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-26/issue-none/Limits-of-random-
walks-with-distributionally-robust-transition-probabilities/
10.1214/21-ECP393.full.

Bertoin:2000:CMC

[Ber00] Jean Bertoin. The convex minorant of the Cauchy process. Electronic Communications in Probability, 5:5:51–5:55, 2000. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1017.

Bertoin:2010:TTS

[Ber10] Jean Bertoin. A two-time-scale phenomenon in a fragmentationcoagulation process. Electronic Communications in Probability, 15:23:253–23:262, 2010. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1552.

Berestycki:2017:EAG

[Ber17] Nathanaël Berestycki. An elementary approach to Gaussian multiplicative chaos. Electronic Communications in Probability, 22 (??):27:1–27:12, ???? 2017. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1494554429.

Bergqvist:2011:RPR

[BF11] Göran Bergqvist and Peter Forrester. Rank probabilities for real random N × N × 2 tensors. Electronic Communications in Probability, 16:55:630–55:637, 2011. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1655.

Benjamini:2016:SRC

[BFGG^{+}16] Itai Benjamini, Eric Foxall, Ori Gurel-Gurevich, Matthew Junge, and Harry Kesten. Site recurrence for coalescing random walk. Electronic Communications in Probability, 21(??):47:1–47:12, ???? 2016. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1466450523.

Beckman:2019:FMT

[BFJ^{+}19] Erin Beckman, Natalie Frank, Yufeng Jiang, Matthew Junge, and Si Tang. The frog model on trees with drift. Electronic Communications in Probability, 24(??):26:1–26:10, ???? 2019. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1559354659.

Borodin:2009:MDB

[BFP^{+}09] Alexei Borodin, Patrik Ferrari, Michael Prahofer, Tomohiro Sasamoto, and Jon Warren. Maximum of Dyson Brownian motion and non-colliding systems with a boundary. Electronic Communications in Probability, 14:47:486–47:494, 2009. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1503.

Bjorklund:2015:POD

[BFRH15] Johan Björklund, Victor Falgas-Ravry, and Cecilia Holmgren. On percolation in one-dimensional stable Poisson graphs. Electronic Communications in Probability, 20(??):50:1–50:6, ???? 2015. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/3958.

Benjamini:2013:ARR

[BFT13] Itai Benjamini, Hilary Finucane, and Romain Tessera. Algebraically recurrent random walks on groups. Electronic Communications in Probability, 18:28:1–28:8, 2013. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/2519.

Beffara:2010:SLP

[BFV10] Vincent Beffara, Sacha Friedli, and Yvan Velenik. Scaling limit of the prudent walk. Electronic Communications in Probability, 15:5:44–5:58, 2010. CODEN ???? ISSN 1083-589X. URL http:

//ecp.ejpecp.org/article/view/1527.

Bourgade:2007:EFP

[BFY07] Paul Bourgade, Takahiko Fujita, and Marc Yor. Euler’s formulae for ζ(2n) and products of Cauchy variables. Electronic Communications in Probability, 12:9:73–9:80, 2007. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1244.

Biroli:2020:LDL

[BG20] Giulio Biroli and Alice Guionnet. Large deviations for the largest eigenvalues and eigenvectors of spiked Gaussian random matrices. Electronic Communications in Probability, 25(??):1– 13, 2020. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-25/issue-none/Large-deviations-
for-the-largest-eigenvalues-and-eigenvectors-of-spiked/
10.1214/20-ECP343.full.

Bayraktar:2021:SEB

[BG21] Erhan Bayraktar and Gaoyue Guo. Strong equivalence between metrics of Wasserstein type. Electronic Communications in Probability, 26(??):1–13, 2021. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-26/issue-
none/Strong-equivalence-between-metrics-of-Wasserstein-
type/10.1214/21-ECP383.full.

Bertail:2008:EBM

[BGHK08] Patrice Bertail, Emmanuelle Gautherat, and Hugo Harari-Kermadec. Exponential bounds for multivariate self-normalized sums. Electronic Communications in Probability, 13:57:628– 57:640, 2008. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1430.

Benaych-Georges:2014:LEE

[BGP14] Florent Benaych-Georges and Sandrine Péché. Largest eigenvalues and eigenvectors of band or sparse random matrices. Electronic Communications in Probability, 19:4:1–4:9, 2014. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3027.

Bojdecki:2007:SEF

[BGT07] Tomasz Bojdecki, Luis Gorostiza, and Anna Talarczyk. Some extensions of fractional Brownian motion and sub-fractional Brownian motion related to particle systems. Electronic Communications in Probability, 12:17:161–17:172, 2007. CODEN

????  ISSN 1083-589X. URL http://ecp.ejpecp.org/
article/view/1272.

Bojdecki:2010:PSQ

[BGT10] Tomasz Bojdecki, Luis Gorostiza, and Anna Talarczyk. Particle systems with quasi-homogeneous initial states and their occupation time fluctuations. Electronic Communications in Probability, 15:18:191–18:202, 2010. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1547.

Bednorz:2021:TRL

[BGT21] Witold Bednorz, Grzegorz Głowienko, and Anna Talarczyk. Time regularity of Lévy-type evolution in Hilbert spaces and of some α-stable processes. Electronic Communications in Probability, 26(??):1–13, 2021. CODEN ???? ISSN 1083-589X.

Benaych-Georges:2018:ENN

[BGZ18] Florent Benaych-Georges and Ofer Zeitouni. Eigenvectors of non normal random matrices. Electronic Communications in Probability, 23(??):70:1–70:12, ???? 2018. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1539309734.

Balanca:2012:SIO

[BH12] Paul Balança and Erick Herbin. A set-indexed Ornstein– Uhlenbeck process. Electronic Communications in Probability, 17:39:1–39:14, 2012. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1903.

Bocharov:2016:LDR

[BH16] Sergey Bocharov and Simon C. Harris. Limiting distribution of the rightmost particle in catalytic branching Brownian motion. Electronic Communications in Probability, 21(??):70:1– 70:12, ???? 2016. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1475601101.

Bernstein:2022:CEC

[BHJR22] Emma Bernstein, Clare Hamblen, Matthew Junge, and Lily Reeves. Chase-escape on the configuration model. Electronic Communications in Probability, 27(??):1–14, 2022. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
journals/electronic-communications-in-probability/volume-
27/issue-none/Chase-escape-on-the-configuration-model/
10.1214/22-ECP470.full.

Bethuelsen:2021:QIP

[BHM21] Stein Andreas Bethuelsen, Christian Hirsch, and Christian Mönch. Quenched invariance principle for random walks on dynamically averaging random conductances. Electronic Communications in Probability, 26(??):1–13, 2021. CODEN ???? ISSN 1083-589X.

Beiglbock:2022:PSM

[BHN22] Mathias Beiglböck, David Hobson, and Dominykas Norgilas. The potential of the shadow measure. Electronic Communications in Probability, 27(??):1–12, 2022. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-27/issue-
none/The-potential-of-the-shadow-measure/10.1214/22-
ECP457.full.

Ben-Hamou:2018:CSR

[BHP18] Anna Ben-Hamou and Yuval Peres. Cutoff for a stratified random walk on the hypercube. Electronic Communications in Probability, 23(??):32:1–32:10, ???? 2018. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1527300061.

Bose:2010:SNC

[BHS10] Arup Bose, Rajat Hazra, and Koushik Saha. Spectral norm of circulant type matrices with heavy tailed entries. Electronic Communications in Probability, 15:29:299–29:313, 2010. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1554.

Bose:2011:HIH

[BHS11] Arup Bose, Rajat Hazra, and Koushik Saha. Half independence and half cumulants. Electronic Communications in Probability, 16:37:405–37:422, 2011. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1651.

Benaim:2018:UFC

[BHS18] Michel Benaïm, Tobias Hurth, and Edouard Strickler. A user-friendly condition for exponential ergodicity in randomly switched environments. Electronic Communications in Probability, 23(??):44:1–44:12, ???? 2018. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1532505675.

Buraczewski:2015:FLT

[BI15] Dariusz Buraczewski and Alexander Iksanov. Functional limit theorems for divergent perpetuities in the contractive case. Electronic Communications in Probability, 20(??):10:1–10:14, ???? 2015. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/3915.

Bianchi:2013:UIS

[Bia13] Luigi Bianchi. Uniqueness for an inviscid stochastic dyadic model on a tree. Electronic Communications in Probability, 18:8:1– 8:12, 2013. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/2382.

Biggins:2004:LDM

[Big04] J. D. Biggins. Large deviations for mixtures. Electronic Communications in Probability, 9:7:60–7:71, 2004. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/1106.

Birkner:2004:CWD

[Bir04] Matthias Birkner. A condition for weak disorder for directed polymers in random environment. Electronic Communications in Probability, 9:3:22–3:25, 2004. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1104.

Bossy:2018:PAL

[BJ18] Mireille Bossy and Jean-François Jabir. Particle approximation for Lagrangian Stochastic Models with specular boundary condition. Electronic Communications in Probability, 23(??):15:1– 15:14, ???? 2018. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1519722245.

Berard:2020:CPC

[BJ20] Jean Bérard and Nicolas Juillet. A coupling proof of convex ordering for compound distributions. Electronic Communications in Probability, 25(??):45:1–45:9, ???? 2020. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1593569165.

Bruckerhoff:2022:IMO

[BJ22] Martin Brückerhoff and Nicolas Juillet. Instability of martingale optimal transport in dimension. Electronic Communications in Probability, 27(??):1–10, 2022. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-27/issue-
none/Instability-of-martingale-optimal-transport-in-
dimension-d2/10.1214/22-ECP463.full.

Bjorner:2009:NRF

[Bjo09] Anders Bjorner. Note: Random-to-front shuffles on trees. Electronic Communications in Probability, 14:4:36–4:41, 2009. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1445.

Bjornberg:2015:LCR

[Bjö15] Jakob E. Björnberg. Large cycles in random permutations related to the Heisenberg model. Electronic Communications in Probability, 20(??):55:1–55:11, ???? 2015. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/4328.

Broman:2017:EPT

[BJT17] Erik I. Broman, Johan Jonasson, and Johan Tykesson. The existence phase transition for two Poisson random fractal models. Electronic Communications in Probability, 22(??):21:1–21:8, ???? 2017. CODEN ???? ISSN 1083-589X. URL http://

projecteuclid.org/euclid.ecp/1490839345.

Blath:2011:SEC

[BK11] Jochen Blath and Noemi Kurt. Survival and extinction of caring double-branching annihilating random walk. Electronic Communications in Probability, 16:26:271–26:282, 2011. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1631.

Berend:2013:CMM

[BK13] Daniel Berend and Aryeh Kontorovich. On the concentration of the missing mass. Electronic Communications in Probability, 18: 3:1–3:7, 2013. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/2359.

Brzezniak:2022:ESE

[BKP22] Zdzisław Brzeźniak, Tomasz Komorowski, and Szymon Peszat. Ergodicity for stochastic equations of Navier–Stokes type. Electronic Communications in Probability, 27(??):1–10, 2022. CO-DEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/journals/electronic-communications-in-probability/
volume-27/issue-none/Ergodicity-for-stochastic-equations-
of-NavierStokes-type/10.1214/21-ECP443.full.

Benjamini:2006:RWW

[BKR06] Itai Benjamini, Gady Kozma, and Dan Romik. Random walks with k-wise independent increments. Electronic Communications in Probability, 11:10:100–10:107, 2006. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1201.

Bansaye:2016:TPF

[BKS16] Vincent Bansaye, Thomas G. Kurtz, and Florian Simatos. Tightness for processes with fixed points of discontinuities and applications in varying environment. Electronic Communications in Probability, 21(??):81:1–81:9, ???? 2016. CODEN ????

ISSN 1083-589X. URL http://projecteuclid.org/euclid.
ecp/1480647626.

Balan:2010:ECC

[BL10] Raluca Balan and Sana Louhichi. Explicit conditions for the convergence of point processes associated to stationary arrays. Electronic Communications in Probability, 15:39:428–39:441, 2010. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1563.

Baik:2013:AAP

[BL13] Jinho Baik and Zhipeng Liu. On the average of the Airy process and its time reversal. Electronic Communications in Probability, 18:89:1–89:10, 2013. CODEN ???? ISSN 1083-589X. URL http:

//ecp.ejpecp.org/article/view/2952.

Bednorz:2008:RPC

[BLL08] Witold Bednorz, Krzysztof Latuszynski, and Rafal Latala. A regeneration proof of the Central Limit Theorem for uniformly ergodic Markov chains. Electronic Communications in Probability, 13:9:85–9:98, 2008. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1354.

Bar-Lev:2016:WDM

[BLL16] Shaul K. Bar-Lev and Gérard Letac. When does the minimum of a sample of an exponential family belong to an exponential family? Electronic Communications in Probability, 21(??):6:1– 6:8, ???? 2016. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1454514626.

Benaim:2012:QES

[BLMZ12] Michel Benaïm, Stéphane Le Borgne, Florent Malrieu, and Pierre-André Zitt. Quantitative ergodicity for some switched dynamical systems. Electronic Communications in Probability, 17: 56:1–56:14, 2012. CODEN ???? ISSN 1083-589X. URL http:/

/ecp.ejpecp.org/article/view/1932.

Blondel:2023:NAS

[Blo23] Oriane Blondel. A note on the antisymmetry in the speed of a random walk in reversible dynamic random environment. Electronic Communications in Probability, 28(??):1– 6, 2023. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-28/issue-none/A-note-on-the-antisymmetry-
in-the-speed-of-a/10.1214/23-ECP514.full.

Bandeira:2017:MPL

[BLR17] Afonso S. Bandeira, Asad Lodhia, and Philippe Rigollet. Marčenko–Pastur law for Kendall’s tau. Electronic Communications in Probability, 22(??):32:1–32:7, ???? 2017. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1496455233.

Boedihardjo:2015:UFD

[BLY15] Horatio Boedihardjo, Terry Lyons, and Danyu Yang. Uniform factorial decay estimates for controlled differential equations. Electronic Communications in Probability, 20(??):94:1– 94:11, ???? 2015. CODEN ???? ISSN 1083-589X. URL http:/

/ecp.ejpecp.org/article/view/4124.

Bertacchi:2018:SME

[BLZ18] Daniela Bertacchi, Jüri Lember, and Fabio Zucca. A stochastic model for the evolution of species with random fitness. Electronic Communications in Probability, 23(??):88:1–88:13, ???? 2018. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1543028983.

Bodineau:2005:UPL

[BM05] Thierry Bodineau and James Martin. A universality property for last-passage percolation paths close to the axis. Electronic Communications in Probability, 10:11:105–11:112, 2005. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1139.

Bertoin:2018:BMC

[BM18] Jean Bertoin and Bastien Mallein. Biggins’ martingale convergence for branching Lévy processes. Electronic Communications in Probability, 23(??):83:1–83:12, ???? 2018. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1540433049.

Boudabra:2020:RGT

[BM20] Maher Boudabra and Greg Markowsky. Remarks on Gross’ technique for obtaining a conformal Skorohod embedding of planar Brownian motion. Electronic Communications in Probability, 25 (??):20:1–20:13, ???? 2020. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1582945213.

Backhausz:2022:RGM

[BM22] Ágnes Backhausz and Tamás F. Móri. A random graph of moderate density. Electronic Communications in Probability, 27(??): 1–12, 2022. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-27/issue-none/A-random-graph-of-
moderate-density/10.1214/21-ECP444.full.

Bjornberg:2020:CRP

[BMMU20] Jakob E. Björnberg, Cécile Mailler, Peter Mörters, and Daniel Ueltschi. Characterising random partitions by random colouring. Electronic Communications in Probability, 25(??):4:1–4:12, ???? 2020. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1578906086.

Baldi:2007:CIG

[BMV07] Paolo Baldi, Domenico Marinucci, and Veeravalli Varadarajan. On the characterization of isotropic Gaussian fields on homogeneous spaces of compact groups. Electronic Communications in Probability, 12:29:291–29:302, 2007. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1316.

Breton:2008:EBN

[BN08] Jean-Christophe Breton and Ivan Nourdin. Error bounds on the non-normal approximation of Hermite power variations of fractional Brownian motion. Electronic Communications in Probability, 13:46:482–46:493, 2008. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1415.

Bell:2017:NLT

[BN17] Denis Bell and David Nualart. Noncentral limit theorem for the generalized Hermite process. Electronic Communications in Probability, 22(??):66:1–66:13, ???? 2017. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1511427621.

Boscain:2020:EBM

[BN20] Ugo Boscain and Robert W. Neel. Extensions of Brownian motion to a family of grushin-type singularities. Electronic Communications in Probability, 25(??):29:1–29:12, ???? 2020. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
euclid.ecp/1586332822.

Boufoussi:2003:SDF

[BO03] Brahim Boufoussi and Youssef Ouknine. On a SDE driven by a fractional Brownian motion and with monotone drift. Electronic Communications in Probability, 8:14:122–14:134, 2003. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1084.

Bobkov:2008:NDM

[Bob08] Sergey Bobkov. A note on the distributions of the maximum of linear Bernoulli processes. Electronic Communications in Probability, 13:26:266–26:271, 2008. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1375.

Borovkov:2010:DBM

[Bor10] Konstantin Borovkov. On the distribution of the Brownian motion process on its way to hitting zero. Electronic Communications in Probability, 15:26:281–26:285, 2010. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1555.

Bordenave:2011:SSP

[Bor11] Charles Bordenave. On the spectrum of sum and product of non-Hermitian random matrices. Electronic Communications in Probability, 16:10:104–10:113, 2011. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1606.

Bordenave:2013:ERM

[Bor13] Charles Bordenave. On Euclidean random matrices in high dimension. Electronic Communications in Probability, 18:25:1– 25:8, 2013. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/2340.

Borovkov:2022:NRB

[Bor22] Konstantin Borovkov. A note on recovering the Brownian motion component from a Lévy process. Electronic Communications in Probability, 27(??):1–6, 2022. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-27/issue-
none/A-note-on-recovering-the-Brownian-motion-component-
from-a/10.1214/22-ECP477.full.

Bottcher:2011:CFP

[Böt11] Björn Böttcher. On the construction of Feller processes with unbounded coefficients. Electronic Communications in Probability, 16:48:545–48:555, 2011. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1652.

Bourguin:2016:VVS

[Bou16] Solesne Bourguin. Vector-valued semicircular limits on the free Poisson chaos. Electronic Communications in Probability, 21(??):

55:1–55:11, ???? 2016. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1472830237.

Biggins:2009:LDR

[BP09] J. D. Biggins and D. B. Penman. Large deviations in randomly coloured random graphs. Electronic Communications in Probability, 14:29:290–29:301, 2009. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1478.

Best:2010:ASM

[BP10] Katharina Best and Peter Pfaffelhuber. The Aldous–Shields model revisited with application to cellular ageing. Electronic Communications in Probability, 15:43:475–43:488, 2010. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1581.

Burdzy:2020:BPR

[BP20] Krzysztof Burdzy and Jim Pitman. Bounds on the probability of radically different opinions. Electronic Communications in Probability, 25(??):14:1–14:12, ???? 2020. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1581325211.

Benjamini:2021:HSA

[BP21] Itai Benjamini and Christoforos Panagiotis. Hyperbolic self avoiding walk. Electronic Communications in Probability, 26 (??):1–5, 2021. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-26/issue-none/Hyperbolic-self-
avoiding-walk/10.1214/21-ECP388.full.

Bertoin:1999:CBP

[BPR99] Jean Bertoin, Jim Pitman, and Juan Ruiz de Chavez. Constructions of a Brownian path with a given minimum. Electronic Communications in Probability, 4:5:31–5:37, 1999. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1003.

Berti:2013:SRT

[BPR13] Patrizia Berti, Luca Pratelli, and Pietro Rigo. A Skorohod representation theorem without separability. Electronic Communications in Probability, 18:80:1–80:12, 2013. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/2793.

Berti:2015:GLS

[BPR15] Patrizia Berti, Luca Pratelli, and Pietro Rigo. Gluing lemmas and Skorohod representations. Electronic Communications in Probability, 20(??):53:1–53:11, ???? 2015. CODEN

????  ISSN 1083-589X. URL http://ecp.ejpecp.org/
article/view/3870.

Berard:2007:CLT

[BR07] Jean Berard and Alejandro Ramirez. Central Limit Theorem for the excited random walk in dimension d ≥ 2. Electronic Communications in Probability, 12:30:303–30:314, 2007. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1317.

Bruggeman:2016:ODD

[BR16] Cameron Bruggeman and Johannes Ruf. A one-dimensional diffusion hits points fast. Electronic Communications in Probability, 21(??):22:1–22:7, ???? 2016. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1457617913.

Bardina:2010:WAF

[BRT10] Xavier Bardina, Carles Rovira, and Samy Tindel. Weak approximation of fractional SDEs: the Donsker setting. Electronic Communications in Probability, 15:30:314–30:329, 2010. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1561.

Benjamini:1996:PBM

[BS96] Itai Benjamini and Oded Schramm. Percolation beyond Z^{d}, many questions and a few answers. Electronic Communications in Probability, 1:8:71–8:82, 1996. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/978.

Bose:2007:APR

[BS07a] Arup Bose and Arnab Sen. On asymptotic properties of the rank of a special random adjacency matrix. Electronic Communications in Probability, 12:20:200–20:205, 2007. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1266.

Bose:2007:SNR

[BS07b] Arup Bose and Arnab Sen. Spectral norm of random large dimensional noncentral Toeplitz and Hankel matrices. Electronic Communications in Probability, 12:3:21–3:27, 2007. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1243.

Bertoin:2016:LES

[BS16] Jean Bertoin and Robin Stephenson. Local explosion in selfsimilar growth-fragmentation processes. Electronic Communications in Probability, 21(??):66:1–66:12, ???? 2016. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1473854582.

Basu:2017:KII

[BS17] Deepan Basu and Artem Sapozhnikov. Kesten’s incipient infinite cluster and quasi-multiplicativity of crossing probabilities. Electronic Communications in Probability, 22(??):26:1– 26:12, ???? 2017. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1494036081.

Basdevant:2018:ASA

[BS18a] A.-L. Basdevant and A. Singh. Almost-sure asymptotics for the number of heaps inside a random sequence. Electronic Communications in Probability, 23(??):17:1–17:8, ???? 2018. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
euclid.ecp/1520391724.

Benjamini:2018:WDR

[BS18b] Itai Benjamini and Alexander Shamov. Where does a random process hit a fractal barrier? Electronic Communications in Probability, 23(??):25:1–25:5, ???? 2018. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1524881133.

Barbier:2022:MSS

[BS22] Jean Barbier and Manuel Sáenz. Marginals of a spherical spin glass model with correlated disorder. Electronic Communications in Probability, 27(??):1–12, 2022. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-27/issue-
none/Marginals-of-a-spherical-spin-glass-model-with-
correlated-disorder/10.1214/22-ECP489.full.

Bourguin:2011:CTG

[BT11] Solesne Bourguin and Ciprian Tudor. Cramér theorem for gamma random variables. Electronic Communications in Probability, 16:34:365–34:378, 2011. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1639.

Boucheron:2012:CIO

[BT12] Stéphane Boucheron and Maud Thomas. Concentration inequalities for order statistics. Electronic Communications in Probability, 17:51:1–51:12, 2012. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/2210.

Benjamini:2017:FPP

[BT17] Itai Benjamini and Romain Tessera. First passage percolation on a hyperbolic graph admits bi-infinite geodesics. Electronic Communications in Probability, 22(??):14:1–14:8, ???? 2017. CO-DEN ???? ISSN 1083-589X. URL http://projecteuclid.

org/euclid.ecp/1487062816.

Bercu:2019:NIC

[BT19] Bernard Bercu and Taieb Touati. New insights on concentration inequalities for self-normalized martingales. Electronic Communications in Probability, 24(??):63:1–63:12, ???? 2019. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
euclid.ecp/1570845629.

Bowditch:2020:SBR

[BT20] Adam Bowditch and Yuki Tokushige. The speed of a biased random walk on a Galton–Watson tree is analytic. Electronic Communications in Probability, 25(??):1–11, 2020. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
journals/electronic-communications-in-probability/volume-
25/issue-none/The-speed-of-a-biased-random-walk-on-a-
Galton/10.1214/20-ECP344.full. See retraction notice [?].

Bowditch:2021:RSB

[BT21] Adam Bowditch and Yuki Tokushige. Retraction of: The speed of a biased random walk on a Galton–Watson tree is analytic. Electronic Communications in Probability, 26(??):??, 2021. CO-DEN ???? ISSN 1083-589X. See [BT20].

Buckley:2013:AHK

[Buc13] Stephen Buckley. Anomalous heat kernel behaviour for the dynamic random conductance model. Electronic Communications in Probability, 18:1:1–1:11, 2013. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/2525.

Buonaguidi:2021:DRQ

[Buo21] Bruno Buonaguidi. On the dimension reduction in the quickest detection problem for diffusion processes with exponential penalty for the delay. Electronic Communications in Probability, 26(??):1–12, 2021. CODEN ???? ISSN 1083-589X.

Burdzy:2019:SBH

[Bur19] Krzysztof Burdzy. Simultaneous boundary hitting by coupled reflected Brownian motions. Electronic Communications in Probability, 24(??):22:1–22:12, ???? 2019. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1555034600.

Butez:2017:LDB

[But17] Raphaël Butez. Large deviations for biorthogonal ensembles and variational formulation for the Dykema–Haagerup distribution. Electronic Communications in Probability, 22(??):37:1– 37:11, ???? 2017. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1499068820.

Butez:2018:LRR

[But18] Raphaël Butez. The largest root of random Kac polynomials is heavy tailed. Electronic Communications in Probability, 23(??): 20:1–20:9, ???? 2018. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1521079421.

Backhoff-Veraguas:2020:WMR

[BVBP20] Julio Backhoff-Veraguas, Mathias Beiglböck, and Gudmund Pammer. Weak monotone rearrangement on the line. Electronic Communications in Probability, 25(??):18:1–18:16, ???? 2020. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1582254306.

Benjamini:2003:ERW

[BW03] Itai Benjamini and David Wilson. Excited random walk. Electronic Communications in Probability, 8:9:86–9:92, 2003. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1072.

Burdzy:2004:GO

[BW04] Krzysztof Burdzy and David White. A Gaussian oscillator. Electronic Communications in Probability, 9:10:92–10:95, 2004. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1113.

Burdzy:2008:MPP

[BW08] Krzysztof Burdzy and David White. Markov processes with product-form stationary distribution. Electronic Communications in Probability, 13:56:614–56:627, 2008. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/1428.

Bertoin:2001:SSS

[BY01] Jean Bertoin and Marc Yor. On subordinators, self-similar Markov processes and some factorizations of the exponential variable. Electronic Communications in Probability, 6:10:95–10:106, 2001. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1039.

Bertoin:2013:PJI

[BY13] Jean Bertoin and Marc Yor. Pure jump increasing processes and the change of variables formula. Electronic Communications in Probability, 18:41:1–41:7, 2013. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/2700.

Benjamini:2007:MAP

[BYZ07] Itai Benjamini, Ariel Yadin, and Ofer Zeitouni. Maximal arithmetic progressions in random subsets. Electronic Communications in Probability, 12:35:365–35:376, 2007. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/

view/1321. See erratum [BYZ12].

Benjamini:2012:EMA

[BYZ12] Itai Benjamini, Ariel Yadin, and Ofer Zeitouni. Erratum: Maximal arithmetic progressions in random subsets. Electronic Communications in Probability, 17:18:1, 2012. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2014. See [BYZ07].

Balan:2006:SAM

[BZ06] Raluca Balan and Ingrid-Mona Zamfirescu. Strong approximation for mixing sequences with infinite variance. Electronic Communications in Probability, 11:2:11–2:23, 2006. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1175.

Bayraktar:2016:RBM

[BZ16] Erhan Bayraktar and Yuchong Zhang. A rank-based mean field game in the strong formulation. Electronic Communications in Probability, 21(??):72:1–72:12, ???? 2016. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1476841494.

Butez:2017:ULD

[BZ17] Raphaël Butez and Ofer Zeitouni. Universal large deviations for Kac polynomials. Electronic Communications in Probability, 22 (??):6:1–6:10, ???? 2017. CODEN ???? ISSN 1083-589X. URL

http://projecteuclid.org/euclid.ecp/1483952415.

Borovkov:2018:MDS

[BZ18] Konstantin Borovkov and Mikhail Zhitlukhin. On the maximum of the discretely sampled fractional Brownian motion with small Hurst parameter. Electronic Communications in Probability, 23 (??):65:1–65:8, ???? 2018. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1537257726.

Can:2015:CPO

[Can15] Van Hao Can. Contact process on one-dimensional long range percolation. Electronic Communications in Probability, 20(??): 93:1–93:11, ???? 2015. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/4461.

Carlsson:2005:SNT

[Car05] Niclas Carlsson. Some notes on topological recurrence. Electronic Communications in Probability, 10:9:82–9:93, 2005. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1137.

Caravenna:2018:MCR

[Car18] Francesco Caravenna. On the maximum of conditioned random walks and tightness for pinning models. Electronic Communications in Probability, 23(??):69:1–69:13, ???? 2018. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1539309733.

Carfagnini:2022:AGC

[Car22] Marco Carfagnini. An application of the Gaussian correlation inequality to the small deviations for a Kolmogorov diffusion. Electronic Communications in Probability, 27(??):1– 7, 2022. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-27/issue-none/An-application-of-
the-Gaussian-correlation-inequality-to-the-small/10.
1214/22-ECP459.full.

Caravenna:2010:LDP

[CB10] Francesco Caravenna and Martin Borecki. Localization for (1 + 1)-dimensional pinning models with (∇ + Δ)-interaction. Electronic Communications in Probability, 15:48:534–48:548, 2010. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1584.

Carmona:1998:FBM

[CC98] Philippe Carmona and Laure Coutin. Fractional Brownian motion and the Markov property. Electronic Communications in Probability, 3:12:95–12:107, 1998. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/998.

Chiu:2018:PQV

[CC18] Henry Chiu and Rama Cont. On pathwise quadratic variation for càdlàg functions. Electronic Communications in Probability, 23(??):85:1–85:12, ???? 2018. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1542942174.

Chen:2019:LTT

[CCGS19] Dayue Chen, Peng Chen, Nina Gantert, and Dominik Schmid. Limit theorems for the tagged particle in exclusion processes on regular trees. Electronic Communications in Probability, 24(??): 2:1–2:10, ???? 2019. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1548299047.

Chiarini:2015:NEP

[CCH15] Alberto Chiarini, Alessandra Cipriani, and Rajat Subhra Hazra. A note on the extremal process of the supercritical Gaussian Free Field. Electronic Communications in Probability, 20(??): 74:1–74:10, ???? 2015. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/4332.

Carmona:2013:MFF

[CD13] René Carmona and François Delarue. Mean field forwardbackward stochastic differential equations. Electronic Communications in Probability, 18:68:1–68:15, 2013. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/2446.

Cator:2017:SAS

[CD17a] Eric Cator and Henk Don. Self-averaging sequences which fail to converge. Electronic Communications in Probability, 22(??): 16:1–16:12, ???? 2017. CODEN ???? ISSN 1083-589X. URL

http://projecteuclid.org/euclid.ecp/1487386904.

Cerf:2017:MCR

[CD17b] Raphaël Cerf and Joseba Dalmau. A Markov chain representation of the normalized Perron–Frobenius eigenvector. Electronic Communications in Probability, 22(??):52:1–52:6, ???? 2017. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1507860209.

Coutin:2020:DTW

[CD20] Laure Coutin and Laurent Decreusefond. Donsker’s theorem in Wasserstein-1 distance. Electronic Communications in Probability, 25(??):27:1–27:13, ???? 2020. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1585620027.

Caravenna:2013:GSI

[CdH13] Francesco Caravenna and Frank den Hollander. A general smoothing inequality for disordered polymers. Electronic Communications in Probability, 18:76:1–76:15, 2013. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/2874.

Chang:2017:SIM

[CDNX17] Jiawei Chang, Nick Duffield, Hao Ni, and Weijun Xu. Signature inversion for monotone paths. Electronic Communications in Probability, 22(??):42:1–42:11, ???? 2017. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1502762748.

Couronne:2011:CSP

[CEG11] Olivier Couronné, Nathanaël Enriquez, and Lucas Gerin. Construction of a short path in high-dimensional first passage percolation. Electronic Communications in Probability, 16:3:22–3:28, 2011. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1595.

Chassagneux:2011:NEU

[CEK11] Jean François Chassagneux, Romuald Elie, and Idris Kharroubi. A note on existence and uniqueness for solutions of multidimensional reflected BSDEs. Electronic Communications in Probability, 16:12:120–12:128, 2011. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1614.

Cerf:2014:TTF

[Cer14] Raphaël Cerf. The travel time in a finite box in supercritical Bernoulli percolation. Electronic Communications in Probability, 19:20:1–20:9, 2014. CODEN ???? ISSN 1083-589X. URL http:

//ecp.ejpecp.org/article/view/3015.

Cetin:2012:FAM

[Çet12] Umut Çetin. Filtered Azéma martingales. Electronic Communications in Probability, 17:62:1–62:13, 2012. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/2310.

Christensen:2020:NNS

[CF20] Sören Christensen and Simon Fischer. Note on the (non- )smoothness of discrete time value functions in optimal stopping. Electronic Communications in Probability, 25(??):1–10, 2020. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-25/issue-none/Note-on-the-non-
smoothness-of-discrete-time-value-functions/10.1214/
20-ECP335.full.

Caravenna:2005:CAB

[CG05] Francesco Caravenna and Giambattista Giacomin. On constrained annealed bounds for pinning and wetting models. Electronic Communications in Probability, 10:18:179–18:189, 2005. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1150.

Cerf:2015:LBR

[CG15] Raphaël Cerf and Matthias Gorny. A lower bound on the relative entropy with respect to a symmetric probability. Electronic Communications in Probability, 20(??):5:1–5:5, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3920.

Chen:2019:BPC

[CGP19] Xinxin Chen and Nadine Guillotin-Plantard. Branching processes in correlated random environment. Electronic Communications in Probability, 24(??):71:1–71:13, ???? 2019. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
euclid.ecp/1573528176.

Castell:2013:OSE

[CGPPS13] Fabienne Castell, Nadine Guillotin-Plantard, Françoise Pène, and Bruno Schapira. On the one-sided exit problem for stable processes in random scenery. Electronic Communications in Probability, 18:33:1–33:7, 2013. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/2444.

Cattiaux:2010:PIC

[CGR10] Patrick Cattiaux, Arnaud Guillin, and Cyril Roberto. Poincaré inequality and the L^{p} convergence of semi-groups. Electronic Communications in Probability, 15:25:270–25:280, 2010. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1559.

Chen:2021:NPL

[CGS21] Xinxin Chen, Christophe Garban, and Atul Shekhar. A new proof of Liggett’s theorem for non-interacting Brownian motions. Electronic Communications in Probability, 26(??):1–12, 2021. CODEN ???? ISSN 1083-589X.

Carmona:1996:SSO

[CGXM96] Rene Carmona, Stanislav Grishin, Lin Xu, and Stanislav Molchanov. Surface stretching for Ornstein–Uhlenbeck velocity fields. Electronic Communications in Probability, 2:1:1–1:11, 1996. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/980.

Carmona:1997:SSO

[CGXM97] Rene A. Carmona, Stanislav Grishin, Lin Xu, and Stanislav Molchanov. Surface stretching for Ornstein Uhlenbeck velocity fields. Electronic Communications in Probability, 2:1:1–1:11, 1997. CODEN ???? ISSN 1083-589X.

Csaki:2004:IPR

[CH04] Endre Csaki and Yueyun Hu. Invariance principles for ranked excursion lengths and heights. Electronic Communications in Probability, 9:2:14–2:21, 2004. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1103.

Couzinie:2021:WRP

[CH21] Yannick Couzinié and Christian Hirsch. Weakly reinforced Pólya urns on countable networks. Electronic Communications in Probability, 26(??):1–10, 2021. CODEN ???? ISSN 1083-589X.

Chakrabarty:2010:CLT

[Cha10] Arijit Chakrabarty. Central Limit Theorem for truncated heavy tailed Banach valued random vectors. Electronic Communications in Probability, 15:33:346–33:364, 2010. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/1564.

Chang:2015:CRL

[Cha15] Yinshan Chang. A convergence result on the lengths of Markovian loops. Electronic Communications in Probability, 20(??): 73:1–73:11, ???? 2015. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/4263.

Chan:2020:RCM

[Cha20] Swee Hong Chan. A rotor configuration with maximum escape rate. Electronic Communications in Probability, 25(??): 19:1–19:5, ???? 2020. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1582686118.

Chen:2017:MPL

[Che17] Joe P. Chen. The moving particle lemma for the exclusion process on a weighted graph. Electronic Communications in Probability, 22(??):47:1–47:13, ???? 2017. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1506931447.

Chen:2021:ATT

[Che21a] Wei-Kuo Chen. On the Almeida–Thouless transition line in the Sherrington–Kirkpatrick model with centered Gaussian external field. Electronic Communications in Probability, 26(??):1–9, 2021. CODEN ???? ISSN 1083-589X.

Chen:2021:MLE

[Che21b] Xiaohui Chen. Maximum likelihood estimation of potential energy in interacting particle systems from single-trajectory data.

Electronic Communications in Probability, 26(??):1–13, 2021. CODEN ???? ISSN 1083-589X.

Chen:2021:PAS

[Che21c] Yu-Ting Chen. Precise asymptotics of some meeting times arising from the voter model on large random regular graphs. Electronic Communications in Probability, 26(??):1–13, 2021. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
journals/electronic-communications-in-probability/volume-
26/issue-none/Precise-asymptotics-of-some-meeting-times-
arising-from-the-voter/10.1214/21-ECP373.full.

Capitaine:1997:MRS

[CHL97] Mireille Capitaine, Elton P. Hsu, and Michel Ledoux. Martingale representation and a simple proof of logarithmic Sobolev inequalities on path spaces. Electronic Communications in Probability, 2:7:71–7:81, 1997. CODEN ???? ISSN 1083-589X. URL http:

//ecp.ejpecp.org/article/view/986.

Cortines:2021:MSE

[CHL21] Aser Cortines, Lisa Hartung, and Oren Louidor. More on the structure of extreme level sets in branching Brownian motion. Electronic Communications in Probability, 26(??):1–14, 2021. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-26/issue-none/More-on-the-structure-
of-extreme-level-sets-in-branching/10.1214/20-ECP369.
full.

Choi:2018:HTM

[Cho18] Michael C. H. Choi. Hitting time and mixing time bounds of Stein’s factors. Electronic Communications in Probability, 23 (??):6:1–6:10, ???? 2018. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1518663615.

Cannings:2013:RWA

[CJ13] Chris Cannings and Jonathan Jordan. Random walk attachment graphs. Electronic Communications in Probability, 18:77:1– 77:5, 2013. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/2518.

Cotar:2018:EDR

[CJK18] Codina Cotar, Benedikt Jahnel, and Christof Külske. Extremal decomposition for random Gibbs measures: from general metastates to metastates on extremal random Gibbs measures. Electronic Communications in Probability, 23(??):95:1– 95:12, ???? 2018. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1545102491.

Chigansky:2008:DBM

[CK08] Pavel Chigansky and Fima Klebaner. Distribution of the Brownian motion on its way to hitting zero. Electronic Communications in Probability, 13:58:641–58:648, 2008. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1432.

Chigansky:2012:CPA

[CK12] Pavel Chigansky and Fima Klebaner. Compound Poisson approximation for triangular arrays with application to threshold estimation. Electronic Communications in Probability, 17:29:1– 29:10, 2012. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/2009.

Cox:2014:MEM

[CK14] Alexander Cox and Martin Klimmek. From minimal embeddings to minimal diffusions. Electronic Communications in Probability, 19:46:1–46:13, 2014. CODEN ???? ISSN 1083-589X. URL http:

//ecp.ejpecp.org/article/view/2889.

Candellero:2018:CBM

[CK18a] Elisabetta Candellero and Wilfrid S. Kendall. Coupling of Brownian motions in Banach spaces. Electronic Communications in Probability, 23(??):9:1–9:13, ???? 2018. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1519182084.

Cygan:2018:RML

[CK18b] Wojciech Cygan and Judith Kloas. On recurrence of the multidimensional Lindley process. Electronic Communications in Probability, 23(??):4:1–4:14, ???? 2018. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1518426010.

Csaki:1999:CEB

[CKS99] Endre Csáki, Davar Khoshnevisan, and Zhan Shi. Capacity estimates, boundary crossings and the Ornstein–Uhlenbeck process in Wiener space. Electronic Communications in Probability, 4: 13:103–13:109, 1999. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1011.

Chigansky:2006:RPF

[CL06] Pavel Chigansky and Robert Liptser. On a role of predictor in the filtering stability. Electronic Communications in Probability, 11:13:129–13:140, 2006. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1205.

Chatterjee:2009:OAS

[CL09] Sourav Chatterjee and Michel Ledoux. An observation about submatrices. Electronic Communications in Probability, 14: 48:495–48:500, 2009. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1504.

Chen:2014:GPU

[CL14] Jun Chen and Cyrille Lucas. A generalized Pólya’s urn with graph based interactions: convergence at linearity. Electronic Communications in Probability, 19(??):67:1–67:13, ???? 2014. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/3094.

Clark:2014:OVR

[Cla14] Jeremy Clark. Optimizing a variable-rate diffusion to hit an infinitesimal target at a set time. Electronic Communications in Probability, 19:48:1–48:19, 2014. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/2846.

Conforti:2015:BMC

[CLMR15] Giovanni Conforti, Christian Léonard, Rüdiger Murr, and Sylvie Rœlly. Bridges of Markov counting processes. reciprocal classes and duality formulas. Electronic Communications in Probability, 20(??):18:1–18:12, ???? 2015. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/3697.

Chobanyan:2005:SLL

[CLS05] Sergei Chobanyan, Shlomo Levental, and Habib Salehi. Strong law of large numbers under a general moment condition. Electronic Communications in Probability, 10:22:218–22:222, 2005. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1156.

Cammarota:2012:MVS

[CM12a] Valentina Cammarota and Peter Mörters. On the most visited sites of planar Brownian motion. Electronic Communications in Probability, 17:15:1–15:9, 2012. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1809.

Cranston:2012:CIS

[CM12b] Michael Cranston and Stanislav Molchanov. On a concentration inequality for sums of independent isotropic vectors. Electronic Communications in Probability, 17:27:1–27:8, 2012. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2063.

Chleboun:2013:MTB

[CM13] Paul Chleboun and Fabio Martinelli. Mixing time bounds for oriented kinetically constrained spin models. Electronic Communications in Probability, 18:60:1–60:9, 2013. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/2516.

Chaumont:2018:SPE

[CM18a] Loïc Chaumont and Jacek Małecki. Short proofs in extrema of spectrally one sided Lévy processes. Electronic Communications in Probability, 23(??):55:1–55:12, ???? 2018. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1535767266.

Cortines:2018:GES

[CM18b] Aser Cortines and Bastien Mallein. The genealogy of an exactly solvable Ornstein–Uhlenbeck type branching process with selection. Electronic Communications in Probability, 23(??):98:1– 98:13, ???? 2018. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1545102494.

Curien:2018:HFP

[CM18c] Nicolas Curien and Cyril Marzouk. How fast planar maps get swallowed by a peeling process. Electronic Communications in Probability, 23(??):18:1–18:11, ???? 2018. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1520391725.

Chen:2022:TNB

[CM22] Xinxin Chen and Bastien Mallein. Total number of births on the negative half-line of the binary branching Brownian motion in the boundary case. Electronic Communications in Probability, 27 (??):1–11, 2022. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-27/issue-none/Total-number-of-
births-on-the-negative-half-line-of/10.1214/22-ECP449.
full.

Contreras:2023:LPG

[CMT23] Daniel Contreras, Sébastien Martineau, and Vincent Tassion. Locality of percolation for graphs with polynomial growth. Electronic Communications in Probability, 28(??):1– 9, 2023. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-28/issue-none/Locality-of-percolation-
for-graphs-with-polynomial-growth/10.1214/22-ECP508.
full.

Campese:2016:MGA

[CNPP16] Simon Campese, Ivan Nourdin, Giovanni Peccati, and Guillaume Poly. Multivariate Gaussian approximations on Markov chaoses. Electronic Communications in Probability, 21(??):48:1– 48:9, ???? 2016. CODEN ???? ISSN 1083-589X. URL https:/

/projecteuclid.org/euclid.ecp/1467399737.

Comman:2008:SWT

[Com08] Henri Comman. Stone–Weierstrass type theorems for large deviations. Electronic Communications in Probability, 13:22:225– 22:240, 2008. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1370.

Conforti:2016:BMC

[Con16] Giovanni Conforti. Bridges of Markov counting processes: quantitative estimates. Electronic Communications in Probability, 21 (??):19:1–19:13, ???? 2016. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1456499563.

Coupier:2011:MGS

[Cou11] David Coupier. Multiple geodesics with the same direction. Electronic Communications in Probability, 16:46:517–46:527, 2011. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1656.

Conlon:2005:HND

[CP05] Joseph Conlon and Ian Pilizzotto. On homogenization of non-divergence form partial difference equations. Electronic Communications in Probability, 10:13:125–13:135, 2005. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1141.

Curien:2011:RLM

[CP11] Nicolas Curien and Yuval Peres. Random laminations and multitype branching processes. Electronic Communications in Probability, 16:39:435–39:446, 2011. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1641.

Czuppon:2014:SLR

[CP14] Peter Czuppon and Peter Pfaffelhuber. Some limit results for Markov chains indexed by trees. Electronic Communications in Probability, 19(??):77:1–77:11, ???? 2014. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3601.

Can:2017:CTM

[CP17a] Van Hao Can and Viet-Hung Pham. A Cramér type moderate deviation theorem for the critical Curie–Weiss model. Electronic Communications in Probability, 22(??):62:1–62:12, ???? 2017. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1510736419.

Cuneo:2017:RRS

[CP17b] Noé Cuneo and Christophe Poquet. On the relaxation rate of short chains of rotors interacting with Langevin thermostats. Electronic Communications in Probability, 22(??):35:1–35:8, ???? 2017. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1498010648.

Cancrini:2019:PCB

[CP19a] Nicoletta Cancrini and Gustavo Posta. Propagation of chaos for a balls into bins model. Electronic Communications in Probability, 24(??):1:1–1:9, ???? 2019. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1546571102.

Cheliotis:2019:SME

[CP19b] Dimitris Cheliotis and Nickos Papadatos. On sequential maxima of exponential sample means, with an application to ruin probability. Electronic Communications in Probability, 24(??): 74:1–74:7, ???? 2019. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1575363644.

Cancrini:2020:MTR

[CP20] Nicoletta Cancrini and Gustavo Posta. Mixing time for the Repeated Balls into Bins dynamics. Electronic Communications in Probability, 25(??):1–14, 2020. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-25/issue-
none/Mixing-time-for-the-Repeated-Balls-into-Bins-dynamics/
10.1214/20-ECP338.full.

Chiclana:2022:NCS

[CP22] Rafael Chiclana and Yuval Peres. No cutoff in spherically symmetric trees. Electronic Communications in Probability, 27(??): 1–11, 2022. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-27/issue-none/No-cutoff-in-Spherically-
symmetric-trees/10.1214/22-ECP468.full.

Cichomski:2023:CPB

[CP23] Stanisław Cichomski and Fedor Petrov. A combinatorial proof of the Burdzy–Pitman conjecture. Electronic Communications in Probability, 28(??):1–7, 2023. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-28/issue-
none/A-combinatorial-proof-of-the-BurdzyPitman-conjecture/
10.1214/23-ECP512.full.

Cator:2012:IIC

[CPS12] Eric Cator, Leandro Pimentel, and Marcio Souza. Influence of the initial condition in equilibrium last-passage percolation models. Electronic Communications in Probability, 17:7:1–7:7, 2012. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1727.

Catuogno:2005:GSD

[CR05] Pedro Catuogno and Paulo Ruffino. Geometry of stochastic delay differential equations. Electronic Communications in Probability, 10:19:190–19:195, 2005. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1151.

Crane:2013:CMB

[Cra13] Harry Crane. Consistent Markov branching trees with discrete edge lengths. Electronic Communications in Probability, 18:73:1– 73:14, 2013. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/2872.

Criens:2021:DYW

[Cri21] David Criens. A dual Yamada -Watanabe theorem for lévy driven stochastic differential equations. Electronic Communications in Probability, 26(??):1–10, 2021. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-26/issue-
none/A-dual-YamadaWatanabe-theorem-for-L%c3%a9vy-driven-
stochastic-differential-equations/10.1214/21-ECP384.
full.

Corwin:2014:EAL

[CS14] Ivan Corwin and Xin Sun. Ergodicity of the Airy line ensemble. Electronic Communications in Probability, 19:49:1–49:11, 2014. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3504.

Cerny:2016:MTR

[ČS16] Jiří Černý and Artem Sapozhnikov. Mixing time for the random walk on the range of the random walk on tori. Electronic Communications in Probability, 21(??):26:1–26:10, ???? 2016. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1457617917.

Chen:2013:NFP

[CSC13] Yu-Ting Chen, Yuan-Chung Sheu, and Ming-Chi Chang. A note on first passage functionals for hyper-exponential jump-diffusion processes. Electronic Communications in Probability, 18:2:1–2:8, 2013. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/2017.

Cranston:1999:LEI

[CSS99] Michael Cranston, Michael Scheutzow, and David Steinsaltz. Linear expansion of isotropic Brownian flows. Electronic Communications in Probability, 4:12:91–12:101, 1999. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/1010.

Cox:2007:SRT

[CV07] Sonja Cox and Mark Veraar. Some remarks on tangent martingale difference sequences in L^{1}-spaces. Electronic Communications in Probability, 12:40:421–40:433, 2007. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/1328.

Cui:2013:ABN

[CY13] Jing Cui and Litan Yan. Asymptotic behavior for neutral stochastic partial differential equations with infinite delays. Electronic Communications in Probability, 18:45:1–45:12, 2013. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2858.

Cai:2021:SRR

[CZ21] Zhenhao Cai and Yuan Zhang. Some rigorous results on the phase transition of finitary random interlacements. Electronic Communications in Probability, 26(??):1–11, 2021. CODEN ???? ISSN 1083-589X.

Daly:2013:CPA

[Dal13] Fraser Daly. Compound Poisson approximation with association or negative association via Stein’s method. Electronic Communications in Probability, 18:30:1–30:12, 2013. CODEN

????  ISSN 1083-589X. URL http://ecp.ejpecp.org/
article/view/2751.

Daly:2017:MFS

[Dal17] Fraser Daly. On magic factors in Stein’s method for compound Poisson approximation. Electronic Communications in Probability, 22(??):67:1–67:10, ???? 2017. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1511427622.

Duheille-Bienvenue:2003:CLT

[DBGP03] Frédérique Duheille-Bienvenue and Nadine Guillotin-Plantard. Central Limit Theorems for the products of random matrices sampled by a random walk. Electronic Communications in Probability, 8:5:43–5:50, 2003. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1068.

delBarrio:2013:BPS

[dBJP13] Eustasio del Barrio, Arnold Janssen, and Markus Pauly. The m(n) out of k(n) bootstrap for partial sums of St. Petersburg type games. Electronic Communications in Probability, 18:91:1– 91:10, 2013. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/2772.

deBuyer:2015:DDE

[dBM15] Paul de Buyer and Jean-Christophe Mourrat. Diffusive decay of the environment viewed by the particle. Electronic Communications in Probability, 20(??):23:1–23:12, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3998.

Duminil-Copin:2013:LWC

[DC13] Hugo Duminil-Copin. Limit of the Wulff Crystal when approaching criticality for site percolation on the triangular lattice. Electronic Communications in Probability, 18:93:1–93:9, 2013. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3163.

Diop:2015:ASN

[DC15] Mamadou Abdoul Diop and Tomàs Caraballo. Asymptotic stability of neutral stochastic functional integro-differential equations with impulses. Electronic Communications in Probability, 20(??):1:1–1:13, ???? 2015. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/3036.

Dacunha-Castelle:2006:DLM

[DCF06] Didier Dacunha-Castelle and Lisandro Fermin. Disaggregation of long memory processes on C^{∞} class. Electronic Communications in Probability, 11:4:35–4:44, 2006. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1133.

Duminil-Copin:2013:CID

[DCLYY13] Hugo Duminil-Copin, Cyrille Lucas, Ariel Yadin, and Amir Yehudayoff. Containing internal diffusion limited aggregation. Electronic Communications in Probability, 18:50:1–50:8, 2013. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/2862.

Dirr:2010:LP

[DDG^{+}10] Nicolas Dirr, Patrick Dondl, Geoffrey Grimmett, Alexander Holroyd, and Michael Scheutzow. Lipschitz percolation. Electronic Communications in Probability, 15:2:14–2:21, 2010. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1521.

DaPrato:2007:MKP

[DDT07] Giuseppe Da Prato, Arnaud Debussche, and Luciano Tubaro. A modified Kardar–Parisi–Zhang model. Electronic Communications in Probability, 12:42:442–42:453, 2007. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/1333.

deLaFortelle:2006:SFL

[de 06] Arnaud de La Fortelle. Yule process sample path asymptotics. Electronic Communications in Probability, 11:20:193– 20:199, 2006. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1215.

Dubach:2023:DRO

[DE23] Guillaume Dubach and László Erdős. Dynamics of a rank-one perturbation of a Hermitian matrix. Electronic Communications in Probability, 28(??):1–13, 2023. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-28/issue-
none/Dynamics-of-a-rank-one-perturbation-of-a-Hermitian-
matrix/10.1214/23-ECP516.full.

DeBlassie:2007:CLL

[DeB07] Dante DeBlassie. The chance of a long lifetime for Brownian motion in a horn-shaped domain. Electronic Communications in Probability, 12:14:134–14:139, 2007. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1263.

Defosseux:2011:GLU

[Def11] Manon Defosseux. Generalized Laguerre unitary ensembles and an interacting particles model with a wall. Electronic Communications in Probability, 16:6:59–6:69, 2011. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/1602.

Defosseux:2012:IPM

[Def12] Manon Defosseux. Interacting particle models and the Pieri-type formulas : the symplectic case with non equal weights. Electronic Communications in Probability, 17:32:1–32:12, 2012. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2193.

Deijfen:2009:SRG

[Dei09] Maria Deijfen. Stationary random graphs with prescribed iid degrees on a spatial Poisson process. Electronic Communications in Probability, 14:8:81–8:89, 2009. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1448.

Delyon:2010:CIS

[Del10] Bernard Delyon. Concentration inequalities for the spectral measure of random matrices. Electronic Communications in Probability, 15:49:549–49:562, 2010. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1585.

Dembo:1996:MDM

[Dem96] Amir Dembo. Moderate deviations for martingales with bounded jumps. Electronic Communications in Probability, 1:3:11–3:17, 1996. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/973.

Demni:2011:KRV

[Dem11] Nizar Demni. Kanter random variable and positive free stable distributions. Electronic Communications in Probability, 16: 14:137–14:149, 2011. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1608.

Dembin:2020:AIP

[Dem20] Barbara Dembin. Anchored isoperimetric profile of the infinite cluster in supercritical bond percolation is Lipschitz continuous. Electronic Communications in Probability, 25(??):34:1– 34:13, ???? 2020. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1588125634.

Dereudre:2016:VPG

[Der16] David Dereudre. Variational principle for Gibbs point processes with finite range interaction. Electronic Communications in Probability, 21(??):10:1–10:11, ???? 2016. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1455560034.

Dawson:2016:LDH

[DF16] Donald A. Dawson and Shui Feng. Large deviations for homozygosity. Electronic Communications in Probability, 21(??): 83:1–83:8, ???? 2016. CODEN ???? ISSN 1083-589X. URL

http://projecteuclid.org/euclid.ecp/1481187621.

Dembo:2018:SEO

[DFK18] Amir Dembo, Ryoki Fukushima, and Naoki Kubota. Slowdown estimates for one-dimensional random walks in random environment with holding times. Electronic Communications in Probability, 23(??):89:1–89:12, ???? 2018. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1543028986.

Devroye:2000:PSQ

[DFN00] Luc Devroye, James Fill, and Ralph Neininger. Perfect simulation from the Quicksort limit distribution. Electronic Communications in Probability, 5:12:95–12:99, 2000. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/1024.

Dedecker:2015:SCI

[DG15] Jérôme Dedecker and Sébastien Gouëzel. Subgaussian concentration inequalities for geometrically ergodic Markov chains. Electronic Communications in Probability, 20(??):64:1–64:12, ???? 2015. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/3966.

Ding:2017:FPP

[DG17] Jian Ding and Subhajit Goswami. First passage percolation on the exponential of two-dimensional branching random walk. Electronic Communications in Probability, 22(??):69:1–69:14, ???? 2017. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1514451624.

Dionigi:2021:SSB

[DGdHM21] Pierfrancesco Dionigi, Diego Garlaschelli, Frank den Hollander, and Michel Mandjes. A spectral signature of breaking of ensemble equivalence for constrained random graphs. Electronic Communications in Probability, 26(??):1–15, 2021. CODEN ???? ISSN 1083-589X.

Devillers:2013:MVR

[DGG^{+}13] Olivier Devillers, Marc Glisse, Xavier Goaoc, Guillaume Moroz, and Matthias Reitzner. The monotonicity of f-vectors of random polytopes. Electronic Communications in Probability, 18:23:1– 23:8, 2013. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/2469.

Dyszewski:2020:LDM

[DGH20] Piotr Dyszewski, Nina Gantert, and Thomas Höfelsauer. Large deviations for the maximum of a branching random walk with stretched exponential tails. Electronic Communications in Probability, 25(??):1–13, 2020. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-25/issue-
none/Large-deviations-for-the-maximum-of-a-branching-
random-walk/10.1214/20-ECP353.full.

Damek:2019:ACM

[DGK19] Ewa Damek, Nina Gantert, and Konrad Kolesko. Absolute continuity of the martingale limit in branching processes in random environment. Electronic Communications in Probability, 24(??): 42:1–42:13, ???? 2019. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1562119370.

Dolinsky:2020:NCM

[DGO20] Yan Dolinsky, Benjamin Gottesman, and Gurel-Gurevich Ori. A note on costs minimization with stochastic target constraints. Electronic Communications in Probability, 25(??):11:1–11:12, ???? 2020. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1580353229.

Depperschmidt:2011:MMM

[DGP11] Andrej Depperschmidt, Andreas Greven, and Peter Pfaffelhuber. Marked metric measure spaces. Electronic Communications in Probability, 16:17:174–17:188, 2011. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1615.

Debussche:2011:AFE

[DHI11] Arnaud Debussche, Michael Hoegele, and Peter Imkeller. Asymptotic first exit times of the Chafee–Infante equation with small heavy-tailed Lévy noise. Electronic Communications in Probability, 16:21:213–21:225, 2011. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1622.

denHollander:2014:LDP

[dHP14] Frank den Hollander and Julien Poisat. Large deviation principles for words drawn from correlated letter sequences. Electronic Communications in Probability, 19:12:1–12:16, 2014. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2681.

DalBorgo:2018:ARS

[DHR18] Martina Dal Borgo, Emma Hovhannisyan, and Alain Rouault. Asymptotic results in solvable two-charge models. Electronic Communications in Probability, 23(??):16:1–16:12, ???? 2018. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1519722246.

Dembo:2014:MIW

[DHS14] Amir Dembo, Ruojun Huang, and Vladas Sidoravicius. Monotone interaction of walk and graph: recurrence versus transience. Electronic Communications in Probability, 19(??):76:1– 76:12, ???? 2014. CODEN ???? ISSN 1083-589X. URL http:/

/ecp.ejpecp.org/article/view/3607.

Dietert:2015:CGF

[Die15] Helge Dietert. Characterisation of gradient flows on finite state Markov chains. Electronic Communications in Probability, 20 (??):29:1–29:8, ???? 2015. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/3521.

Deijfen:2006:SRG

[DJ06] Maria Deijfen and Johan Jonasson. Stationary random graphs on Z with prescribed iid degrees and finite mean connections. Electronic Communications in Probability, 11:33:336–33:346, 2006. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1239.

Doney:2012:LDR

[DJ12] Ronald Doney and Elinor Jones. Large deviation results for random walks conditioned to stay positive. Electronic Communications in Probability, 17:38:1–38:11, 2012. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/2282.

Devroye:2014:PNF

[DJ14] Luc Devroye and Svante Janson. Protected nodes and fringe subtrees in some random trees. Electronic Communications in Probability, 19:6:1–6:10, 2014. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/3048.

Djilali:2018:OTL

[Dji18] Ait-Aoudia Djilali. Occupation time of Lévy processes with jumps rational Laplace transforms. Electronic Communications in Probability, 23(??):68:1–68:13, ???? 2018. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1539137386.

DeMarco:2016:TEN

[DJR16] Stefano De Marco, Antoine Jacquier, and Patrick Roome. Two examples of non strictly convex large deviations. Electronic Communications in Probability, 21(??):38:1–38:12, ???? 2016. CO-DEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1462368222.

Durrett:2020:CCE

[DJT20] Rick Durrett, Matthew Junge, and Si Tang. Coexistence in chase–escape. Electronic Communications in Probability, 25(??): 22:1–22:14, ???? 2020. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1583290992.

Dolgopyat:2012:SLR

[DK12] Dmitry Dolgopyat and Elena Kosygina. Scaling limits of recurrent excited random walks on integers. Electronic Communications in Probability, 17:35:1–35:14, 2012. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/2213.

Damek:2018:RTS

[DK18] Ewa Damek and Bartosz Kołodziejek. A renewal theorem and supremum of a perturbed random walk. Electronic Communications in Probability, 23(??):82:1–82:13, ???? 2018. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1540346606.

Dygert:2019:BPD

[DKJ^{+}19] Brittany Dygert, Christoph Kinzel, Matthew Junge, Annie Raymond, Erik Slivken, and Jennifer Zhu. The bullet problem with discrete speeds. Electronic Communications in Probability, 24 (??):27:1–27:11, ???? 2019. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1559700463.

Damron:2016:CFV

[DKNS16] Michael Damron, Hana Kogan, Charles M. Newman, and Vladas Sidoravicius. Coarsening with a frozen vertex. Electronic Communications in Probability, 21(??):9:1–9:4, ???? 2016. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
euclid.ecp/1455560033.

Dahmer:2014:KTL

[DKW14] Iulia Dahmer, Robert Knobloch, and Anton Wakolbinger. The Kingman tree length process has infinite quadratic variation. Electronic Communications in Probability, 19(??):87:1–87:12, ???? 2014. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/3318.

Dumbgen:2008:EBA

[DL08] Lutz Dümbgen and Christoph Leuenberger. Explicit bounds for the approximation error in Benford’s law. Electronic Communications in Probability, 13:10:99–10:112, 2008. CO-DEN ???? ISSN 1083-589X. URL http://arxiv.org/abs/

0705.4488;   http://ecp.ejpecp.org/article/view/1358;
 http://weber.math.washington.edu/~ejpecp/ECP/index.
php.

Dolgopyat:2009:NPA

[DL09a] Dmitry Dolgopyat and Carlangelo Liverani. Non-perturbative approach to random walk in Markovian environment. Electronic Communications in Probability, 14:24:245–24:251, 2009. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1467.

Duquesne:2009:RRI

[DL09b] Thomas Duquesne and Jean-François Le Gall. On the re-rooting invariance property of Lévy trees. Electronic Communications in Probability, 14:31:317–31:326, 2009. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1484.

Dembo:2018:LDP

[DL18] Amir Dembo and Eyal Lubetzky. A large deviation principle for the Erdős–Rényi uniform random graph. Electronic Communications in Probability, 23(??):79:1–79:13, ???? 2018. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1540346603.

delaPena:2009:EIS

[dlPP09] Victor de la Peña and Guodong Pang. Exponential inequalities for self-normalized processes with applications. Electronic Communications in Probability, 14:37:372–37:381, 2009. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1490.

Dieker:2009:RBM

[DM09] A. B. Dieker and J. Moriarty. Reflected Brownian motion in a wedge: sum-of-exponential stationary densities. Electronic Communications in Probability, 14:1:1–1:16, 2009. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1437.

Ding:2014:MUM

[DM14] Jian Ding and Elchanan Mossel. Mixing under monotone censoring. Electronic Communications in Probability, 19:45:1–45:6, 2014. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/3157.

Damek:2018:ACC

[DM18] Ewa Damek and Sebastian Mentemeier. Absolute continuity of complex martingales and of solutions to complex smoothing equations. Electronic Communications in Probability, 23(??): 60:1–60:12, ???? 2018. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1536718013.

Diel:2021:LCS

[DM21] Roland Diel and Dieter Mitsche. On the largest component of subcritical random hyperbolic graphs. Electronic Communications in Probability, 26(??):1–14, 2021. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-26/issue-
none/On-the-largest-component-of-subcritical-random-
hyperbolic-graphs/10.1214/21-ECP380.full.

Dominguez-Molina:2013:CRH

[DMPARA13] J. Armando Dominguez-Molina, Víctor Pérez-Abreu, and Alfonso Rocha-Arteaga. Covariation representations for Hermitian Lévy process ensembles of free infinitely divisible distributions. Electronic Communications in Probability, 18:6:1–6:14, 2013. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/2113.

Darses:2007:DPC

[DN07] Sébastien Darses and Ivan Nourdin. Dynamical properties and characterization of gradient drift diffusions. Electronic Communications in Probability, 12:37:390–37:400, 2007. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1324.

DOvidio:2010:ESF

[D’O10] Mirko D’Ovidio. Explicit solutions to fractional differential equations via generalized gamma convolution. Electronic Communications in Probability, 15:42:457–42:474, 2010. CODEN

????  ISSN 1083-589X. URL http://ecp.ejpecp.org/
article/view/1570.

Dohmen:2013:LBP

[Doh13] Klaus Dohmen. Lower bounds for the probability of a union via chordal graphs. Electronic Communications in Probability, 18: 70:1–70:4, 2013. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/2357.

Dokuchaev:2015:DFE

[Dok15] Nikolai Dokuchaev. On the dependence of the first exit times on the fluctuations of the domain boundary. Electronic Communications in Probability, 20(??):80:1–80:3, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/4531.

Dolinsky:2014:HGO

[Dol14] Yan Dolinsky. Hedging of game options under model uncertainty in discrete time. Electronic Communications in Probability, 19: 19:1–19:11, 2014. CODEN ???? ISSN 1083-589X. URL http:/

/ecp.ejpecp.org/article/view/2714.

Doney:2020:RRT

[Don20] Ron Doney. The remainder in the renewal theorem. Electronic Communications in Probability, 25(??):5:1–5:8, ???? 2020. CO-DEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1580180423. See erratum [Don22].

Doney:2022:ERR

[Don22] Ron Doney. Erratum: The remainder in the renewal theorem. Electronic Communications in Probability, 27(??):1– 5, 2022. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-27/issue-none/Erratum-The-remainder-
in-the-renewal-theorem/10.1214/22-ECP456.full.  See

[Don20].

DOvidio:2016:SDR

[DOS16] Mirko D’Ovidio, Enzo Orsingher, and Ludmila Sakhno. Spectral densities related to some fractional stochastic differential equations. Electronic Communications in Probability, 21(??):18:1– 18:15, ???? 2016. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1456412898.

Ding:2013:SMT

[DP13] Jian Ding and Yuval Peres. Sensitivity of mixing times. Electronic Communications in Probability, 18:88:1–88:6, 2013. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2765.

Dobler:2014:RFM

[DP14] Christian Döbler and Lorenz Pfeifroth. Recurrence for the frog model with drift on Z^{d}. Electronic Communications in Probability, 19(??):79:1–79:13, ???? 2014. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/3740.

Dobler:2018:FMT

[DP18a] Christian Döbler and Giovanni Peccati. Fourth moment theorems on the Poisson space: analytic statements via product formulae. Electronic Communications in Probability, 23(??):91:1– 91:12, ???? 2018. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1544843114.

Drame:2018:AGC

[DP18b] Ibrahima Dramé and Étienne Pardoux. Approximation of a generalized continuous-state branching process with interaction. Electronic Communications in Probability, 23(??):73:1–73:14, ???? 2018. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1539763346.

Depperschmidt:2015:SLD

[DPS15] Andrej Depperschmidt, Peter Pfaffelhuber, and Annika Scheuringer. Some large deviations in Kingman’s coalescent. Electronic Communications in Probability, 20(??):7:1–7:14, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3107.

Delattre:2012:TFS

[DR12] Sylvain Delattre and Mathieu Rosenbaum. Testing the finiteness of the support of a distribution: a statistical look at Tsirelson’s equation. Electronic Communications in Probability, 17:25:1– 25:7, 2012. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1834.

Deijfen:2020:ISF

[DR20] Maria Deijfen and Sebastian Rosengrenï. The initial set in the frog model is irrelevant. Electronic Communications in Probability, 25(??):1–7, 2020. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-25/issue-
none/The-initial-set-in-the-frog-model-is-irrelevant/
10.1214/20-ECP329.full.

Dembo:2006:LMD

[DS06] Amir Dembo and Qi-Man Shao. Large and moderate deviations for Hotelling’s T^{2}-statistics. Electronic Communications in Probability, 11:15:149–15:159, 2006. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1209.

Doring:2010:ART

[DS10] Leif Döring and Mladen Savov. An application of renewal theorems to exponential moments of local times. Electronic Communications in Probability, 15:24:263–24:269, 2010. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1558.

Duy:2015:MSM

[DS15] Khanh Trinh Duy and Tomoyuki Shirai. The mean spectral measures of random Jacobi matrices related to Gaussian beta ensembles. Electronic Communications in Probability, 20(??): 68:1–68:13, ???? 2015. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/4252.

Deng:2016:CBF

[DS16] Chang-Song Deng and René L. Schilling. Complete Bernstein functions and subordinators with nested ranges. A note on a paper by P. Marchal. Electronic Communications in Probability, 21(??):78:1–78:5, ???? 2016. CODEN ???? ISSN 1083-589X.

URL http://projecteuclid.org/euclid.ecp/1480388671.

Duong:2018:VFP

[DT18] Manh Hong Duong and Julian Tugaut. The Vlasov–Fokker– Planck equation in non-convex landscapes: convergence to equilibrium. Electronic Communications in Probability, 23(??):19:1– 19:10, ???? 2018. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1521079417.

Dubedat:2003:ST

[Dub03] Julien Dubédat. SLE and triangles. Electronic Communications in Probability, 8:4:28–4:42, 2003. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1067.

Dubach:2023:EFC

[Dub23] Guillaume Dubach. Explicit formulas concerning eigenvectors of weakly non-unitary matrices. Electronic Communications in Probability, 28(??):1–11, 2023. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-28/issue-
none/Explicit-formulas-concerning-eigenvectors-of-weakly-
non-unitary-matrices/10.1214/22-ECP507.full.

Ducatez:2019:FBR

[Duc19] Raphael Ducatez. A forward-backward random process for the spectrum of 1D Anderson operators. Electronic Communications in Probability, 24(??):69:1–69:13, ???? 2019. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1573095648.

Duerre:2006:UMD

[Due06] Maximilian Duerre. Uniqueness of multi-dimensional infinite volume self-organized critical forest-fire models. Electronic Communications in Probability, 11:31:304–31:315, 2006. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1229.

Dunlap:2020:CPA

[Dun20] Alexander Dunlap. The continuum parabolic Anderson model with a half-laplacian and periodic noise. Electronic Communications in Probability, 25(??):1–14, 2020. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-25/issue-
none/The-continuum-parabolic-Anderson-model-with-a-half-
Laplacian-and/10.1214/20-ECP342.full.

Duquesne:2009:EPH

[Duq09] Thomas Duquesne. An elementary proof of Hawkes’s conjecture on Galton–Watson trees. Electronic Communications in Probability, 14:15:151–15:164, 2009. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1454.

Duraj:2014:HFK

[Dur14] Jetlir Duraj. On harmonic functions of killed random walks in convex cones. Electronic Communications in Probability, 19(??): 80:1–80:10, ???? 2014. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/3219.

Dallaporta:2011:NCL

[DV11] Sandrine Dallaporta and Van Vu. A note on the Central Limit Theorem for the eigenvalue counting function of Wigner matrices. Electronic Communications in Probability, 16:30:214–30:322, 2011. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1634.

Deijfen:2023:CDB

[DV23] Maria Deijfen and Timo Vilkas. Competition on Z^{d} driven by branching random walk. Electronic Communications in Probability, 28(??):1–11, 2023. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-28/issue-
none/Competition-on-Zd-driven-by-branching-random-walk/
10.1214/23-ECP521.full.

Denisov:2012:MAS

[DW12] Denis Denisov and Vitali Wachtel. Martingale approach to subexponential asymptotics for random walks. Electronic Communications in Probability, 17:6:1–6:9, 2012. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/1757.

Dey:2015:NMC

[DW15] Partha S Dey and Edward C Waymire. On normalized multiplicative cascades under strong disorder. Electronic Communications in Probability, 20(??):32:1–32:13, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3936.

Denisov:2016:ULT

[DW16] Denis Denisov and Vitali Wachtel. Universality of local times of killed and reflected random walks. Electronic Communications in Probability, 21(??):1:1–1:11, ???? 2016. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1454514621.

Dong:2011:IMS

[DXZ11] Zhao Dong, Lihu Xu, and Xicheng Zhang. Invariant measures of stochastic 2D Navier–Stokes equation driven by α-stable processes. Electronic Communications in Probability, 16:59:678– 59:688, 2011. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1664.

Dadoun:2021:MCM

[DY21] Benjamin Dadoun and Pierre Youssef. Maximal correlation and monotonicity of free entropy and of Stein discrepancy. Electronic Communications in Probability, 26(??):1–10, 2021. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-26/issue-none/Maximal-correlation-
and-monotonicity-of-free-entropy-and-of-Stein/10.1214/
21-ECP391.full.

Dembo:1996:TAS

[DZ96] Amir Dembo and Ofer Zeitouni. Transportation approach to some concentration inequalities in product spaces. Electronic Communications in Probability, 1:9:83–9:90, 1996. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/979.

Dassios:2013:ESH

[DZ13] Angelos Dassios and Hongbiao Zhao. Exact simulation of Hawkes process with exponentially decaying intensity. Electronic Communications in Probability, 18:62:1–62:13, 2013. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2717.

Deng:2019:CGF

[DZ19] Xiaoxue Deng and Ping Zhong. Continuity and growth of free multiplicative convolution semigroups. Electronic Communications in Probability, 24(??):4:1–4:11, ???? 2019. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1548817631.

Dzindzalieta:2013:ELF

[Dzi13] Dainius Dzindzalieta. Extremal Lipschitz functions in the deviation inequalities from the mean. Electronic Communications in Probability, 18:66:1–66:5, 2013. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/2814.

Easo:2022:WAG

[Eas22] Philip Easo. The wired arboreal gas on regular trees. Electronic Communications in Probability, 27(??):1–10, 2022. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
journals/electronic-communications-in-probability/volume-
27/issue-none/The-wired-arboreal-gas-on-regular-trees/
10.1214/22-ECP460.full.

Evans:2015:WDS

[EHW15] Steven N. Evans, Alexandru Hening, and Eric S. Wayman. When do skew-products exist? Electronic Communications in Probability, 20(??):54:1–54:14, ???? 2015. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/4040.

Eisenbaum:2021:SPI

[Eis21] Nathalie Eisenbaum. Some percolations involving the Gaussian free fields. Electronic Communications in Probability, 26(??): 1–8, 2021. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-26/issue-none/Some-percolations-
involving-the-Gaussian-free-fields/10.1214/21-ECP379.
full.

Ejsmont:2012:LLP

[Ejs12] Wiktor Ejsmont. Laha–Lukacs properties of some free processes. Electronic Communications in Probability, 17:13:1–13:8, 2012. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1865.

Ejsmont:2013:NCF

[Ejs13] Wiktor Ejsmont. Noncommutative characterization of free Meixner processes. Electronic Communications in Probability, 18:22:1–22:12, 2013. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/2356.

Eisenbaum:2008:PGG

[EK08] Nathalie Eisenbaum and Andreas Kyprianou. On the parabolic generator of a general one-dimensional Lévy process. Electronic Communications in Probability, 13:20:198–20:209, 2008. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1366.

Ernst:2022:LFN

[EKR22] Philip A. Ernst, Abram M. Kagan, and L. C. G. Rogers. The least favorable noise. Electronic Communications in Probability, 27(??):1–11, 2022. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-27/issue-
none/The-least-favorable-noise/10.1214/22-ECP467.full.

Etore:2014:DAP

[EM14] Pierre Etoré and Ester Mariucci. L_{1}-distance for additive processes with time-homogeneous Lévy measures. Electronic Communications in Probability, 19:57:1–57:10, 2014. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/3678.

Enriquez:2016:AEE

[EM16] Nathanaël Enriquez and Laurent Ménard. Asymptotic expansion of the expected spectral measure of Wigner matrices. Electronic Communications in Probability, 21(??):58:1–58:11, ???? 2016. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1473186615.

Essifi:2021:UEI

[EM21] Rim Essifi and Sami Mustapha. Upper estimates for inhomogeneous random walks confined to the positive orthant. Electronic Communications in Probability, 26(??):1–14, 2021. CODEN ???? ISSN 1083-589X.

Erbar:2015:LDW

[EMR15] Matthias Erbar, Jan Maas, and Michiel Renger. From large deviations to Wasserstein gradient flows in multiple dimensions. Electronic Communications in Probability, 20(??):89:1–89:12, ???? 2015. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/4315.

Emrah:2016:LSI

[Emr16] Elnur Emrah. Limit shapes for inhomogeneous corner growth models with exponential and geometric weights. Electronic Communications in Probability, 21(??):42:1–42:16, ???? 2016. CO-DEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1463683990.

Enriquez:2020:SCR

[EN20] Nathanaël Enriquez and Nathan Noiry. A solvable class of renewal processes. Electronic Communications in Probability, 25 (??):1–14, 2020. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-25/issue-none/A-solvable-class-
of-renewal-processes/10.1214/20-ECP348.full.

Evans:1998:EIS

[EP98] Steven N. Evans and Yuval Peres. Eventual intersection for sequences of Lévy processes. Electronic Communications in Probability, 3:3:21–3:27, 1998. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/989.

Englander:2017:SAB

[EP17] János Engländer and Yuval Peres. Survival asymptotics for branching random walks in IID environments. Electronic Communications in Probability, 22(??):29:1–29:12, ???? 2017. CO-DEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1495677741.

Eckhoff:2009:UMM

[ER09] Maren Eckhoff and Silke Rolles. Uniqueness of the mixing measure for a random walk in a random environment on the positive integers. Electronic Communications in Probability, 14:3:31– 3:35, 2009. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1441.

Evans:2019:RRS

[ER19] Steven N. Evans and Daniel Raban. Rotatable random sequences in local fields. Electronic Communications in Probability, 24(??): 37:1–37:12, ???? 2019. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1561169053.

Erickson:2016:DTI

[Eri16] K. Bruce Erickson. Deterministic time intervals on which a class of persistent processes are away from their origins. Electronic Communications in Probability, 21(??):62:1–62:12, ???? 2016. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1473685926.

Erny:2021:CCS

[Ern21] Xavier Erny. A convergence criterion for systems of point processes from the convergence of their stochastic intensities. Electronic Communications in Probability, 26(??):1–10, 2021. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-26/issue-none/A-convergence-criterion-
for-systems-of-point-processes-from-the/10.1214/21-ECP372.
full.

Eichelsbacher:2009:MDT

[ES09] Peter Eichelsbacher and Jens Sommerauer. Moderate deviations for traces of words in a mult-matrix model. Electronic Communications in Probability, 14:56:572–56:586, 2009. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1515.

Erdos:2016:FFW

[ES16] László Erdős and Dominik Schröder. Fluctuations of functions of Wigner matrices. Electronic Communications in Probability, 21(??):86:1–86:15, ???? 2016. CODEN ???? ISSN 1083-589X.

URL http://projecteuclid.org/euclid.ecp/1483347665.

Ertul:2021:SDC

[ES21] Anatole Ertul and Assaf Shapira. Self-diffusion coefficient in the Kob–Andersen model. Electronic Communications in Probability, 26(??):1–12, 2021. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-26/issue-

none/Self-diffusion-coefficient-in-the-Kob-Andersen-
model/10.1214/20-ECP370.full.

Es-Sarhir:2009:HIF

[ESvRS09] Abdelhadi Es-Sarhir, Max-K. von Renesse, and Michael Scheutzow. Harnack inequality for functional SDEs with bounded memory. Electronic Communications in Probability, 14:54:560–54:565, 2009. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1513.

Enriquez:2008:RSS

[ESY08] Nathanael Enriquez, Christophe Sabot, and Marc Yor. Renewal series and square-root boundaries for Bessel processes. Electronic Communications in Probability, 13:59:649–59:652, 2008. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1436.

Ethier:2014:PPD

[Eth14] Stewart N. Ethier. A property of Petrov’s diffusion. Electronic Communications in Probability, 19(??):65:1–65:4, ???? 2014. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/3684.

Evans:2006:ENZ

[Eva06] Steven Evans. The expected number of zeros of a random system of p-adic polynomials. Electronic Communications in Probability, 11:29:278–29:290, 2006. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1230.

Eisenbaum:2009:OIF

[EW09] Nathalie Eisenbaum and Alexander Walsh. An optimal Itô formula for Lévy processes. Electronic Communications in Probability, 14:20:202–20:209, 2009. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1469.

Evans:2017:RST

[EW17] Steven N. Evans and Anton Wakolbinger. Radix sort trees in the large. Electronic Communications in Probability, 22(??):68:1– 68:13, ???? 2017. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1512097482.

Evans:1999:IES

[EZ99] Steven N. Evans and Xiaowen Zhou. Identifiability of exchangeable sequences with identically distributed partial sums. Electronic Communications in Probability, 4:2:9–2:13, 1999. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1000.

Fang:2015:UEB

[Fan15] Xiao Fang. A universal error bound in the CLT for counting monochromatic edges in uniformly colored graphs. Electronic Communications in Probability, 20(??):21:1–21:6, ???? 2015. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/3707.

Fan:2016:DAL

[Fan16] Wai-Tong Louis Fan. Discrete approximations to local times for reflected diffusions. Electronic Communications in Probability, 21(??):16:1–16:12, ???? 2016. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1456238572.

Fargason:1998:PDB

[Far98] Chad Fargason. Percolation dimension of Brownian motion in R^{3}. Electronic Communications in Probability, 3:7:51–7:63, 1998. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/993.

Fathi:2018:SSF

[Fat18] Max Fathi. A sharp symmetrized form of Talagrand’s transportentropy inequality for the Gaussian measure. Electronic Communications in Probability, 23(??):81:1–81:9, ???? 2018. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
euclid.ecp/1540346605.

Friedli:2004:LRP

[FdLS04] Sacha Friedli, Benoîte Borge de Lima, and Vladas Sidoravicius. On long range percolation with heavy tails. Electronic Communications in Probability, 9:19:175–19:177, 2004. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1122.

Feyel:2007:NCS

[FdM07] Denis Feyel, Arnaud de La Pradelle, and Gabriel Mokobodzki. A non-commutative sewing lemma. Electronic Communications in Probability, 13:3:24–3:34, 2007. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1345.

Feng:2023:HDP

[Fen23] Shui Feng. Hierarchical Dirichlet process and relative entropy. Electronic Communications in Probability, 28(??):1– 12, 2023. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-28/issue-none/Hierarchical-Dirichlet-
process-and-relative-entropy/10.1214/23-ECP511.full.

Feray:2020:CLT

[Fér20] Valentin Féray. On the central limit theorem for the two-sided descent statistics in Coxeter groups. Electronic Communications in Probability, 25(??):28:1–28:6, ???? 2020. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1585620028.

Fittipaldi:2012:SAC

[FF12] Maria Fittipaldi and Joaquin Fontbona T. On SDE associated with continuous-state branching processes conditioned to never be extinct. Electronic Communications in Probability, 17:49:1– 49:13, 2012. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1972.

Friedland:2013:SOR

[FG13] Omer Friedland and Ohad Giladi. A simple observation on random matrices with continuous diagonal entries. Electronic Communications in Probability, 18:53:1–53:7, 2013. CODEN

????  ISSN 1083-589X. URL http://ecp.ejpecp.org/
article/view/2633.

Feldheim:2021:PTB

[FGG21] Ohad N. Feldheim and Ori Gurel-Gurevich. The power of thinning in balanced allocation. Electronic Communications in Probability, 26(??):1–8, 2021. CODEN ???? ISSN 1083-589X.

Fan:2012:LDE

[FGL12] Xiequan Fan, Ion Grama, and Quansheng Liu. Large deviation exponential inequalities for supermartingales. Electronic Communications in Probability, 17:59:1–59:8, 2012. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2318.

Fontbona:2010:MOT

[FGM10] Joaquin Fontbona, Hélène Guérin, and Sylvie Méléard. Measurability of optimal transportation and strong coupling of martingale measures. Electronic Communications in Probability, 15: 13:124–13:133, 2010. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1534.

Fernandez:2011:RMA

[FGM11] Roberto Fernandez, Sandro Gallo, and Gregory Maillard. Regular g-measures are not always Gibbsian. Electronic Communications in Probability, 16:64:732–64:740, 2011. CODEN

????  ISSN 1083-589X. URL http://ecp.ejpecp.org/
article/view/1681.

Fill:2019:TLQ

[FH19] James Allen Fill and Wei-Chun Hung. On the tails of the limiting QuickSort density. Electronic Communications in Probability, 24 (??):7:1–7:11, ???? 2019. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1550113298.

Favero:2022:ABS

[FH22] Martina Favero and Henrik Hult. Asymptotic behaviour of sampling and transition probabilities in coalescent models under selection and parent dependent mutations. Electronic Communications in Probability, 27(??):1–13, 2022. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-27/issue-
none/Asymptotic-behaviour-of-sampling-and-transition-
probabilities-in-coalescent-models/10.1214/22-ECP472.
full.

Foxall:2018:CRW

[FHJ18] Eric Foxall, Tom Hutchcroft, and Matthew Junge. Coalescing random walk on unimodular graphs. Electronic Communications in Probability, 23(??):62:1–62:10, ???? 2018. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1536804170.

Fukasawa:2021:NDS

[FI21] Masaaki Fukasawa and Mitsumasa Ikeda. A new discretization scheme for one dimensional stochastic differential equations using time change method. Electronic Communications in Probability, 26(??):1–12, 2021. CODEN ???? ISSN 1083-589X.

Fitzsimmons:2000:SFM

[Fit00] P. Fitzsimmons. Strict fine maxima. Electronic Communications in Probability, 5:11:91–11:94, 2000. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1023.

Fitzsimmons:2006:ERE

[Fit06] Patrick Fitzsimmons. On the existence of recurrent extensions of self-similar Markov processes. Electronic Communications in Probability, 11:25:230–25:241, 2006. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1222.

Fill:2000:CSF

[FJ00] James Fill and Svante Janson. A characterization of the set of fixed points of the Quicksort transformation. Electronic Communications in Probability, 5:9:77–9:84, 2000. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1021.

Fukushima:2021:LDR

[FJ21] Ryoki Fukushima and Stefan Junk. On large deviation rate functions for a continuous-time directed polymer in weak disorder. Electronic Communications in Probability, 26(??):1– 10, 2021. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-26/issue-none/On-large-deviation-
rate-functions-for-a-continuous-time-directed/10.1214/
21-ECP378.full.

Friesen:2020:BBM

[FJR20] Martin Friesen, Peng Jin, and Barbara Rüdiger. On the boundary behavior of multi-type continuous-state branching processes with immigration. Electronic Communications in Probability, 25 (??):1–14, 2020. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-25/issue-none/On-the-boundary-
behavior-of-multi-type-continuous-state-branching/10.
1214/20-ECP364.full.

Fodor:2023:RDP

[FKV23] Ferenc Fodor, Péter Kevei, and Viktor Vígh. On random disc-polygons in a disc-polygon. Electronic Communications in Probability, 28(??):1–11, 2023. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-28/issue-
none/On-random-disc-polygons-in-a-disc-polygon/10.1214/
23-ECP515.full.

Fuchs:2020:DAE

[FLP20] Michael Fuchs, Chih-Hong Lee, and Ariel R. Paningbatan. Distributional analysis of the extra-clustering model with uniformly generated phylogenetic trees. Electronic Communications in Probability, 25(??):13:1–13:13, ???? 2020. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1581130925.

Flury:2008:NBL

[Flu08] Markus Flury. A note on the ballistic limit of random motion in a random potential. Electronic Communications in Probability, 13:38:393–38:400, 2008. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1394.

Frikha:2012:CBS

[FM12a] Noufel Frikha and Stéphane Menozzi. Concentration bounds for stochastic approximations. Electronic Communications in Probability, 17:47:1–47:15, 2012. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1952.

Frikha:2012:ECB

[FM12b] Noufel Frikha and Stéphane Menozzi. Erratum: Concentration bounds for stochastic approximation. Electronic Communications in Probability, 17:60:1–60:2, 2012. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/2495.

Fukasawa:2021:RCM

[FM21] Masaaki Fukasawa and Kazuki Matsushita. Realized cumulants for martingales. Electronic Communications in Probability, 26 (??):1–10, 2021. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-26/issue-none/Realized-cumulants-
for-martingales/10.1214/21-ECP382.full.

Fotsa–Mbogne:2017:NFD

[FMP17] David Jaures Fotsa-Mbogne and Etienne Pardoux. Nonlinear filtering with degenerate noise. Electronic Communications in Probability, 22(??):44:1–44:14, ???? 2017. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1502762750.

Forsstrom:2020:TST

[For20] Palö Forsström Forsström. A tame sequence of transitive Boolean functions. Electronic Communications in Probability, 25(??): 1–8, 2020. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-25/issue-none/A-tame-sequence-
of-transitive-Boolean-functions/10.1214/20-ECP366.full.

Forsstrom:2021:WSB

[For21] Malin P. Forsström. When are sequences of Boolean functions tame? Electronic Communications in Probability, 26(??):1–13, 2021. CODEN ???? ISSN 1083-589X.

Foucart:2013:ISW

[Fou13] Clément Foucart. The impact of selection in the Λ–Wright– Fisher model. Electronic Communications in Probability, 18: 72:1–72:10, 2013. CODEN ???? ISSN 1083-589X. URL http:/

/ecp.ejpecp.org/article/view/2838. See erratum [Fou14].

Foucart:2014:EIS

[Fou14] Clément Foucart. Erratum: The impact of selection in the Λ– Wright–Fisher Model. Electronic Communications in Probability, 19:15:1–15:3, 2014. CODEN ???? ISSN 1083-589X. URL http:

//ecp.ejpecp.org/article/view/3351. See [Fou13].

Fournier:2011:SSH

[FP11] Nicolas Fournier and Jacques Printems. Stability of the stochastic heat equation in L^{1}([0, 1]). Electronic Communications in Probability, 16:32:337–32:352, 2011. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1636.

Forman:2020:MSI

[FPRW20] Noah Forman, Soumik Pal, Douglas Rizzolo, and Matthias Winkel. Metrics on sets of interval partitions with diversity. Electronic Communications in Probability, 25(??):38:1–38:16, ???? 2020. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1591668057.

Fuhrman:2016:RNM

[FPZ16] Marco Fuhrman, Huyên Pham, and Federica Zeni. Representation of non-Markovian optimal stopping problems by constrained BSDEs with a single jump. Electronic Communications in Probability, 21(??):3:1–3:7, ???? 2016. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1454514623.

Fradon:2013:SDE

[Fra13] Myriam Fradon. Stochastic differential equations on domains defined by multiple constraints. Electronic Communications in Probability, 18:26:1–26:13, 2013. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/2730.

Freund:2012:ASA

[Fre12] Fabian Freund. Almost sure asymptotics for the number of types for simple Ξ-coalescents. Electronic Communications in Probability, 17:3:1–3:11, 2012. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1704.

Fontes:2019:MRW

[FS19] Luiz Renato Fontes and Rinaldo B. Schinazi. Metastability of a random walk with catastrophes. Electronic Communications in Probability, 24(??):70:1–70:8, ???? 2019. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1573268934.

Fierro:2007:SSA

[FT07] Raul Fierro and Soledad Torres. A stochastic scheme of approximation for ordinary differential equations. Electronic Communications in Probability, 13:1:1–1:9, 2007. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/1341.

Fukushima:2009:LTQ

[Fuk09] Ryoki Fukushima. From the Lifshitz tail to the quenched survival asymptotics in the trapping problem. Electronic Communications in Probability, 14:42:435–42:446, 2009. CODEN

????  ISSN 1083-589X. URL http://ecp.ejpecp.org/
article/view/1497.

Funaki:2007:DSL

[Fun07] Tadahisa Funaki. Dichotomy in a scaling limit under Wiener measure with density. Electronic Communications in Probability, 12:18:173–18:183, 2007. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1271.

Fresen:2014:CRP

[FV14] Daniel Fresen and Richard Vitale. Concentration of random polytopes around the expected convex hull. Electronic Communications in Probability, 19:59:1–59:8, 2014. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/3376.

Ferrari:2021:UTD

[FV21] Patrik L. Ferrari and Bálint Vet. Upper tail decay of KPZ models with Brownian initial conditions. Electronic Communications in Probability, 26(??):1–14, 2021. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-26/issue-
none/Upper-tail-decay-of-KPZ-models-with-Brownian-initial-
conditions/10.1214/21-ECP385.full.

Federico:2016:CTR

[FvdHH16] Lorenzo Federico, Remco van der Hofstad, and Tim Hulshof. Connectivity threshold for random subgraphs of the Hamming graph. Electronic Communications in Probability, 21(??):27:1– 27:8, ???? 2016. CODEN ???? ISSN 1083-589X. URL https:/

/projecteuclid.org/euclid.ecp/1457978024.

Fagnola:2000:MSQ

[FW00] Franco Fagnola and Stephen Wills. Mild solutions of quantum stochastic differential equations. Electronic Communications in Probability, 5:17:158–17:171, 2000. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1029.

Freij:2010:POS

[FW10] Ragnar Freij and Johan Wästlund. Partially ordered secretaries. Electronic Communications in Probability, 15:45:504– 45:507, 2010. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1579.

Fang:2017:RSG

[FW17] Shizan Fang and Bo Wu. Remarks on spectral gaps on the Riemannian path space. Electronic Communications in Probability, 22(??):19:1–19:13, ???? 2017. CODEN ???? ISSN 1083-589X.

URL http://projecteuclid.org/euclid.ecp/1489457081.

Feng:2002:LDQ

[FX02] Shui Feng and Jie Xiong. Large deviations and quasi-potential of a Fleming–Viot process. Electronic Communications in Probability, 7:2:13–2:25, 2002. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1043.

Feng:2018:CRS

[FXA18] Renjie Feng, Xingcheng Xu, and Robert J. Adler. Critical radius and supremum of random spherical harmonics (II). Electronic Communications in Probability, 23(??):50:1–50:11, ???? 2018. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1535767261.

Fuchs:2015:RSD

[FY15] Michael Fuchs and Pei-Duo Yu. Rumor source detection for rumor spreading on random increasing trees. Electronic Communications in Probability, 20(??):2:1–2:12, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3743.

Fang:2010:CMD

[FZ10] Ming Fang and Ofer Zeitouni. Consistent minimal displacement of branching random walks. Electronic Communications in Probability, 15:11:106–11:118, 2010. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1533.

Gaal:2014:LRO

[Gaá14] Alexisz Gaál. Long-range order in a hard disk model in statistical mechanics. Electronic Communications in Probability, 19:9:1– 9:9, 2014. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/3047.

Galicza:2020:PVN

[Gal20] Pál Galicza. Pivotality versus noise stability for monotone transitive functions. Electronic Communications in Probability, 25 (??):17:1–17:6, ???? 2020. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1582167853.

Ganesan:2014:DCP

[Gan14] Ghurumuruhan Ganesan. Disjoint crossings, positive speed and deviation estimates for first passage percolation. Electronic Communications in Probability, 19:52:1–52:8, 2014. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3490.

Ganesan:2021:SWM

[Gan21] Ghurumuruhan Ganesan. Strong and weighted matchings in inhomogeneous random graphs. Electronic Communications in Probability, 26(??):1–12, 2021. CODEN ???? ISSN 1083-589X.

Gao:2003:MML

[Gao03] Fuchang Gao. The mean of a maximum likelihood estimator associated with the Brownian bridge. Electronic Communications in Probability, 8:1:1–1:5, 2003. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1064.

Gao:2008:EEM

[Gao08] Fuchang Gao. Entropy estimate for k-monotone functions via small ball probability of integrated Brownian motions. Electronic Communications in Probability, 13:12:121–12:130, 2008. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1357.

Gassiat:2019:MPR

[Gas19] Paul Gassiat. On the martingale property in the rough Bergomi model. Electronic Communications in Probability, 24(??):33:1– 33:9, ???? 2019. CODEN ???? ISSN 1083-589X. URL https:/

/projecteuclid.org/euclid.ecp/1560477645.

Gauthier:2016:SAD

[Gau16] Carl-Erik Gauthier. Self attracting diffusions on a sphere and application to a periodic case. Electronic Communications in Probability, 21(??):53:1–53:12, ???? 2016. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1470081068.

Georgiou:2010:SER

[Geo10] Nicos Georgiou. Soft edge results for longest increasing paths on the planar lattice. Electronic Communications in Probability, 15: 1:1–1:13, 2010. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1519.

Gine:2004:SNC

[GG04] Evarist Giné and Friedrich Götze. On standard normal convergence of the multivariate Student t-statistic for symmetric random vectors. Electronic Communications in Probability, 9: 17:162–17:171, 2004. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1120.

Ghosh:2011:ASB

[GG11] Subhankar Ghosh and Larry Goldstein. Applications of size biased couplings for concentration of measures. Electronic Communications in Probability, 16:7:70–7:83, 2011. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/1605.

Geiss:2014:FSF

[GG14] Stefan Geiss and Emmanuel Gobet. Fractional smoothness of functionals of diffusion processes under a change of measure. Electronic Communications in Probability, 19:33:1–33:14, 2014. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2786.

Gerhold:2020:LDR

[GG20] Stefan Gerhold and Christoph Gerstenecker. Large deviations related to the law of the iterated logarithm for Itô diffusions. Electronic Communications in Probability, 25(??):16:1–16:11, ???? 2020. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1581995086.

Gurel-Gurevich:2010:FAR

[GGA10] Ori Gurel-Gurevich and Gideon Amir. On fixation of activated random walks. Electronic Communications in Probability, 15: 12:119–12:123, 2010. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1536.

Gurel-Gurevich:2017:RME

[GGNS17] Ori Gurel-Gurevich, Asaf Nachmias, and Juan Souto. Recurrence of multiply-ended planar triangulations. Electronic Communications in Probability, 22(??):5:1–5:6, ???? 2017. CODEN ????

ISSN 1083-589X. URL http://projecteuclid.org/euclid.
ecp/1483671681.

Gurel-Gurevich:2014:LCR

[GGPZ14] Ori Gurel-Gurevich, Yuval Peres, and Ofer Zeitouni. Localization for controlled random walks and martingales. Electronic Communications in Probability, 19:23:1–23:8, 2014. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3081.

Gantert:2018:LDM

[GH18a] Nina Gantert and Thomas Höfelsauer. Large deviations for the maximum of a branching random walk. Electronic Communications in Probability, 23(??):34:1–34:12, ???? 2018. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1528358639.

Gu:2018:CEP

[GH18b] Yu Gu and Jingyu Huang. Chaos expansion of 2D parabolic Anderson model. Electronic Communications in Probability, 23 (??):26:1–26:10, ???? 2018. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1524881134.

Griffiths:2020:UAM

[GH20] Robert Griffiths and Kais Hamza. A universal approach to matching marginals and sums. Electronic Communications in Probability, 25(??):1–12, 2020. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-25/issue-
none/A-universal-approach-to-matching-marginals-and-
sums/10.1214/20-ECP357.full.

Georgakopoulos:2021:TIR

[GH21] Agelos Georgakopoulos and John Haslegrave. A time-invariant random graph with splitting events. Electronic Communications in Probability, 26(??):1–15, 2021. CODEN ???? ISSN 1083-589X.

Gulisashvili:2016:PHB

[GHJ16] Archil Gulisashvili, Blanka Horvath, and Antoine Jacquier. On the probability of hitting the boundary for Brownian motions on the SABR plane. Electronic Communications in Probability, 21(??):75:1–75:13, ???? 2016. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1477600776.

Gao:2017:PST

[GHJL17] Qiang Gao, Mingshang Hu, Xiaojun Ji, and Guomin Liu. Product space for two processes with independent increments under nonlinear expectations. Electronic Communications in Probability, 22(??):11:1–11:12, ???? 2017. CODEN ???? ISSN

1083-589X. URL http://projecteuclid.org/euclid.ecp/
1485421234.

Ghosh:2016:PMR

[Gho16] Subhroshekhar Ghosh. Palm measures and rigidity phenomena in point processes. Electronic Communications in Probability, 21(??):85:1–85:14, ???? 2016. CODEN ???? ISSN 1083-589X.

URL http://projecteuclid.org/euclid.ecp/1481770936.

Ghosh:2022:LDR

[Gho22] Partha Pratim Ghosh. Large deviations for the rightmost position of a last progeny modified branching random walk. Electronic Communications in Probability, 27(??):1–13, 2022. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-27/issue-none/Large-deviations-
for-the-right-most-position-of-a-last/10.1214/22-ECP446.
full.

Garban:2018:NMG

[GHSS18] Christophe Garban, Nina Holden, Avelio Sepúlveda, and Xin Sun. Negative moments for Gaussian multiplicative chaos on fractal sets. Electronic Communications in Probability, 23(??): 100:1–100:10, ???? 2018. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1545188960.

Gufler:2022:TER

[GIK22] Stephan Gufler, Jan Lukas Igelbrink, and Nicola Kistler. TAP equations are repulsive. Electronic Communications in Probability, 27(??):1–7, 2022. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-27/issue-
none/TAP-equations-are-repulsive/10.1214/22-ECP505.full.

Grigorova:2020:SVN

[GIOQ20] Miryana Grigorova, Peter Imkeller, Youssef Ouknine, and Marie-Claire Quenez. On the strict value of the non-linear optimal stopping problem. Electronic Communications in Probability, 25 (??):1–9, 2020. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-25/issue-none/On-the-strict-value-
of-the-non-linear-optimal-stopping/10.1214/20-ECP328.
full.

Gao:2009:DIM

[GJ09a] Fuqing Gao and Hui Jiang. Deviation inequalities and moderate deviations for estimators of parameters in an Ornstein– Uhlenbeck process with linear drift. Electronic Communications in Probability, 14:21:210–21:223, 2009. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1466.

Goncalves:2009:DFZ

[GJ09b] Patricia Goncalves and Milton Jara. Density fluctuations for a zero-range process on the percolation cluster. Electronic Communications in Probability, 14:38:382–38:395, 2009. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1491.

Goldberg:2012:CRM

[GJ12] Leslie Goldberg and Mark Jerrum. A counterexample to rapid mixing of the Ge–Stefankovic process. Electronic Communications in Probability, 17:5:1–5:6, 2012. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1712.

Groisman:2018:FPQ

[GJ18] Pablo Groisman and Matthieu Jonckheere. Front propagation and quasi-stationary distributions for one-dimensional Lévy processes. Electronic Communications in Probability, 23(??):93:1– 93:11, ???? 2018. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1544843116.

Gavalakis:2021:ITP

[GK21] Lampros Gavalakis and Ioannis Kontoyiannis. An informationtheoretic proof of a finite de Finetti theorem. Electronic Communications in Probability, 26(??):1–5, 2021. CODEN ???? ISSN 1083-589X.

Gu:2022:GFR

[GK22] Yu Gu and Tomasz Komorowski. Gaussian fluctuations of replica overlap in directed polymers. Electronic Communications in Probability, 27(??):1–12, 2022. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-27/issue-
none/Gaussian-fluctuations-of-replica-overlap-in-directed-
polymers/10.1214/22-ECP476.full.

Gobet:2003:CGB

[GKH03] Emmanuel Gobet and Arturo Kohatsu-Higa. Computation of Greeks for barrier and lookback options using Malliavin calculus. Electronic Communications in Probability, 8:6:51–6:62, 2003. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1069.

Glenz:2018:HPB

[GKS18] Constantin Glenz, Nicola Kistler, and Marius A. Schmidt. High points of branching Brownian motion and McKean’s martingale in the Bovier–Hartung extremal process. Electronic Communications in Probability, 23(??):86:1–86:12, ???? 2018. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1542942175.

Garg:2021:SPI

[GKS21] Ankit Garg, Tarun Kathuria, and Nikhil Srivastava. Scalar Poincaré implies matrix Poincaré. Electronic Communications in Probability, 26(??):1–4, 2021. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-26/issue-
none/Scalar-Poincar%c3%a9-implies-matrix-Poincar%c3%
a9/10.1214/21-ECP371.full.

Gobet:2008:SEC

[GL08] Emmanuel Gobet and Céline Labart. Sharp estimates for the convergence of the density of the Euler scheme in small time. Electronic Communications in Probability, 13:35:352– 35:363, 2008. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1393.

Guntuboyina:2009:CSM

[GL09] Adityanand Guntuboyina and Hannes Leeb. Concentration of the spectral measure of large Wishart matrices with dependent entries. Electronic Communications in Probability, 14:33:334– 33:342, 2009. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1483.

Gobet:2014:OJV

[GL14] Emmanuel Gobet and Nicolas Landon. Optimization of joint p-variations of Brownian semimartingales. Electronic Communications in Probability, 19:34:1–34:14, 2014. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/2975.

Glazman:2015:CCW

[Gla15] Alexander Glazman. Connective constant for a weighted selfavoiding walk on Z^{2}. Electronic Communications in Probability, 20(??):86:1–86:13, ???? 2015. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/3844.

Galane:2023:SLC

[GŁM23] Lesiba Ch. Galane, Rafał M. Łochowski, and Farai J. Mhlanga. On SDEs with Lipschitz coefficients, driven by continuous, model-free martingales. Electronic Communications in Probability, 28(??):1–12, 2023. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-28/issue-
none/On-SDEs-with-Lipschitz-coefficients-driven-by-continuous-
model-free/10.1214/23-ECP520.full.

Gloria:2014:WIS

[Glo14] Antoine Gloria. When are increment-stationary random point sets stationary? Electronic Communications in Probability, 19: 29:1–29:14, 2014. CODEN ???? ISSN 1083-589X. URL http:/

/ecp.ejpecp.org/article/view/3288.

Gheissari:2018:CIP

[GLP18] Reza Gheissari, Eyal Lubetzky, and Yuval Peres. Concentration inequalities for polynomials of contracting Ising models. Electronic Communications in Probability, 23(??):76:1–76:12, ???? 2018. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1539914642.

Gao:2014:CPG

[GLY14] Fuchang Gao, Zhenxia Liu, and Xiangfeng Yang. Conditional persistence of Gaussian random walks. Electronic Communications in Probability, 19(??):70:1–70:9, ???? 2014. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3587.

Gnedin:2012:EUI

[GM12] Alexander Gnedin and Alexander Marynych. Exponential– uniform identities related to records. Electronic Communications in Probability, 17:26:1–26:5, 2012. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/2011.

Garet:2013:CBR

[GM13] Olivier Garet and Régine Marchand. The critical branching random walk in a random environment dies out. Electronic Communications in Probability, 18:9:1–9:15, 2013. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2438.

Guillin:2016:OLD

[GM16] Arnaud Guillin and Pierre Monmarché. Optimal linear drift for the speed of convergence of an hypoelliptic diffusion. Electronic Communications in Probability, 21(??):74:1–74:14, ???? 2016. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1477600775. See erratum [GM17].

Guillin:2017:EOL

[GM17] Arnaud Guillin and Pierre Monmarché. Erratum: Optimal linear drift for the speed of convergence of an hypoelliptic diffusion. Electronic Communications in Probability, 22(??):15:1– 15:2, ???? 2017. CODEN ???? ISSN 1083-589X. URL http://

projecteuclid.org/euclid.ecp/1487062817. See [GM16].

Garet:2021:PFP

[GM21] Olivier Garet and Régine Marchand. Percolation and firstpassage percolation on oriented graphs. Electronic Communications in Probability, 26(??):1–14, 2021. CODEN ???? ISSN 1083-589X.

Gaudio:2022:SAE

[GM22] Julia Gaudio and Elchanan Mossel. Shotgun assembly of Erdős–Rényi random graphs. Electronic Communications in Probability, 27(??):1–14, 2022. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-27/issue-
none/Shotgun-assembly-of-Erd%c5%91s-R%c3%a9nyi-random-
graphs/10.1214/22-ECP445.full.

Gromoll:2018:QCL

[GMP18] H. Christian Gromoll, Mark W. Meckes, and Leonid Petrov. Quenched central limit theorem in a corner growth setting. Electronic Communications in Probability, 23(??):101:1–101:12, ???? 2018. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1545188961.

Grejo:2016:FSI

[GMRC16] Carolina Grejo, Fábio Machado, and Alejandro Roldán-Correa. The fitness of the strongest individual in the subcritical GMS model. Electronic Communications in Probability, 21(??):12:1– 12:5, ???? 2016. CODEN ???? ISSN 1083-589X. URL https:/

/projecteuclid.org/euclid.ecp/1455560036.

Goldschmidt:2008:FRP

[GMS08] Christina Goldschmidt, James Martin, and Dario Spano. Fragmenting random permutations. Electronic Communications in Probability, 13:44:461–44:474, 2008. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1402.

Garet:2015:CAS

[GMT15] Olivier Garet, Regine Marchand, and Marie Théret. Continuity of the asymptotic shape of the supercritical contact process. Electronic Communications in Probability, 20(??):92:1–92:11, ???? 2015. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/4103.

Georgiou:2018:RIP

[GMW18] Nicholas Georgiou, Aleksandar Mijatović, and Andrew R. Wade. A radial invariance principle for non-homogeneous random walks. Electronic Communications in Probability, 23(??):56:1–56:11, ???? 2018. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1536718009.

Gaans:2006:IMS

[GN06] Onno Gaans and Jan Neerven. Invariant measures for stochastic Cauchy problems with asymptotically unstable drift semigroup. Electronic Communications in Probability, 11:3:24–3:34, 2006. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1184.

Gantert:2014:RES

[GN14] Nina Gantert and Jan Nagel. Reconstructing the environment seen by a RWRE. Electronic Communications in Probability, 19: 26:1–26:9, 2014. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/3013.

Gnedin:2008:CRP

[Gne08] Alexander Gnedin. Corners and records of the Poisson process in quadrant. Electronic Communications in Probability, 13:18:187– 18:193, 2008. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1351.

Gnedin:2010:SSM

[Gne10] Alexander Gnedin. A species sampling model with finitely many types. Electronic Communications in Probability, 15:8:79–8:88, 2010. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1532.

Gnedin:2021:SSF

[Gne21] Alexander Gnedin. On sequential selection and a first passage problem for the Poisson process. Electronic Communications in Probability, 26(??):1–13, 2021. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-26/issue-
none/On-sequential-selection-and-a-first-passage-problem-
for-the/10.1214/21-ECP377.full.

Goldstein:2022:ZBE

[Gol22] Larry Goldstein. Zero bias enhanced Stein couplings. Electronic Communications in Probability, 27(??):1–13, 2022. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
journals/electronic-communications-in-probability/volume-
27/issue-none/Zero-bias-enhanced-Stein-couplings/10.
1214/22-ECP504.full.

Goreac:2015:NGT

[Gor15] Dan Goreac. A note on general Tauberian-type results for controlled stochastic dynamics. Electronic Communications in Probability, 20(??):90:1–90:12, ???? 2015. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/4142.

Gouezel:2018:GNS

[Gou18] Sébastien Gouëzel. Growth of normalizing sequences in limit theorems for conservative maps. Electronic Communications in Probability, 23(??):99:1–99:11, ???? 2018. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1545188959.

Gozlan:2006:ICT

[Goz06] Nathael Gozlan. Integral criteria for transportation cost inequalities. Electronic Communications in Probability, 11:7:64–7:77, 2006. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1198.

Giacomin:2001:RTS

[GP01] Giambattista Giacomin and Gustavo Posta. On recurrent and transient sets of inhomogeneous symmetric random walks. Electronic Communications in Probability, 6:4:39–4:53, 2001. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1033.

Graversen:2011:RUC

[GP11] Svend-Erik Graversen and Jan Pedersen. Representations of Urbanik’s classes and multiparameter Ornstein–Uhlenbeck processes. Electronic Communications in Probability, 16:20:200– 20:212, 2011. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1621.

Gunderson:2014:LBB

[GP14] Karen Gunderson and Michał Przykucki. Lower bounds for bootstrap percolation on Galton–Watson trees. Electronic Communications in Probability, 19:41:1–41:7, 2014. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/3315.

Gwynne:2019:BDG

[GP19] Ewain Gwynne and Joshua Pfeffer. Bounds for distances and geodesic dimension in Liouville first passage percolation. Electronic Communications in Probability, 24(??):56:1–56:12, ???? 2019. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1568361882.

Guillotin-Plantard:2013:QLD

[GPHS13] Nadine Guillotin-Plantard, Yueyun Hu, and Bruno Schapira. The quenched limiting distributions of a one-dimensional random walk in random scenery. Electronic Communications in Probability, 18:85:1–85:7, 2013. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/2918.

Guillotin-Plantard:2008:FLT

[GPL08] Nadine Guillotin-Plantard and Arnaud Le Ny. A functional limit theorem for a 2d-random walk with dependent marginals. Electronic Communications in Probability, 13:34:337–34:351, 2008. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1386.

Guillotin-Plantard:2014:QFC

[GPPdS14] Nadine Guillotin-Plantard, Julien Poisat, and Renato Soares dos Santos. A quenched functional central limit theorem for planar random walks in random sceneries. Electronic Communications in Probability, 19:3:1–3:9, 2014. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/3002.

Grabchak:2015:TUT

[Gra15] Michael Grabchak. Three upsilon transforms related to tempered stable distributions. Electronic Communications in Probability, 20(??):82:1–82:10, ???? 2015. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/4366.

Graf:2016:CHD

[Gra16] Robert Graf. Critical heights of destruction for a forest-fire model on the half-plane. Electronic Communications in Probability, 21 (??):39:1–39:10, ???? 2016. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1462901217.

Griffin:2002:TSS

[Gri02] Philip Griffin. Tightness of the Student t-statistic. Electronic Communications in Probability, 7:18:181–18:190, 2002. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1059.

Gripenberg:2011:WCG

[Gri11] Gustaf Gripenberg. White and colored Gaussian noises as limits of sums of random dilations and translations of a single function. Electronic Communications in Probability, 16:45:507– 45:516, 2011. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1650.

Gross:2019:CSE

[Gro19] Renan Gross. A conformal Skorokhod embedding. Electronic Communications in Probability, 24(??):68:1–68:11, ???? 2019.

CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1573095647.

Gantert:2014:LDW

[GRR14] Nina Gantert, Kavita Ramanan, and Franz Rembart. Large deviations for weighted sums of stretched exponential random variables. Electronic Communications in Probability, 19:40:1– 40:14, 2014. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/3266.

Guiol:2003:MSD

[GRS03] Herve Guiol, Krishnamurthi Ravishankar, and Ellen Saada. Microscopic structure of a decreasing shock for the asymmetric kstep exclusion process. Electronic Communications in Probability, 8:19:170–19:178, 2003. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1080.

Grube:2023:SSM

[Gru23] Sebastian Grube. Strong solutions to McKean–Vlasov SDEs with coefficients of nemytskii-type. Electronic Communications in Probability, 28(??):1–13, 2023. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-28/issue-
none/Strong-solutions-to-McKeanVlasov-SDEs-with-coefficients-
of-Nemytskii-type/10.1214/23-ECP519.full.

Goldstein:2009:BEB

[GS09] Larry Goldstein and Qi-Man Shao. Berry–Esséen bounds for projections of coordinate symmetric random vectors. Electronic Communications in Probability, 14:46:474–46:485, 2009. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1502.

Goreac:2012:NLM

[GS12] Dan Goreac and Oana Silvia Serea. A note on linearization methods and dynamic programming principles for stochastic discontinuous control problems. Electronic Communications in Probability, 17:12:1–12:12, 2012. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1844.

Geiss:2021:SLD

[GS21] Sarah Geiss and Michael Scheutzow. Sharpness of Lenglart’s domination inequality and a sharp monotone version. Electronic Communications in Probability, 26(??):1–8, 2021. CODEN ???? ISSN 1083-589X.

Groeneboom:2011:TMB

[GT11] Piet Groeneboom and Nico Temme. The tail of the maximum of Brownian motion minus a parabola. Electronic Communications in Probability, 16:41:458–41:466, 2011. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1645.

Gabrysch:2018:GWI

[GT18] Katja Gabrysch and Erik Thörnblad. The greedy walk on an inhomogeneous Poisson process. Electronic Communications in Probability, 23(??):14:1–14:11, ???? 2018. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1519722244.

Guillotin:1999:EOM

[Gui99] Nadine Guillotin. Edge occupation measure for a reversible Markov chain. Electronic Communications in Probability, 4: 11:87–11:90, 1999. CODEN ???? ISSN 1083-589X. URL http:

//ecp.ejpecp.org/article/view/1009.

Giraudo:2014:CEC

[GV14] Davide Giraudo and Dalibor Volny. A counter example to central limit theorem in Hilbert spaces under a strong mixing condition. Electronic Communications in Probability, 19:62:1–62:12, 2014. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3249.

Gao:2021:AEP

[GWZ21] Fuqing Gao, Yujing Wang, and Youzhou Zhou. Asymptotic expansions and precise deviations in the Kingman coalescent. Electronic Communications in Probability, 26(??):1– 11, 2021. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-26/issue-none/Asymptotic-expansions-
and-precise-deviations-in-the-Kingman-coalescent/10.
1214/21-ECP375.full.

Guionnet:2000:CSM

[GZ00] Alice Guionnet and Ofer Zeitouni. Concentration of the spectral measure for large matrices. Electronic Communications in Probability, 5:14:119–14:136, 2000. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1026.

Garcia-Zelada:2019:CCG

[GZ19] David García-Zelada. Concentration for Coulomb gases on compact manifolds. Electronic Communications in Probability, 24 (??):12:1–12:18, ???? 2019. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1553133701.

Hutzenthaler:2007:GRS

[HA07] Martin Hutzenthaler and Roland Alkemper. Graphical representation of some duality relations in stochastic population models. Electronic Communications in Probability, 12:21:206–21:220, 2007. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1283.

Hafouta:2018:SMN

[Haf18] Yeor Hafouta. Stein’s method for nonconventional sums. Electronic Communications in Probability, 23(??):38:1–38:14, ???? 2018. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1528509622.

Haggstrom:2002:MRH

[Häg02] Olle Häggström. A monotonicity result for hard-core and Widom–Rowlinson models on certain d-dimensional lattices. Electronic Communications in Probability, 7:7:67–7:78, 2002. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1048.

Hajek:2014:MIS

[Haj14] Bruce Hajek. A maximal inequality for supermartingales. Electronic Communications in Probability, 19:55:1–55:10, 2014. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3237.

Hajri:2015:FAT

[Haj15] Hatem Hajri. On flows associated to Tanaka’s SDE and related works. Electronic Communications in Probability, 20(??):16:1– 16:12, ???? 2015. CODEN ???? ISSN 1083-589X. URL http:/

/ecp.ejpecp.org/article/view/4058.

Hammond:2005:CEP

[Ham05] Alan Hammond. Critical exponents in percolation via lattice animals. Electronic Communications in Probability, 10:6:45–6:59, 2005. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1131.

Handa:1998:LBT

[Han98] Kenji Handa. A lower bound for time correlation of lattice gases. Electronic Communications in Probability, 4:1:1–1:8, 1998. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/999.

Handa:1999:LBT

[Han99] Kenji Handa. A lower bound for time correlation of lattice gases. Electronic Communications in Probability, 4:1–8, 1999. CODEN ???? ISSN 1083-589X.

Handa:2005:SFS

[Han05] Kenji Handa. Sampling formulae for symmetric selection. Electronic Communications in Probability, 10:23:223–23:234, 2005. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1159.

Han:2021:GTT

[Han21] Xiyue Han. A Gladyshev theorem for trifractional Brownian motion and n-th order fractional Brownian motion. Electronic Communications in Probability, 26(??):1–12, 2021. CODEN ???? ISSN 1083-589X.

Hara:2004:FDD

[Har04] Keisuke Hara. Finite dimensional determinants as characteristic functions of quadratic Wiener functionals. Electronic Communications in Probability, 9:4:26–4:35, 2004. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/1091.

Hardy:2012:NLD

[Har12] Adrien Hardy. A note on large deviations for 2D Coulomb gas with weakly confining potential. Electronic Communications in Probability, 17:19:1–19:12, 2012. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1818.

Hariya:2014:CBL

[Har14] Yuu Hariya. A connection of the Brascamp–Lieb inequality with Skorokhod embedding. Electronic Communications in Probability, 19:61:1–61:12, 2014. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/3025.

Hariya:2016:PIG

[Har16] Yuu Hariya. A pathwise interpretation of the Gorin–Shkolnikov identity. Electronic Communications in Probability, 21(??):52:1– 52:6, ???? 2016. CODEN ???? ISSN 1083-589X. URL https:/

/projecteuclid.org/euclid.ecp/1469557026.

Hartarsky:2021:BKC

[Har21] Ivailo Hartarsky. Bisection for kinetically constrained models revisited. Electronic Communications in Probability, 26(??):1– 10, 2021. CODEN ???? ISSN 1083-589X.

Hashorva:2005:BCB

[Has05] Enkelejd Hashorva. Boundary crossings of Brownian motion. Electronic Communications in Probability, 10:21:207–21:217, 2005. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1155.

Hajri:2017:ASF

[HCA17] Hatem Hajri, Mine Caglar, and Marc Arnaudon. Application of stochastic flows to the sticky Brownian motion equation. Electronic Communications in Probability, 22(??):3:1–3:10, ???? 2017. CODEN ???? ISSN 1083-589X. URL http://

projecteuclid.org/euclid.ecp/1483585772.

Hofmann-Credner:2008:WTR

[HCS08] Katrin Hofmann-Credner and Michael Stolz. Wigner theorems for random matrices with dependent entries: Ensembles associated to symmetric spaces and sample covariance matrices. Electronic Communications in Probability, 13:39:401–39:414, 2008. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1395.

Huang:2020:CPP

[HD20] Xiangying Huang and Rick Durrett. The contact process on periodic trees. Electronic Communications in Probability, 25(??): 24:1–24:12, ???? 2020. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1585101897. See

corrigendum [HD23].

Huang:2023:CCP

[HD23] Xiangying Huang and Rick Durrett. Corrigendum to: The contact process on periodic trees. Electronic Communications in Probability, 28(??):1–8, 2023. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-28/issue-
none/Corrigendum-to-The-contact-process-on-periodic-
trees/10.1214/23-ECP518.full. See [HD20].

Hermon:2017:CRG

[Her17] Jonathan Hermon. Cutoff for Ramanujan graphs via degree inflation. Electronic Communications in Probability, 22(??):45:1– 45:10, ???? 2017. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1502762751.

Harris:2007:SPB

[HH07] John Harris and Simon Harris. Survival probabilities for branching Brownian motion with absorption. Electronic Communications in Probability, 12:10:81–10:92, 2007. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/1259.

Hernandez-Hernandez:2022:GBD

[HHJ22] Ma. Elena Hernández-Hernández and Saul D. Jacka. A generalisation of the Burkholder–Davis–Gundy inequalities. Electronic Communications in Probability, 27(??):1–8, 2022. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
journals/electronic-communications-in-probability/volume-
27/issue-none/A-generalisation-of-the-Burkholder-Davis-
Gundy-inequalities/10.1214/22-ECP493.full.

Holroyd:2018:FDC

[HHL18] Alexander E. Holroyd, Tom Hutchcroft, and Avi Levy. Finitely dependent cycle coloring. Electronic Communications in Probability, 23(??):64:1–64:12, ???? 2018. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1536977437.

Hu:2016:IFS

[HHN16] Yaozhong Hu, Jingyu Huang, and David Nualart. On the intermittency front of stochastic heat equation driven by colored noises. Electronic Communications in Probability, 21(??):21:1– 21:13, ???? 2016. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1456840982.

Henderson:2018:OSS

[HHZ18] Vicky Henderson, David Hobson, and Matthew Zeng. Optimal stopping and the sufficiency of randomized threshold strategies. Electronic Communications in Probability, 23(??):30:1– 30:11, ???? 2018. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1525312854.

Hildebrand:2006:CDG

[Hil06] Martin Hildebrand. On the Chung–Diaconis–Graham random process. Electronic Communications in Probability, 11:34:347– 34:356, 2006. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1237.

Hillion:2012:CEA

[Hil12] Erwan Hillion. Concavity of entropy along binomial convolutions. Electronic Communications in Probability, 17:4:1–4:9, 2012. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1707.

Haslegrave:2018:NCP

[HJ18] John Haslegrave and Jonathan Jordan. Non-convergence of proportions of types in a preferential attachment graph with three co-existing types. Electronic Communications in Probability, 23 (??):54:1–54:12, ???? 2018. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1535767265.

Hinsen:2020:PTC

[HJCW20] Alexander Hinsen, Benedikt Jahnel, Elie Cali, and Jean-Philippe Wary. Phase transitions for chase-escape models on Poisson– Gilbert graphs. Electronic Communications in Probability, 25 (??):25:1–25:14, ???? 2020. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1585188174.

Hirsch:2022:SPT

[HJM22] Christian Hirsch, Benedikt Jahnel, and Stephen Muirhead. Sharp phase transition for Cox percolation. Electronic Communications in Probability, 27(??):1–13, 2022. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-27/issue-
none/Sharp-phase-transition-for-Cox-percolation/10.1214/
22-ECP487.full.

Hashorva:2012:ISD

[HJT12] Enkelejd Hashorva, Lanpeng Ji, and Zhongquan Tan. On the infinite sums of deflated Gaussian products. Electronic Communications in Probability, 17:31:1–31:8, 2012. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/1921.

Hirsch:2020:LLD

[HJT20] Christian Hirsch, Benedikt Jahnel, and András Tóbiás. Lower large deviations for geometric functionals. Electronic Communications in Probability, 25(??):41:1–41:12, ???? 2020. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
euclid.ecp/1592272819.

Hubalek:2011:CSR

[HK11] Friedrich Hubalek and Alexey Kuznetsov. A convergent series representation for the density of the supremum of a stable process. Electronic Communications in Probability, 16:8:84–8:95, 2011. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1601.

Hackmann:2013:NSR

[HK13] Daniel Hackmann and Alexey Kuznetsov. A note on the series representation for the density of the supremum of a stable process. Electronic Communications in Probability, 18:42:1–42:5, 2013. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/2757.

Hasebe:2014:FSD

[HK14] Takahiro Hasebe and Alexey Kuznetsov. On free stable distributions. Electronic Communications in Probability, 19:56:1–56:12, 2014. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/3443.

Hartung:2015:GPC

[HK15] Lisa Hartung and Anton Klimovsky. The glassy phase of the complex branching Brownian motion energy model. Electronic Communications in Probability, 20(??):78:1–78:15, ???? 2015. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/4360.

Huang:2016:SIG

[HK16] Ruojun Huang and Takashi Kumagai. Stability and instability of Gaussian heat kernel estimates for random walks among time-dependent conductances. Electronic Communications in Probability, 21(??):5:1–5:11, ???? 2016. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1454514625.

Huang:2017:MLB

[HK17] Jingyu Huang and Davar Khoshnevisan. On the multifractal local behavior of parabolic stochastic PDEs. Electronic Communications in Probability, 22(??):49:1–49:11, ???? 2017. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
euclid.ecp/1506931449.

Huang:2018:EFD

[HKST18] Ruojun Huang, Daniel Kious, Vladas Sidoravicius, and Pierre Tarrès. Explicit formula for the density of local times of Markov jump processes. Electronic Communications in Probability, 23 (??):90:1–90:7, ???? 2018. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1544843113.

Hsu:2012:TIS

[HKZ12a] Daniel Hsu, Sham Kakade, and Tong Zhang. Tail inequalities for sums of random matrices that depend on the intrinsic dimension. Electronic Communications in Probability, 17:14:1–14:13, 2012. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1869.

Hsu:2012:TIQ

[HKZ12b] Daniel Hsu, Sham Kakade, and Tong Zhang. A tail inequality for quadratic forms of subGaussian random vectors. Electronic Communications in Probability, 17:52:1–52:6, 2012. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2079.

He:2013:NSL

[HL13] Hui He and Nana Luan. A note on the scaling limits of contour functions of Galton–Watson trees. Electronic Communications in Probability, 18:79:1–79:13, 2013. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/2781.

Hairer:2015:SCC

[HL15a] Martin Hairer and Cyril Labbé. A simple construction of the continuum parabolic Anderson model on R^{2}. Electronic Communications in Probability, 20(??):43:1–43:11, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/4038.

Holroyd:2015:SDC

[HL15b] Alexander E. Holroyd and Thomas M. Liggett. Symmetric 1-dependent colorings of the integers. Electronic Communications in Probability, 20(??):31:1–31:8, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/4070.

Havet:2020:QMI

[HLMV20] Antoine Havet, Matthieu Lerasle, Eric Moulines, and Elodie Vernet. A quantitative McDiarmid’s inequality for geometrically ergodic Markov chains. Electronic Communications in Probability, 25(??):15:1–15:11, ???? 2020. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1581325212.

Hu:2013:NDS

[HLN13] Yaozhong Hu, Fei Lu, and David Nualart. Non-degeneracy of some Sobolev pseudo-norms of fractional Brownian motion. Electronic Communications in Probability, 18:84:1–84:8, 2013. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2986.

Huang:2021:NOR

[HLSX21] Xiangyu Huang, Yong Liu, Vladas Sidoravicius, and Kainan Xiang. A note on once reinforced random walk on ladder Z ×{0, 1}. Electronic Communications in Probability, 26(??):1–12, 2021. CODEN ???? ISSN 1083-589X.

Hu:2015:IEN

[HLWZ15] Mingshang Hu, Hanwu Li, Falei Wang, and Guoqiang Zheng. Invariant and ergodic nonlinear expectations for g-diffusion processes. Electronic Communications in Probability, 20(??):30:1– 30:15, ???? 2015. CODEN ???? ISSN 1083-589X. URL http:/

/ecp.ejpecp.org/article/view/3886.

Haggstrom:2009:STD

[HM09] Olle Häggström and Péter Mester. Some two-dimensional finite energy percolation processes. Electronic Communications in Probability, 14:5:42–5:54, 2009. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1446.

Hamana:2014:APD

[HM14] Yuji Hamana and Hiroyuki Matsumoto. Asymptotics of the probability distributions of the first hitting times of Bessel processes.

Electronic Communications in Probability, 19:5:1–5:5, 2014. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3215.

Holroyd:2015:PAB

[HM15] Alexander E Holroyd and James B Martin. Poisson allocations with bounded connected cells. Electronic Communications in Probability, 20(??):69:1–69:8, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3853.

Houdre:2016:CDL

[HM16] Christian Houdré and Heinrich Matzinger. Closeness to the diagonal for longest common subsequences in random words. Electronic Communications in Probability, 21(??):36:1–36:19, ???? 2016. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1461781966.

Hambly:2022:MVE

[HM22] Ben Hambly and Julian Meier. McKean–Vlasov equations with positive feedback through elastic stopping times. Electronic Communications in Probability, 27(??):1–13, 2022. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
journals/electronic-communications-in-probability/volume-
27/issue-none/McKean--Vlasov-equations-with-positive-
feedback-through-elastic-stopping-times/10.1214/22-ECP482.
full.

Hariya:2023:TDE

[HM23] Yuu Hariya and Yohei Matsumura. On two-dimensional extensions of Bougerol’s identity in law. Electronic Communications in Probability, 28(??):1–7, 2023. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-28/issue-
none/On-two-dimensional-extensions-of-Bougerols-identity-
in-law/10.1214/23-ECP510.full.

Hambly:2001:PTS

[HMO01] B. Hambly, James Martin, and Neil O’Connell. Pitman’s 2M −X theorem for skip-free random walks with Markovian increments. Electronic Communications in Probability, 6:7:73–7:77, 2001. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1036.

Huss:2015:RRG

[HMSH15] Wilfried Huss, Sebastian Müller, and Ecaterina Sava-Huss. Rotor-routing on Galton–Watson trees. Electronic Communications in Probability, 20(??):49:1–49:12, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/4000.

Hu:2009:SIR

[HN09] Yaozhong Hu and David Nualart. Stochastic integral representation of the L^{2} modulus of Brownian local time and a central limit theorem. Electronic Communications in Probability, 14: 51:529–51:539, 2009. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1511.

Hu:2010:CLT

[HN10] Yaozhong Hu and David Nualart. Central limit theorem for the third moment in space of the Brownian local time increments. Electronic Communications in Probability, 15:36:396– 36:410, 2010. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1573.

Heil:2011:RLB

[HN11] Hadrian Heil and Makoto Nakashima. A remark on localization for branching random walks in random environment. Electronic Communications in Probability, 16:31:323–31:336, 2011. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1603.

Hoepfner:2009:EYW

[Hoe09] Reinhard Hoepfner. An extension of the Yamada–Watanabe condition for pathwise uniqueness to stochastic differential equations with jumps. Electronic Communications in Probability, 14: 43:447–43:456, 2009. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1499.

Holmes:2009:SLS

[Hol09] Mark Holmes. The scaling limit of senile reinforced random walk. Electronic Communications in Probability, 14:10:104– 10:115, 2009. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1449.

Holmes:2015:SMS

[Hol15] Mark P. Holmes. On strict monotonicity of the speed for excited random walks in one dimension. Electronic Communications in Probability, 20(??):41:1–41:7, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/4212.

Holderrieth:2021:CPD

[Hol21] Peter Holderrieth. Cores for piecewise-deterministic Markov processes used in Markov chain Monte Carlo. Electronic Communications in Probability, 26(??):1–12, 2021. CODEN ???? ISSN 1083-589X.

Hong:2019:IHC

[Hon19] Jieliang Hong. Improved Hölder continuity near the boundary of one-dimensional super-Brownian motion. Electronic Communications in Probability, 24(??):28:1–28:12, ???? 2019. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
euclid.ecp/1559700464.

Hong:2021:EMC

[Hon21] Jieliang Hong. An exit measure construction of the total local time of super-Brownian motion. Electronic Communications in Probability, 26(??):1–9, 2021. CODEN ???? ISSN 1083-589X.

Hooghiemstra:1999:OTB

[Hoo99] Gerard Hooghiemstra. On the occupation time of Brownian excursion. Electronic Communications in Probability, 4:8:61–8:64, 1999. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1006.

Hough:2009:TTR

[Hou09] Robert Hough. Tesselation of a triangle by repeated barycentric subdivision. Electronic Communications in Probability, 14: 27:270–27:277, 2009. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1471.

Holroyd:2003:TMP

[HP03] Alexander Holroyd and Yuval Peres. Trees and matchings from point processes. Electronic Communications in Probability, 8: 3:17–3:27, 2003. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1066.

Hutchcroft:2015:CRW

[HP15] Tom Hutchcroft and Yuval Peres. Collisions of random walks in reversible random graphs. Electronic Communications in Probability, 20(??):63:1–63:6, ???? 2015. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/4330.

Holroyd:2014:WNS

[HPS14] Alexander E. Holroyd, Yuval Peres, and Jeffrey E. Steif. Wald for non-stopping times: the rewards of impatient prophets. Electronic Communications in Probability, 19(??):78:1–78:9, ???? 2014. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/3609.

Haggstrom:2007:VCM

[HR07] Olle Häggström and Jeffrey Rosenthal. On variance conditions for Markov chain CLTs. Electronic Communications in Probability, 12:43:454–43:464, 2007. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1336.

Harris:2014:SLL

[HR14] Simon Harris and Matthew Roberts. A strong law of large numbers for branching processes: almost sure spine events. Electronic Communications in Probability, 19:27:1–27:6, 2014. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2641.

Hahn:2011:TCG

[HRKU11] Marjorie Hahn, Jelena Ryvkina, Kei Kobayashi, and Sabir Umarov. On time-changed Gaussian processes and their associated Fokker–Planck–Kolmogorov equations. Electronic Communications in Probability, 16:15:150–15:164, 2011. CODEN

????  ISSN 1083-589X. URL http://ecp.ejpecp.org/
article/view/1620.

Habeck:2020:SDI

[HRS20] Michael Habeck, Daniel Rudolf, and Björn Sprungk. Stability of doubly-intractable distributions. Electronic Communications in Probability, 25(??):1–13, 2020. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-25/issue-
none/Stability-of-doubly-intractable-distributions/10.
1214/20-ECP341.full.

Hu:2009:NDP

[HS09] Yueyun Hu and Qi-Man Shao. A note on directed polymers in Gaussian environments. Electronic Communications in Probability, 14:50:518–50:528, 2009. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1509.

Hasebe:2011:JCN

[HS11] Takahiro Hasebe and Hayato Saigo. Joint cumulants for natural independence. Electronic Communications in Probability, 16:44:491–44:506, 2011. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1647.

Huss:2012:TRR

[HS12] Wilfried Huss and Ecaterina Sava. Transience and recurrence of rotor-router walks on directed covers of graphs. Electronic Communications in Probability, 17:41:1–41:13, 2012. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2096.

Hartarsky:2022:SBP

[HS22] Ivailo Hartarsky and Réka Szabó. Subcritical bootstrap percolation via Toom contours. Electronic Communications in Probability, 27(??):1–13, 2022. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-27/issue-
none/Subcritical-bootstrap-percolation-via-Toom-contours/
10.1214/22-ECP496.full.

Hu:2015:MDB

[HSY15] Yueyun Hu, Zhan Shi, and Marc Yor. The maximal drawdown of the Brownian meander. Electronic Communications in Probability, 20(??):39:1–39:6, ???? 2015. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/3945.

Hobson:2005:DBC

[HT05] Tim Hobson and Rodge Tribe. On the duality between coalescing Brownian particles and the heat equation driven by Fisher– Wright noise. Electronic Communications in Probability, 10: 14:136–14:145, 2005. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1143.

Huang:2020:RSS

[HT20] Han Huang and Konstantin Tikhomirov. A remark on the smallest singular value of powers of Gaussian matrices. Electronic Communications in Probability, 25(??):10:1–10:8, ???? 2020. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1580353228.

Huang:2017:SHE

[Hua17] Jingyu Huang. On stochastic heat equation with measure initial data. Electronic Communications in Probability, 22(??):40:1– 40:6, ???? 2017. CODEN ???? ISSN 1083-589X. URL https:/

/projecteuclid.org/euclid.ecp/1502416901.

Huang:2018:CMB

[Hua18] Brice Huang. Convergence of maximum bisection ratio of sparse random graphs. Electronic Communications in Probability, 23 (??):51:1–51:10, ???? 2018. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1535767262.

Huang:2019:GTI

[Hua19] Ruojun Huang. Growing in time IDLA cluster is recurrent. Electronic Communications in Probability, 24(??):31:1–31:11, ???? 2019. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1560391477.

Huang:2021:APC

[Hua21] Yichao Huang. Another probabilistic construction of Φ_{2}^{2n}. Electronic Communications in Probability, 26(??):1–13, 2021. CO-DEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/journals/electronic-communications-in-probability/
volume-26/issue-none/Another-probabilistic-construction-
of-%ce%a622n/10.1214/21-ECP389.full.

Huesmann:2016:TCE

[Hue16] Martin Huesmann. Transport cost estimates for random measures in dimension one. Electronic Communications in Probability, 21(??):46:1–46:10, ???? 2016. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1464709607.

Huss:2008:IDL

[Hus08] Wilfried Huss. Internal diffusion-limited aggregation on nonamenable graphs. Electronic Communications in Probability, 13:

27:272–27:279, 2008. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1374.

Hutzenthaler:2011:SBD

[Hut11] Martin Hutzenthaler. Supercritical branching diffusions in random environment. Electronic Communications in Probability, 16:69:781–69:791, 2011. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1685.

Hutchcroft:2018:HWB

[Hut18] Tom Hutchcroft. The Hammersley–Welsh bound for self-avoiding walk revisited. Electronic Communications in Probability, 23(??): 5:1–5:8, ???? 2018. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1518426011.

Holmes:2008:EIA

[HvdHS08] Mark Holmes, Remco van der Hofstad, and Gordon Slade. An extension of the inductive approach to the lace expansion. Electronic Communications in Probability, 13:29:291–29:301, 2008. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1377.

Han:2019:ELB

[HW19] Yanjun Han and Guanyang Wang. Expectation of the largest bet size in the Labouchere system. Electronic Communications in Probability, 24(??):11:1–11:10, ???? 2019. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1550826036.

Hariya:2022:BDF

[HW22] Yuu Hariya and Sou Watanabe. The Boué–Dupuis formula and the exponential hypercontractivity in the Gaussian space. Electronic Communications in Probability, 27(??):1– 13, 2022. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-27/issue-none/The-Bou%c3%a9Dupuis-
formula-and-the-exponential-hypercontractivity-in-the-
Gaussian/10.1214/22-ECP461.full.

Hingant:2021:QSD

[HY21] Erwan Hingant and Romain Yvinec. Quasi-stationary distribution and metastability for the stochastic Becker–Döring model.

Electronic Communications in Probability, 26(??):1–14, 2021. CODEN ???? ISSN 1083-589X.

He:2019:SCL

[HYZ19] Guoman He, Ge Yang, and Yixia Zhu. Some conditional limiting theorems for symmetric Markov processes with tightness property. Electronic Communications in Probability, 24(??):60:1– 60:11, ???? 2019. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1569895735.

Huang:2007:NIP

[HZ07] Wei Huang and Li-Xin Zhang. A note on the invariance principle of the product of sums of random variables. Electronic Communications in Probability, 12:6:51–6:56, 2007. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1255.

Iksanov:2018:FLT

[IK18] Alexander Iksanov and Zakhar Kabluchko. A functional limit theorem for the profile of random recursive trees. Electronic Communications in Probability, 23(??):87:1–87:13, ???? 2018. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1542942176.

Ilienko:2019:CPP

[Ili19] Andrii Ilienko. Convergence of point processes associated with coupon collector’s and Dixie cup problems. Electronic Communications in Probability, 24(??):51:1–51:9, ???? 2019. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1568253713.

Iksanov:2007:PPW

[IM07] Alex Iksanov and Martin Möhle. A probabilistic proof of a weak limit law for the number of cuts needed to isolate the root of a random recursive tree. Electronic Communications in Probability, 12:4:28–4:35, 2007. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1253.

Iksanov:2010:EMF

[IM10] Alexander Iksanov and Matthias Meiners. Exponential moments of first passage times and related quantities for random walks. Electronic Communications in Probability, 15:34:365– 34:375, 2010. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1569.

Isaev:2016:BHC

[IM16] Mikhail Isaev and Brendan D. McKay. On a bound of Hoeffding in the complex case. Electronic Communications in Probability, 21(??):14:1–14:7, ???? 2016. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1455717136.

Ignatiouk-Robert:2010:MBR

[IR10] Irina Ignatiouk-Robert. T -Martin boundary of reflected random walks on a half-space. Electronic Communications in Probability, 15:15:149–15:161, 2010. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1541.

Ichiba:2017:YAC

[IS17] Tomoyuki Ichiba and Andrey Sarantsev. Yet another condition for absence of collisions for competing Brownian particles. Electronic Communications in Probability, 22(??):8:1–8:7, ???? 2017. CODEN ???? ISSN 1083-589X. URL http://projecteuclid.

org/euclid.ecp/1484363135.

Istas:2005:SHF

[Ist05] Jacques Istas. Spherical and hyperbolic fractional Brownian motion. Electronic Communications in Probability, 10:26:254– 26:262, 2005. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1166.

Istas:2006:FFI

[Ist06] Jacques Istas. On fractional fields indexed by metric spaces. Electronic Communications in Probability, 11:26:242–26:251, 2006. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1223.

Jaber:2017:SIC

[Jab17] Eduardo Abi Jaber. Stochastic invariance of closed sets for jumpdiffusions with non-Lipschitz coefficients. Electronic Communications in Probability, 22(??):53:1–53:15, ???? 2017. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
euclid.ecp/1507860210.

Jacquot:2014:LGA

[Jac14] Stephanie S. M. Jacquot. Large gaps asymptotics for the 1dimensional random Schrödinger operator. Electronic Communications in Probability, 19(??):82:1–82:12, ???? 2014. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2724.

Jackson:2023:GEQ

[Jac23] Joe Jackson. Global existence for quadratic FBSDE systems and application to stochastic differential games. Electronic Communications in Probability, 28(??):1–14, 2023. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-28/issue-
none/Global-existence-for-quadratic-FBSDE-systems-and-
application-to-stochastic/10.1214/23-ECP513.full.

Jansons:1996:END

[Jan96] Kalvis M. Jansons. Excursions into a new duality relation for diffusion processes. Electronic Communications in Probability, 1:7:65–7:69, 1996. CODEN ???? ISSN 1083-589X. URL http:

//ecp.ejpecp.org/article/view/977.

Jansons:1997:DTS

[Jan97] Kalvis M. Jansons. The distribution of time spent by a standard excursion above a given level, with applications to ring polymers near a discontinuity in potential. Electronic Communications in Probability, 2:5:53–5:58, 1997. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/984.

Janson:2009:SRM

[Jan09] Svante Janson. Standard representation of multivariate functions on a general probability space. Electronic Communications in Probability, 14:34:343–34:346, 2009. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1477.

Janson:2013:MLM

[Jan13] Svante Janson. Moments of the location of the maximum of Brownian motion with parabolic drift. Electronic Communications in Probability, 18:15:1–15:8, 2013. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/2330.

Janson:2015:TLQ

[Jan15] Svante Janson. On the tails of the limiting Quicksort distribution. Electronic Communications in Probability, 20(??):81:1– 81:7, ???? 2015. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/4525.

Janjigian:2019:UTL

[Jan19a] Christopher Janjigian. Upper tail large deviations in Brownian directed percolation. Electronic Communications in Probability, 24(??):45:1–45:10, ???? 2019. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1562292106.

Janson:2019:RRP

[Jan19b] Svante Janson. Random replacements in Pólya urns with infinitely many colours. Electronic Communications in Probability, 24(??):23:1–23:11, ???? 2019. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1556589626.

Janson:2020:INS

[Jan20] Svante Janson. On the independence number of some random trees. Electronic Communications in Probability, 25(??): 1–14, 2020. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-25/issue-none/On-the-independence-
number-of-some-random-trees/10.1214/20-ECP345.full.

Janson:2021:GSC

[Jan21] Svante Janson. On general subtrees of a conditioned Galton Watson tree. Electronic Communications in Probability, 26(??): 1–12, 2021. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-26/issue-none/On-general-subtrees-
of-a-conditioned-GaltonWatson-tree/10.1214/21-ECP392.
full.

Janson:2004:CMI

[JC04] Svante Janson and Philippe Chassaing. The center of mass of the ISE and the Wiener index of trees. Electronic Communications in Probability, 9:20:178–20:187, 2004. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1088.

Jegaraj:2009:STA

[Jeg09] Terence Jegaraj. Small time asymptotics of Ornstein–Uhlenbeck densities in Hilbert spaces. Electronic Communications in Probability, 14:53:552–53:559, 2009. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1510.

Jara:2019:SSS

[JF19] Milton Jara and Gregorio R. Moreno Flores. Scaling of the sasamoto-spohn model in equilibrium. Electronic Communications in Probability, 24(??):3:1–3:12, ???? 2019. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1548817627.

Johnson:2016:CDF

[JJ16] Tobias Johnson and Matthew Junge. The critical density for the frog model is the degree of the tree. Electronic Communications in Probability, 21(??):82:1–82:12, ???? 2016. CODEN ????

ISSN 1083-589X. URL http://projecteuclid.org/euclid.
ecp/1480734227.

Jarai:2004:BDH

[JK04] Antal Jarai and Harry Kesten. A bound for the distribution of the hitting time of arbitrary sets by random walk. Electronic Communications in Probability, 9:16:152–16:161, 2004. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1119.

Jiang:2008:DRF

[JK08] Thomas Jiang and Kun-Lin Kuo. Distribution of a random functional of a Ferguson–Dirichlet process over the unit sphere. Electronic Communications in Probability, 13:49:518–49:525, 2008. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1416.

Jansen:2013:GRC

[JK13] Sabine Jansen and Noemi Kurt. Graphical representation of certain moment dualities and application to population models with balancing selection. Electronic Communications in Probability, 18:14:1–14:15, 2013. CODEN ???? ISSN 1083-589X. URL http:

//ecp.ejpecp.org/article/view/2194.

Jamneshan:2017:MQB

[JKL17] Asgar Jamneshan, Michael Kupper, and Peng Luo. Multidimensional quadratic BSDEs with separated generators. Electronic Communications in Probability, 22(??):58:1–58:10, ???? 2017. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1508292096.

Jagers:2008:GBP

[JL08] Peter Jagers and Andreas Lagerås. General branching processes conditioned on extinction are still branching processes. Electronic Communications in Probability, 13:51:540–51:547, 2008. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1419.

Jung:2018:DLS

[JL18] Paul Jung and Jaehun Lee. Delocalization and limiting spectral distribution of Erdős–Rényi graphs with constant expected degree. Electronic Communications in Probability, 23(??):92:1– 92:13, ???? 2018. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1544843115.

Jalowy:2021:RRR

[JL21] Jonas Jalowy and Matthias Löwe. Reconstructing a recurrent random environment from a single trajectory of a Random Walk in Random Environment with errors. Electronic Communications in Probability, 26(??):1–12, 2021. CODEN ???? ISSN 1083-589X.

Jonasson:2015:RMD

[JM15] Johan Jonasson and Benjamin J. Morris. Rapid mixing of dealer shuffles and clumpy shuffles. Electronic Communications in Probability, 20(??):20:1–20:11, ???? 2015. CODEN

????  ISSN 1083-589X. URL http://ecp.ejpecp.org/
article/view/3682.

Jonasson:2004:OSR

[Jon04] Johan Jonasson. On the optimal strategy in a random game. Electronic Communications in Probability, 9:14:132– 14:139, 2004. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1100.

Jonasson:2013:BIP

[Jon13] Johan Jonasson. The BK inequality for pivotal sampling a.k.a. the Srinivasan sampling process. Electronic Communications in Probability, 18:35:1–35:6, 2013. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/2045.

Jost:2007:NET

[Jos07] Céline Jost. A note on ergodic transformations of self-similar Volterra Gaussian processes. Electronic Communications in Probability, 12:25:259–25:266, 2007. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1298.

Jourdain:2012:EPI

[Jou12] Benjamin Jourdain. Equivalence of the Poincaré inequality with a transport-chi-square inequality in dimension one. Electronic Communications in Probability, 17:43:1–43:12, 2012. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2115.

Jakobsen:2019:IOC

[JPR19] Espen R. Jakobsen, Athena Picarelli, and Christoph Reisinger. Improved order 1/4 convergence for piecewise constant policy approximation of stochastic control problems. Electronic Communications in Probability, 24(??):59:1–59:10, ???? 2019. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
euclid.ecp/1568793626.

Jones:2011:CHT

[JR11] Owen Jones and David Rolls. A characterisation of, and hypothesis test for, continuous local martingales. Electronic Communications in Probability, 16:56:638–56:651, 2011. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1673.

Johnson:2019:SFM

[JR19] Tobias Johnson and Leonardo T. Rolla. Sensitivity of the frog model to initial conditions. Electronic Communications in Probability, 24(??):29:1–29:9, ???? 2019. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1559700465.

Jonasson:2000:CTP

[JS00] Johan Jonasson and Oded Schramm. On the cover time of planar graphs. Electronic Communications in Probability, 5:10:85– 10:90, 2000. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1022.

Johndrow:2018:FMM

[JS18] James Johndrow and Aaron Smith. Fast mixing of Metropolis– Hastings with unimodal targets. Electronic Communications in Probability, 23(??):71:1–71:9, ???? 2018. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1539655259.

Jain:2022:RDR

[JSS22] Vishesh Jain, Ashwin Sah, and Mehtaab Sawhney. Rank deficiency of random matrices. Electronic Communications in Probability, 27(??):1–9, 2022. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-27/issue-
none/Rank-deficiency-of-random-matrices/10.1214/22-ECP455.
full.

Jabir:2018:MFL

[JTT18] Jean-François Jabir, Denis Talay, and Milica Tomašević. Meanfield limit of a particle approximation of the one-dimensional parabolic-parabolic Keller–Segel model without smoothing. Electronic Communications in Probability, 23(??):84:1–84:14, ???? 2018. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1540865165.

Julien:2019:RWS

[Jul19] Brémont Julien. Random walk in a stratified independent random environment. Electronic Communications in Probability, 24 (??):47:1–47:15, ???? 2019. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1567649074.

Jung:2011:IFS

[Jun11a] Paul Jung. Indicator fractional stable motions. Electronic Communications in Probability, 16:16:165–16:173, 2011. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1611.

Junglen:2011:QBA

[Jun11b] Stefan Junglen. Quantization balls and asymptotics of quantization radii for probability distributions with radial exponential tails. Electronic Communications in Probability, 16:27:283– 27:295, 2011. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1629.

Janssen:2009:ESM

[JV09] A. J. E. M. Janssen and J. S. H. Van Leeuwaarden. Equidistant sampling for the maximum of a Brownian motion with drift on a finite horizon. Electronic Communications in Probability, 14: 14:143–14:150, 2009. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1453.

Kabluchko:2019:IDH

[Kab19] Zakhar Kabluchko. An infinite-dimensional helix invariant under spherical projections. Electronic Communications in Probability, 24(??):25:1–25:13, ???? 2019. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1559354658.

Kahn:2003:ITC

[Kah03] Jeff Kahn. Inequality of two critical probabilities for percolation. Electronic Communications in Probability, 8:21:184–21:187, 2003. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1099.

Kargin:2007:PNC

[Kar07] Vladislav Kargin. A proof of a non-commutative Central Limit Theorem by the Lindeberg method. Electronic Communications in Probability, 12:5:36–5:50, 2007. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1250.

Kargin:2008:AGS

[Kar08] Vladislav Kargin. On asymptotic growth of the support of free multiplicative convolutions. Electronic Communications in Probability, 13:40:415–40:421, 2008. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1396.

Kargin:2009:SRT

[Kar09] Vladislav Kargin. Spectrum of random Toeplitz matrices with band structure. Electronic Communications in Probability, 14: 40:412–40:423, 2009. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1492.

Kazashi:2018:DMR

[Kaz18] Yoshihito Kazashi. Discrete maximal regularity of an implicit Euler–Maruyama scheme with non-uniform time discretisation for a class of stochastic partial differential equations. Electronic Communications in Probability, 23(??):29:1–29:14, ???? 2018. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1524881137.

Khan:2005:LLR

[KDN05] Tämur Khan, Luc Devroye, and Ralph Neininger. A limit law for the root value of minimax trees. Electronic Communications in Probability, 10:28:273–28:281, 2005. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1168.

Kasprzak:2017:NBP

[KDV17] Mikołaj J. Kasprzak, Andrew B. Duncan, and Sebastian J. Vollmer. Note on A. Barbour’s paper on Stein’s method for diffusion approximations. Electronic Communications in Probability, 22(??):23:1–23:8, ???? 2017. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1492221618.

Kendall:2004:GEP

[Ken04] Wilfrid Kendall. Geometric ergodicity and perfect simulation. Electronic Communications in Probability, 9:15:140–15:151, 2004. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1117.

Kendall:2009:BCC

[Ken09] Wilfrid Kendall. Brownian couplings, convexity, and shy-ness. Electronic Communications in Probability, 14:7:66–7:80, 2009. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1417.

Kersting:2017:RTM

[Ker17] Götz Kersting. On recurrence and transience of multivariate near-critical stochastic processes. Electronic Communications in Probability, 22(??):7:1–7:12, ???? 2017. CODEN ????

ISSN 1083-589X. URL http://projecteuclid.org/euclid.
ecp/1484190065.

Kesten:1996:NCT

[Kes96] Harry Kesten. On the non-convexity of the time constant in firstpassage percolation. Electronic Communications in Probability, 1:1:1–1:6, 1996. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/971.

Kevei:2016:NKG

[Kev16] Péter Kevei. A note on the Kesten–Grincevičius–Goldie theorem. Electronic Communications in Probability, 21(??):51:1– 51:12, ???? 2016. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1469557025.

Kulske:2009:SEB

[KF09] Christof Külske and Marco Formentin. A symmetric entropy bound on the non-reconstruction regime of Markov chains on Galton–Watson trees. Electronic Communications in Probability, 14:57:587–57:596, 2009. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1516.

Kiesel:1997:SLS

[Kie97] Rüdiger Kiesel. Strong laws and summability for sequences of ϕ-mixing random variables in Banach spaces. Electronic Communications in Probability, 2:3:27–3:41, 1997. CODEN ???? ISSN

1083-589X. URL http://ecp.ejpecp.org/article/view/982.

Kifer:2015:ERL

[Kif15] Yuri Kifer. An Erdős–Rényi law for nonconventional sums. Electronic Communications in Probability, 20(??):83:1–83:8, ???? 2015. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/4613. See erratum [Kif16].

Kifer:2016:EER

[Kif16] Yuri Kifer. Erratum: An Erdős–Rényi law for nonconventional sums. Electronic Communications in Probability, 21(??):33:1, ???? 2016. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1461251161. See [Kif15].

Kink:2008:MZS

[Kin08] Peter Kink. A martingale on the zero-set of a holomorphic function. Electronic Communications in Probability, 13:55:606– 55:613, 2008. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1425.

Kiss:2014:LDB

[Kis14] Demeter Kiss. Large deviation bounds for the volume of the largest cluster in 2d critical percolation. Electronic Communications in Probability, 19:31:1–31:11, 2014. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/3438.

Karatzas:2015:ODC

[KK15] Ioannis Karatzas and Constantinos Kardaras. Optional decomposition for continuous semimartingales under arbitrary filtrations. Electronic Communications in Probability, 20(??):59:1– 59:10, ???? 2015. CODEN ???? ISSN 1083-589X. URL http:/

/ecp.ejpecp.org/article/view/4090.

Karczewska:2014:SVE

[KL14] Anna Karczewska and Carlos Lizama. Stochastic Volterra equations under perturbations. Electronic Communications in Probability, 19:28:1–28:14, 2014. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/3365.

Klebaner:2002:OPW

[Kle02] Fima Klebaner. Option price when the stock is a semimartingale. Electronic Communications in Probability, 7:8:79–8:83, 2002. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1049.

Klimmek:2012:WPC

[Kli12a] Martin Klimmek. The Wronskian parametrises the class of diffusions with a given distribution at a random time. Electronic Communications in Probability, 17:50:1–50:8, 2012. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1976.

Klimovsky:2012:HDG

[Kli12b] Anton Klimovsky. High-dimensional Gaussian fields with isotropic increments seen through spin glasses. Electronic Communications in Probability, 17:17:1–17:14, 2012. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/1994.

Kume:2018:LBS

[KLL18] Alfred Kume, Fabrizio Leisen, and Antonio Lijoiï. Limiting behaviour of the stationary search cost distribution driven by a generalized gamma process. Electronic Communications in Probability, 23(??):11:1–11:10, ???? 2018. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1519354834.

Klebaner:2015:LDP

[KLM15] Fima Klebaner, Artem Logachev, and Anatoli Mogulski. Large deviations for processes on half-line. Electronic Communications in Probability, 20(??):75:1–75:14, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/4130.

Khoshnevisan:2005:EVA

[KLS05] Davar Khoshnevisan, David Levin, and Zhan Shi. An extremevalue analysis of the LIL for Brownian motion. Electronic Communications in Probability, 10:20:196–20:206, 2005. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1154.

Knopfel:2021:LDP

[KLS21] Holger Knöpfel, Matthias Löwe, and Holger Sambale. Large deviations, a phase transition, and logarithmic Sobolev inequalities in the block spin Potts model. Electronic Communications in Probability, 26(??):1–14, 2021. CODEN ???? ISSN 1083-589X.

Konarovskyi:2019:DKD

[KLvR19] Vitalii Konarovskyi, Tobias Lehmann, and Max-K. von Renesse. Dean–Kawasaki dynamics: ill-posedness vs. triviality. Electronic Communications in Probability, 24(??):8:1–8:9, ???? 2019. CO-DEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1550113299.

Kontoyiannis:2006:MCC

[KM06] Ioannis Kontoyiannis and Mokshay Madiman. Measure concentration for compound Poisson distributions. Electronic Communications in Probability, 11:5:45–5:57, 2006. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/1190.

Kabluchko:2008:ECR

[KM08] Zakhar Kabluchko and Axel Munk. Exact convergence rate for the maximum of standardized Gaussian increments. Electronic Communications in Probability, 13:30:302–30:310, 2008. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1380.

Khorunzhiy:2009:UBE

[KM09] Oleksiy Khorunzhiy and Jean-François Marckert. Uniform bounds for exponential moment of maximum of a Dyck path. Electronic Communications in Probability, 14:32:327–32:333, 2009. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1486.

Kolesko:2017:CCM

[KM17] Konrad Kolesko and Matthias Meiners. Convergence of complex martingales in the branching random walk: the boundary. Electronic Communications in Probability, 22(??):18:1– 18:14, ???? 2017. CODEN ???? ISSN 1083-589X. URL http:/

/projecteuclid.org/euclid.ecp/1488337251.

Kammoun:2020:PIR

[KM20] Slim Kammoun Kammoun and Mylène Maïda. A product of invariant random permutations has the same small cycle structure as uniform. Electronic Communications in Probability, 25(??): 1–14, 2020. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-25/issue-none/A-product-of-invariant-
random-permutations-has-the-same-small/10.1214/20-ECP334.
full.

Kondo:2006:SPE

[KMiS06] Hitoshi Kondo, Makoto Maejima, and Ken iti Sato. Some properties of exponential integrals of Lévy processes and examples. Electronic Communications in Probability, 11:30:291– 30:303, 2006. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1232.

Kotecky:2016:RIP

[KMU16] Roman Kotecký, Piotr Miłoś, and Daniel Ueltschi. The random interchange process on the hypercube. Electronic Communications in Probability, 21(??):4:1–4:9, ???? 2016. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1454514624.

Kiro:2019:RZS

[KN19] Avner Kiro and Alon Nishry. Rigidity for zero sets of Gaussian entire functions. Electronic Communications in Probability, 24 (??):30:1–30:9, ???? 2019. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1559700466.

Kowalski:2015:CLF

[KNN15] Emmanuel Kowalski, Joseph Najnudel, and Ashkan Nikeghbali. A characterization of limiting functions arising in mod-* convergence. Electronic Communications in Probability, 20(??):79:1– 79:11, ???? 2015. CODEN ???? ISSN 1083-589X. URL http:/

/ecp.ejpecp.org/article/view/4381.

Konig:2001:ELP

[KO01] Wolfgang König and Neil O’Connell. Eigenvalues of the Laguerre process as non-colliding squared Bessel processes. Electronic Communications in Probability, 6:11:107–11:114, 2001. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1040.

Kuelske:2006:SFL

[KO06] Christof Kuelske and Enza Orlandi. A simple fluctuation lower bound for a disordered massless random continuous spin model in d = 2. Electronic Communications in Probability, 11:21:200– 21:205, 2006. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1218.

Kordzakhia:2005:EMH

[Kor05] George Kordzakhia. The escape model on a homogeneous tree. Electronic Communications in Probability, 10:12:113– 12:124, 2005. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1140.

Kosters:2008:SOC

[Kös08] Holger Kösters. On the second-order correlation function of the characteristic polynomial of a real symmetric Wigner matrix. Electronic Communications in Probability, 13:42:435– 42:447, 2008. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1400.

Kovchegov:2009:OPB

[Kov09] Yevgeniy Kovchegov. Orthogonality and probability: beyond nearest neighbor transitions. Electronic Communications in Probability, 14:9:90–9:103, 2009. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1447.

Kovchegov:2010:OPM

[Kov10] Yevgeniy Kovchegov. Orthogonality and probability: mixing times. Electronic Communications in Probability, 15:6:59–6:67, 2010. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1525.

Krishnapur:2004:RGW

[KP04] Manjunath Krishnapur and Yuval Peres. Recurrent graphs where two independent random walks collide finitely often. Electronic Communications in Probability, 9:8:72–8:81, 2004. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1111.

Kukla:2015:SDB

[KP15] Jonas Kukla and Helmut Pitters. A spectral decomposition for the Bolthausen–Sznitman coalescent and the Kingman coalescent. Electronic Communications in Probability, 20(??):87:1– 87:13, ???? 2015. CODEN ???? ISSN 1083-589X. URL http:/

/ecp.ejpecp.org/article/view/4612.

Kobayashi:2022:LTS

[KP22] Kei Kobayashi and Hyunchul Park. Large-time and small-time behaviors of the spectral heat content for time-changed stable processes. Electronic Communications in Probability, 27(??): 1–11, 2022. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-27/issue-none/Large-time-and-small-
time-behaviors-of-the-spectral-heat/10.1214/22-ECP478.
full.

Kwapien:1996:PCB

[KPS96] S. Kwapien, M. Pycia, and W. Schachermayer. A proof of a conjecture of Bobkov and Houdré. Electronic Communications in Probability, 1:2:7–2:10, 1996. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/972.

Kardaras:2019:PSB

[KR19] Constantinos Kardaras and Johannes Ruf. Projections of scaled Bessel processes. Electronic Communications in Probability, 24 (??):43:1–43:11, ???? 2019. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1562119371.

Krikun:2007:CAP

[Kri07] Maxim Krikun. Connected allocation to Poisson points in R^{2}. Electronic Communications in Probability, 12:15:140–15:145, 2007. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1268.

Krizmanic:2014:FWC

[Kri14] Danijel Krizmanic. On functional weak convergence for partial sum processes. Electronic Communications in Probability, 19:

60:1–60:12, 2014. CODEN ???? ISSN 1083-589X. URL http:/

/ecp.ejpecp.org/article/view/3686.

Kriechbaum:2021:STB

[Kri21] Xaver Kriechbaum. Subsequential tightness for branching random walk in random environment. Electronic Communications in Probability, 26(??):1–12, 2021. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-26/issue-
none/Subsequential-tightness-for-branching-random-walk-
in-random-environment/10.1214/21-ECP386.full.

Kaj:1997:SAS

[KS97] Ingemar Kaj and Serik Sagitov. Superprocess approximation for a spatially homogeneous branching walk. Electronic Communications in Probability, 2:6:59–6:70, 1997. CODEN ???? ISSN

1083-589X. URL http://ecp.ejpecp.org/article/view/985.

Kessler:2002:IER

[KS02] David Kessler and Jeremy Schiff. Inclusion-Exclusion Redux. Electronic Communications in Probability, 7:9:85–9:96, 2002. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1050.

Kovchegov:2003:LSL

[KS03] Yevgeniy Kovchegov and Scott Sheffield. Linear speed large deviations for percolation clusters. Electronic Communications in Probability, 8:20:179–20:183, 2003. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1098.

Kozlova:2005:NOT

[KS05a] Marina Kozlova and Paavo Salminen. A note on occupation times of stationary processes. Electronic Communications in Probability, 10:10:94–10:104, 2005. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1138.

Kyprianou:2005:NSO

[KS05b] Andreas Kyprianou and Budhi Surya. On the Novikov–Shiryaev optimal stopping problems in continuous time. Electronic Communications in Probability, 10:15:146–15:154, 2005. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1144.

Kuwada:2007:COK

[KS07] Kazumasa Kuwada and Karl-Theodor Sturm. A counterexample for the optimality of Kendall–Cranston coupling. Electronic Communications in Probability, 12:8:66–8:72, 2007. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1160.

Kuhn:2009:NSI

[KS09] Christoph Kühn and Maximilian Stroh. A note on stochastic integration with respect to optional semimartingales. Electronic Communications in Probability, 14:19:192–19:201, 2009. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1465.

Konig:2010:RWC

[KS10] Wolfgang König and Patrick Schmid. Random walks conditioned to stay in Weyl chambers of type C and D. Electronic Communications in Probability, 15:27:286–27:296, 2010. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1560.

Kolb:2014:EEK

[KS14] Martin Kolb and Mladen Savov. Exponential ergodicity of killed Lévy processes in a finite interval. Electronic Communications in Probability, 19:30:1–30:9, 2014. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/3006.

Kifer:2018:NRM

[KS18] Yuri Kifer and Sasha Sodin. Nonconventional random matrix products. Electronic Communications in Probability, 23(??): 37:1–37:12, ???? 2018. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1528509621. See

erratum [KS19].

Kifer:2019:ENR

[KS19] Yuri Kifer and Sasha Sodin. Erratum: Nonconventional random matrix products. Electronic Communications in Probability, 24 (??):6:1, ???? 2019. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1549530018. See

[KS18].

Karatzas:2011:OST

[KSS11] Ioannis Karatzas, Albert Shiryaev, and Mykhaylo Shkolnikov. On the one-sided Tanaka equation with drift. Electronic Communications in Probability, 16:58:664–58:677, 2011. CODEN

????  ISSN 1083-589X. URL http://ecp.ejpecp.org/
article/view/1665.

Konig:2012:LDL

[KSW12] Wolfgang König, Michele Salvi, and Tilman Wolff. Large deviations for the local times of a random walk among random conductances. Electronic Communications in Probability, 17:10:1– 10:11, 2012. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1820.

Khoshnevisan:2006:NFP

[KSY06] Davar Khoshnevisan, Paavo Salminen, and Marc Yor. A note on a.s. finiteness of perpetual integral functionals of diffusions. Electronic Communications in Probability, 11:11:108–11:117, 2006. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1203.

Katori:2003:NBM

[KT03] Makoto Katori and Hideki Tanemura. Noncolliding Brownian motions and Harish-Chandra formula. Electronic Communications in Probability, 8:13:112–13:121, 2003. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/1076.

Kuhn:2011:OPN

[KT11] Christoph Kühn and Marc Teusch. Optional processes with non-exploding realized power variation along stopping times are làglàd. Electronic Communications in Probability, 16:1:1–1:8, 2011. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1591.

Katori:2013:CBM

[KT13] Makoto Katori and Hideki Tanemura. Complex Brownian motion representation of the Dyson model. Electronic Communications in Probability, 18:4:1–4:16, 2013. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/2554.

Krishnan:2017:IGS

[KTA17] Sunder Ram Krishnan, Jonathan E. Taylor, and Robert J. Adler. The intrinsic geometry of some random manifolds. Electronic Communications in Probability, 22(??):1:1–1:12, ???? 2017. CO-

DEN ???? ISSN 1083-589X. URL http://projecteuclid.org/
euclid.ecp/1483585770.

Kusuoka:2017:RTP

[KTT17] Seiichiro Kusuoka, Hiroshi Takahashi, and Yozo Tamura. Recurrence and transience properties of multi-dimensional diffusion processes in selfsimilar and semi-selfsimilar random environments. Electronic Communications in Probability, 22(??): 4:1–4:11, ???? 2017. CODEN ???? ISSN 1083-589X. URL

http://projecteuclid.org/euclid.ecp/1483585773.

Kuan:2016:IPS

[Kua16] Jeffrey Kuan. An interacting particle system with geometric jump rates near a partially reflecting boundary. Electronic Communications in Probability, 21(??):76:1–76:15, ???? 2016. CO-

DEN ???? ISSN 1083-589X. URL http://projecteuclid.org/
euclid.ecp/1479178838.

Kuan:2021:SNM

[Kua21] Jeffrey Kuan. A short note on Markov duality in multi-species higher spin stochastic vertex models. Electronic Communications in Probability, 26(??):1–11, 2021. CODEN ???? ISSN 1083-589X.

Kuba:2011:ACC

[Kub11] Markus Kuba. Analysis of a class of cannibal urns. Electronic Communications in Probability, 16:51:583–51:599, 2011. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1669.

Kulczycki:2016:MCS

[Kul16] Tadeusz Kulczycki. Mid-concavity of survival probability for isotropic Lévy processes. Electronic Communications in Probability, 21(??):29:1–29:9, ???? 2016. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1459878296.

Kurtz:2014:WSS

[Kur14] Thomas Kurtz. Weak and strong solutions of general stochastic models. Electronic Communications in Probability, 19:58:1– 58:16, 2014. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/2833.

Kuznetsov:2000:USG

[Kuz00] Sergei Kuznetsov. On uniqueness of a solution of Lu = u^{α} with given trace. Electronic Communications in Probability, 5:15:137– 15:147, 2000. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1027.

Koudou:2011:WDM

[KV11] Angelo Koudou and Pierre Vallois. Which distributions have the Matsumoto–Yor property? Electronic Communications in Probability, 16:49:556–49:566, 2011. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1663.

Klicnarova:2013:ZWC

[KV13] Jana Klicnarova and Dalibor Volny. On Zhao–Woodroofe’s condition for martingale approximation. Electronic Communications in Probability, 18:36:1–36:8, 2013. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/2780.

Kotowski:2015:NLG

[KV15] Michał Kotowski and Bálint Virág. Non-Liouville groups with return probability exponent at most 1/2. Electronic Communications in Probability, 20(??):12:1–12:12, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3774.

Kiss:2022:MSE

[KV22] József Kiss and Bálint Vető. Moments of the superdiffusive elephant random walk with general step distribution. Electronic Communications in Probability, 27(??):1–12, 2022. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-27/issue-none/Moments-of-the-superdiffusive-
elephant-random-walk-with-general-step/10.1214/22-ECP485.
full.

Kumagai:2013:FMD

[KZ13] Takashi Kumagai and Ofer Zeitouni. Fluctuations of maxima of discrete Gaussian free fields on a class of recurrent graphs. Electronic Communications in Probability, 18:75:1–75:12, 2013.

CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2632.

Kovac:2019:VEM

[KZK19] Vjekoslav Kovač and Pavel Zorin-Kranich. Variational estimates for martingale paraproducts. Electronic Communications in Probability, 24(??):48:1–48:14, ???? 2019. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1568253710.

Labbe:2013:QSD

[Lab13] Cyril Labbé. Quasi-stationary distributions associated with explosive CSBP. Electronic Communications in Probability, 18: 57:1–57:13, 2013. CODEN ???? ISSN 1083-589X. URL http:/

/ecp.ejpecp.org/article/view/2508.

Lacoin:2010:MAD

[Lac10] Hubert Lacoin. The martingale approach to disorder irrelevance for pinning models. Electronic Communications in Probability, 15:38:418–38:427, 2010. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1572.

Lacoin:2015:PCC

[Lac15] Hubert Lacoin. A product chain without cutoff. Electronic Communications in Probability, 20(??):19:1–19:9, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3765.

Lacker:2018:SFP

[Lac18] Daniel Lacker. On a strong form of propagation of chaos for McKean–Vlasov equations. Electronic Communications in Probability, 23(??):45:1–45:11, ???? 2018. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1532657017.

Lageraas:2007:PMC

[Lag07] Andreas Lagerås. A population model for Λ-coalescents with neutral mutations. Electronic Communications in Probability, 12:2:9–2:20, 2007. CODEN ???? ISSN 1083-589X. URL http:

//ecp.ejpecp.org/article/view/1245.

Lalley:2003:SCL

[Lal03] Steven Lalley. Strict convexity of the limit shape in first-passage percolation. Electronic Communications in Probability, 8:15:135– 15:141, 2003. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1089.

Lam:2022:REA

[Lâm22] Phúc Lâm. On the rate of escape or approach to the origin of a random string. Electronic Communications in Probability, 27 (??):1–13, 2022. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-27/issue-none/On-the-rate-of-escape-
or-approach-to-the-origin/10.1214/22-ECP451.full.

Landa:2022:SPR

[Lan22] Boris Landa. Scaling positive random matrices: concentration and asymptotic convergence. Electronic Communications in Probability, 27(??):1–13, 2022. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-27/issue-
none/Scaling-positive-random-matrices-concentration-
and-asymptotic-convergence/10.1214/22-ECP502.full.

Latala:2008:BBP

[Lat08] Rafal Latala. On the boundedness of Bernoulli processes over thin sets. Electronic Communications in Probability, 13:17:175– 17:186, 2008. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1362.

Laurent:2013:SNN

[Lau13] Stéphane Laurent. Standardness and nonstandardness of nextjump time filtrations. Electronic Communications in Probability, 18:56:1–56:11, 2013. CODEN ???? ISSN 1083-589X. URL http:

//ecp.ejpecp.org/article/view/2766.

Laukkarinen:2017:NMS

[Lau17] Eija Laukkarinen. A note on Malliavin smoothness on the Lévy space. Electronic Communications in Probability, 22(??):34:1– 34:12, ???? 2017. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1498010647.

Laulin:2020:MAP

[Lau20] Lucile Laulin. A martingale approach for Pólya urn processes. Electronic Communications in Probability, 25(??):39:1–39:13, ???? 2020. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1591840901.

Laulin:2022:ISA

[Lau22] Lucile Laulin. Introducing smooth amnesia to the memory of the Elephant Random Walk. Electronic Communications in Probability, 27(??):1–12, 2022. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-27/issue-
none/Introducing-smooth-amnesia-to-the-memory-of-the-
Elephant-Random/10.1214/22-ECP495.full.

Lawler:1996:DFP

[Law96] Gregory Lawler. The dimension of the frontier of planar Brownian motion. Electronic Communications in Probability, 1:5:29– 5:47, 1996. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/975.

Lawler:1998:LEW

[Law98] Gregory F. Lawler. Loop-erased walks intersect infinitely often in four dimensions. Electronic Communications in Probability, 3:5:35–5:42, 1998. CODEN ???? ISSN 1083-589X. URL http:

//ecp.ejpecp.org/article/view/991.

Lawi:2008:HLP

[Law08] Stephan Lawi. Hermite and Laguerre polynomials and matrixvalued stochastic processes. Electronic Communications in Probability, 13:8:67–8:84, 2008. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1353.

Lawler:2014:PPL

[Law14] Gregory Lawler. The probability that planar loop-erased random walk uses a given edge. Electronic Communications in Probability, 19:51:1–51:13, 2014. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/2908.

Leonetti:2022:MDA

[LC22] Paolo Leonetti and Amir Khorrami Chokami. The maximum domain of attraction of multivariate extreme value distributions is small. Electronic Communications in Probability, 27 (??):1–8, 2022. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-27/issue-none/The-maximum-domain-
of-attraction-of-multivariate-extreme-value-distributions/
10.1214/22-ECP501.full.

Liu:2021:HCS

[LD21] Ruoyang Liu and Kai Du. Hölder continuity of solutions to the Dirichlet problem for SPDEs with spatially correlated noise. Electronic Communications in Probability, 26(??):1–13, 2021. CODEN ???? ISSN 1083-589X.

LePrince:2008:RBD

[Le 08] Vincent Le Prince. A relation between dimension of the harmonic measure, entropy and drift for a random walk on a hyperbolic space. Electronic Communications in Probability, 13: 5:45–5:53, 2008. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1350.

LeJan:2019:MRN

[Le 19] Yves Le Jan. On Markovian random networks. Electronic Communications in Probability, 24(??):40:1–40:7, ???? 2019. CO-DEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1562119368.

Ledger:2016:SMT

[Led16] Sean Ledger. Skorokhod’s M1 topology for distribution-valued processes. Electronic Communications in Probability, 21(??): 34:1–34:11, ???? 2016. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1461251162.

Ledesma:2017:HFA

[Led17] Diego S. Ledesma. A heat flow approach to the Godbillon–Vey class. Electronic Communications in Probability, 22(??):2:1–2:6, ???? 2017. CODEN ???? ISSN 1083-589X. URL http://

projecteuclid.org/euclid.ecp/1483585771.

Lejay:2011:SSP

[Lej11] Antoine Lejay. Simulation of a stochastic process in a discontinuous layered medium. Electronic Communications in Probability, 16:67:764–67:774, 2011. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1686.

Lopez-Garcia:2009:CDL

[LG09] Marcos Lopez-Garcia. Characterization of distributions with the length-bias scaling property. Electronic Communications in Probability, 14:18:186–18:191, 2009. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1458.

Lopez-Garcia:2020:KCH

[LG20] Marcos López-García. Krein condition and the Hilbert transform. Electronic Communications in Probability, 25(??):1–7, 2020. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-25/issue-none/Krein-condition-
and-the-Hilbert-transform/10.1214/20-ECP351.full.

Li:1999:GCI

[Li99] Wenbo V. Li. A Gaussian correlation inequality and its applications to small ball probabilities. Electronic Communications in Probability, 4:14:111–14:118, 1999. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1012.

Li:2014:UIH

[Li14] Zhongyang Li. Uniqueness of the infinite homogeneous cluster in the 1–2 model. Electronic Communications in Probability, 19: 22:1–22:8, 2014. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/3105.

Li:2016:HB

[Li16] Xue-Mei Li. On hypoelliptic bridge. Electronic Communications in Probability, 21(??):24:1–24:12, ???? 2016. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1457617915.

Li:2017:SCB

[Li17] Xue-Mei Li. On the semi-classical Brownian Bridge Measure. Electronic Communications in Probability, 22(??):38:1–38:15, ???? 2017. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1501833630.

Li:2020:ASF

[Li20] Zhongyang Li. Asymptotics of Schur functions on almost staircase partitions. Electronic Communications in Probability, 25 (??):1–13, 2020. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-25/issue-none/Asymptotics-of-Schur-
functions-on-almost-staircase-partitions/10.1214/20-
ECP332.full.

Li:2021:MPP

[Li21] Linjun Li. On the Manhattan pinball problem. Electronic Communications in Probability, 26(??):1–11, 2021. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
journals/electronic-communications-in-probability/volume-
26/issue-none/On-the-Manhattan-pinball-problem/10.1214/
21-ECP394.full.

Limic:1999:BLP

[Lim99] Vlada Limic. On the behavior of LIFO preemptive resume queues in heavy traffic. Electronic Communications in Probability, 5: 2:13–2:27, 1999. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1014.

Lindvall:1999:STS

[Lin99] Torgny Lindvall. On Strassen’s theorem on stochastic domination. Electronic Communications in Probability, 4:7:51–7:59, 1999. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1005.

Lin:2009:ASL

[Lin09] Fuming Lin. An almost sure limit theorem for the maxima of strongly dependent Gaussian sequences. Electronic Communications in Probability, 14:22:224–22:231, 2009. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1461.

Lin:2019:MDS

[Lin19] Yier Lin. Markov duality for stochastic six vertex model. Electronic Communications in Probability, 24(??):67:1–67:17, ???? 2019. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1573095646.

Liu:2015:GID

[Liu15] Yuan Liu. Gaussian integrability of distance function under the Lyapunov condition. Electronic Communications in Probability, 20(??):9:1–9:10, ???? 2015. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/3838.

Liu:2022:WGB

[Liu22] Zhipeng Liu. When the geodesic becomes rigid in the directed landscape. Electronic Communications in Probability, 27(??):

1–13, 2022. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-27/issue-none/When-the-geodesic-
becomes-rigid-in-the-directed-landscape/10.1214/22-ECP484.
full.

Lalley:2008:OCM

[LK08] Steven Lalley and George Kordzakhia. An oriented competition model on Z_{+}^{2}. Electronic Communications in Probability, 13: 52:548–52:561, 2008. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1422.

Liu:2007:SLT

[LL07] Wei-Dong Liu and Zheng-Yan Lin. Some LIL type results on the partial sums and trimmed sums with multidimensional indices. Electronic Communications in Probability, 12:22:221– 22:233, 2007. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1286.

Liang:2015:WMM

[LL15] Xingang Liang and Quansheng Liu. Weighted moments for Mandelbrot’s martingales. Electronic Communications in Probability, 20(??):85:1–85:12, ???? 2015. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/4443.

Lanchier:2020:PCM

[LL20] Nicolas Lanchier and Hsin-Lun Li. Probability of consensus in the multivariate deffuant model on finite connected graphs. Electronic Communications in Probability, 25(??):1– 12, 2020. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-25/issue-none/Probability-of-consensus-
in-the-multivariate-Deffuant-model-on-finite/10.1214/
20-ECP359.full.

Lalley:2009:GIH

[LLN09] Steven Lalley, Gregory Lawler, and Hariharan Narayanan. Geometric interpretation of half-plane capacity. Electronic Communications in Probability, 14:55:566–55:571, 2009. CODEN

????  ISSN 1083-589X. URL http://ecp.ejpecp.org/
article/view/1517.

LeGall:2006:OMS

[LM06] Jean-François Le Gall and Mathieu Merle. On the occupation measure of super-Brownian motion. Electronic Communications in Probability, 11:27:252–27:265, 2006. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1225.

Lin:2017:SOB

[LM17] Yier Lin and Bastien Mallein. Second order behavior of the block counting process of beta coalescents. Electronic Communications in Probability, 22(??):61:1–61:8, ???? 2017. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
euclid.ecp/1510736418.

Lugosi:2003:NRC

[LMK03] Gábor Lugosi, Shahar Mendelson, and Vladimir Koltchinskii. A note on the richness of convex hulls of VC classes. Electronic Communications in Probability, 8:18:167–18:169, 2003. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1097.

Lin:2022:CLT

[LMV22] Han-Mai Lin, Florence Merlevède, and Dalibor Volný. On the central limit theorem for stationary random fields under L^{1}-projective condition. Electronic Communications in Probability, 27(??):1–12, 2022. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-27/issue-
none/On-the-central-limit-theorem-for-stationary-random-
fields-under/10.1214/22-ECP486.full.

Loehr:2013:EGP

[Loe13] Wolfgang Loehr. Equivalence of Gromov–Prohorov- and Gromov’s \underline{□}_{λ}-metric on the space of metric measure spaces. Electronic Communications in Probability, 18:17:1–17:10, 2013. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2268.

Londono:2004:STN

[Lon04] Jaime Londono. State tameness: A new approach for credit constrains. Electronic Communications in Probability, 9:1:1–1:13, 2004. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1102.

Louis:2004:EPE

[Lou04] Pierre-Yves Louis. Ergodicity of PCA: Equivalence between spatial and temporal mixing conditions. Electronic Communications in Probability, 9:13:119–13:131, 2004. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1116.

Lewis:1999:CM

[LP99] Thomas M. Lewis and Geoffrey Pritchard. Correlation measures. Electronic Communications in Probability, 4:10:77–10:85, 1999. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1008.

Luschgy:2008:MEL

[LP08] Harald Luschgy and Gilles Pagès. Moment estimates for Lévy processes. Electronic Communications in Probability, 13:41:422– 41:434, 2008. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1397.

Li:2012:SSJ

[LP12] Zenghu Li and Fei Pu. Strong solutions of jump-type stochastic equations. Electronic Communications in Probability, 17:33:1– 33:13, 2012. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1915.

Lawler:2019:WGC

[LP19] Gregory F. Lawler and Petr Panov. Weighted graphs and complex Gaussian free fields. Electronic Communications in Probability, 24(??):38:1–38:9, ???? 2019. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1561169054.

Laketa:2022:SHD

[LPN22] Petra Laketa, Dušan Pokorný, and Stanislav Nagy. Simple halfspace depth. Electronic Communications in Probability, 27(??): 1–12, 2022. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-27/issue-none/Simple-halfspace-
depth/10.1214/22-ECP503.full.

Lechner:2015:UEA

[LPP15] Richard Lechner, Markus Passenbrunner, and Joscha Prochno. Uniform estimates for averages of order statistics of matrices. Electronic Communications in Probability, 20(??):27:1– 27:12, ???? 2015. CODEN ???? ISSN 1083-589X. URL http:/

/ecp.ejpecp.org/article/view/3992.

Lawler:2015:CBT

[LR15] Gregory F Lawler and Mohammad Abbas Rezaei. Up-to-constants bounds on the two-point Green’s function for SLE curves. Electronic Communications in Probability, 20(??):45:1– 45:13, ???? 2015. CODEN ???? ISSN 1083-589X. URL http:/

/ecp.ejpecp.org/article/view/4246.

Lohr:2016:BFM

[LR16] Wolfgang Löhr and Thomas Rippl. Boundedly finite measures: separation and convergence by an algebra of functions. Electronic Communications in Probability, 21(??):60:1–60:16, ???? 2016. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1473424721.

Lampert:2013:MSR

[LS13a] Amichai Lampert and Assaf Shapira. On maximizing the speed of a random walk in fixed environments. Electronic Communications in Probability, 18:40:1–40:9, 2013. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/2431.

Ley:2013:SDA

[LS13b] Christophe Ley and Yvik Swan. Stein’s density approach and information inequalities. Electronic Communications in Probability, 18:7:1–7:14, 2013. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/2578.

Lie:2018:QIC

[LS18a] Han Cheng Lie and T. J. Sullivan. Quasi-invariance of countable products of Cauchy measures under non-unitary dilations. Electronic Communications in Probability, 23(??):8:1–8:6, ???? 2018. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1519182083.

Lowe:2018:FBS

[LS18b] Matthias Löwe and Kristina Schubert. Fluctuations for block spin Ising models. Electronic Communications in Probability, 23 (??):53:1–53:12, ???? 2018. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1535767264.

Leobacher:2015:ESC

[LST15] Gunther Leobacher, Michaela Szölgyenyi, and Stefan Thonhauser. On the existence of solutions of a class of SDEs with discontinuous drift and singular diffusion. Electronic Communications in Probability, 20(??):6:1–6:14, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3149.

Li:1999:RPF

[LSY99] Zenghu Li, Tokuzo Shiga, and Lihua Yao. A reversibility problem for Fleming–Viot processes. Electronic Communications in Probability, 4:9:65–9:76, 1999. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1007.

Liu:2011:EUI

[LT11] Wei Liu and Jonas Toelle. Existence and uniqueness of invariant measures for stochastic evolution equations with weakly dissipative drifts. Electronic Communications in Probability, 16: 40:447–40:457, 2011. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1643.

Lalley:2020:ODE

[LT20] Steven P. Lalley and Si Tang. Occupation densities of ensembles of branching random walks. Electronic Communications in Probability, 25(??):12:1–12:13, ???? 2020. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1580461331.

Ledger:2018:RWR

[LTV18] Sean Ledger, Bálint Tóth, and Benedek Valkó. Random walk on the randomly-oriented Manhattan lattice. Electronic Communications in Probability, 23(??):43:1–43:11, ???? 2018. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
euclid.ecp/1532505674.

Luo:2014:UDF

[Luo14] Dejun Luo. Uniqueness of degenerate Fokker–Planck equations with weakly differentiable drift whose gradient is given by a singular integral. Electronic Communications in Probability, 19: 42:1–42:14, 2014. CODEN ???? ISSN 1083-589X. URL http:/

/ecp.ejpecp.org/article/view/3407.

Lupu:2016:LPD

[Lup16] Titus Lupu. Loop percolation on discrete half-plane. Electronic Communications in Probability, 21(??):30:1–30:9, ???? 2016. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1459878297.

Lifshits:2020:CSS

[LV20] Mikhail Lifshits and Vladislav Vysotsky. On the completion of Skorokhod space. Electronic Communications in Probability, 25 (??):1–10, 2020. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-25/issue-none/On-the-completion-
of-Skorokhod-space/10.1214/20-ECP346.full.

Liao:2005:MRB

[LW05] Ming Liao and Longmin Wang. Motion of a rigid body under random perturbation. Electronic Communications in Probability, 10:24:235–24:243, 2005. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1163.

Liu:2009:IRF

[LW09] Wei Liu and Liming Wu. Identification of the rate function for large deviations of an irreducible Markov chain. Electronic Communications in Probability, 14:52:540–52:551, 2009. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1512.

Lacker:2015:TIM

[LW15] Daniel Lacker and Kevin Webster. Translation invariant mean field games with common noise. Electronic Communications in Probability, 20(??):42:1–42:13, ???? 2015. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/3822.

Lupu:2016:NIR

[LW16] Titus Lupu and Wendelin Werner. A note on Ising random currents, Ising-FK, loop-soups and the Gaussian free field. Electronic Communications in Probability, 21(??):13:1–13:7, ???? 2016. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1455717135.

Li:2015:FTB

[LX15] Fang Li and Lihu Xu. Finite time blowup of the stochastic shadow Gierer–Meinhardt system. Electronic Communications in Probability, 20(??):65:1–65:13, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/4298.

Lee:2019:LNE

[LX19] Cheuk Yin Lee and Yimin Xiao. Local nondeterminism and the exact modulus of continuity for stochastic wave equation. Electronic Communications in Probability, 24(??):52:1–52:8, ???? 2019. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1568253714.

Liu:2022:GSS

[LX22] Qingwei Liu and Aihua Xia. Geometric sums, size biasing and zero biasing. Electronic Communications in Probability, 27(??): 1–13, 2022. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-27/issue-none/Geometric-sums-size-
biasing-and-zero-biasing/10.1214/22-ECP462.full.

Liu:2013:ASS

[LY13] Hanbing Liu and Juan Yang. An approximation scheme of stochastic Stokes equations. Electronic Communications in Probability, 18:21:1–21:10, 2013. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/2374.

Lyons:2016:RPI

[LY16] Terry Lyons and Danyu Yang. Recovering the pathwise Itô solution from averaged Stratonovich solutions. Electronic Communications in Probability, 21(??):7:1–7:18, ???? 2016. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1454682823.

Li:2022:LMD

[LY22] Yuqiang Li and Qiang Yao. Large and moderate deviations for record numbers in some non-nearest neighbor random walks. Electronic Communications in Probability, 27(??): 1–13, 2022. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-27/issue-none/Large-and-moderate-

deviations-for-record-numbers-in-some-nonnearest/10.
1214/22-ECP497.full.

Lyons:2018:NTT

[Lyo18] Russell Lyons. A note on tail triviality for determinantal point processes. Electronic Communications in Probability, 23(??): 72:1–72:3, ???? 2018. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1539763345.

Lyons:2019:EBM

[Lyo19] Russell Lyons. Exit boundaries of multidimensional SDEs. Electronic Communications in Probability, 24(??):24:1–24:2, ???? 2019. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1558339220.

Li:2023:ASC

[LZ23] Jingyu Li and Yong Zhang. Almost sure central limit theorems for stochastic wave equations. Electronic Communications in Probability, 28(??):1–12, 2023. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-28/issue-
none/Almost-sure-central-limit-theorems-for-stochastic-
wave-equations/10.1214/23-ECP517.full.

Marchal:2017:SGB

[MA17] Olivier Marchal and Julyan Arbel. On the sub-Gaussianity of the beta and Dirichlet distributions. Electronic Communications in Probability, 22(??):54:1–54:14, ???? 2017. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1507860211.

Machida:2002:FAA

[Mac02] Motoya Machida. Fill’s algorithm for absolutely continuous stochastically monotone kernels. Electronic Communications in Probability, 7:15:141–15:155, 2002. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1056.

Maillard:2013:NSP

[Mai13] Pascal Maillard. A note on stable point processes occurring in branching Brownian motion. Electronic Communications in Probability, 18:5:1–5:9, 2013. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/2390.

Major:2006:MVH

[Maj06] Peter Major. A multivariate version of Hoeffding’s inequality. Electronic Communications in Probability, 11:24:220–24:229, 2006. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1221.

Makhnin:2008:FPE

[Mak08] Oleg Makhnin. Filtering and parameter estimation for a jump stochastic process with discrete observations. Electronic Communications in Probability, 13:21:210–21:224, 2008. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1363.

Mallein:2015:MDD

[Mal15] Bastien Mallein. Maximal displacement in the d-dimensional branching Brownian motion. Electronic Communications in Probability, 20(??):76:1–76:12, ???? 2015. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/4216.

Malyshkin:2020:SPA

[Mal20] Yury Malyshkin. Sublinear preferential attachment combined with a growing number of choices. Electronic Communications in Probability, 25(??):1–12, 2020. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-25/issue-
none/Sublinear-preferential-attachment-combined-with-
a-growing-number-of-choices/10.1214/20-ECP368.full.

Manstavicius:2005:NMP

[Man05] Martynas Manstavicius. A non-Markovian process with unbounded p-variation. Electronic Communications in Probability, 10:3:17–3:28, 2005. CODEN ???? ISSN 1083-589X. URL http:

//ecp.ejpecp.org/article/view/1128.

Marchand:2014:RBP

[MAPS14] Éric Marchand, Djilali Ait Aoudia, François Perron, and Latifa Ben Hadj Slimene. On runs, bivariate Poisson mixtures and distributions that arise in Bernoulli arrays. Electronic Communications in Probability, 19:8:1–8:12, 2014. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/3152.

Marchal:1998:BBT

[Mar98] Philippe Marchal. The best bounds in a theorem of Russell Lyons. Electronic Communications in Probability, 3:11:91–11:94, 1998. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/997.

Marchal:1999:LER

[Mar99] Philippe Marchal. Loop-erased random walks, spanning trees and Hamiltonian cycles. Electronic Communications in Probability, 5:4:39–4:50, 1999. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1016.

Marchal:2005:MCS

[Mar05] Philippe Marchal. Measure concentration for stable laws with index close to 2. Electronic Communications in Probability, 10: 4:29–4:35, 2005. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1129.

Marchal:2009:STE

[Mar09] Philippe Marchal. Small time expansions for transition probabilities of some Lévy processes. Electronic Communications in Probability, 14:13:132–13:142, 2009. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1452.

Markstrom:2010:CPN

[Mar10] Klas Markström. Closure properties and negatively associated measures violating the van den Berg–Kesten inequality. Electronic Communications in Probability, 15:41:449–41:456, 2010. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1575.

Markowsky:2011:EET

[Mar11] Greg Markowsky. On the expected exit time of planar Brownian motion from simply connected domains. Electronic Communications in Probability, 16:57:652–57:663, 2011. CODEN

????  ISSN 1083-589X. URL http://ecp.ejpecp.org/
article/view/1653.

Marcus:2014:MGD

[Mar14] Michael B. Marcus. Multivariate gamma distributions. Electronic Communications in Probability, 19(??):86:1–86:10, ???? 2014. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/3794.

Martineau:2017:SCC

[Mar17] Sébastien Martineau. The set of connective constants of Cayley graphs contains a Cantor space. Electronic Communications in Probability, 22(??):12:1–12:4, ???? 2017. CODEN ????

ISSN 1083-589X. URL http://projecteuclid.org/euclid.
ecp/1485507643.

Marchina:2018:CIS

[Mar18] Antoine Marchina. Comparison inequalities for suprema of bounded empirical processes. Electronic Communications in Probability, 23(??):33:1–33:7, ???? 2018. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1528358638.

Martineau:2022:CPV

[Mar22] Sébastien Martineau. On coprime percolation, the visibility graphon, and the local limit of the GCD profile. Electronic Communications in Probability, 27(??):1–14, 2022. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
journals/electronic-communications-in-probability/volume-
27/issue-none/On-coprime-percolation-the-visibility-
graphon-and-the-local-limit/10.1214/21-ECP381.full.

Maslanka:2018:TAM

[Maś18] Mariusz Maślanka. Tail asymptotics of maximums on trees in the critical case. Electronic Communications in Probability, 23(??): 48:1–48:11, ???? 2018. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1532678499.

Maurelli:2020:NES

[Mau20] Mario Maurelli. Non-explosion by Stratonovich noise for ODEs. Electronic Communications in Probability, 25(??):1– 10, 2020. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-25/issue-none/Non-explosion-by-
Stratonovich-noise-for-ODEs/10.1214/20-ECP347.full.

McVinish:2008:OPE

[McV08] Ross McVinish. Optimising prediction error among completely monotone covariance sequences. Electronic Communications in Probability, 13:11:113–11:120, 2008. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1355.

McVinish:2022:DOP

[McV22] Ross McVinish. Dominating occupancy processes by the independent site approximation. Electronic Communications in Probability, 27(??):1–10, 2022. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-27/issue-
none/Dominating-occupancy-processes-by-the-independent-
site-approximation/10.1214/22-ECP499.full.

Mordecki:2021:TSO

[ME21] Ernesto Mordecki and Facundo Oliú Eguren. Two-sided optimal stopping for lévy processes. Electronic Communications in Probability, 26(??):1–12, 2021. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-26/issue-
none/Two-sided-optimal-stopping-for-L%c3%a9vy-processes/
10.1214/21-ECP376.full.

Meckes:2007:SNR

[Mec07] Mark Meckes. On the spectral norm of a random Toeplitz matrix. Electronic Communications in Probability, 12:31:315– 31:325, 2007. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1313.

Meckes:2009:QAG

[Mec09] Elizabeth Meckes. Quantitative asymptotics of graphical projection pursuit. Electronic Communications in Probability, 14: 17:176–17:185, 2009. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1457.

Menozzi:2011:PTM

[Men11] Stephane Menozzi. Parametrix techniques and martingale problems for some degenerate Kolmogorov equations. Electronic Communications in Probability, 16:23:234–23:250, 2011. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1619.

Mentemeier:2013:NKC

[Men13] Sebastian Mentemeier. A note on Kesten’s Choquet–Deny lemma. Electronic Communications in Probability, 18:65:1–65:7, 2013. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/2629. See erratum [Men14].

Mentemeier:2014:ENK

[Men14] Sebastian Mentemeier. Erratum: “A note on Kesten’s Choquet– Deny lemma”. Electronic Communications in Probability, 19: 35:1–35:2, 2014. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/3381. See [Men13].

Mendelson:2018:CNR

[Men18] Shahar Mendelson. Column normalization of a random measurement matrix. Electronic Communications in Probability, 23(??): 13:1–13:8, ???? 2018. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1519722243.

Mackey:2016:EMS

[MG16a] Lester Mackey and Jackson Gorham. Erratum: Multivariate Stein factors for a class of strongly log-concave distributions. Electronic Communications in Probability, 21(??):80:1– 80:2, ???? 2016. CODEN ???? ISSN 1083-589X. URL http://

projecteuclid.org/euclid.ecp/1480647625. See [MG16b].

Mackey:2016:MSF

[MG16b] Lester Mackey and Jackson Gorham. Multivariate Stein factors for a class of strongly log-concave distributions. Electronic Communications in Probability, 21(??):56:1–56:14, ???? 2016. CO-DEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1472830238. See erratum [MG16a].

Moon:2013:HBV

[MHC13] Hee-Jin Moon, Chang-Ho Han, and Yong-Kab Choi. How big are the l^{∞}-valued random fields? Electronic Communications in Probability, 18:61:1–61:9, 2013. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/2417.

Michna:2013:EFS

[Mic13] Zbigniew Michna. Explicit formula for the supremum distribution of a spectrally negative stable process. Electronic Communications in Probability, 18:10:1–10:6, 2013. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/2236.

Michelen:2019:CPI

[Mic19] Marcus Michelen. Critical percolation and the incipient infinite cluster on Galton–Watson trees. Electronic Communications in Probability, 24(??):10:1–10:13, ???? 2019. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1550480494.

Miermont:2008:SSL

[Mie08] Grégory Miermont. On the sphericity of scaling limits of random planar quadrangulations. Electronic Communications in Probability, 13:24:248–24:257, 2008. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1368.

Mikami:2002:OCA

[Mik02] Toshio Mikami. Optimal control for absolutely continuous stochastic processes and the mass transportation problem. Electronic Communications in Probability, 7:20:199–20:213, 2002. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1061.

Millan:2008:RGL

[Mil08] Juan Carlos Pardo Millan. On the rate of growth of Lévy processes with no positive jumps conditioned to stay positive. Electronic Communications in Probability, 13:47:494–47:506, 2008. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1414.

Mercier:2015:PMR

[MLV15] Sabine Mercier, Agnès Lagnoux, and Pierre Vallois. Probability that the maximum of the reflected Brownian motion over a finite interval [0, t] is achieved by its last zero before t. Electronic Communications in Probability, 20(??):62:1–62:9, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/4279.

Meckes:2013:SMP

[MM13] Elizabeth Meckes and Mark Meckes. Spectral measures of powers of random matrices. Electronic Communications in Probability, 18:78:1–78:13, 2013. CODEN ???? ISSN 1083-589X. URL http:

//ecp.ejpecp.org/article/view/2551.

Maillard:2021:BCE

[MM21] Pascal Maillard and Bastien Mallein. On the branching convolution equation E = Z ⊗ E. Electronic Communications in Probability, 26(??):1–12, 2021. CODEN ???? ISSN 1083-589X.

Mijatovic:2018:PSB

[MMB18] Aleksandar Mijatović, Veno Mramor, and Gerónimo Uribe Bravo. Projections of spherical Brownian motion. Electronic Communications in Probability, 23(??):52:1–52:12, ???? 2018. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1535767263.

Maas:2008:COF

[MN08] Jan Maas and Jan Neerven. A Clark–Ocone formula in UMD Banach spaces. Electronic Communications in Probability, 13: 15:151–15:164, 2008. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1361.

Maejima:2009:NNC

[MN09] Makoto Maejima and Genta Nakahara. A note on new classes of infinitely divisible distributions on R^{d}. Electronic Communications in Probability, 14:36:358–36:371, 2009. CODEN

????  ISSN 1083-589X. URL http://ecp.ejpecp.org/
article/view/1487.

Ma:2020:IPA

[MNX20] Nicholas Ma, David Nualart, and Panqiu Xia. Intermittency for the parabolic Anderson model of Skorohod type driven by a rough noise. Electronic Communications in Probability, 25(??): 48:1–48:10, ???? 2020. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1594713622.

Maples:2012:NCR

[MNZ12] Kenneth Maples, Ashkan Nikeghbali, and Dirk Zeindler. On the number of cycles in a random permutation. Electronic Communications in Probability, 17:20:1–20:13, 2012. CODEN

????  ISSN 1083-589X. URL http://ecp.ejpecp.org/
article/view/1934.

Mandler:2022:WFR

[MO22] Christian Mandler and Ludger Overbeck. The weak functional representation of historical martingales. Electronic Communications in Probability, 27(??):1–12, 2022. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-27/issue-
none/The-weak-functional-representation-of-historical-
martingales/10.1214/22-ECP492.full.

Mohle:2011:CPD

[Möh11] Martin Möhle. Coalescent processes derived from some compound Poisson population models. Electronic Communications in Probability, 16:50:567–50:582, 2011. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1654.

Mohle:2018:SDB

[Möh18] Martin Möhle. A spectral decomposition for the block counting process and the fixation line of the beta(3,1)-coalescent. Electronic Communications in Probability, 23(??):102:1–102:15, ???? 2018. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1545447666.

Mohle:2019:SDS

[Möh19] Martin Möhle. A spectral decomposition for a simple mutation model. Electronic Communications in Probability, 24(??):15:1– 15:14, ???? 2019. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1553133704.

Montenegro:2007:SEV

[Mon07] Ravi Montenegro. Sharp edge, vertex, and mixed Cheeger type inequalities for finite Markov kernels. Electronic Communications in Probability, 12:36:377–36:389, 2007. CODEN

????  ISSN 1083-589X. URL http://ecp.ejpecp.org/
article/view/1269.

Morandin:2005:RBP

[Mor05] Francesco Morandin. A resummed branching process representation for a class of nonlinear ODEs. Electronic Communications in Probability, 10:1:1–1:6, 2005. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1126.

Morris:2008:SGI

[Mor08] Ben Morris. Spectral gap for the interchange process in a box. Electronic Communications in Probability, 13:31:311– 31:318, 2008. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1381.

Marinucci:2013:MSC

[MP13] Domenico Marinucci and Giovanni Peccati. Mean-square continuity on homogeneous spaces of compact groups. Electronic Communications in Probability, 18:37:1–37:10, 2013. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2400.

Malyshkin:2014:PCC

[MP14a] Yury Malyshkin and Elliot Paquette. The power of choice combined with preferential attachment. Electronic Communications in Probability, 19:43:1–43:13, 2014. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/3461.

Mohle:2014:SDB

[MP14b] Martin Möhle and Helmut Pitters. A spectral decomposition for the block counting process of the Bolthausen–Sznitman coalescent. Electronic Communications in Probability, 19:47:1–47:11, 2014. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/3464.

Mijatovic:2016:JAD

[MP16] Aleksandar Mijatović and Martijn Pistorius. Joint asymptotic distribution of certain path functionals of the reflected process. Electronic Communications in Probability, 21(??):43:1– 43:18, ???? 2016. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1464033418.

Mariano:2020:CSE

[MP20] Phanuel Mariano and Hugo Panzo. Conformal Skorokhod embeddings and related extremal problems. Electronic Communications in Probability, 25(??):42:1–42:11, ???? 2020. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1592446014.

Michna:2015:DSS

[MPP15] Zbigniew Michna, Zbigniew Palmowski, and Martijn Pistorius. The distribution of the supremum for spectrally asymmetric Lévy processes. Electronic Communications in Probability, 20 (??):24:1–24:10, ???? 2015. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/2999.

Mogulskii:2014:LDE

[MPY14] Anatoli Mogulskii, Eugene Pechersky, and Anatoli Yambartsev. Large deviations for excursions of non-homogeneous Markov processes. Electronic Communications in Probability, 19:36:1–36:8, 2014. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/3289.

Marcus:2001:NID

[MR01] Michael Marcus and Jan Rosinski. L^{1}-norm of infinitely divisible random vectors and certain stochastic integrals. Electronic Communications in Probability, 6:2:15–2:29, 2001. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1031.

Marcus:2008:IDG

[MR08] Michael Marcus and Jay Rosen. Infinite divisibility of Gaussian squares with non-zero means. Electronic Communications in Probability, 13:36:364–36:376, 2008. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1389.

Merkl:2011:CIE

[MR11] Franz Merkl and Silke Rolles. Correlation inequalities for edgereinforced random walk. Electronic Communications in Probability, 16:66:753–66:763, 2011. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1683.

Meiners:2013:ANA

[MR13] Raphael Meiners and Anselm Reichenbachs. On the accuracy of the normal approximation for the free energy in the random energy model. Electronic Communications in Probability, 18: 12:1–12:11, 2013. CODEN ???? ISSN 1083-589X. URL http:/

/ecp.ejpecp.org/article/view/2377.

Marcus:2015:NSC

[MR15a] Michael B. Marcus and Jay S. Rosen. Necessary and sufficient conditions for the continuity of permanental processes associated with transient Lévy processes. Electronic Communications in Probability, 20(??):57:1–57:6, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/4183.

Morgado:2015:ETD

[MR15b] Leandro Batista Morgado and Paulo R. C. Ruffino. Extension of time for decomposition of stochastic flows in spaces with complementary foliations. Electronic Communications in Probability, 20 (??):38:1–38:9, ???? 2015. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/3762.

Marcus:2022:LUM

[MR22a] Michael B. Marcus and Jay Rosen. Local and uniform moduli of continuity of chi-square processes. Electronic Communications in Probability, 27(??):1–10, 2022. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-27/issue-
none/Local-and-uniform-moduli-of-continuity-of-chisquare-
processes/10.1214/22-ECP471.full.

Monmarche:2022:OLS

[MR22b] Pierre Monmarché and Mouad Ramil. Overdamped limit at stationarity for non-equilibrium Langevin diffusions. Electronic Communications in Probability, 27(??):1–8, 2022. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
journals/electronic-communications-in-probability/volume-
27/issue-none/Overdamped-limit-at-stationarity-for-non-
equilibrium-Langevin-diffusions/10.1214/22-ECP447.full.

Miranda:2011:GCL

[MS11] Yuri Mejia Miranda and Gordon Slade. The growth constants of lattice trees and lattice animals in high dimensions. Electronic Communications in Probability, 16:13:129–13:136, 2011. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1612.

Meckes:2019:ETR

[MS19a] Elizabeth Meckes and Kathryn Stewart. On the eigenvalues of truncations of random unitary matrices. Electronic Communications in Probability, 24(??):57:1–57:12, ???? 2019. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1568361883.

Mordecki:2019:OSO

[MS19b] Ernesto Mordecki and Paavo Salminen. Optimal stopping of oscillating Brownian motion. Electronic Communications in Probability, 24(??):50:1–50:12, ???? 2019. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1568253712.

Mustapha:2019:DHF

[MS19c] Sami Mustapha and Mohamed Sifi. Discrete harmonic functions in Lipschitz domains. Electronic Communications in Probability, 24(??):58:1–58:15, ???? 2019. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1568793625.

Maejima:2010:CMI

[MU10] Makoto Maejima and Yohei Ueda. Compositions of mappings of infinitely divisible distributions with applications to finding the limits of some nested subclasses. Electronic Communications in Probability, 15:21:227–21:239, 2010. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1557.

Mijatovic:2012:CIF

[MU12] Aleksandar Mijatovic and Mikhail Urusov. Convergence of integral functionals of one-dimensional diffusions. Electronic Communications in Probability, 17:61:1–61:13, 2012. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/1825.

Muirhead:2015:TSL

[Mui15] Stephen Muirhead. Two-site localisation in the bouchaud trap model with slowly varying traps. Electronic Communications in Probability, 20(??):25:1–25:15, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3723.

Muller:2008:RBM

[Mül08] Sebastian Müller. Recurrence for branching Markov chains. Electronic Communications in Probability, 13:54:576–54:605, 2008. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1424.

Marynych:2014:WCN

[MV14] Alexander Marynych and Glib Verovkin. Weak convergence of the number of zero increments in the random walk with barrier. Electronic Communications in Probability, 19(??):74:1– 74:11, ???? 2014. CODEN ???? ISSN 1083-589X. URL http:/

/ecp.ejpecp.org/article/view/3641.

Mueller:2009:CBS

[MW09] Carl Mueller and Zhixin Wu. A connection between the stochastic heat equation and fractional Brownian motion, and a simple proof of a result of Talagrand. Electronic Communications in Probability, 14:6:55–6:65, 2009. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1403. See erra-

tum [MW12].

Mueller:2012:ECB

[MW12] Carl Mueller and Zhixin Wu. Erratum: “A connection between the stochastic heat equation and fractional Brownian motion, and a simple proof of a result of Talagrand”. Electronic Communications in Probability, 17:8:1–8:10, 2012. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1774. See [MW09].

Molchanov:2016:CHL

[MW16] Ilya Molchanov and Florian Wespi. Convex hulls of Lévy processes. Electronic Communications in Probability, 21(??):69:1– 69:11, ???? 2016. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1475266872.

Ma:2011:TII

[MWW11] Yutao Ma, Ran Wang, and Liming Wu. Transportationinformation inequalities for continuum Gibbs measures. Electronic Communications in Probability, 16:52:600–52:613, 2011. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1670.

Mikulevicius:2019:SCM

[MX19] R. Mikulevičius and Fanhui Xu. On some càdlàguity moment estimates of stochastic jump processes. Electronic Communications in Probability, 24(??):77:1–77:11, ???? 2019. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1576811275.

Matsumoto:1999:SCP

[MY99] Hiroyuki Matsumoto and Marc Yor. Some changes of probabilities related to a geometric Brownian motion version of Pitman’s 2M − X theorem. Electronic Communications in Probability, 4: 3:15–3:23, 1999. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1001.

Madan:2012:MWI

[MY12] Dilip Madan and Marc Yor. Moments of Wiener integrals for subordinators. Electronic Communications in Probability, 17: 55:1–55:8, 2012. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/2206.

Masuda:2013:EEI

[MY13] Hiroki Masuda and Nakahiro Yoshida. Edgeworth expansion for the integrated Lévy driven Ornstein–Uhlenbeck process. Electronic Communications in Probability, 18:94:1–94:10, 2013. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2726.

McDonald:2020:EFS

[MY20] Curtis McDonald and Serdar Yüksel. Exponential filter stability via Dobrushin’s coefficient. Electronic Communications in Probability, 25(??):1–13, 2020. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-25/issue-
none/Exponential-filter-stability-via-Dobrushins-coefficient/
10.1214/20-ECP333.full.

Ma:2021:BBC

[MYZ21] Shaojuan Ma, Xu Yang, and Xiaowen Zhou. Boundary behaviors for a class of continuous-state nonlinear branching processes in critical cases. Electronic Communications in Probability, 26 (??):1–10, 2021. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-26/issue-none/Boundary-behaviors-
for-a-class-of-continuous-state-nonlinear-branching/
10.1214/21-ECP374.full.

Mason:2005:APW

[MZ05a] David Mason and Joel Zinn. Acknowledgment of priority: When Does a Randomly Weighted Self-normalized Sum Converge in Distribution? (Elect. Comm. in Probab. 10 (2005), 70–81). Electronic Communications in Probability, 10:30:297, 2005. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1170. See [MZ05b].

Mason:2005:WDR

[MZ05b] David Mason and Joel Zinn. When does a randomly weighted self-normalized sum converge in distribution? Electronic Communications in Probability, 10:8:70–8:81, 2005. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/

view/1135. See acknowledgment of priority [MZ05a].

Ma:2014:LSP

[MZ14a] Yutao Ma and Zhengliang Zhang. Logarithmic Sobolev and Poincaré inequalities for the circular Cauchy distribution. Electronic Communications in Probability, 19:10:1–10:9, 2014. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3071.

Mariani:2014:RVS

[MZ14b] Mauro Mariani and Lorenzo Zambotti. A renewal version of the Sanov theorem. Electronic Communications in Probability, 19 (??):69:1–69:13, ???? 2014. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/3325.

Magda:2018:MAR

[MZ18] Peligrad Magda and Na Zhang. Martingale approximations for random fields. Electronic Communications in Probability, 23(??): 28:1–28:9, ???? 2018. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1524881136.

Nakajima:2019:DNR

[Nak19] Shuta Nakajima. Divergence of non-random fluctuation in First Passage Percolation. Electronic Communications in Probability, 24(??):65:1–65:13, ???? 2019. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1572509099.

Neeman:2014:MVN

[Nee14] Joe Neeman. A multidimensional version of noise stability. Electronic Communications in Probability, 19(??):72:1–72:10, ???? 2014. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/3005.

Nehring:2014:CPP

[Neh14] Benjamin Nehring. A characterization of the Poisson process revisited. Electronic Communications in Probability, 19(??):68:1– 68:5, ???? 2014. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/3622.

Neunhauserer:2011:FEP

[Neu11] Jörg Neunhäuserer. A family of exceptional parameters for nonuniform self-similar measures. Electronic Communications in Probability, 16:19:192–19:199, 2011. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1618.

Nguyen:2022:ETF

[Ngu22] Nhu N. Nguyen. Exponential tightness of a family of Skorohod integrals. Electronic Communications in Probability, 27(??): 1–14, 2022. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-27/issue-none/Exponential-tightness-
of-a-family-of-Skorohod-integrals/10.1214/21-ECP442.
full.

Nicolas:2006:SSC

[Nic06] Fournier Nicolas. Standard stochastic coalescence with sum kernels. Electronic Communications in Probability, 11:14:141– 14:148, 2006. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1206.

Nie:2022:CIP

[Nie22] Florian Nie. The compact interface property for the stochastic heat equation with seed bank. Electronic Communications in Probability, 27(??):1–15, 2022. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-27/issue-
none/The-compact-interface-property-for-the-stochastic-
heat-equation-with/10.1214/22-ECP465.full.

Nikula:2020:SDL

[Nik20] Miika Nikula. Small deviations in lognormal Mandelbrot cascades. Electronic Communications in Probability, 25(??):7:1– 7:12, ???? 2020. CODEN ???? ISSN 1083-589X. URL https:/

/projecteuclid.org/euclid.ecp/1580202245.

Noba:2020:GSF

[Nob20] Kei Noba. Generalized scale functions of standard processes with no positive jumps. Electronic Communications in Probability, 25 (??):8:1–8:12, ???? 2020. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1580266866.

Nourdin:2011:YAP

[Nou11] Ivan Nourdin. Yet another proof of the Nualart–Peccati criterion. Electronic Communications in Probability, 16:42:467– 42:481, 2011. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1642.

Nachmias:2012:NAC

[NP12a] Asaf Nachmias and Yuval Peres. Non-amenable Cayley graphs of high girth have p_{c} < p_{u} and mean-field exponents. Electronic Communications in Probability, 17:57:1–57:8, 2012. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2139.

Nourdin:2012:CLS

[NP12b] Ivan Nourdin and Guillaume Poly. Convergence in law in the second Wiener/Wigner chaos. Electronic Communications in Probability, 17:36:1–36:12, 2012. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/2023. See erra-

tum [NP12c].

Nourdin:2012:ECL

[NP12c] Ivan Nourdin and Guillaume Poly. Erratum: Convergence in law in the second Wiener/Wigner chaos. Electronic Communications in Probability, 17:54:1–54:3, 2012. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/2383. See

[NP12b].

Nechita:2013:RPQ

[NP13] Ion Nechita and Clément Pellegrini. Random pure quantum states via unitary Brownian motion. Electronic Communications in Probability, 18:27:1–27:13, 2013. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/2426.

Najnudel:2020:FCC

[NP20] Joseph Najnudel and Jim Pitman. Feller coupling of cycles of permutations and Poisson spacings in inhomogeneous Bernoulli trials. Electronic Communications in Probability, 25(??):1– 11, 2020. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-25/issue-none/Feller-coupling-
of-cycles-of-permutations-and-Poisson-spacings-in/10.
1214/20-ECP352.full.

Nourdin:2019:BEB

[NPY19] Ivan Nourdin, Giovanni Peccati, and Xiaochuan Yang. Berry– Esseen bounds in the Breuer–Major CLT and Gebelein’s inequality. Electronic Communications in Probability, 24(??):34:1–34:12, ???? 2019. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1561169050.

Neuman:2018:FBM

[NR18] Eyal Neuman and Mathieu Rosenbaum. Fractional Brownian motion with zero Hurst parameter: a rough volatility viewpoint. Electronic Communications in Probability, 23(??):61:1– 61:12, ???? 2018. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1536718014.

Nagy:2021:OLT

[NR21] Lóránt Nagy and Miklós Rásonyi. Optimal long-term investment in illiquid markets when prices have negative memory. Electronic Communications in Probability, 26(??):1–12, 2021. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-26/issue-none/Optimal-long-term-
investment-in-illiquid-markets-when-prices-have/10.1214/
21-ECP387.full.

Nualart:2013:JCA

[NS13] David Nualart and Jason Swanson. Joint convergence along different subsequences of the signed cubic variation of fractional Brownian motion II. Electronic Communications in Probability, 18:81:1–81:11, 2013. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/2840.

Nualart:2018:MBS

[Nua18] Eulalia Nualart. Moment bounds for some fractional stochastic heat equations on the ball. Electronic Communications in Probability, 23(??):41:1–41:12, ???? 2018. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1532505672.

Nutz:2012:PCS

[Nut12] Marcel Nutz. Pathwise construction of stochastic integrals. Electronic Communications in Probability, 17:24:1–24:7, 2012. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2099.

Ng:2015:MDC

[NW15] Stephen Ng and Meg Walters. A method to derive concentration of measure bounds on Markov chains. Electronic Communications in Probability, 20(??):95:1–95:13, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3817.

Nualart:2013:CLT

[NX13] David Nualart and Fangjun Xu. Central limit theorem for an additive functional of the fractional Brownian motion II. Electronic Communications in Probability, 18:74:1–74:10, 2013. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2761.

Ni:2015:CEC

[NX15] Hao Ni and Weijun Xu. Concentration and exact convergence rates for expected Brownian signatures. Electronic Communications in Probability, 20(??):8:1–8:11, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3636.

Nikeghbali:2009:BFR

[NY09] Ashkan Nikeghbali and Marc Yor. The Barnes G function and its relations with sums and products of generalized Gamma convolution variables. Electronic Communications in Probability, 14:39:396–39:411, 2009. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1488.

Nagahata:2010:NDS

[NY10] Yukio Nagahata and Nobuo Yoshida. A note on the diffusive scaling limit for a class of linear systems. Electronic Communications in Probability, 15:7:68–7:78, 2010. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1530.

Nualart:2020:SES

[NZ20] David Nualart and Guangqu Zheng. Spatial ergodicity of stochastic wave equations in dimensions 1, 2 and 3. Electronic Communications in Probability, 25(??):1–11, 2020. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
journals/electronic-communications-in-probability/volume-
25/issue-none/Spatial-ergodicity-of-stochastic-wave-
equations-in-dimensions-1-2/10.1214/20-ECP361.full.

Ocafrain:2020:PRC

[Oça20] William Oçafrain. Polynomial rate of convergence to the Yaglom limit for Brownian motion with drift. Electronic Communications in Probability, 25(??):35:1–35:12, ???? 2020. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1588903421.

Orsingher:2012:PRF

[OD12] Enzo Orsingher and Mirko D’Ovidio. Probabilistic representation of fundamental solutions to \frac{∂u}{∂t} = κ_{m} \frac{∂^{m}u}{∂x^{m}}. Electronic Communications in Probability, 17:34:1–34:12, 2012. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1885.

Orenshtein:2016:ZOL

[OdS16] Tal Orenshtein and Renato Soares dos Santos. Zero-one law for directional transience of one-dimensional random walks in dynamic random environments. Electronic Communications in Probability, 21(??):15:1–15:11, ???? 2016. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1455897057.

Okada:2014:LZT

[Oka14] Izumi Okada. Last zero time or maximum time of the winding number of Brownian motions. Electronic Communications in Probability, 19(??):64:1–64:8, ???? 2014. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/3485.

Oliveira:2010:SRH

[Oli10a] Roberto Oliveira. Sums of random Hermitian matrices and an inequality by Rudelson. Electronic Communications in Probability, 15:19:203–19:212, 2010. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1544.

Olivier:2010:DIS

[Oli10b] Wintenberger Olivier. Deviation inequalities for sums of weakly dependent time series. Electronic Communications in Probability, 15:44:489–44:503, 2010. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1577.

Olofsson:2018:BOS

[Olo18] Marcus Olofsson. A Brownian optimal switching problem under incomplete information. Electronic Communications in Probability, 23(??):67:1–67:12, ???? 2018. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1537840941.

ORourke:2012:NMP

[O’R12] Sean O’Rourke. A note on the Marchenko–Pastur law for a class of random matrices with dependent entries. Electronic Communications in Probability, 17:28:1–28:13, 2012. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2020.

Oraby:2007:SLH

[Ora07] Tamer Oraby. The spectral laws of Hermitian block-matrices with large random blocks. Electronic Communications in Probability, 12:44:465–44:476, 2007. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1335.

Orenshtein:2021:RIP

[Ore21] Tal Orenshtein. Rough invariance principle for delayed regenerative processes. Electronic Communications in Probability, 26 (??):1–13, 2021. CODEN ???? ISSN 1083-589X.

Ouyang:2022:QSN

[ORV22] Cheng Ouyang and William Roberson-Vickery. Quasi-sure non-self-intersection for rough differential equations driven by fractional Brownian motion. Electronic Communications in Probability, 27(??):1–12, 2022. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-27/issue-
none/Quasi-sure-non-self-intersection-for-rough-differential-
equations-driven/10.1214/22-ECP454.full.

Ondrejat:2013:EPM

[OS13] Martin Ondrejat and Jan Seidler. On existence of progressively measurable modifications. Electronic Communications in Probability, 18:20:1–20:6, 2013. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/2548.

Orbanz:2016:BLG

[OS16] Peter Orbanz and Balazs Szegedy. Borel liftings of graph limits. Electronic Communications in Probability, 21(??):65:1–65:4, ???? 2016. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1473854581.

Osekowski:2008:SIB

[Ose08] Adam Osekowski. Sharp inequality for bounded submartingales and their differential subordinates. Electronic Communications in Probability, 13:61:660–61:675, 2008. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1433.

Osekowski:2009:SMI

[Ose09] Adam Osekowski. Sharp maximal inequality for martingales and stochastic integrals. Electronic Communications in Probability, 14:2:17–2:30, 2009. CODEN ???? ISSN 1083-589X. URL http:

//ecp.ejpecp.org/article/view/1438.

Osekowski:2010:STI

[Ose10] Adam Osekowski. Sharp tail inequalities for nonnegative submartingales and their strong differential subordinates. Electronic Communications in Probability, 15:46:508–46:521, 2010. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1582.

Osekowski:2011:RAD

[Ose11] Adam Osekowski. On relaxing the assumption of differential subordination in some martingale inequalities. Electronic Communications in Probability, 16:2:9–2:21, 2011. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/1593.

Osekowski:2014:MWT

[Ose14] Adam Osekowski. Maximal weak-type inequality for stochastic integrals. Electronic Communications in Probability, 19:24:1– 24:13, 2014. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/3151.

Osekowski:2016:WMI

[Osȩ16] Adam Osȩkowski. Weighted maximal inequality for differentially subordinate martingales. Electronic Communications in Probability, 21(??):23:1–23:10, ???? 2016. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1457617914.

Osekowski:2020:RIN

[Ose20] Adam Osekowski. A ratio inequality for nonnegative martingales and their differential subordinates. Electronic Communications in Probability, 25(??):21:1–21:12, ???? 2020. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
euclid.ecp/1582945214.

Ostrovsky:2013:TBB

[Ost13] Dmitry Ostrovsky. Theory of Barnes Beta distributions. Electronic Communications in Probability, 18:59:1–59:16, 2013. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2445.

Ostrovsky:2014:NDR

[Ost14] Dmitry Ostrovsky. A note on the S_{2}(δ) distribution and the Riemann Xi function. Electronic Communications in Probability, 19(??):85:1–85:13, ???? 2014. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/3608.

Otobe:2009:TGD

[Oto09] Yoshiki Otobe. A type of Gauss’ divergence formula on Wiener spaces. Electronic Communications in Probability, 14:44:457– 44:463, 2009. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1498.

Ouimet:2018:PDS

[Oui18] Frédéric Ouimet. Poisson–Dirichlet statistics for the extremes of a randomized Riemann zeta function. Electronic Communications in Probability, 23(??):46:1–46:15, ???? 2018. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1532657018.

ODonnell:2013:SKS

[OW13] Ryan O’Donnell and Karl Wimmer. Sharpness of KKL on Schreier graphs. Electronic Communications in Probability, 18: 18:1–18:12, 2013. CODEN ???? ISSN 1083-589X. URL http:/

/ecp.ejpecp.org/article/view/1961.

Osekowski:2022:WTE

[OW22] Adam Osekowski and Mateusz Wojtas. Weak-type estimates for martingale maximal functions. Electronic Communications in Probability, 27(??):1–11, 2022. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-27/issue-
none/Weak-type-estimates-for-martingale-maximal-functions/
10.1214/22-ECP494.full.

Owo:2015:RBS

[Owo15] Jean-Marc Owo. Reflected backward stochastic differential equations driven by countable Brownian motions with continuous coefficients. Electronic Communications in Probability, 20(??): 26:1–26:11, ???? 2015. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/3771.

OConnell:2001:RNC

[OY01] Neil O’Connell and Marc Yor. A representation for non-colliding random walks. Electronic Communications in Probability, 7:1:1– 1:12, 2001. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1042.

Oz:2020:VSB

[Öz20] Mehmet Öz. On the volume of the shrinking branching Brownian sausage. Electronic Communications in Probability, 25(??):37:1– 37:12, ???? 2020. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1589335619.

Ozawa:2020:EPC

[Oza20] Narutaka Ozawa. An entropic proof of cutoff on Ramanujan graphs. Electronic Communications in Probability, 25(??):1– 8, 2020. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-25/issue-none/An-entropic-proof-
of-cutoff-on-Ramanujan-graphs/10.1214/20-ECP358.full.

Pal:2008:SB

[Pal08] Soumik Pal. Symmetrization of Bernoulli. Electronic Communications in Probability, 13:19:194–19:197, 2008. CODEN

????  ISSN 1083-589X. URL http://ecp.ejpecp.org/
article/view/1364.

Panchenko:2001:NTC

[Pan01] Dmitriy Panchenko. A note on Talagrand’s concentration inequality. Electronic Communications in Probability, 6:5:55–5:65, 2001. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1034.

Panchenko:2002:SEI

[Pan02] Dmitriy Panchenko. Some extensions of an inequality of Vapnik and Chervonenkis. Electronic Communications in Probability, 7: 6:55–6:65, 2002. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1047.

Panchenko:2005:QAP

[Pan05] Dmitriy Panchenko. A question about the Parisi functional. Electronic Communications in Probability, 10:16:155–16:166, 2005. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1145.

Panchenko:2007:NTP

[Pan07] Dmitriy Panchenko. A note on Talagrand’s positivity principle. Electronic Communications in Probability, 12:38:401–38:410, 2007. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1326.

Panchenko:2008:DPF

[Pan08] Dmitry Panchenko. On differentiability of the Parisi formula. Electronic Communications in Probability, 13:23:241– 23:247, 2008. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1365.

Panchenko:2010:DSR

[Pan10] Dmitry Panchenko. On the Dovbysh–Sudakov representation result. Electronic Communications in Probability, 15:31:330– 31:338, 2010. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1562.

Panzo:2021:IFL

[Pan21] Hugo Panzo. Independent factorization of the last zero arcsine law for Bessel processes with drift. Electronic Communications in Probability, 26(??):1–11, 2021. CODEN ???? ISSN 1083-589X.

Parczewski:2017:DTT

[Par17] Peter Parczewski. Donsker-type theorems for correlated geometric fractional Brownian motions and related processes. Electronic Communications in Probability, 22(??):55:1–55:13, ???? 2017. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1507860212.

Patie:2007:TSE

[Pat07] Pierre Patie. Two-sided exit problem for a spectrally negative αstable Ornstein–Uhlenbeck process and the Wright’s generalized hypergeometric functions. Electronic Communications in Probability, 12:16:146–16:160, 2007. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1265.

Peccati:2004:WCO

[Pec04] Giovanni Peccati. Weak convergence to Ocone martingales: a remark. Electronic Communications in Probability, 9:18:172– 18:174, 2004. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1121. See [vZ02].

Peccati:2007:GAM

[Pec07] Giovanni Peccati. Gaussian approximations of multiple integrals. Electronic Communications in Probability, 12:34:350– 34:364, 2007. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1322.

Peche:2019:NPW

[Péc19] S. Péché. A note on the Pennington–Worah distribution. Electronic Communications in Probability, 24(??):66:1–66:7, ???? 2019. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1572509100.

Peligrad:2020:CAF

[Pel20] Magda Peligrad. On the CLT for additive functionals of Markov chains. Electronic Communications in Probability, 25(??):40:1– 40:10, ???? 2020. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1592272818.

Peligrad:2023:CSM

[Pel23] Magda Peligrad. On the CLT for stationary Markov chains with trivial tail sigma field. Electronic Communications in Probability, 28(??):1–7, 2023. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-28/issue-none/On-the-CLT-for-stationary-
Markov-chains-with-trivial-tail/10.1214/23-ECP509.full.

Penrose:2018:NTV

[Pen18] Mathew D. Penrose. Non-triviality of the vacancy phase transition for the Boolean model. Electronic Communications in Probability, 23(??):49:1–49:8, ???? 2018. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1533002443.

Penrose:2022:GCS

[Pen22] Mathew D. Penrose. Giant component of the soft random geometric graph. Electronic Communications in Probability, 27(??):

1–10, 2022. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-27/issue-none/Giant-component-
of-the-soft-random-geometric-graph/10.1214/22-ECP491.
full.

Perruchaud:2022:KDB

[Per22] Pierre Perruchaud. Kinetic Dyson Brownian motion. Electronic Communications in Probability, 27(??):1–12, 2022. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
journals/electronic-communications-in-probability/volume-
27/issue-none/Kinetic-Dyson-Brownian-motion/10.1214/
22-ECP480.full.

Peskir:2008:LHT

[Pes08] Goran Peskir. The law of the hitting times to points by a stable Lévy process with no negative jumps. Electronic Communications in Probability, 13:60:653–60:659, 2008. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1431.

Pete:2008:NPI

[Pet08] Gabor Pete. A note on percolation on Z^{d}: isoperimetric profile via exponential cluster repulsion. Electronic Communications in Probability, 13:37:377–37:392, 2008. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1390.

Petrov:2010:RSP

[Pet10] Leonid Petrov. Random strict partitions and determinantal point processes. Electronic Communications in Probability, 15:16:162– 16:175, 2010. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1542.

Peterson:2015:STO

[Pet15] Jonathon Peterson. Strong transience of one-dimensional random walk in a random environment. Electronic Communications in Probability, 20(??):67:1–67:10, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/4352.

Pilipenko:2017:FLT

[Pil17] Andrey Pilipenko. A functional limit theorem for excited random walks. Electronic Communications in Probability, 22(??):39:1– 39:9, ???? 2017. CODEN ???? ISSN 1083-589X. URL https:/

/projecteuclid.org/euclid.ecp/1502244193.

Pimentel:2006:TCC

[Pim06] Leandro Pimentel. The time constant and critical probabilities in percolation models. Electronic Communications in Probability, 11:16:160–16:167, 2006. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1210.

Pinelis:2016:MSI

[Pin16] Iosif Pinelis. On a multidimensional spherically invariant extension of the Rademacher–Gaussian comparison. Electronic Communications in Probability, 21(??):67:1–67:5, ???? 2016. CO-DEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1474301119.

Pinsky:2017:SCB

[Pin17] Ross G. Pinsky. Some connections between permutation cycles and Touchard polynomials and between permutations that fix a set and covers of multisets. Electronic Communications in Probability, 22(??):17:1–17:9, ???? 2017. CODEN ????

ISSN 1083-589X. URL http://projecteuclid.org/euclid.
ecp/1487386905.

Pinsky:2019:KCO

[Pin19] Ross Pinsky. Kemeny’s constant for one-dimensional diffusions. Electronic Communications in Probability, 24(??):36:1– 36:5, ???? 2019. CODEN ???? ISSN 1083-589X. URL https:/

/projecteuclid.org/euclid.ecp/1561169052.

Pogany:2015:RD

[PN15] Tibor K Pogany and Saralees Nadarajah. On the result of Doney. Electronic Communications in Probability, 20(??):58:1–58:4, ???? 2015. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/4081.

Pogany:2016:APR

[PN16] Tibor K. Pogány and Saralees Nadarajah. Acknowledgement of priority: On the result of Doney. Electronic Communications in Probability, 21(??):32:1, ???? 2016. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1460643503.

Polaczyk:2019:CES

[Pol19] Bartłomiej Polaczyk. Concentration of the empirical spectral distribution of random matrices with dependent entries. Electronic Communications in Probability, 24(??):78:1–78:15, ???? 2019. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1576897461.

Posfai:2009:EMC

[Pos09] Anna Posfai. An extension of Mineka’s coupling inequality. Electronic Communications in Probability, 14:45:464–45:473, 2009. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1501.

Pavlyukevich:2020:GPP

[PP20] Ilya Pavlyukevich and Andrey Pilipenko. Generalized Peano problem with Lévy noise. Electronic Communications in Probability, 25(??):1–14, 2020. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-25/issue-
none/Generalized-Peano-problem-with-L%c3%a9vy-noise/
10.1214/20-ECP365.full.

Peres:2011:RTE

[PR11] Yuval Peres and Sebastien Roch. Reconstruction on trees: Exponential moment bounds for linear estimators. Electronic Communications in Probability, 16:24:251–24:261, 2011. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1630.

Perkowski:2012:CM

[PR12a] Nicolas Perkowski and Johannes Ruf. Conditioned martingales. Electronic Communications in Probability, 17:48:1–48:12, 2012. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1955.

Procaccia:2012:CEI

[PR12b] Eviatar Procaccia and Ron Rosenthal. Concentration estimates for the isoperimetric constant of the supercritical percolation cluster. Electronic Communications in Probability, 17:30:1– 30:11, 2012. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/2185.

Pitman:2015:BGT

[PR15] Jim Pitman and Miklos Z. Racz. Beta-gamma tail asymptotics. Electronic Communications in Probability, 20(??):84:1–84:7, ???? 2015. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/4545.

Prokaj:2018:LMD

[PR18] Vilmos Prokaj and Johannes Ruf. Local martingales in discrete time. Electronic Communications in Probability, 23(??):31:1– 31:11, ???? 2018. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1525312855.

Pratelli:2019:ECP

[PR19] Luca Pratelli and Pietro Rigo. On the existence of continuous processes with given one-dimensional distributions. Electronic Communications in Probability, 24(??):46:1–46:9, ???? 2019. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1566957630.

Privault:2009:MIS

[Pri09] Nicolas Privault. Moment identities for Skorohod integrals on the Wiener space and applications. Electronic Communications in Probability, 14:11:116–11:121, 2009. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1450.

Privault:2015:SAI

[Pri15] Nicolas Privault. Stein approximation for Itô and Skorohod integrals by Edgeworth type expansions. Electronic Communications in Probability, 20(??):35:1–35:10, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3827.

Price:2017:IHK

[Pri17] Thomas McMurray Price. An inequality for the heat kernel on an Abelian Cayley graph. Electronic Communications in Probability, 22(??):57:1–57:8, ???? 2017. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1508292095.

Profeta:2018:SPS

[Pro18] Christophe Profeta. On the supremum of products of symmetric stable processes. Electronic Communications in Probability, 23 (??):97:1–97:13, ???? 2018. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1545102493.

Possamai:2013:RSM

[PRT13] Dylan Possamaï, Guillaume Royer, and Nizar Touzi. On the robust superhedging of measurable claims. Electronic Communications in Probability, 18:95:1–95:13, 2013. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/2739.

Peche:2008:LBS

[PS08] Sandrine Peche and Alexander Soshnikov. On the lower bound of the spectral norm of symmetric random matrices with independent entries. Electronic Communications in Probability, 13: 28:280–28:290, 2008. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1376.

Polito:2016:GSF

[PS16] Federico Polito and Enrico Scalas. A generalization of the spacefractional Poisson process and its connection to some Lévy processes. Electronic Communications in Probability, 21(??):20:1– 20:14, ???? 2016. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1456840981.

Podder:2017:GWP

[PS17] Moumanti Podder and Joel Spencer. Galton–Watson probability contraction. Electronic Communications in Probability, 22(??): 20:1–20:16, ???? 2017. CODEN ???? ISSN 1083-589X. URL

http://projecteuclid.org/euclid.ecp/1489651213.

Pal:2019:NTC

[PS19] Soumik Pal and Andrey Sarantsev. A note on transportation cost inequalities for diffusions with reflections. Electronic Communications in Probability, 24(??):21:1–21:11, ???? 2019. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
euclid.ecp/1554429763.

Peres:2020:ARD

[PS20] Yuval Peres and Andrew Swan. Analyticity for rapidly determined properties of Poisson Galton–Watson trees. Electronic Communications in Probability, 25(??):43:1–43:8, ???? 2020. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1592553612.

Peled:2013:GLF

[PSY13] Ron Peled, Wojciech Samotij, and Amir Yehudayoff. Grounded Lipschitz functions on trees are typically flat. Electronic Communications in Probability, 18:55:1–55:9, 2013. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/2796.

Procaccia:2011:GRI

[PT11] Eviatar Procaccia and Johan Tykesson. Geometry of the random interlacement. Electronic Communications in Probability, 16:47:528–47:544, 2011. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1660.

Puhalskii:2019:LDL

[Puh19] Anatolii Puhalskii. Large deviations of the long term distribution of a non Markov process. Electronic Communications in Probability, 24(??):35:1–35:11, ???? 2019. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1561169051.

Popov:2005:RWA

[PV05] Serguei Popov and Marina Vachkovskaia. Random walk attracted by percolation clusters. Electronic Communications in Probability, 10:27:263–27:272, 2005. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1167.

Puckette:1996:SCD

[PW96] Emily E. Puckette and Wendelin Werner. Simulations and conjectures for disconnection exponents. Electronic Communications in Probability, 1:6:49–6:64, 1996. CODEN ???? ISSN

1083-589X. URL http://ecp.ejpecp.org/article/view/976.

Pardoux:2011:BML

[PW11] Etienne Pardoux and Anton Wakolbinger. From Brownian motion with a local time drift to Feller’s branching diffusion with logistic growth. Electronic Communications in Probability, 16: 63:720–63:731, 2011. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1679.

Pitman:2018:SBP

[PW18] Jim Pitman and Matthias Winkel. Squared Bessel processes of positive and negative dimension embedded in Brownian local times. Electronic Communications in Probability, 23(??):74:1– 74:13, ???? 2018. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1539763347.

Pitman:2017:ETP

[PY17] Jim Pitman and Yuri Yakubovich. An ergodic theorem for partially exchangeable random partitions. Electronic Communications in Probability, 22(??):64:1–64:10, ???? 2017. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1511406072.

Pyc:2016:TDH

[PŻ16] Andrzej Pyć and Tomasz Żak. Transition density of a hyperbolic Bessel process. Electronic Communications in Probability, 21 (??):50:1–50:10, ???? 2016. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1469557024.

Procaccia:2018:CPD

[PZ18] Eviatar B. Procaccia and Yuan Zhang. On covering paths with 3 dimensional random walk. Electronic Communications in Probability, 23(??):57:1–57:11, ???? 2018. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1536718010.

Qin:2017:IBM

[QM17] Chuan Qin and Ben Morris. Improved bounds for the mixing time of the random-to-random shuffle. Electronic Communications in Probability, 22(??):22:1–22:7, ???? 2017. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1491271396.

Quastel:2011:LBP

[QR11] Jeremy Quastel and Daniel Remenik. Local Brownian property of the narrow wedge solution of the KPZ equation. Electronic Communications in Probability, 16:62:712–62:719, 2011. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1678.

Rassoul-Agha:2005:ZOL

[RA05] Firas Rassoul-Agha. On the zero–one law and the law of large numbers for random walk in mixing random environment. Electronic Communications in Probability, 10:5:36–5:44, 2005. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1130.

Ramadas:2014:MNV

[Ram14] Harishchandra Ramadas. Mixing of the noisy voter model. Electronic Communications in Probability, 19:17:1–17:17, 2014. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2968.

Rao:2019:HIM

[Rao19] Shravas Rao. A Hoeffding inequality for Markov chains. Electronic Communications in Probability, 24(??):14:1–14:11, ???? 2019. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1553133703.

Rao:2021:LOP

[Rao21] Shravas Rao. The Littlewood–Offord problem for Markov chains. Electronic Communications in Probability, 26(??):1–11, 2021. CODEN ???? ISSN 1083-589X.

Raschel:2010:GFM

[Ras10] Kilian Raschel. Green functions and Martin compactification for killed random walks related to SU(3). Electronic Communications in Probability, 15:17:176–17:190, 2010. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1543.

Rath:2015:SPP

[Rát15] Balázs Ráth. A short proof of the phase transition for the vacant set of random interlacements. Electronic Communications in Probability, 20(??):3:1–3:11, ???? 2015. CODEN

????  ISSN 1083-589X. URL http://ecp.ejpecp.org/
article/view/3734.

Rath:2018:MGF

[Rát18] Balázs Ráth. A moment-generating formula for Erdős–Rényi component sizes. Electronic Communications in Probability, 23 (??):24:1–24:14, ???? 2018. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1524708114.

Rey-Bellet:2015:VRI

[RBS15] Luc Rey-Bellet and Konstantinos Spiliopoulos. Variance reduction for irreversible Langevin samplers and diffusion on graphs.

Electronic Communications in Probability, 20(??):15:1–15:16, ???? 2015. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/3855.

Reitzner:2013:PPP

[Rei13] Matthias Reitzner. Poisson point processes: large deviation inequalities for the convex distance. Electronic Communications in Probability, 18:96:1–96:7, 2013. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/2851.

Ren:2022:OPP

[Ren22] Panpan Ren. Order preservation and positive correlation for nonlinear Fokker–Planck equation. Electronic Communications in Probability, 27(??):1–12, 2022. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-27/issue-
none/Order-preservation-and-positive-correlation-for-
nonlinear-Fokker--Planck-equation/10.1214/22-ECP466.
full.

Ressel:2001:SAU

[Res01] Paul Ressel. Subdiagonal and almost uniform distributions. Electronic Communications in Probability, 7:10:97–10:100, 2001. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1051.

Revelle:2003:HKA

[Rev03] David Revelle. Heat kernel asymptotics on the lamplighter group. Electronic Communications in Probability, 8:16:142– 16:154, 2003. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1092.

Reygner:2015:CSD

[Rey15] Julien Reygner. Chaoticity of the stationary distribution of rankbased interacting diffusions. Electronic Communications in Probability, 20(??):60:1–60:20, ???? 2015. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/4063.

Riedle:2018:SCL

[Rie18] Markus Riedle. Stable cylindrical Lévy processes and the stochastic Cauchy problem. Electronic Communications in Probability, 23(??):36:1–36:12, ???? 2018. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1528358641.

Rincon:1998:EDD

[Rin98] L. A. Rincon. Estimates for the derivative of diffusion semigroups. Electronic Communications in Probability, 3:8:65–8:74, 1998. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/994.

Rio:2011:ACM

[Rio11] Emmanuel Rio. Asymptotic constants for minimal distance in the central limit theorem. Electronic Communications in Probability, 16:9:96–9:103, 2011. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1609.

Rio:2013:EHA

[Rio13a] Emmanuel Rio. Extensions of the Hoeffding–Azuma inequalities. Electronic Communications in Probability, 18:54:1–54:6, 2013. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/2690.

Rio:2013:MCI

[Rio13b] Emmanuel Rio. On McDiarmid’s concentration inequality. Electronic Communications in Probability, 18:44:1–44:11, 2013. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2659.

Rio:2015:EIW

[Rio15] Emmanuel Rio. Exponential inequalities for weighted sums of bounded random variables. Electronic Communications in Probability, 20(??):77:1–77:10, ???? 2015. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/4204.

Rio:2017:ACF

[Rio17] Emmanuel Rio. About the constants in the Fuk–Nagaev inequalities. Electronic Communications in Probability, 22(??): 28:1–28:12, ???? 2017. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1494921750.

Rio:2018:ADI

[Rio18] Emmanuel Rio. About Doob’s inequality, entropy and Tchebichef. Electronic Communications in Probability, 23(??):

78:1–78:12, ???? 2018. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1540346602.

Raad:2020:SHP

[RL20] Mads Bonde Raad and Eva Löcherbach. Stability for Hawkes processes with inhibition. Electronic Communications in Probability, 25(??):33:1–33:9, ???? 2020. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1588039621.

Ruffino:2016:GSD

[RM16] Paulo R. Ruffino and Leandro Morgado. Geometry of stochastic delay differential equations with jumps in manifolds. Electronic Communications in Probability, 21(??):37:1–37:9, ???? 2016. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1461868666.

Roch:2005:BFM

[Roc05] Sébastien Roch. Bounding fastest mixing. Electronic Communications in Probability, 10:29:282–29:296, 2005. CODEN

????  ISSN 1083-589X. URL http://ecp.ejpecp.org/
article/view/1169.

Roitershtein:2005:LSL

[Roi05] Alexander Roitershtein. A log-scale limit theorem for onedimensional random walks in random environments. Electronic Communications in Probability, 10:25:244–25:253, 2005. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1164.

Rokhlin:2007:MSP

[Rok07] Dmitry Rokhlin. Martingale selection problem and asset pricing in finite discrete time. Electronic Communications in Probability, 12:1:1–1:8, 2007. CODEN ???? ISSN 1083-589X. URL http:/

/ecp.ejpecp.org/article/view/1240.

Rokhlin:2014:SPM

[Rok14] Dmitry B. Rokhlin. Stochastic Perron’s method for optimal control problems with state constraints. Electronic Communications in Probability, 19(??):73:1–73:15, ???? 2014. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3616.

Rokhlin:2015:CLT

[Rok15] Dmitry B. Rokhlin. Central limit theorem under variance uncertainty. Electronic Communications in Probability, 20(??):66:1– 66:10, ???? 2015. CODEN ???? ISSN 1083-589X. URL http:/

/ecp.ejpecp.org/article/view/4341.

Rosenthal:2002:QCR

[Ros02] Jeffrey Rosenthal. Quantitative convergence rates of Markov chains: A simple account. Electronic Communications in Probability, 7:13:123–13:128, 2002. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1054.

Rossignol:2008:TPP

[Ros08] Raphaël Rossignol. Threshold phenomena on product spaces: Bkkkl revisited (once more). Electronic Communications in Probability, 13:4:35–4:44, 2008. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1344.

Rosati:2020:KRS

[Ros20] Tommaso Cornelis Rosati. Killed rough super-Brownian motion. Electronic Communications in Probability, 25(??):44:1– 44:12, ???? 2020. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1592877699.

Roberts:1997:GEH

[RR97] Gareth O. Roberts and Jeffrey S. Rosenthal. Geometric ergodicity and hybrid Markov chains. Electronic Communications in Probability, 2:2:13–2:25, 1997. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/981.

Rasonyi:2014:CTP

[RR14] Miklós Rásonyi and Andrea Meireles Rodrigues. Continuoustime portfolio optimisation for a behavioural investor with bounded utility on gains. Electronic Communications in Probability, 19:37:1–37:13, 2014. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/2990.

Rosenthal:2015:SBC

[RR15] Jeffrey S. Rosenthal and Peter Rosenthal. Spectral bounds for certain two-factor non-reversible MCMC algorithms. Electronic Communications in Probability, 20(??):91:1–91:10, ???? 2015.

CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/4528.

Ramirez:2011:HET

[RRZ11] Jose Ramirez, Brian Rider, and Ofer Zeitouni. Hard edge tail asymptotics. Electronic Communications in Probability, 16: 65:741–65:752, 2011. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1682.

Rueschendorf:2006:ETB

[RS06] Ludger Rueschendorf and Eva-Maria Schopp. Exponential tail bounds for max-recursive sequences. Electronic Communications in Probability, 11:28:266–28:277, 2006. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1227.

Rao:2007:MFR

[RS07] N. Raj Rao and Roland Speicher. Multiplication of free random variables and the S-transform: the case of vanishing mean. Electronic Communications in Probability, 12:24:248–24:258, 2007. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1274.

Raimond:2011:IDG

[RS11a] Olivier Raimond and Bruno Schapira. Internal DLA generated by cookie random walks on Z. Electronic Communications in Probability, 16:43:483–43:490, 2011. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1646.

Rath:2011:TRI

[RS11b] Balazs Rath and Artem Sapozhnikov. On the transience of random interlacements. Electronic Communications in Probability, 16:35:379–35:391, 2011. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1637.

Regev:2016:CMR

[RS16] Oded Regev and Igor Shinkar. A counterexample to monotonicity of relative mass in random walks. Electronic Communications in Probability, 21(??):8:1–8:8, ???? 2016. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1454682824.

Roy:2021:CDT

[RS21] Rishideep Roy and Kumarjit Saha. Coexistence in discrete time multi-type competing frog models. Electronic Communications in Probability, 26(??):1–9, 2021. CODEN ???? ISSN 1083-589X.

Ren:2018:SDC

[RSS18] Yan-Xia Ren, Renming Song, and Zhenyao Sun. A 2-spine decomposition of the critical Galton–Watson tree and a probabilistic proof of Yaglom’s theorem. Electronic Communications in Probability, 23(??):42:1–42:12, ???? 2018. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1532505673.

Rolla:2008:LPP

[RT08] Leonardo Rolla and Augusto Teixeira. Last passage percolation in macroscopically inhomogeneous media. Electronic Communications in Probability, 13:13:131–13:139, 2008. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1287.

Rudolf:2013:PHR

[RU13] Daniel Rudolf and Mario Ullrich. Positivity of hit-and-run and related algorithms. Electronic Communications in Probability, 18:49:1–49:8, 2013. CODEN ???? ISSN 1083-589X. URL http:

//ecp.ejpecp.org/article/view/2507.

Ruf:2015:MPC

[Ruf15a] Johannes Ruf. The martingale property in the context of stochastic differential equations. Electronic Communications in Probability, 20(??):34:1–34:10, ???? 2015. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/3449.

Ruffino:2015:AAP

[Ruf15b] Paulo R. C. Ruffino. Application of an averaging principle on foliated diffusions: topology of the leaves. Electronic Communications in Probability, 20(??):28:1–28:5, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3715.

Ruf:2017:PCL

[Ruf17] Johannes Ruf. Piecewise constant local martingales with bounded numbers of jumps. Electronic Communications in Probability, 22(??):31:1–31:5, ???? 2017. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1496282536.

Rudelson:2013:HWI

[RV13] Mark Rudelson and Roman Vershynin. Hanson–Wright inequality and sub-Gaussian concentration. Electronic Communications in Probability, 18:82:1–82:9, 2013. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/2865.

Rath:2017:TSV

[RV17] Balázs Ráth and Daniel Valesin. On the threshold of spread-out voter model percolation. Electronic Communications in Probability, 22(??):50:1–50:12, ???? 2017. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1507255233.

Rempala:2002:APS

[RW02] Grzegorz Rempala and Jacek Wesolowski. Asymptotics for products of sums and U-statistics. Electronic Communications in Probability, 7:5:47–5:54, 2002. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1046.

Ruggiero:2009:CRI

[RW09] Matteo Ruggiero and Stephen Walker. Countable representation for infinite dimensional diffusions derived from the twoparameter Poisson–Dirichlet process. Electronic Communications in Probability, 14:49:501–49:517, 2009. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/1508.

Roch:2021:SCR

[RW21] Sebastien Roch and Kun-Chieh Wang. Sufficient condition for root reconstruction by parsimony on binary trees with general weights. Electronic Communications in Probability, 26(??):1–13, 2021. CODEN ???? ISSN 1083-589X.

Rosenberg:2020:CLT

[RX20] Steven Rosenberg and Jie Xu. Central limit theorems on compact metric spaces. Electronic Communications in Probability, 25 (??):1–10, 2020. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-25/issue-none/Central-limit-theorems-
on-compact-metric-spaces/10.1214/20-ECP336.full.

Ryznar:1998:UUB

[RŻ98] Michał Ryznar and Tomasz Żak. Uniform upper bound for a stable measure of a small ball. Electronic Communications in Probability, 3:9:75–9:78, 1998. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/995.

Roberts:2013:IPR

[RZ13] Matthew Roberts and Lee Zhao. Increasing paths in regular trees. Electronic Communications in Probability, 18:87:1–87:10, 2013. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/2784.

Sabanis:2013:NTE

[Sab13] Sotirios Sabanis. A note on tamed Euler approximations. Electronic Communications in Probability, 18:47:1–47:10, 2013. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2824.

Sabot:2021:PLV

[Sab21] Christophe Sabot. Polynomial localization of the 2D-vertex reinforced jump process. Electronic Communications in Probability, 26(??):1–9, 2021. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-26/issue-none/Polynomial-localization-
of-the-{$2$D}-Vertex-Reinforced-Jump-Process/10.1214/
20-ECP356.full.

Saintier:2007:GST

[Sai07] Nicolas Saintier. A general stochastic target problem with jump diffusion and an application to a hedging problem for large investors. Electronic Communications in Probability, 12:12:106– 12:119, 2007. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1261.

Salez:2015:MPE

[Sal15] Justin Salez. The Mézard–Parisi equation for matchings in pseudo-dimension d > 1. Electronic Communications in Probability, 20(??):13:1–13:7, ???? 2015. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/3791.

Salazar:2022:BET

[Sal22] Mauricio Salazar. On a Berry–Esseen type limit theorem for Boolean convolution. Electronic Communications in Probability, 27(??):1–10, 2022. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-27/issue-
none/On-a-Berry--Esseen-type-limit-theorem-for-Boolean-
convolution/10.1214/22-ECP448.full.

Samee:2010:PSF

[Sam10] Farman Samee. On the principle of smooth fit for killed diffusions. Electronic Communications in Probability, 15:9:89–9:98, 2010. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1531.

Sandric:2013:TCC

[San13] Nikola Sandrić. A transience condition for a class of onedimensional symmetric Lévy processes. Electronic Communications in Probability, 18:71:1–71:13, 2013. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/2802.

Sapozhnikov:2010:UBE

[Sap10] Artem Sapozhnikov. Upper bound on the expected size of the intrinsic ball. Electronic Communications in Probability, 15: 28:297–28:298, 2010. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1553.

Sapozhnikov:2011:IIC

[Sap11] Artem Sapozhnikov. The incipient infinite cluster does not stochastically dominate the invasion percolation cluster in two dimensions. Electronic Communications in Probability, 16:68:775– 68:780, 2011. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1684.

Sarkar:2021:NLW

[Sar21] Sourav Sarkar. A note on the local weak limit of a sequence of expander graphs. Electronic Communications in Probability, 26 (??):1–6, 2021. CODEN ???? ISSN 1083-589X.

Saumard:2019:BLC

[Sau19] Adrien Saumard. Bi-log-concavity: some properties and some remarks towards a multi-dimensional extension. Electronic Communications in Probability, 24(??):61:1–61:8, ???? 2019. CO-DEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1569895736.

Savov:2014:RS

[Sav14] Mladen Svetoslavov Savov. On the range of subordinators. Electronic Communications in Probability, 19(??):84:1–84:10, ???? 2014. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/3629.

Spitzner:2007:AVF

[SB07] Dan Spitzner and Thomas Boucher. Asymptotic variance of functionals of discrete-time Markov chains via the Drazin inverse. Electronic Communications in Probability, 12:13:120– 13:133, 2007. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1262.

Sami:2015:DHF

[SBS15] Mustapha Sami, Aymen Bouaziz, and Mohamed Sifi. Discrete harmonic functions on an orthant in Z^{d}. Electronic Communications in Probability, 20(??):52:1–52:13, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/4249.

Sheu:2009:NBM

[SC09] Yuan-Chung Sheu and Yu-Ting Chen. A note on r-balayages of matrix-exponential Lévy processes. Electronic Communications in Probability, 14:16:165–16:175, 2009. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1456.

Schweinsberg:1999:NSC

[Sch99] Jason Schweinsberg. A necessary and sufficient condition for the lambda-coalescent to come down from infinity. Electronic Communications in Probability, 5:1:1–1:11, 1999. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1013.

Schramm:2001:PF

[Sch01] Oded Schramm. A percolation formula. Electronic Communications in Probability, 6:12:115–12:120, 2001. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/1041.

Schlemm:2009:FPP

[Sch09] Eckhard Schlemm. First-passage percolation on width-two stretches with exponential link weights. Electronic Communications in Probability, 14:41:424–41:434, 2009. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/1493.

Schapira:2012:LVR

[Sch12] Bruno Schapira. A 0-1 law for vertex-reinforced random walks on Z with weight of order k^{α}, α ∈ [0, 1/2). Electronic Communications in Probability, 17:22:1–22:8, 2012. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/2084.

Schubert:2016:SDR

[Sch16] Kristina Schubert. Spectral density for random matrices with independent skew-diagonals. Electronic Communications in Probability, 21(??):40:1–40:12, ???? 2016. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1463081069.

Scheutzow:2020:CGC

[Sch20] Michael Scheutzow. Couplings, generalized couplings and uniqueness of invariant measures. Electronic Communications in Probability, 25(??):1–7, 2020. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-25/issue-
none/Couplings-generalized-couplings-and-uniqueness-
of-invariant-measures/10.1214/20-ECP363.full.

Sen:2016:LCS

[Sen16] Sanchayan Sen. On the largest component in the subcritical regime of the Bohman–Frieze process. Electronic Communications in Probability, 21(??):64:1–64:15, ???? 2016. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1473854580.

Sepanski:2003:LIL

[Sep03] Steven Sepanski. A law of the iterated logarithm for the sample covariance matrix. Electronic Communications in Probability, 8: 7:63–7:76, 2003. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1070.

Sava-Huss:2014:ETR

[SHH14] Ecaterina Sava-Huss and Wilfried Huss. Erratum: “Transience and recurrence of rotor-router walks on directed covers of graphs”. Electronic Communications in Probability, 19(??):

71:1–71:6, ???? 2014. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/3848.

Silvestri:2020:IDC

[Sil20] Vittoria Silvestri. Internal DLA on cylinder graphs: fluctuations and mixing. Electronic Communications in Probability, 25(??): 1–14, 2020. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-25/issue-none/Internal-DLA-on-
cylinder-graphs-fluctuations-and-mixing/10.1214/20-ECP339.
full.

Simon:2000:SME

[Sim00] Thomas Simon. Support of a Marcus equation in dimension 1. Electronic Communications in Probability, 5:16:149–16:157, 2000. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1028.

Simon:2011:MSP

[Sim11] Thomas Simon. A multiplicative short proof for the unimodality of stable densities. Electronic Communications in Probability, 16:54:623–54:629, 2011. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1672.

Simm:2017:RSP

[Sim17] Nick Simm. On the real spectrum of a product of Gaussian matrices. Electronic Communications in Probability, 22(??):41:1– 41:11, ???? 2017. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1502416902.

Singh:2014:RVR

[Sin14] Arvind Singh. Recurrence for vertex-reinforced random walks on Z with weak reinforcements. Electronic Communications in Probability, 19:14:1–14:6, 2014. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/3242.

Siorpaes:2014:DAP

[Sio14] Pietro Siorpaes. On a dyadic approximation of predictable processes of finite variation. Electronic Communications in Probability, 19:21:1–21:12, 2014. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/2972.

Sirbu:2014:NSF

[Sir14] Mihai Sirbu. A note on the strong formulation of stochastic control problems with model uncertainty. Electronic Communications in Probability, 19(??):81:1–81:10, ???? 2014. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3436.

Schmidt:2015:DRE

[SK15] Marius Alexander Schmidt and Nicola Kistler. From Derrida’s random energy model to branching random walks: from 1 to 3. Electronic Communications in Probability, 20(??):47:1–47:12, ???? 2015. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/4189.

Sokol:2013:ONT

[Sok13] Alexander Sokol. Optimal novikov-type criteria for local martingales with jumps. Electronic Communications in Probability, 18: 39:1–39:8, 2013. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/2312.

Soshnikov:2004:PSL

[Sos04] Alexander Soshnikov. Poisson statistics for the largest eigenvalues of Wigner random matrices with heavy tails. Electronic Communications in Probability, 9:9:82–9:91, 2004. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1112.

Sepanski:2000:WLL

[SP00] Steven Sepanski and Zhidong Pan. A weak law of large numbers for the sample covariance matrix. Electronic Communications in Probability, 5:8:73–8:76, 2000. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1020.

Spinu:2013:PLB

[Spi13] Florin Spinu. The probability law of the Brownian motion normalized by its range. Electronic Communications in Probability, 18:46:1–46:8, 2013. CODEN ???? ISSN 1083-589X. URL http:

//ecp.ejpecp.org/article/view/2568.

Spruill:2007:ADC

[Spr07] Marcus Spruill. Asymptotic distribution of coordinates on high dimensional spheres. Electronic Communications in Probability, 12:23:234–23:247, 2007. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1294.

Squillace:2021:DCC

[Squ21] Joseph Squillace. On the dependence of the component counting process of a uniform random variable. Electronic Communications in Probability, 26(??):1–12, 2021. CODEN ???? ISSN 1083-589X.

Steif:2006:SRP

[SS06] Jeffrey Steif and Aidan Sudbury. Some results for poisoning in a catalytic model. Electronic Communications in Probability, 11: 17:168–17:177, 2006. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1211.

Sturm:2008:TVM

[SS08] Anja Sturm and Jan Swart. Tightness of voter model interfaces. Electronic Communications in Probability, 13:16:165– 16:174, 2008. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1360.

Scheutzow:2021:CMC

[SS21] Michael Scheutzow and Dominik Schindler. Convergence of Markov chain transition probabilities. Electronic Communications in Probability, 26(??):1–13, 2021. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-26/issue-
none/Convergence-of-Markov-chain-transition-probabilities/
10.1214/21-ECP395.full.

Sanz-Sole:2015:ACS

[SSS15] Marta Sanz-Solé and André Süss. Absolute continuity for SPDEs with irregular fundamental solution. Electronic Communications in Probability, 20(??):14:1–14:11, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3831.

Schramm:1999:TCH

[ST99] Oded Schramm and Boris Tsirelson. Trees, not cubes: Hypercontractivity, cosiness, and noise stability. Electronic Communications in Probability, 4:6:39–6:49, 1999. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1004.

Sidoravicius:2017:NOD

[ST17] Vladas Sidoravicius and Laurent Tournier. Note on a onedimensional system of annihilating particles. Electronic Communications in Probability, 22(??):59:1–59:9, ???? 2017. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
euclid.ecp/1508292097.

Sousi:2020:CSM

[ST20] Perla Sousi and Sam Thomas. Chen–Stein method for the uncovered set of random walk on Z_{n}^{d} for d ≥ 3. Electronic Communications in Probability, 25(??):1–11, 2020. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-25/issue-
none/ChenStein-method-for-the-uncovered-set-of-random-
walk-on/10.1214/20-ECP331.full.

Stenflo:2008:PSL

[Ste08] Örjan Stenflo. Perfect sampling from the limit of deterministic products of stochastic matrices. Electronic Communications in Probability, 13:45:474–45:481, 2008. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1409.

Stenlund:2013:LLT

[Ste13] Mikko Stenlund. A local limit theorem for random walks in balanced environments. Electronic Communications in Probability, 18:19:1–19:13, 2013. CODEN ???? ISSN 1083-589X. URL http:

//ecp.ejpecp.org/article/view/2336.

Stucki:2013:CPG

[Stu13] Kaspar Stucki. Continuum percolation for Gibbs point processes. Electronic Communications in Probability, 18:67:1–67:10, 2013. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2837.

Stupfler:2016:WCK

[Stu16] Gilles Stupfler. On the weak convergence of the kernel density estimator in the uniform topology. Electronic Communications in Probability, 21(??):17:1–17:13, ???? 2016. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1456412897.

Subramanian:2012:DCP

[Sub12] Sneha Subramanian. On the distribution of critical points of a polynomial. Electronic Communications in Probability, 17:37:1– 37:9, 2012. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/2040.

Song:2004:SBG

[SV04] Renming Song and Zoran Vondracek. Sharp bounds for Green and jumping functions of subordinate killed Brownian motions. Electronic Communications in Probability, 9:11:96–11:105, 2004. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1114.

Song:2008:RBS

[SV08] Renming Song and Zoran Vondracek. On the relationship between subordinate killed and killed subordinate processes. Electronic Communications in Probability, 13:33:325–33:336, 2008. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1388.

Sayit:2011:AFM

[SV11a] Hasanjan Sayit and Frederi Viens. Arbitrage-free models in markets with transaction costs. Electronic Communications in Probability, 16:53:614–53:622, 2011. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1671.

Sen:2011:ACL

[SV11b] Arnab Sen and Balint Virag. Absolute continuity of the limiting eigenvalue distribution of the random Toeplitz matrix. Electronic Communications in Probability, 16:61:706–61:711, 2011. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1675.

Szabo:2016:SEC

[SV16] Réka Szabó and Daniel Valesin. From survival to extinction of the contact process by the removal of a single edge. Electronic Communications in Probability, 21(??):54:1–54:8, ???? 2016. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1472144932.

Soucaliuc:2002:NRB

[SW02] Florin Soucaliuc and Wendelin Werner. A note on reflecting Brownian motions. Electronic Communications in Probability, 7:12:117–12:122, 2002. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1053.

Savov:2010:RIL

[SW10] Mladen Savov and Matthias Winkel. Right inverses of Lévy processes: the excursion measure in the general case. Electronic Communications in Probability, 15:51:572–51:584, 2010. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1590.

Sun:2016:SUR

[SW16] Xin Sun and David B. Wilson. Sandpiles and unicycles on random planar maps. Electronic Communications in Probability, 21 (??):57:1–57:12, ???? 2016. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1473186614.

Swart:2001:DSW

[Swa01] Jan Swart. A 2-dimensional SDE whose solutions are not unique. Electronic Communications in Probability, 6:6:67–6:71, 2001. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1035.

Swart:2013:NIN

[Swa13] Jan Swart. Noninvadability implies noncoexistence for a class of cancellative systems. Electronic Communications in Probability, 18:38:1–38:12, 2013. CODEN ???? ISSN 1083-589X. URL http:

//ecp.ejpecp.org/article/view/2471.

Song:2018:UHD

[SXY18] Renming Song, Yimin Xiao, and Xiaochuan Yang. Uniform Hausdorff dimension result for the inverse images of stable Lévy processes. Electronic Communications in Probability, 23(??): 75:1–75:10, ???? 2018. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1539914641.

Seidler:2017:NCT

[SŽ17] Jan Seidler and František Žák. A note on continuous-time stochastic approximation in infinite dimensions. Electronic Communications in Probability, 22(??):36:1–36:13, ???? 2017. CO-DEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1498204833.

Sznitman:2012:ITR

[Szn12] Alain-Sol Sznitman. An isomorphism theorem for random interlacements. Electronic Communications in Probability, 17:9:1– 9:9, 2012. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1792.

Sznitman:2019:CVS

[Szn19] Alain-Sol Sznitman. On coupling and “vacant set level set” percolation. Electronic Communications in Probability, 24(??):20:1– 20:12, ???? 2019. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1554170646.

Takeda:2010:FKP

[Tak10] Masayoshi Takeda. Feynman–Kac penalisations of symmetric stable processes. Electronic Communications in Probability, 15: 4:32–4:43, 2010. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1524.

Tamas:2007:DDN

[Tam07] Móri Tamás. Degree distribution nearby the origin of a preferential attachment graph. Electronic Communications in Probability, 12:27:276–27:282, 2007. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1299.

Taniguchi:2006:QWF

[Tan06] Setsuo Taniguchi. On the quadratic Wiener functional associated with the Malliavin derivative of the square norm of Brownian sample path on interval. Electronic Communications in Probability, 11:1:1–1:10, 2006. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1174.

Tan:2017:EOD

[Tan17] Yan Shuo Tan. Energy optimization for distributions on the sphere and improvement to the Welch bounds. Electronic Communications in Probability, 22(??):43:1–43:12, ???? 2017. CO-DEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1502762749.

Tappe:2013:YWT

[Tap13] Stefan Tappe. The Yamada–Watanabe theorem for mild solutions to stochastic partial differential equations. Electronic Communications in Probability, 18:24:1–24:13, 2013. CODEN

????  ISSN 1083-589X. URL http://ecp.ejpecp.org/
article/view/2392.

Tappe:2015:FIM

[Tap15] Stefan Tappe. Flatness of invariant manifolds for stochastic partial differential equations driven by Lévy processes. Electronic Communications in Probability, 20(??):40:1–40:11, ???? 2015. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/3943.

Tassy:2010:RIG

[Tas10] Martin Tassy. Random interlacements on Galton–Watson trees. Electronic Communications in Probability, 15:50:562–50:571, 2010. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1586.

Tehranchi:2017:IGM

[Teh17] Michael R. Tehranchi. Inequalities for the Gaussian measure of convex sets. Electronic Communications in Probability, 22(??): 51:1–51:7, ???? 2017. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1507536392.

Telcs:2000:TPE

[Tel00] Andras Telcs. Transition probability estimates for reversible Markov chains. Electronic Communications in Probability, 5: 3:29–3:37, 2000. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1015.

Thalmaier:1998:SRH

[Tha98] Anton Thalmaier. Some remarks on the heat flow for functions and forms. Electronic Communications in Probability, 3:6:43– 6:49, 1998. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/992.

Thibault:2020:NAM

[Thi20] Pautrel Thibault. New asymptotics for the mean number of zeros of random trigonometric polynomials with strongly dependent Gaussian coefficients. Electronic Communications in Probability, 25(??):36:1–36:13, ???? 2020. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1588903422.

Thorisson:2016:CDF

[Tho16] Hermann Thorisson. Convergence in density in finite time windows and the Skorohod representation. Electronic Communications in Probability, 21(??):63:1–63:9, ???? 2016. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1473685927.

Timar:2004:TGF

[Tim04] Adam Timar. Tree and grid factors of general point processes. Electronic Communications in Probability, 9:6:53–6:59, 2004. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1073.

Timar:2019:OES

[Tim19] Ádám Timár. One-ended spanning trees in amenable unimodular graphs. Electronic Communications in Probability, 24(??):72:1– 72:12, ???? 2019. CODEN ???? ISSN 1083-589X. URL https:

//projecteuclid.org/euclid.ecp/1573528177.

Tkocz:2011:GMD

[Tko11] Tomasz Tkocz. Gaussian measures of dilations of convex rotationally symmetric sets in C^{n}. Electronic Communications in Probability, 16:5:38–5:49, 2011. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1599.

Tkocz:2013:NTP

[Tko13] Tomasz Tkocz. A note on the tensor product of two random unitary matrices. Electronic Communications in Probability, 18: 16:1–16:7, 2013. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/2484.

Tevzadze:2006:EME

[TM06] Revaz Tevzadze and Mikhael Mania. An exponential martingale equation. Electronic Communications in Probability, 11:22:206– 22:216, 2006. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1220.

Treilhard:2015:CIM

[TM15] John Treilhard and Abdol-Reza Mansouri. Concentration inequalities via Malliavin calculus with applications. Electronic Communications in Probability, 20(??):36:1–36:14, ???? 2015.

CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/3931.

Todino:2022:LBE

[Tod22] Anna Paola Todino. Limiting behavior for the excursion area of band-limited spherical random fields. Electronic Communications in Probability, 27(??):1–12, 2022. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-27/issue-
none/Limiting-behavior-for-the-excursion-area-of-band-
limited-spherical/10.1214/22-ECP488.full.

Toth:2013:CTM

[Tót13] Bálint Tóth. Comment on a theorem of M. Maxwell and M. Woodroofe. Electronic Communications in Probability, 18:13:1– 13:4, 2013. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/2366.

Tough:2022:SAP

[Tou22] Oliver Tough. Stochastic approximation of the paths of killed Markov processes conditioned on survival. Electronic Communications in Probability, 27(??):1–13, 2022. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-27/issue-
none/Stochastic-approximation-of-the-paths-of-killed-
Markov-processes-conditioned/10.1214/22-ECP475.full.

Trevisan:2013:BRM

[Tre13a] Dario Trevisan. BV-regularity for the Malliavin derivative of the maximum of the Wiener process. Electronic Communications in Probability, 18:29:1–29:9, 2013. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/2314.

Trevisan:2013:ZNL

[Tre13b] Dario Trevisan. Zero noise limits using local times. Electronic Communications in Probability, 18:31:1–31:7, 2013. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2587.

Treszczotko:2018:RWD

[Tre18] Łukasz Treszczotko. Random walks in doubly random scenery. Electronic Communications in Probability, 23(??):66:1–66:11, ???? 2018. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1537495700.

Trinh:2019:CLT

[Tri19] Khanh Duy Trinh. On central limit theorems in stochastic geometry for add-one cost stabilizing functionals. Electronic Communications in Probability, 24(??):76:1–76:15, ???? 2019. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
euclid.ecp/1576206122.

Tropp:2011:FIM

[Tro11] Joel Tropp. Freedman’s inequality for matrix martingales. Electronic Communications in Probability, 16:25:262–25:270, 2011. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1624.

Tsai:2018:SDA

[Tsa18] Li-Cheng Tsai. Stationary distributions of the Atlas model. Electronic Communications in Probability, 23(??):10:1–10:10, ???? 2018. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1519354833.

Tsirelson:2013:URT

[Tsi13] Boris Tsirelson. From uniform renewal theorem to uniform large and moderate deviations for renewal-reward processes. Electronic Communications in Probability, 18:52:1–52:13, 2013. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2719.

Turnbull:2020:CRG

[TT20] Shane Turnbull and Amanda Turner. Coexistence in a random growth model with competition. Electronic Communications in Probability, 25(??):26:1–26:14, ???? 2020. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1585274664.

Tucci:2011:API

[Tuc11] Gabriel Tucci. Asymptotic products of independent Gaussian random matrices with correlated entries. Electronic Communications in Probability, 16:33:353–33:364, 2011. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/1635.

Tudor:2009:HRS

[Tud09] Ciprian Tudor. Hsu–Robbins and Spitzer’s theorems for the variations of fractional Brownian motion. Electronic Communications in Probability, 14:28:278–28:289, 2009. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/1481.

Tugaut:2016:SPK

[Tug16] Julian Tugaut. A simple proof of a Kramers’ type law for selfstabilizing diffusions. Electronic Communications in Probability, 21(??):11:1–11:7, ???? 2016. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1455560035.

Tracy:2003:SDE

[TW03] Craig Tracy and Harold Widom. A system of differential equations for the Airy process. Electronic Communications in Probability, 8:10:93–10:98, 2003. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1074.

Tribe:2012:ODA

[TYZ12] Roger Tribe, Jonathan Yip, and Oleg Zaboronski. One dimensional annihilating and coalescing particle systems as extended Pfaffian point processes. Electronic Communications in Probability, 17:40:1–40:7, 2012. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/2133.

Tribe:2015:EOD

[TYZ15] Roger Tribe, Siu Kwan Yip, and Oleg Zaboronski. Erratum: One dimensional particle systems as extended Pfaffian point processes. Electronic Communications in Probability, 20(??):46:1– 46:2, ???? 2015. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/4302.

Uchiyama:2015:MNS

[Uch15] Kôhei Uchiyama. The mean number of sites visited by a random walk pinned at a distant point. Electronic Communications in Probability, 20(??):17:1–17:9, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/4027.

Uchiyama:2018:LHR

[Uch18] Kôhei Uchiyama. On the ladder heights of random walks attracted to stable laws of exponent 1. Electronic Communications in Probability, 23(??):23:1–23:12, ???? 2018. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1522375382.

Uemura:2007:EJT

[Uem07] Toshihiro Uemura. On an extension of jump-type symmetric Dirichlet forms. Electronic Communications in Probability, 12: 7:57–7:65, 2007. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1256.

Unterberger:2010:MES

[Unt10] Jeremie Unterberger. Moment estimates for solutions of linear stochastic differential equations driven by analytic fractional Brownian motion. Electronic Communications in Probability, 15: 37:411–37:417, 2010. CODEN ???? ISSN 1083-589X. URL http:

//ecp.ejpecp.org/article/view/1574.

Vadlamani:2006:GGU

[VA06] Sreekar Vadlamani and Robert Adler. Global geometry under isotropic Brownian flows. Electronic Communications in Probability, 11:19:182–19:192, 2006. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1212.

Valdivia:2017:ILG

[Val17] Arturo Valdivia. Information loss on Gaussian Volterra process. Electronic Communications in Probability, 22(??):60:1– 60:5, ???? 2017. CODEN ???? ISSN 1083-589X. URL https:/

/projecteuclid.org/euclid.ecp/1508896983.

Valdivia:2019:CFF

[Val19] Arturo Valdivia. Closed-form formulas for the distribution of the jumps of doubly-stochastic Poisson processes. Electronic Communications in Probability, 24(??):13:1–13:12, ???? 2019. CO-DEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1553133702.

Van:2007:SLL

[Van07] Thanh Le Van. On the strong law of large numbers for ddimensional arrays of random variables. Electronic Communications in Probability, 12:41:434–41:441, 2007. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/1331.

VanHandel:2008:DTN

[Van08] Ramon Van Handel. Discrete time nonlinear filters with informative observations are stable. Electronic Communications in Probability, 13:53:562–53:575, 2008. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1423.

vanBatenburg:2015:DII

[vB15] Wouter Cames van Batenburg. The dimension of the incipient infinite cluster. Electronic Communications in Probability, 20 (??):33:1–33:10, ???? 2015. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/3570.

vandenBerg:2012:SLC

[vdBC12] Jacob van den Berg and Rene Conijn. On the size of the largest cluster in 2d critical percolation. Electronic Communications in Probability, 17:58:1–58:13, 2012. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/2263.

vandenBerg:2013:GBS

[vdBC13] Jacob van den Berg and Rene Conijn. The gaps between the sizes of large clusters in 2d critical percolation. Electronic Communications in Probability, 18:92:1–92:9, 2013. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3065.

vandenBerg:2016:ENC

[vdBC16] J. van den Berg and R. P. Conijn. The expected number of critical percolation clusters intersecting a line segment. Electronic Communications in Probability, 21(??):28:1–28:10, ???? 2016. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1458324257.

vandenBerg:2020:LBP

[vdBD20] J. van den Berg and H. Don. A lower bound for point-to-point connection probabilities in critical percolation. Electronic Communications in Probability, 25(??):47:1–47:9, ???? 2020. CO-DEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1593569167.

vandenBerg:2010:ERD

[vdBHH10] Jacob van den Berg, Marcelo Hilário, and Alexander Holroyd. Escape of resources in a distributed clustering process. Electronic Communications in Probability, 15:40:442–40:448, 2010.

CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/1567.

vandenBerg:2007:SPI

[vdBJV07] Jacob van den Berg, Antal Jarai, and Balint Vagvolgyi. The size of a pond in 2d invasion percolation. Electronic Communications in Probability, 12:39:411–39:420, 2007. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1327.

vandenBerg:2012:PPB

[vdBKN12] Jacob van den Berg, Demeter Kiss, and Pierre Nolin. A percolation process on the binary tree where large finite clusters are frozen. Electronic Communications in Probability, 17:2:1– 2:11, 2012. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1694.

vandenBerg:2017:BRB

[vdBN17] Jacob van den Berg and Pierre Nolin. Boundary rules and breaking of self-organized criticality in 2D frozen percolation. Electronic Communications in Probability, 22(??):65:1–65:15, ???? 2017. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1511406073.

vanderHofstad:2009:LLT

[vdHKM09] Remco van der Hofstad, Wouter Kager, and Tobias Müller. A local limit theorem for the critical random graph. Electronic Communications in Probability, 14:12:122–12:131, 2009. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1451.

Venker:2013:PSR

[Ven13] Martin Venker. Particle systems with repulsion exponent β and random matrices. Electronic Communications in Probability, 18: 83:1–83:12, 2013. CODEN ???? ISSN 1083-589X. URL http:/

/ecp.ejpecp.org/article/view/2864.

Vidmar:2014:NTF

[Vid14] Matija Vidmar. A note on the times of first passage for ‘nearly right-continuous’ random walks. Electronic Communications in Probability, 19(??):75:1–75:7, ???? 2014. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3735.

Valesin:2013:SAA

[VM13] Daniel Valesin and Thomas Mountford. Supercriticality of an annealed approximation of Boolean networks. Electronic Communications in Probability, 18:32:1–32:12, 2013. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/2479.

Vollering:2020:MPR

[Völ20] Florian Völlering. Markov process representation of semigroups whose generators include negative rates. Electronic Communications in Probability, 25(??):1–7, 2020. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-25/issue-
none/Markov-process-representation-of-semigroups-whose-
generators-include-negative-rates/10.1214/20-ECP349.
full.

Vovk:2008:CTT

[Vov08] Vladimir Vovk. Continuous-time trading and the emergence of volatility. Electronic Communications in Probability, 13:32:319– 32:324, 2008. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1383.

Vandenberg-Rodes:2010:LTP

[VR10] Alexander Vandenberg-Rodes. A limit theorem for particle current in the symmetric exclusion process. Electronic Communications in Probability, 15:22:240–22:252, 2010. CODEN

????  ISSN 1083-589X. URL http://ecp.ejpecp.org/
article/view/1550.

vonSoosten:2019:RCW

[vSW19] Per von Soosten and Simone Warzel. Random characteristics for Wigner matrices. Electronic Communications in Probability, 24(??):75:1–75:12, ???? 2019. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1576119990.

Vakeroudis:2012:SPW

[VY12a] Stavros Vakeroudis and Marc Yor. A scaling proof for Walsh’s Brownian motion extended arc-sine law. Electronic Communications in Probability, 17:63:1–63:9, 2012. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/2319.

Vakeroudis:2012:SID

[VY12b] Stavros Vakeroudis and Marc Yor. Some infinite divisibility properties of the reciprocal of planar Brownian motion exit time from a cone. Electronic Communications in Probability, 17:23:1–23:9, 2012. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/2090.

Vysotsky:2021:WRF

[Vys21] Vladislav Vysotsky. When is the rate function of a random vector strictly convex? Electronic Communications in Probability, 26 (??):1–11, 2021. CODEN ???? ISSN 1083-589X.

vanZanten:2002:COM

[vZ02] Harry van Zanten. Continuous Ocone martingales as weak limits of rescaled martingales. Electronic Communications in Probability, 7:21:215–21:222, 2002. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1062. See remark and

counterexample [Pec04].

vanZanten:2008:REG

[vZ08] Harry van Zanten. A remark on the equivalence of Gaussian processes. Electronic Communications in Probability, 13:6:54– 6:59, 2008. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1348.

VanNeerven:2011:MIS

[VZ11] Jan Van Neerven and Jiahui Zhu. A maximal inequality for stochastic convolutions in 2-smooth Banach spaces. Electronic Communications in Probability, 16:60:689–60:705, 2011. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1677.

Wagner:2016:BHP

[Wag16] Vanja Wagner. Boundary Harnack principle for the absolute value of a one-dimensional subordinate Brownian motion killed at 0. Electronic Communications in Probability, 21(??):84:1– 84:12, ???? 2016. CODEN ???? ISSN 1083-589X. URL http:/

/projecteuclid.org/euclid.ecp/1481684430.

Wainrib:2013:DAP

[Wai13] Gilles Wainrib. Double averaging principle for periodically forced stochastic slow-fast systems. Electronic Communications in Probability, 18:51:1–51:12, 2013. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1975.

Wang:2002:SCC

[Wan02] Hao Wang. State classification for a class of interacting superprocesses with location dependent branching. Electronic Communications in Probability, 7:16:157–16:167, 2002. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1057.

Wang:2009:FEO

[Wan09] Jian Wang. First eigenvalue of one-dimensional diffusion processes. Electronic Communications in Probability, 14:23:232– 23:244, 2009. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1464.

Wang:2014:CIG

[Wan14] Neng-Yi Wang. Concentration inequalities for Gibbs sampling under d_{l_{2}}-metric. Electronic Communications in Probability, 19 (??):63:1–63:11, ???? 2014. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/3502.

Wang:2015:HDB

[Wan15] Minmin Wang. Height and diameter of Brownian tree. Electronic Communications in Probability, 20(??):88:1–88:15, ???? 2015. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/4193.

Wang:2017:CRR

[Wan17] Neng-Yi Wang. Convergence rates of the random scan Gibbs sampler under the Dobrushin’s uniqueness condition. Electronic Communications in Probability, 22(??):56:1–56:7, ???? 2017. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1507860213.

Wang:2018:OSM

[Wan18] Yanhui Wang. Order statistics of the moduli of the eigenvalues of product random matrices from polynomial ensembles. Electronic Communications in Probability, 23(??):21:1–21:14, ???? 2018. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1522375377.

Wang:2021:CMB

[Wan21] Minmin Wang. k-cut model for the Brownian continuum random tree. Electronic Communications in Probability, 26(??):1– 11, 2021. CODEN ???? ISSN 1083-589X.

Wang:2022:DDS

[Wan22] Feng-Yu Wang. Distribution dependent SDEs for Navier–Stokes type equations. Electronic Communications in Probability, 27 (??):1–12, 2022. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-27/issue-none/Distribution-dependent-
SDEs-for-Navier--Stokes-type-equations/10.1214/22-ECP479.
full.

Warren:1999:RDA

[War99] Jon Warren. On a result of David Aldous concerning the trees in a conditioned excursion. Electronic Communications in Probability, 4:4:25–4:29, 1999. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1002.

Wastlund:2008:RMP

[Wäs08] Johan Wästlund. Random matching problems on the complete graph. Electronic Communications in Probability, 13:25:258– 25:265, 2008. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1372.

Wastlund:2009:EPL

[Wäs09] Johan Wästlund. An easy proof of the ζ(2) limit in the random assignment problem. Electronic Communications in Probability, 14:26:261–26:269, 2009. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1475.

Watbled:2012:SAF

[Wat12] Frédérique Watbled. Sharp asymptotics for the free energy of 1+1 dimensional directed polymers in an infinitely divisible environment. Electronic Communications in Probability, 17:53:1– 53:9, 2012. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/2221.

Wu:2019:HMN

[WCS19] Zhixin Wu, Arijit Chakrabarty, and Gennady Samorodnitsky. High minima of non-smooth Gaussian processes. Electronic Communications in Probability, 24(??):53:1–53:12, ???? 2019. CO-DEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1568253715.

Weerasinghe:2006:CSG

[Wee06] Ananda Weerasinghe. A controller and a stopper game with degenerate variance control. Electronic Communications in Probability, 11:9:89–9:99, 2006. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1202.

Weininger:2003:PCI

[Wei03] Nicholas Weininger. Positive correlation for increasing events with disjoint dependencies does not imply positive correlation for all increasing events. Electronic Communications in Probability, 8:11:99–11:101, 2003. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1078.

Werner:1996:BDE

[Wer96] Wendelin Werner. Bounds for disconnection exponents. Electronic Communications in Probability, 1:4:19–4:28, 1996. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/974.

Wang:2019:CCM

[WH19] Xiaoqiang Wang and Chunmao Huang. Convergence of complex martingale for a branching random walk in a time random environment. Electronic Communications in Probability, 24(??): 41:1–41:14, ???? 2019. CODEN ???? ISSN 1083-589X. URL

https://projecteuclid.org/euclid.ecp/1562119369.

Wilson:2003:MTR

[Wil03] David Wilson. Mixing time of the Rudvalis shuffle. Electronic Communications in Probability, 8:8:77–8:85, 2003. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1071.

Windisch:2008:RWD

[Win08] David Windisch. Random walk on a discrete torus and random interlacements. Electronic Communications in Probability, 13: 14:140–14:150, 2008. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1359.

Winkel:2020:DSI

[Win20] Matthias Winkel. Diffusions on a space of interval partitions: construction from Bertoin’s BES_{0}(d), d ∈ (0, 1). Electronic Communications in Probability, 25(??):1–13, 2020. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
journals/electronic-communications-in-probability/volume-
25/issue-none/Diffusions-on-a-space-of-interval-partitions-
-constructionfrom-Bertoins/10.1214/20-ECP355.full.

Wojtylak:2012:CSR

[Woj12] Michał Wojtylak. On a class of H-selfadjont random matrices with one eigenvalue of nonpositive type. Electronic Communications in Probability, 17:45:1–45:14, 2012. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/2148.

Wojtylak:2014:SPC

[WP14] Michał Wojtylak and Patryk Pagacz. On the spectral properties of a class of H-selfadjoint random matrices and the underlying combinatorics. Electronic Communications in Probability, 19: 7:1–7:14, 2014. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/3066.

Wu:2008:LDP

[WY08] Liming Wu and Nian Yao. Large deviation principles for Markov processes via Phi–Sobolev inequalities. Electronic Communications in Probability, 13:2:10–2:23, 2008. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/1342.

Wang:2013:WAF

[WYY13] Zhi Wang, Litan Yan, and Xianye Yu. Weak approximation of the fractional Brownian sheet from random walks. Electronic Communications in Probability, 18:90:1–90:13, 2013. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2878.

Wu:2019:STM

[WZ19] Wei Wu and Ofer Zeitouni. Subsequential tightness of the maximum of two dimensional Ginzburg–Landau fields. Electronic Communications in Probability, 24(??):19:1–19:12, ???? 2019. CODEN ???? ISSN 1083-589X. URL https://projecteuclid.

org/euclid.ecp/1553306561.

Xiao:2020:PIH

[XFT20] Lishun Xiao, Shengjun Fan, and Dejian Tian. Probabilistic interpretation of HJB equations by the representation theorem for generators of BSDEs. Electronic Communications in Probability, 25(??):30:1–30:10, ???? 2020. CODEN ???? ISSN 1083-589X.

URL https://projecteuclid.org/euclid.ecp/1586419347.

Xie:2020:FWL

[Xie20] Yongjia Xie. Functional weak limit of random walks in cooling random environment. Electronic Communications in Probability, 25(??):1–14, 2020. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-25/issue-
none/Functional-weak-limit-of-random-walks-in-cooling-
random-environment/10.1214/20-ECP360.full.

Xiong:2004:LTB

[Xio04] Jie Xiong. Long-term behavior for superprocesses over a stochastic flow. Electronic Communications in Probability, 9:5:36–5:52, 2004. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1081.

Yadin:2009:REM

[Yad09] Ariel Yadin. Rate of escape of the mixer chain. Electronic Communications in Probability, 14:35:347–35:357, 2009. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1474.

Yang:2006:SPH

[Yan06] Ming Yang. A short proof of the Hausdorff dimension formula for Lévy processes. Electronic Communications in Probability, 11:23:217–23:219, 2006. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1199.

Yang:2007:TMP

[Yan07] Ming Yang. On a theorem in multi-parameter potential theory. Electronic Communications in Probability, 12:26:267–26:275, 2007. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1293.

Yang:2022:REB

[Yan22] Danyu Yang. A remainder estimate for branched rough differential equations. Electronic Communications in Probability, 27 (??):1–12, 2022. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/journals/electronic-communications-
in-probability/volume-27/issue-none/A-remainder-estimate-
for-branched-rough-differential-equations/10.1214/22-
ECP473.full.

Yao:2014:LLN

[Yao14] Chang-Long Yao. Law of large numbers for critical first-passage percolation on the triangular lattice. Electronic Communications in Probability, 19:18:1–18:14, 2014. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/3268.

Yaskov:2014:LBS

[Yas14] Pavel Yaskov. Lower bounds on the smallest eigenvalue of a sample covariance matrix. Electronic Communications in Probability, 19(??):83:1–83:10, ???? 2014. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/3807.

Yaskov:2015:SLB

[Yas15] Pavel Yaskov. Sharp lower bounds on the least singular value of a random matrix without the fourth moment condition. Electronic Communications in Probability, 20(??):44:1–44:9, ???? 2015. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.

org/article/view/4089.

Yaskov:2016:NSC

[Yas16] Pavel Yaskov. Necessary and sufficient conditions for the Marchenko–Pastur theorem. Electronic Communications in Probability, 21(??):73:1–73:8, ???? 2016. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/euclid.ecp/
1477327674.

Yang:2013:SPG

[YE13] Zhe Yang and Robert Elliott. Some properties of generalized anticipated backward stochastic differential equations. Electronic Communications in Probability, 18:63:1–63:10, 2013. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2415.

Yin:2015:LDE

[Yin15] Mei Yin. Large deviations and exact asymptotics for constrained exponential random graphs. Electronic Communications in Probability, 20(??):56:1–56:14, ???? 2015. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/4010.

Yan:2015:ARP

[YLW15] Litan Yan, Yumiao Li, and Di Wu. Approximating the Rosenblatt process by multiple Wiener integrals. Electronic Communications in Probability, 20(??):11:1–11:16, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/3517.

Yor:2015:GMW

[Yor15] Marc Yor. A Gaussian martingale which is the sum of two independent Gaussian non-semimartingales. Electronic Communications in Probability, 20(??):70:1–70:5, ???? 2015. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/4034.

Yang:2016:CCC

[YRE16] Zhe Yang, Dimbinirina Ramarimbahoaka, and Robert J. Elliott. Comparison and converse comparison theorems for backward stochastic differential equations with Markov chain noise. Electronic Communications in Probability, 21(??):25:1–25:10, ???? 2016. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1457617916.

Yukich:2008:LTM

[Yuk08] Joseph Yukich. Limit theorems for multi-dimensional random quantizers. Electronic Communications in Probability, 13: 48:507–48:517, 2008. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1418.

Yen:2013:IVM

[YY13] Ju-Yi Yen and Marc Yor. Illustration of various methods for solving partly Skorokhod’s embedding problem. Electronic Communications in Probability, 18:48:1–48:5, 2013. CODEN ????

ISSN 1083-589X. URL http://ecp.ejpecp.org/article/
view/2178.

Yan:2018:HID

[YY18] Litan Yan and Xiuwei Yin. Harnack inequality and derivative formula for stochastic heat equation with fractional noise. Electronic Communications in Probability, 23(??):35:1–35:11, ???? 2018. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1528358640.

Zerner:2002:NBL

[Zer02] Martin Zerner. A non-ballistic law of large numbers for random walks in I.I.D. random environment. Electronic Communications in Probability, 7:19:191–19:197, 2002. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1060.

Zerner:2006:RTE

[Zer06] Martin Zerner. Recurrence and transience of excited random walks on Z^{d} and strips. Electronic Communications in Probability, 11:12:118–12:128, 2006. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1200.

Zerner:2007:ZOL

[Zer07] Martin Zerner. The zero-one law for planar random walks in i.i.d. random environments revisited. Electronic Communications in Probability, 12:32:326–32:335, 2007. CODEN ???? ISSN 1083-

589X. URL http://ecp.ejpecp.org/article/view/1314.

Zhao:2012:UAE

[Zha12] James Zhao. Universality of asymptotically Ewens measures on partitions. Electronic Communications in Probability, 17:16:1– 16:11, 2012. CODEN ???? ISSN 1083-589X. URL http://ecp.

ejpecp.org/article/view/1956.

Zhang:2020:SFP

[Zha20a] Hua Zhang. Strong Feller property and continuous dependence on initial data for one-dimensional stochastic differential equations with Hölder continuous coefficients. Electronic Communications in Probability, 25(??):3:1–3:10, ???? 2020. CODEN ????

ISSN 1083-589X. URL https://projecteuclid.org/euclid.
ecp/1578387617.

Zhang:2020:OEC

[Zha20b] Lingfu Zhang. Optimal exponent for coalescence of finite geodesics in exponential last passage percolation. Electronic Communications in Probability, 25(??):1–14, 2020. CODEN

???? ISSN 1083-589X. URL https://projecteuclid.org/
journals/electronic-communications-in-probability/volume-
25/issue-none/Optimal-exponent-for-coalescence-of-finite-
geodesics-in-exponential-last/10.1214/20-ECP354.full.

Zhang:2022:LAD

[Zha22] Xincheng Zhang. The logarithmic anti-derivative of the baikrains distribution satisfies the KP equation. Electronic Communications in Probability, 27(??):1–12, 2022. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-27/issue-
none/The-logarithmic-anti-derivative-of-the-Baik-Rains-
distribution-satisfies/10.1214/22-ECP469.full.

Zhou:2010:ASF

[Zho10] Xiaowen Zhou. Almost sure finiteness for the total occupation time of an (d, α, β)-superprocess. Electronic Communications in Probability, 15:3:22–3:31, 2010. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1523.

Zhou:2014:CDF

[Zho14] Xiaowen Zhou. On criteria of disconnectedness of Λ–Fleming– Viot support. Electronic Communications in Probability, 19: 53:1–53:16, 2014. CODEN ???? ISSN 1083-589X. URL http:/

/ecp.ejpecp.org/article/view/3208.

Zhou:2020:HL

[Zho20] Zhengye Zhou. Hydrodynamic limit for a d-dimensional open symmetric exclusion process. Electronic Communications in Probability, 25(??):1–8, 2020. CODEN ???? ISSN

1083-589X. URL https://projecteuclid.org/journals/
electronic-communications-in-probability/volume-25/issue-
none/Hydrodynamic-limit-for-a-d-dimensional-open-symmetric-
exclusion-process/10.1214/20-ECP350.full.

Zhu:2014:LDE

[Zhu14] Lingjiong Zhu. On the large deviations for Engel’s, Sylvester’s series and Cantor’s products. Electronic Communications in Probability, 19:2:1–2:9, 2014. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/3194.

Zaidi:2003:SLS

[ZN03] Noureddine Zaïdi and David Nualart. Smoothness of the law of the supremum of the fractional Brownian motion. Electronic Communications in Probability, 8:12:102–12:111, 2003. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1079.